LCIO

"2.7.4"

Generated by Doxygen 1.8.6

Fri Dec 2 2016 11:44:41

Contents

1 <a href="http://lcio.desy.de">LCIO</a> (v02-07-04) 1 1.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Deprecated List 3

3 Namespace Index 5 3.1 Namespace List . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

4 Hierarchical Index 7 4.1 Class Hierarchy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

5 Class Index 13 5.1 Class List . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

6 Namespace Documentation 21

6.1 DATA Namespace Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

6.1.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 6.2 EVENT Namespace Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

6.2.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 6.2.2 Typedef Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

6.2.2.1 CalorimeterHitVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 6.2.2.2 ClusterVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 6.2.2.3 DoubleVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 6.2.2.4 FloatVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 6.2.2.5 IntVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 6.2.2.6 LCObjectVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 6.2.2.7 MCParticleVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 6.2.2.8 ParticleIDVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 6.2.2.9 RawCalorimeterHitVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 6.2.2.10 ReconstructedParticleVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 6.2.2.11 ShortVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 6.2.2.12 StringVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 6.2.2.13 TrackerHitVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 6.2.2.14 TrackStateVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 6.2.2.15 TrackVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 6.2.2.16 VertexVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

6.3 IMPL Namespace Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

6.3.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 6.4 IO Namespace Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

6.4.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 6.5 IOIMPL Namespace Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

6.5.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 6.6 lcio Namespace Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

6.6.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 6.7 lcrtrel Namespace Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

6.7.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 6.7.2 Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

6.7.2.1 set_relation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

6.8 SIO Namespace Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

6.8.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 6.9 UTIL Namespace Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

6.9.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 6.9.2 Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

6.9.2.1 CellIDEncoder_cellID1Bit< EVENT::SimCalorimeterHit > . . . . . . . . . . . . 38 6.9.2.2 header . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 6.9.2.3 lctypename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 6.9.2.4 lctypename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 6.9.2.5 lctypename< EVENT::LCObject > . . . . . . . . . . . . . . . . . . . . . . . . 39

7 Class Documentation 41

7.1 IMPL::AccessChecked Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

7.1.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 7.2 lcrtrel_helper::BiDirectional< From, To > Struct Template Reference . . . . . . . . . . . . . . . . 42

7.2.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 7.3 UTIL::BitField64 Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

7.3.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 7.3.2 Constructor & Destructor Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

7.3.2.1 BitField64 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 7.3.3 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

7.3.3.1 init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

7.4 UTIL::BitFieldValue Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 7.4.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.4.2 Constructor & Destructor Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

7.4.2.1 BitFieldValue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 7.4.3 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

7.4.3.1 name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

7.5 UTIL::BitSet32 Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

7.5.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 7.6 UTIL::LCTime::CalendarTime Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

7.6.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 7.7 CalibrationConstant Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

7.7.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 7.8 EVENT::CalorimeterHit Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

7.8.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 7.8.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

7.8.2.1 getCellID1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 7.8.2.2 getEnergyError . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 7.8.2.3 getPosition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 7.8.2.4 getRawHit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 7.8.2.5 getTime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 7.8.2.6 getType . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

7.9 IMPL::CalorimeterHitImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 7.9.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 7.9.2 Constructor & Destructor Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

7.9.2.1 ∼CalorimeterHitImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 7.9.3 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

7.9.3.1 getCellID1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 7.9.3.2 getEnergyError . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 7.9.3.3 getPosition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 7.9.3.4 getRawHit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 7.9.3.5 getTime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 7.9.3.6 getType . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 7.9.3.7 setCellID1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 7.9.3.8 setEnergyError . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 7.9.3.9 setPosition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

7.10 IOIMPL::CalorimeterHitIOImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

7.10.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 7.11 lcrtrel_helper::can_call_ext< B > Struct Template Reference . . . . . . . . . . . . . . . . . . . . 55 7.12 lcrtrel_helper::can_call_ext< true > Struct Template Reference . . . . . . . . . . . . . . . . . . . 55 7.13 UTIL::CellIDDecoder< T > Class Template Reference . . . . . . . . . . . . . . . . . . . . . . . . 56

7.13.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 7.13.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.13.2.1 operator() . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 7.13.2.2 setDefaultEncoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

7.14 UTIL::CellIDEncoder< T > Class Template Reference . . . . . . . . . . . . . . . . . . . . . . . . 57

7.14.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 7.15 EVENT::Cluster Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

7.15.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 7.15.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60

7.15.2.1 getCalorimeterHits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 7.15.2.2 getDirectionError . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 7.15.2.3 getIPhi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 7.15.2.4 getITheta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 7.15.2.5 getParticleIDs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 7.15.2.6 getSubdetectorEnergies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 7.15.2.7 getType . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

7.16 IMPL::ClusterImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 7.16.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 7.16.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

7.16.2.1 getParticleIDs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 7.16.2.2 getSubdetectorEnergies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 7.16.2.3 getType . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 7.16.2.4 subdetectorEnergies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

7.17 IOIMPL::ClusterIOImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

7.17.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 7.18 UTIL::CollectionParameterMap Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

7.18.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 7.19 lcrtrel_helper::CreationPtrInit< T > Struct Template Reference . . . . . . . . . . . . . . . . . . . 66

7.19.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 7.20 EVENT::DataNotAvailableException Class Reference . . . . . . . . . . . . . . . . . . . . . . . . 66

7.20.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 7.21 lcrtrel_helper::DeleteElements< T > Struct Template Reference . . . . . . . . . . . . . . . . . . 67

7.21.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 7.22 lcrtrel_helper::DeletePtr< T > Struct Template Reference . . . . . . . . . . . . . . . . . . . . . . 68

7.22.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 7.23 IO::EndOfDataException Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

7.23.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 7.24 EVENT::EventException Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

7.24.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 7.25 EVENT::Exception Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

7.25.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 7.26 lcrtrel_helper::FromRelation< U > Struct Template Reference . . . . . . . . . . . . . . . . . . . . 70

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.26.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 7.27 HEPEVTIMPL::HEPEVT Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

7.27.1 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

7.27.1.1 fromHepEvt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

7.28 hepevt1_t Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 7.29 hepevt_t Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 7.30 IO::ILCFactory Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

7.30.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 7.30.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

7.30.2.1 createLCWriter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

7.31 UTIL::ILDCellID0 Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 7.31.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 7.31.2 Member Data Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

7.31.2.1 encoder_string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 7.31.2.2 module . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 7.31.2.3 sensor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 7.31.2.4 subdet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

7.32 UTIL::ILDCellIDEncoder< T > Class Template Reference . . . . . . . . . . . . . . . . . . . . . . 74 7.32.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 7.32.2 Constructor & Destructor Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

7.32.2.1 ILDCellIDEncoder . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

7.33 UTIL::ILDDetID Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

7.33.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 7.34 UTIL::ILDTrkHitQualityBit Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76

7.34.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 7.35 UTIL::ILDTrkHitTypeBit Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76

7.35.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 7.36 Index Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 7.37 SomeClass::Index Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 7.38 UTIL::IndexMap Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

7.38.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 7.38.2 Constructor & Destructor Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

7.38.2.1 IndexMap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 7.38.3 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

7.38.3.1 decode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 7.38.3.2 encode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

7.39 IO::IOException Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

7.39.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 7.40 lcrtrel::LC1To1Relation< U, From, To > Struct Template Reference . . . . . . . . . . . . . . . . . 80

7.40.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.41 lcrtrel::LC1ToNRelation< U, From, To > Struct Template Reference . . . . . . . . . . . . . . . . . 80

7.41.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 7.42 lcrtrel_helper::LCBaseLinkContainerTraits< U, T, I, D, b > Struct Template Reference . . . . . . . 80

7.42.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 7.43 lcrtrel_helper::LCBaseLinkTraits< U, T, I, D, b > Struct Template Reference . . . . . . . . . . . . 81

7.43.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 7.44 lcrtrel_helper::LCBaseTraits< U, T, I, D, b > Struct Template Reference . . . . . . . . . . . . . . . 82

7.44.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 7.45 EVENT::LCCollection Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

7.45.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 7.45.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

7.45.2.1 addElement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 7.45.2.2 getFlag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 7.45.2.3 getNumberOfElements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 7.45.2.4 isDefault . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 7.45.2.5 isSubset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 7.45.2.6 isTransient . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 7.45.2.7 parameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 7.45.2.8 removeElementAt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 7.45.2.9 setFlag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86

7.46 IOIMPL::LCCollectionIOVec Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86

7.46.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 7.47 IMPL::LCCollectionVec Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

7.47.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 7.47.2 Constructor & Destructor Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

7.47.2.1 LCCollectionVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 7.47.3 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

7.47.3.1 addElement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 7.47.3.2 getFlag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 7.47.3.3 isDefault . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 7.47.3.4 isSubset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 7.47.3.5 isTransient . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 7.47.3.6 removeElementAt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 7.47.3.7 setDefault . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 7.47.3.8 setSubset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 7.47.3.9 setTransient . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

7.48 EVENT::LCEvent Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 7.48.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 7.48.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

7.48.2.1 addCollection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.48.2.2 getCollection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 7.48.2.3 getTimeStamp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 7.48.2.4 removeCollection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 7.48.2.5 takeCollection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

7.49 IMPL::LCEventImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 7.49.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 7.49.2 Constructor & Destructor Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . 96

7.49.2.1 ∼LCEventImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 7.49.3 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96

7.49.3.1 addCollection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 7.49.3.2 getCollection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 7.49.3.3 getParameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 7.49.3.4 removeCollection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 7.49.3.5 takeCollection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 7.49.3.6 validateCollectionName . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97

7.50 IOIMPL::LCEventIOImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97

7.50.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 7.51 IO::LCEventListener Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98

7.51.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 7.51.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

7.51.2.1 modifyEvent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 7.51.2.2 processEvent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

7.52 lcrtrel::LCExtension< U, T > Struct Template Reference . . . . . . . . . . . . . . . . . . . . . . . 99

7.52.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 7.53 lcrtrel::LCExtensionList< U, T > Class Template Reference . . . . . . . . . . . . . . . . . . . . . 100

7.53.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 7.54 lcrtrel::LCExtensionVector< U, T > Class Template Reference . . . . . . . . . . . . . . . . . . . 100

7.54.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 7.55 IOIMPL::LCFactory Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

7.55.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 7.55.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

7.55.2.1 createLCReader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 7.55.2.2 getInstance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

7.56 UTIL::LCFixedObject< NINT, NFLOAT, NDOUBLE > Class Template Reference . . . . . . . . . . 102 7.56.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 7.56.2 Constructor & Destructor Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

7.56.2.1 LCFixedObject . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 7.56.3 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

7.56.3.1 obj . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

7.57 EVENT::LCFlag Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.57.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 7.58 IMPL::LCFlagImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

7.58.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 7.59 lcrtrel::LCFloatExtension< U > Struct Template Reference . . . . . . . . . . . . . . . . . . . . . 105 7.60 EVENT::LCFloatVec Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106

7.60.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 7.61 UTIL::LCFourVector< TT > Class Template Reference . . . . . . . . . . . . . . . . . . . . . . . 107

7.61.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 7.61.2 Constructor & Destructor Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . 108

7.61.2.1 LCFourVector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 7.61.2.2 LCFourVector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108

7.61.3 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 7.61.3.1 lcObj . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 7.61.3.2 operator-> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108

7.62 EVENT::LCGenericObject Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 7.62.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 7.62.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110

7.62.2.1 getDataDescription . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 7.62.2.2 getTypeName . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110

7.63 IMPL::LCGenericObjectImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 7.63.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 7.63.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112

7.63.2.1 getDataDescription . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112

7.64 IOIMPL::LCGenericObjectIOImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . 112

7.64.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 7.65 lcrtrel::LCIntExtension< U > Struct Template Reference . . . . . . . . . . . . . . . . . . . . . . . 113

7.65.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 7.66 EVENT::LCIntVec Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114

7.66.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 7.67 EVENT::LCIO Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115

7.67.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 7.68 UTIL::LCIO_LONG< T > Class Template Reference . . . . . . . . . . . . . . . . . . . . . . . . 117 7.69 UTIL::lcio_short< T > Class Template Reference . . . . . . . . . . . . . . . . . . . . . . . . . . 117 7.70 IMPL::LCIOExceptionHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117

7.70.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 7.71 SIO::LCIORandomAccess Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118

7.71.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 7.72 SIO::LCIORandomAccessMgr Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

7.72.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 7.72.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.72.2.1 createFromEventMap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 7.72.2.2 getPosition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121

7.73 UTIL::LCIterator< T > Class Template Reference . . . . . . . . . . . . . . . . . . . . . . . . . . 121 7.73.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 7.73.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122

7.73.2.1 operator() . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122

7.74 lcrtrel::LCNToNRelation< U, From, To > Struct Template Reference . . . . . . . . . . . . . . . . 122

7.74.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 7.75 EVENT::LCObject Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123

7.75.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 7.75.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124

7.75.2.1 clone . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124

7.76 UTIL::LCObjectHandle< T > Class Template Reference . . . . . . . . . . . . . . . . . . . . . . . 124

7.76.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 7.77 lcrtrel::LCOwnedExtension< U, T > Class Template Reference . . . . . . . . . . . . . . . . . . . 125

7.77.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 7.78 lcrtrel::LCOwnedExtensionList< U, T > Class Template Reference . . . . . . . . . . . . . . . . . 125

7.78.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 7.79 lcrtrel::LCOwnedExtensionVector< U, T > Class Template Reference . . . . . . . . . . . . . . . . 126

7.79.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 7.80 EVENT::LCParameters Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126

7.80.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 7.80.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128

7.80.2.1 getFloatVals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 7.80.2.2 getIntVals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 7.80.2.3 getStringVals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128

7.81 IMPL::LCParametersImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 7.81.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 7.81.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130

7.81.2.1 getFloatVals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 7.81.2.2 getIntVals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 7.81.2.3 getStringVals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130

7.82 IO::LCReader Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 7.82.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 7.82.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132

7.82.2.1 close . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 7.82.2.2 getEvents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 7.82.2.3 getNumberOfEvents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 7.82.2.4 getNumberOfRuns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 7.82.2.5 getRuns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.82.2.6 open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 7.82.2.7 open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 7.82.2.8 readEvent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 7.82.2.9 readEvent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 7.82.2.10 readNextEvent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 7.82.2.11 readNextEvent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 7.82.2.12 readNextRunHeader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 7.82.2.13 readNextRunHeader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 7.82.2.14 readRunHeader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 7.82.2.15 readRunHeader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 7.82.2.16 readStream . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 7.82.2.17 readStream . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 7.82.2.18 setReadCollectionNames . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 7.82.2.19 skipNEvents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136

7.82.3 Member Data Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136

7.82.3.1 directAccess . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136

7.83 EVENT::LCRelation Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136

7.83.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 7.84 IMPL::LCRelationImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137

7.84.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 7.85 IOIMPL::LCRelationIOImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139

7.85.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 7.86 UTIL::LCRelationNavigator Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139

7.86.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 7.86.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141

7.86.2.1 addRelation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 7.86.2.2 createLCCollection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 7.86.2.3 getRelatedFromObjects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 7.86.2.4 getRelatedFromWeights . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 7.86.2.5 getRelatedToObjects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 7.86.2.6 getRelatedToWeights . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141

7.87 lcrtrel::LCRTRelations Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 7.87.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 7.87.2 Friends And Related Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . 145

7.87.2.1 set_relation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145

7.88 EVENT::LCRunHeader Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145

7.88.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 7.89 IMPL::LCRunHeaderImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146

7.89.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 7.89.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.89.2.1 getDescription . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147

7.90 IOIMPL::LCRunHeaderIOImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . 148

7.90.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 7.91 IO::LCRunListener Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148

7.91.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 7.91.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149

7.91.2.1 processRunHeader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149

7.92 SIO::LCSIO Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 7.92.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 7.92.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150

7.92.2.1 getValidSIOName . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 7.92.2.2 read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 7.92.2.3 seekStream . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 7.92.2.4 write . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 7.92.2.5 write . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 7.92.2.6 write . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151

7.93 UTIL::LCSplitWriter Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 7.93.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 7.93.2 Constructor & Destructor Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . 153

7.93.2.1 LCSplitWriter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 7.93.3 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153

7.93.3.1 close . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 7.93.3.2 flush . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 7.93.3.3 getCountingString . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 7.93.3.4 getFilename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 7.93.3.5 open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 7.93.3.6 open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 7.93.3.7 setBaseFilename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 7.93.3.8 setCompressionLevel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 7.93.3.9 writeEvent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 7.93.3.10 writeRunHeader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154

7.94 UTIL::LCStdHepRdr Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 7.94.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 7.94.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156

7.94.2.1 getNumberOfEvents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 7.94.2.2 getNumberOfTotalEventsExpected . . . . . . . . . . . . . . . . . . . . . . . . 156 7.94.2.3 readEvent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 7.94.2.4 threeCharge . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 7.94.2.5 updateNextEvent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156

7.95 UTIL::LCStdHepRdrNew Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.95.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 7.95.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157

7.95.2.1 getNumberOfEvents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 7.95.2.2 getNumberOfTotalEventsExpected . . . . . . . . . . . . . . . . . . . . . . . . 157 7.95.2.3 readEvent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 7.95.2.4 threeCharge . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 7.95.2.5 updateNextEvent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158

7.96 EVENT::LCStrVec Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158

7.96.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 7.97 UTIL::LCTime Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159

7.97.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 7.97.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161

7.97.2.1 daysInMonth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 7.97.2.2 daysInYear . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 7.97.2.3 getDateString . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161

7.98 UTIL::LCTokenizer Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 7.98.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 7.98.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162

7.98.2.1 operator() . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162

7.99 UTIL::LCTOOLS Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 7.99.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 7.99.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164

7.99.2.1 dumpEvent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 7.99.2.2 getSimulatorStatusString . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164

7.100 UTIL::LCTypedVector< T > Class Template Reference . . . . . . . . . . . . . . . . . . . . . . . 164

7.100.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 7.101 UTIL::LCWarning Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165

7.101.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 7.101.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166

7.101.2.1 registerWarning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166

7.102 IO::LCWriter Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 7.102.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 7.102.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167

7.102.2.1 close . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 7.102.2.2 flush . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 7.102.2.3 open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 7.102.2.4 open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 7.102.2.5 setCompressionLevel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 7.102.2.6 writeEvent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 7.102.2.7 writeRunHeader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.103 Mass Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 7.104 EVENT::MCParticle Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169

7.104.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 7.104.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171

7.104.2.1 getEndpoint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 7.104.2.2 getMomentum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 7.104.2.3 getSimulatorStatus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 7.104.2.4 getTime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 7.104.2.5 isBackscatter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 7.104.2.6 vertexIsNotEndpointOfParent . . . . . . . . . . . . . . . . . . . . . . . . . . . 172

7.105 IMPL::MCParticleCont Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 7.106 IMPL::MCParticleImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173

7.106.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 7.106.2 Constructor & Destructor Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . 176

7.106.2.1 ∼MCParticleImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 7.106.3 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176

7.106.3.1 getEndpoint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 7.106.3.2 getPDG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 7.106.3.3 getSimulatorStatus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176 7.106.3.4 getTime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 7.106.3.5 setPDG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177

7.107 IOIMPL::MCParticleIOImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177

7.107.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 7.108 MCTruth2Rec Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 7.109 MyUserExtension Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 7.110 lcrtrel_helper::NoDelete Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179

7.110.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 7.111 lcrtrel_helper::objorcont< is_container > Struct Template Reference . . . . . . . . . . . . . . . . 179

7.111.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 7.112 lcrtrel_helper::objorcont< false > Struct Template Reference . . . . . . . . . . . . . . . . . . . . 179

7.112.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 7.113 opts_t Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 7.114 pagerInfo Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 7.115 ParentDaughter Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 7.116 EVENT::ParticleID Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181

7.116.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 7.116.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182

7.116.2.1 getAlgorithmType . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 7.116.2.2 getParameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182

7.117 IMPL::ParticleIDImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.117.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 7.117.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184

7.117.2.1 getAlgorithmType . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 7.117.2.2 getParameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184

7.118 IOIMPL::ParticleIDIOImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184

7.118.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 7.119 ParticleIDs Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 7.120 UTIL::PIDHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186

7.120.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 7.120.2 Constructor & Destructor Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . 187

7.120.2.1 PIDHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 7.120.3 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187

7.120.3.1 getParticleID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 7.120.3.2 setParticleIDUsed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187

7.121 IMPL::PIDSort Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188

7.121.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 7.122 EVENT::RawCalorimeterHit Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188

7.122.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 7.122.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189

7.122.2.1 getCellID1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189 7.122.2.2 getTimeStamp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189

7.123 IMPL::RawCalorimeterHitImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 7.123.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 7.123.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191

7.123.2.1 getCellID1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 7.123.2.2 getTimeStamp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 7.123.2.3 setCellID1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191

7.124 IOIMPL::RawCalorimeterHitIOImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . 191

7.124.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 7.125 EVENT::ReadOnlyException Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192

7.125.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 7.126 EVENT::ReconstructedParticle Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . 193

7.126.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 7.126.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195

7.126.2.1 addParticleID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 7.126.2.2 getClusters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 7.126.2.3 getCovMatrix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 7.126.2.4 getMass . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 7.126.2.5 getParticleIDs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 7.126.2.6 getParticleIDUsed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.126.2.7 getTracks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 7.126.2.8 getType . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 7.126.2.9 isCompound . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196

7.127 IMPL::ReconstructedParticleImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . 197 7.127.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 7.127.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199

7.127.2.1 addParticleID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 7.127.2.2 getParticleIDs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 7.127.2.3 getParticleIDUsed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 7.127.2.4 getType . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200

7.128 IOIMPL::ReconstructedParticleIOImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . 200

7.128.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 7.129 lcrtrel_helper::RelationManySide< U, T > Struct Template Reference . . . . . . . . . . . . . . . . 201

7.129.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 7.130 lcrtrel_helper::RelationOneSide< U, T > Struct Template Reference . . . . . . . . . . . . . . . . 201

7.130.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 7.131 SIO::RunEvent Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202

7.131.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 7.132 SIO::RunEventMap Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202

7.132.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 7.132.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204

7.132.2.1 getPosition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204

7.133 RunEventProcessor Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 204

7.133.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 7.134 EVENT::SimCalorimeterHit Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205

7.134.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 7.134.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207

7.134.2.1 getCellID1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 7.134.2.2 getEnergyCont . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 7.134.2.3 getNMCContributions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 7.134.2.4 getNMCParticles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 7.134.2.5 getParticleCont . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 7.134.2.6 getPDGCont . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 7.134.2.7 getPosition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 7.134.2.8 getTimeCont . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208

7.135 IMPL::SimCalorimeterHitImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 7.135.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 7.135.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210

7.135.2.1 addMCParticleContribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 7.135.2.2 addMCParticleContribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.135.2.3 getCellID0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 7.135.2.4 getCellID1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 7.135.2.5 getEnergyCont . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 7.135.2.6 getNMCContributions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 7.135.2.7 getNMCParticles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 7.135.2.8 getParticleCont . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 7.135.2.9 getPDGCont . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 7.135.2.10 getPosition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 7.135.2.11 getTimeCont . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 7.135.2.12 setCellID1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 7.135.2.13 setPosition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212

7.136 IOIMPL::SimCalorimeterHitIOImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . 212

7.136.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 7.137 lcrtrel_helper::SimplePtrInit Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213

7.137.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 7.138 EVENT::SimTrackerHit Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214

7.138.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 7.138.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215

7.138.2.1 getdEdx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 7.138.2.2 getMCParticle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 7.138.2.3 getPathLength . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215 7.138.2.4 getTime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 215

7.139 IMPL::SimTrackerHitImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 7.139.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 7.139.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218

7.139.2.1 getCellID1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 7.139.2.2 getdEdx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 7.139.2.3 getMCParticle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 7.139.2.4 getPathLength . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 7.139.2.5 getTime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 7.139.2.6 setCellID1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 7.139.2.7 setdEdx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218

7.140 IOIMPL::SimTrackerHitIOImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219

7.140.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 7.141 SIO::SIOCalHitHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220

7.141.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 7.142 SIO::SIOClusterHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220

7.142.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 7.143 SIO::SIOCollectionHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221

7.143.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.143.2 Constructor & Destructor Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . 222

7.143.2.1 SIOCollectionHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222

7.144 SIO::SIOEventHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222

7.144.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 7.145 SIO::SIOFloatVecHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223

7.145.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 7.146 SIO::SIOHandlerMgr Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224

7.146.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 7.146.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224

7.146.2.1 registerHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224

7.147 SIO::SIOIndexHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225

7.147.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 7.148 SIO::SIOIntVecHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225

7.148.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 226 7.149 SIO::SIOLCGenericObjectHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . 226

7.149.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 7.149.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227

7.149.2.1 init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227

7.150 SIO::SIOLCParameters Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228

7.150.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 7.151 SIO::SIOLCRelationHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228

7.151.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 7.152 SIO::SIOObjectHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229

7.152.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 7.152.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231

7.152.2.1 init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 7.152.2.2 readBase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 7.152.2.3 writeBase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231

7.153 SIO::SIOParticleHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231

7.153.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 7.154 SIO::SIORandomAccessHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . 232

7.154.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 7.155 SIO::SIORawCalHitHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233

7.155.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 233 7.156 SIO::SIOReader Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234

7.156.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 7.156.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236

7.156.2.1 close . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 7.156.2.2 getEvents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 7.156.2.3 getNumberOfEvents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.156.2.4 getNumberOfRuns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 7.156.2.5 getRuns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 7.156.2.6 open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 7.156.2.7 open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 7.156.2.8 readEvent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 7.156.2.9 readEvent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 7.156.2.10 readNextEvent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 7.156.2.11 readNextEvent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 7.156.2.12 readNextRunHeader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 7.156.2.13 readNextRunHeader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 7.156.2.14 readRunHeader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 7.156.2.15 readRunHeader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 7.156.2.16 readStream . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 7.156.2.17 readStream . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 7.156.2.18 setReadCollectionNames . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 7.156.2.19 skipNEvents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 240

7.157 SIO::SIOReconstructedParticleHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . 241

7.157.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 7.158 SIO::SIORecords Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241

7.158.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242 7.159 SIO::SIORunHeaderHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . 242

7.159.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243 7.160 SIO::SIOSimCalHitHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 243

7.160.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 7.161 SIO::SIOSimTrackHitHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244

7.161.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 7.162 SIO::SIOStrVecHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245

7.162.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 7.163 SIO::SIOTPCHitHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246

7.163.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 7.164 SIO::SIOTrackerDataHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246

7.164.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 7.165 SIO::SIOTrackerHitHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247

7.165.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 248 7.166 SIO::SIOTrackerHitPlaneHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . 248

7.166.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 248 7.167 SIO::SIOTrackerHitZCylinderHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . 249

7.167.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249 7.168 SIO::SIOTrackerPulseHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . 250

7.168.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.169 SIO::SIOTrackerRawDataHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . 250

7.169.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 7.170 SIO::SIOTrackHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251

7.170.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 252 7.171 SIO::SIOVertexHandler Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 252

7.171.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253 7.171.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253

7.171.2.1 init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253

7.172 SIO::SIOWriter Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253 7.172.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 7.172.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255

7.172.2.1 close . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 7.172.2.2 flush . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 7.172.2.3 open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 7.172.2.4 open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 7.172.2.5 setCompressionLevel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 7.172.2.6 setUpHandlers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 7.172.2.7 writeEvent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 7.172.2.8 writeRunHeader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257

7.173 SomeClass Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 7.174 TEST Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258 7.175 lcrtrel_helper::ToRelation< U > Struct Template Reference . . . . . . . . . . . . . . . . . . . . . 258

7.175.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258 7.176 EVENT::TPCHit Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258

7.176.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 7.176.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259

7.176.2.1 getNRawDataWords . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 7.176.2.2 getRawDataWord . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260

7.177 IMPL::TPCHitImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 7.177.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 7.177.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261

7.177.2.1 getNRawDataWords . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 7.177.2.2 getRawDataWord . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262

7.178 IOIMPL::TPCHitIOImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262

7.178.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 7.179 EVENT::Track Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 263

7.179.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 7.179.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264

7.179.2.1 getChi2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 7.179.2.2 getClosestTrackState . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.179.2.3 getCovMatrix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 7.179.2.4 getD0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 7.179.2.5 getOmega . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 7.179.2.6 getPhi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 7.179.2.7 getRadiusOfInnermostHit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 7.179.2.8 getReferencePoint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 7.179.2.9 getSubdetectorHitNumbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 7.179.2.10 getTanLambda . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 7.179.2.11 getTrackState . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 7.179.2.12 getTrackStates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 7.179.2.13 getType . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 7.179.2.14 getZ0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267

7.180 EVENT::TrackerData Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 7.180.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 7.180.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268

7.180.2.1 getCellID1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 7.180.2.2 getTime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269

7.181 IMPL::TrackerDataImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 7.181.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 7.181.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270

7.181.2.1 chargeValues . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 7.181.2.2 getCellID1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270

7.182 IOIMPL::TrackerDataIOImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270

7.182.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 7.183 EVENT::TrackerHit Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271

7.183.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 272 7.183.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 273

7.183.2.1 getCovMatrix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 273 7.183.2.2 getdEdx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 273 7.183.2.3 getRawHits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 273 7.183.2.4 getType . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 273

7.184 IMPL::TrackerHitImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 273 7.184.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275 7.184.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275

7.184.2.1 getCellID1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275 7.184.2.2 getdEdx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276 7.184.2.3 getRawHits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276 7.184.2.4 getTime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276 7.184.2.5 getType . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276 7.184.2.6 setCellID1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.185 IOIMPL::TrackerHitIOImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276

7.185.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 7.186 EVENT::TrackerHitPlane Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277

7.186.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 7.186.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279

7.186.2.1 getU . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 7.186.2.2 getV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279

7.187 IMPL::TrackerHitPlaneImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 7.187.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 7.187.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281

7.187.2.1 getCellID1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 7.187.2.2 getdEdx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282 7.187.2.3 getRawHits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282 7.187.2.4 getTime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282 7.187.2.5 getType . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282 7.187.2.6 getU . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282 7.187.2.7 getV . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282 7.187.2.8 setCellID1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282

7.188 IOIMPL::TrackerHitPlaneIOImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . 283

7.188.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 7.189 EVENT::TrackerHitZCylinder Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283

7.189.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 7.189.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285

7.189.2.1 getCenter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285

7.190 IMPL::TrackerHitZCylinderImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . 285 7.190.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 7.190.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287

7.190.2.1 getCellID1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 7.190.2.2 getCenter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 7.190.2.3 getdEdx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 7.190.2.4 getRawHits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 288 7.190.2.5 getTime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 288 7.190.2.6 getType . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 288 7.190.2.7 setCellID1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 288

7.191 IOIMPL::TrackerHitZCylinderIOImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . 288

7.191.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 289 7.192 EVENT::TrackerPulse Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 289

7.192.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 290 7.192.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 290

7.192.2.1 getCellID1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 290

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.192.2.2 getCovMatrix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291 7.192.2.3 getTrackerData . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291

7.193 IMPL::TrackerPulseImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291 7.193.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292 7.193.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292

7.193.2.1 getCellID1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 292 7.193.2.2 getCovMatrix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293 7.193.2.3 getTrackerData . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293

7.194 IOIMPL::TrackerPulseIOImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293

7.194.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293 7.195 EVENT::TrackerRawData Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294

7.195.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294 7.195.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295

7.195.2.1 getCellID1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295 7.195.2.2 getTime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295

7.196 IMPL::TrackerRawDataImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295 7.196.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 7.196.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297

7.196.2.1 adcValues . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 7.196.2.2 getCellID1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297

7.197 IOIMPL::TrackerRawDataIOImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . 297

7.197.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 7.198 IMPL::TrackImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298

7.198.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 7.198.2 Constructor & Destructor Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . 300

7.198.2.1 TrackImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 7.198.3 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301

7.198.3.1 getChi2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 7.198.3.2 getClosestTrackState . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 7.198.3.3 getCovMatrix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 7.198.3.4 getD0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 7.198.3.5 getOmega . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 301 7.198.3.6 getPhi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 7.198.3.7 getRadiusOfInnermostHit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 7.198.3.8 getReferencePoint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 7.198.3.9 getSubdetectorHitNumbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 7.198.3.10 getTanLambda . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 302 7.198.3.11 getTrackState . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 7.198.3.12 getTrackStates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 7.198.3.13 getType . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.198.3.14 getZ0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 7.198.3.15 subdetectorHitNumbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303

7.199 IOIMPL::TrackIOImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 304

7.199.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 304 7.200 EVENT::TrackState Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 304

7.200.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 7.200.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306

7.200.2.1 getCovMatrix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 7.200.2.2 getLocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 7.200.2.3 getOmega . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 7.200.2.4 getPhi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 7.200.2.5 getReferencePoint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 306 7.200.2.6 getTanLambda . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307

7.201 IMPL::TrackStateImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 7.201.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 308 7.201.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309

7.201.2.1 getCovMatrix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309 7.201.2.2 getLocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309 7.201.2.3 getOmega . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309 7.201.2.4 getReferencePoint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309

7.202 IOIMPL::TrackStateIOImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309

7.202.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310 7.203 TrkCluLink Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310 7.204 UTIL::UnknownAlgorithm Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 311

7.204.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 311 7.205 SIO::SIORecords::Unpack Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 311

7.205.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 7.206 UserClass Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 7.207 UserStruct Struct Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 7.208 EVENT::Vertex Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312

7.208.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 313 7.208.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314

7.208.2.1 getCovMatrix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 7.208.2.2 isPrimary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314

7.209 IMPL::VertexImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314 7.209.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 316 7.209.2 Member Function Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 316

7.209.2.1 getCovMatrix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 316 7.209.2.2 isPrimary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 316

7.210 IOIMPL::VertexIOImpl Class Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 316

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

xxvi  CONTENTS

7.210.1 Detailed Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Chapter 1

<a href="http://lcio.desy.de">LCIO</a> (v02-07-04)

A persistency framework for linear collider simulation studies.

1.1 Overview

This is the API documentation for the C++ version of LCIO. For general documetation on LCIO see the homepage. There you can also find the API documentation of the Java version.

1.2 Usage

A good starting point to browse this documentation is the Namespace List. As all data entities inherit from E-VENT::LCObject (p. 123), the corresponding class documentation shows the complete hierarchy of all objects that can be stored with LCIO. The namespace IO (p. 27) contains the abstract interfaces for reading and writing data.

2  <a href="http://lcio.desy.de">LCIO</a> (v02-07-04)

Chapter 2

Deprecated List

Member EVENT::SimCalorimeterHit::getNMCParticles (p. 207) () const =0  See Also

getNMCContributions()

Member EVENT::SimTrackerHit::getdEdx (p. 215) () const =0  See Also

getEDep()

Class EVENT::TPCHit (p. 258)  Author

gaede Version

Member EVENT::TrackerHit::getdEdx (p. 273) () const =0  See Also

getEDep()

Member IMPL::SimCalorimeterHitImpl::getNMCParticles (p. 211) () const  See Also

getNMCContributions()

Member IMPL::SimTrackerHitImpl::getdEdx (p. 218) () const  See Also

getEDep()

Member IMPL::SimTrackerHitImpl::setdEdx (p. 218) (float dEdX)  See Also

setEDep()

Member IMPL::TrackerHitImpl::getdEdx (p. 276) () const  See Also

getEDep()

Member IMPL::TrackerHitPlaneImpl::getdEdx (p. 282) () const  See Also

getEDep()

Member IMPL::TrackerHitZCylinderImpl::getdEdx (p. 287) () const  See Also

getEDep()

Member UTIL::LCStdHepRdr::readEvent (p. 156) ()

please use updateEvent() Member UTIL::LCStdHepRdrNew::readEvent (p. 158) ()

please use updateEvent()

Namespace Index

3.1 Namespace List

Here is a list of all documented namespaces with brief descriptions:

DATA The namespace DATA (p. 21) has been removed from LCIO all data entities are now defined in EVENT (p. 21) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

EVENT

The namespace EVENT (p. 21) holds all base interfaces of the event entities . . . . . . . . . . 21

IMPL The namespace IMPL (p. 25) holds the implementations of the base interfaces that are defined in EVENT (p. 21) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

IO

The namespace IO (p. 27) holds the base interfaces for io of data . . . . . . . . . . . . . . . . 27

IOIMPL The IOIMPL (p. 27) namespace holds extensions to the default implementations needed for IO (p. 27) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

lcio The LCIO namespace combines EVENT (p. 21), IO (p. 27), IMPL (p. 25) and UTIL (p. 32) for user convenience . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

lcrtrel The namespace lcrtrel holds classes and functions for runtime extensions and relations see LCRTRelations (p. 142) for more detailed documentation . . . . . . . . . . . . . . . . . . . . 28

SIO

The SIO (p. 30) namespace holds the persistency implementation using SIO (p. 30) . . . . . . 30

UTIL The namespace UTIL (p. 32) holds convenient and support classes and methods for the LCIO objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

6  Namespace Index

Hierarchical Index

4.1 Class Hierarchy

This inheritance list is sorted roughly, but not completely, alphabetically:

IMPL::AccessChecked . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

IMPL::CalorimeterHitImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

IOIMPL::CalorimeterHitIOImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 IMPL::ClusterImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

IOIMPL::ClusterIOImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 IMPL::LCCollectionVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

IOIMPL::LCCollectionIOVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 IMPL::LCEventImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

IOIMPL::LCEventIOImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 IMPL::LCGenericObjectImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .110

IOIMPL::LCGenericObjectIOImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 IMPL::LCParametersImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .128

SIO::SIOLCParameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 IMPL::LCRelationImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .137

IOIMPL::LCRelationIOImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 IMPL::LCRunHeaderImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .146

IOIMPL::LCRunHeaderIOImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 IMPL::MCParticleImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .173

IOIMPL::MCParticleIOImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177 IMPL::ParticleIDImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .183

IOIMPL::ParticleIDIOImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 IMPL::RawCalorimeterHitImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .190

IOIMPL::RawCalorimeterHitIOImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 IMPL::ReconstructedParticleImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .197

IOIMPL::ReconstructedParticleIOImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 IMPL::SimCalorimeterHitImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .208

IOIMPL::SimCalorimeterHitIOImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 IMPL::SimTrackerHitImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .216

IOIMPL::SimTrackerHitIOImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 IMPL::TPCHitImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .260

IOIMPL::TPCHitIOImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 IMPL::TrackerDataImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .269

IOIMPL::TrackerDataIOImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 IMPL::TrackerHitImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .273

IOIMPL::TrackerHitIOImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 276 IMPL::TrackerHitPlaneImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .279

IOIMPL::TrackerHitPlaneIOImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 IMPL::TrackerHitZCylinderImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .285

IOIMPL::TrackerHitZCylinderIOImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 288 IMPL::TrackerPulseImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .291

IOIMPL::TrackerPulseIOImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 293 IMPL::TrackerRawDataImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .295

IOIMPL::TrackerRawDataIOImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 IMPL::TrackImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .298

IOIMPL::TrackIOImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 304 IMPL::TrackStateImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .307

IOIMPL::TrackStateIOImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309 IMPL::VertexImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .314

IOIMPL::VertexIOImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 316 lcrtrel_helper::BiDirectional< From, To > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 lcrtrel_helper::BiDirectional< RelationManySide< FromRelation< U >, From >, RelationManySide< To-

Relation< U >, To > > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 lcrtrel::LCNToNRelation< U, From, To > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .122

lcrtrel_helper::BiDirectional< RelationOneSide< FromRelation< U >, From >, RelationManySide< To-

Relation< U >, To > > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 lcrtrel::LC1ToNRelation< U, From, To > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

lcrtrel_helper::BiDirectional< RelationOneSide< FromRelation< U >, From >, RelationOneSide< To-

Relation< U >, To > > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 lcrtrel::LC1To1Relation< U, From, To > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

binary_function

IMPL::PIDSort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .188 UTIL::BitField64 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

UTIL::CellIDEncoder< T > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

UTIL::ILDCellIDEncoder< T > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 UTIL::BitFieldValue . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 bitset

UTIL::BitSet32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 UTIL::LCTime::CalendarTime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 lcrtrel_helper::can_call_ext< B > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 lcrtrel_helper::can_call_ext< true > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 UTIL::CellIDDecoder< T > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 UTIL::CollectionParameterMap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 lcrtrel_helper::CreationPtrInit< T > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 lcrtrel_helper::DeleteElements< T > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 lcrtrel_helper::DeletePtr< T > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 exception

EVENT::Exception . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

EVENT::EventException . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 EVENT::DataNotAvailableException . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 EVENT::ReadOnlyException . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 IO::IOException . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 IO::EndOfDataException . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 UTIL::UnknownAlgorithm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 311

FloatVec

EVENT::LCFloatVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .106 lcrtrel_helper::FromRelation< U > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 HEPEVTIMPL::HEPEVT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 hepevt1_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 hepevt_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 HepLorentzVector

UTIL::LCFourVector< TT > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .107 IO::ILCFactory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

IOIMPL::LCFactory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .101 UTIL::ILDCellID0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 UTIL::ILDDetID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 UTIL::ILDTrkHitQualityBit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 UTIL::ILDTrkHitTypeBit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 UTIL::IndexMap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 IntVec

EVENT::LCIntVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .114 LC1ToNRelation

TrkCluLink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .310 lcrtrel_helper::LCBaseTraits< U, T, I, D, b > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82

lcrtrel_helper::LCBaseLinkContainerTraits< U, T, I, D, b > . . . . . . . . . . . . . . . . . . . . . . . . 80 lcrtrel_helper::LCBaseLinkTraits< U, T, I, D, b > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

lcrtrel_helper::LCBaseTraits< U, std::list< T ∗ >, CreationPtrInit< std::list< T ∗ > >, DeleteElements<

std::list< T ∗ > >, 1 > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 lcrtrel_helper::LCBaseLinkContainerTraits< U, std::list< T ∗ >, CreationPtrInit< std::list< T ∗ > >,

DeleteElements< std::list< T ∗ > > > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 lcrtrel::LCOwnedExtensionList< U, T > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125

lcrtrel_helper::LCBaseTraits< U, std::list< T ∗ >, CreationPtrInit< std::list< T ∗ > >, DeletePtr< std:-

:list< T ∗ > >, 1 > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 lcrtrel_helper::LCBaseLinkContainerTraits< U, std::list< T ∗ >, CreationPtrInit< std::list< T ∗ > >,

DeletePtr< std::list< T ∗ > > > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 lcrtrel::LCExtensionList< U, T > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100

lcrtrel_helper::LCBaseTraits< U, std::list< T ∗ >, CreationPtrInit< std::list< T ∗ > >, DeletePtr< std:-

:list< T ∗ > >, b > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 lcrtrel_helper::LCBaseLinkContainerTraits< U, std::list< T ∗ >, CreationPtrInit< std::list< T ∗ > >,

DeletePtr< std::list< T ∗ > >, false > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 lcrtrel_helper::RelationManySide< U, T > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201

lcrtrel_helper::LCBaseTraits< U, std::vector< T ∗ >, CreationPtrInit< std::vector< T ∗ > >, Delete-

Elements< std::vector< T ∗ > >, 1 > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 lcrtrel_helper::LCBaseLinkContainerTraits< U, std::vector< T ∗ >, CreationPtrInit< std::vector< T ∗

> >, DeleteElements< std::vector< T ∗ > > > . . . . . . . . . . . . . . . . . . . . . . . . 80 lcrtrel::LCOwnedExtensionVector< U, T > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126

lcrtrel_helper::LCBaseTraits< U, std::vector< T ∗ >, CreationPtrInit< std::vector< T ∗ > >, DeletePtr<

std::vector< T ∗ > >, 1 > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 lcrtrel_helper::LCBaseLinkContainerTraits< U, std::vector< T ∗ >, CreationPtrInit< std::vector< T ∗

> >, DeletePtr< std::vector< T ∗ > > > . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 lcrtrel::LCExtensionVector< U, T > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100

lcrtrel_helper::LCBaseTraits< U, T, SimplePtrInit, DeletePtr< T >, 1 > . . . . . . . . . . . . . . . . . . 82

lcrtrel_helper::LCBaseLinkTraits< U, T, SimplePtrInit, DeletePtr< T > > . . . . . . . . . . . . . . . . 81

lcrtrel::LCOwnedExtension< U, T > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 lcrtrel_helper::LCBaseTraits< U, T, SimplePtrInit, NoDelete, 1 > . . . . . . . . . . . . . . . . . . . . . 82

lcrtrel_helper::LCBaseLinkTraits< U, T > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

lcrtrel::LCExtension< U, T > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 lcrtrel_helper::LCBaseTraits< U, T, SimplePtrInit, NoDelete, b > . . . . . . . . . . . . . . . . . . . . . 82

lcrtrel_helper::LCBaseLinkTraits< U, T, SimplePtrInit, NoDelete, false > . . . . . . . . . . . . . . . . 81

lcrtrel_helper::RelationOneSide< U, T > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 EVENT::LCCollection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

IMPL::LCCollectionVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 EVENT::LCEvent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

IMPL::LCEventImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 IO::LCEventListener . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 LCEventListener RunEventProcessor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .204 RunEventProcessor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .204 RunEventProcessor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .204 RunEventProcessor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .204

LCFixedObject

CalibrationConstant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 EVENT::LCFlag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

IMPL::LCFlagImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .104 LCFloatExtension

Mass . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .169 Mass . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .169

lcrtrel::LCFloatExtension< U > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 LCGenericObject

UTIL::LCFixedObject< NINT, NFLOAT, NDOUBLE > . . . . . . . . . . . . . . . . . . . . . . . . . .102 LCIntExtension

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 SomeClass::Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

lcrtrel::LCIntExtension< U > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 EVENT::LCIO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 UTIL::LCIO_LONG< T > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 UTIL::lcio_short< T > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 IMPL::LCIOExceptionHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 SIO::LCIORandomAccess . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 SIO::LCIORandomAccessMgr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 UTIL::LCIterator< T > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 LCNToNRelation

MCTruth2Rec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .178 ParentDaughter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .181

UTIL::LCObjectHandle< T > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 LCObjectVec

IMPL::LCCollectionVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 LCOwnedExtension

MyUserExtension . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .178 UserStruct . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .312

LCOwnedExtensionVector ParticleIDs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .185 ParticleIDs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .185

EVENT::LCParameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126

IMPL::LCParametersImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .128 IO::LCReader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130

SIO::SIOReader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .234 UTIL::LCRelationNavigator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 lcrtrel::LCRTRelations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142

EVENT::LCObject . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .123

EVENT::CalorimeterHit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 IMPL::CalorimeterHitImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 EVENT::Cluster . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 IMPL::ClusterImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 EVENT::LCFloatVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 EVENT::LCGenericObject . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 IMPL::LCGenericObjectImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 EVENT::LCIntVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 EVENT::LCRelation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 IMPL::LCRelationImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 EVENT::LCRunHeader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 IMPL::LCRunHeaderImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 EVENT::LCStrVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 EVENT::MCParticle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169

IMPL::MCParticleImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 EVENT::ParticleID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181

IMPL::ParticleIDImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 EVENT::RawCalorimeterHit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188

IMPL::RawCalorimeterHitImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 EVENT::ReconstructedParticle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193

IMPL::ReconstructedParticleImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 EVENT::SimCalorimeterHit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205

IMPL::SimCalorimeterHitImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 EVENT::SimTrackerHit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214

IMPL::SimTrackerHitImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216 EVENT::TPCHit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258

IMPL::TPCHitImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 EVENT::Track . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 263

IMPL::TrackImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 EVENT::TrackerData . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267

IMPL::TrackerDataImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 EVENT::TrackerHit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271

EVENT::TrackerHitPlane . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277

IMPL::TrackerHitPlaneImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 EVENT::TrackerHitPlane . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 EVENT::TrackerHitZCylinder . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283

IMPL::TrackerHitZCylinderImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 285 EVENT::TrackerHitZCylinder . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 IMPL::TrackerHitImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 273

EVENT::TrackerPulse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 289

IMPL::TrackerPulseImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291 EVENT::TrackerRawData . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294

IMPL::TrackerRawDataImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295 EVENT::TrackState . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 304

IMPL::TrackStateImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 307 EVENT::Vertex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312

IMPL::VertexImpl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314

IO::LCRunListener . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 LCRunListener

RunEventProcessor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .204 RunEventProcessor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .204 RunEventProcessor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .204 RunEventProcessor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .204

SIO::LCSIO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 UTIL::LCStdHepRdr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 UTIL::LCStdHepRdrNew . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 UTIL::LCTime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 UTIL::LCTokenizer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 UTIL::LCTOOLS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 UTIL::LCWarning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 IO::LCWriter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166

SIO::SIOWriter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .253 UTIL::LCSplitWriter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .151

IMPL::MCParticleCont . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 lcrtrel_helper::NoDelete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 lcrtrel_helper::objorcont< is_container > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 lcrtrel_helper::objorcont< false > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 opts_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 pagerInfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 UTIL::PIDHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 SIO::RunEvent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 SIO::RunEventMap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 lcrtrel_helper::SimplePtrInit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 SIO_block

SIO::SIOCollectionHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .221 SIO::SIOEventHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .222 SIO::SIOIndexHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .225 SIO::SIORandomAccessHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .232 SIO::SIORunHeaderHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .242

SIO::SIOHandlerMgr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 SIO::SIOObjectHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229

SIO::SIOCalHitHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .220 SIO::SIOClusterHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .220 SIO::SIOFloatVecHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .223 SIO::SIOIntVecHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .225 SIO::SIOLCGenericObjectHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .226 SIO::SIOLCRelationHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .228 SIO::SIOParticleHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .231 SIO::SIORawCalHitHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .233 SIO::SIOReconstructedParticleHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .241 SIO::SIOSimCalHitHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .243 SIO::SIOSimTrackHitHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .244 SIO::SIOStrVecHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .245 SIO::SIOTPCHitHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .246 SIO::SIOTrackerDataHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .246 SIO::SIOTrackerHitHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .247 SIO::SIOTrackerHitPlaneHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .248 SIO::SIOTrackerHitZCylinderHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .249 SIO::SIOTrackerPulseHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .250 SIO::SIOTrackerRawDataHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .250 SIO::SIOTrackHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .251 SIO::SIOVertexHandler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .252

SIO::SIORecords . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 SomeClass . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 StringVec

EVENT::LCStrVec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .158 TEST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258 lcrtrel_helper::ToRelation< U > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258 SIO::SIORecords::Unpack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 311 UserClass . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 vector

UTIL::LCTypedVector< T > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .164

Class Index

5.1 Class List

Here are the classes, structs, unions and interfaces with brief descriptions:

IMPL::AccessChecked

Controls access to objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

lcrtrel_helper::BiDirectional< From, To >

Helper class for biderectional relations provides the to and from type . . . . . . . . . . . . . . 42

UTIL::BitField64 A bit field of 64bits that allows convenient declaration and manipulation of sub fields of various widths . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

UTIL::BitFieldValue

Helper class for BitField64 (p. 43) that corresponds to one field value . . . . . . . . . . . . . 45

UTIL::BitSet32 Convenient helper class for setting single bits in a 32bit-field, such as collection flags or hit types

- extends std::bitset<32> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

UTIL::LCTime::CalendarTime

Helper struct that holds the calendar time . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

CalibrationConstant

Example for a simple calibration class based on the LCFixedObject template . . . . . . . . . . 48

EVENT::CalorimeterHit The generic calorimeter hit - created from SimCalorimeterHit (p. 205) or RawCalorimeterHit (p. 188) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

IMPL::CalorimeterHitImpl

Implementation of the real data CalorimeterHit . . . . . . . . . . . . . . . . . . . . . . . . . . 51

IOIMPL::CalorimeterHitIOImpl

Adding stuff needed for io (friend declarations, etc.) . . . . . . . . . . . . . . . . . . . . . . . 54

lcrtrel_helper::can_call_ext< B > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 lcrtrel_helper::can_call_ext< true > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 UTIL::CellIDDecoder< T >

Convenient class for decoding cellIDs from collection parameter LCIO::CellIDEncoding …. 56

UTIL::CellIDEncoder< T >

Specialization for SimTrackerHits that have only one cellID . . . . . . . . . . . . . . . . . . . 57

EVENT::Cluster

The LCIO (p. 115) cluster . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

IMPL::ClusterImpl

Implementation of Cluster . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

IOIMPL::ClusterIOImpl

Adding stuff needed for io (friend declarations, etc.) . . . . . . . . . . . . . . . . . . . . . . . 64

UTIL::CollectionParameterMap Helper class for setting and retrieving an std::map<string,int> as collection parameters …. 65 lcrtrel_helper::CreationPtrInit< T >

Factory for objects of type T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

EVENT::DataNotAvailableException

EventException (p. 69) used for data not available . . . . . . . . . . . . . . . . . . . . . . . 66

lcrtrel_helper::DeleteElements< T >

Delete function for containers of owned objects . . . . . . . . . . . . . . . . . . . . . . . . . 67

lcrtrel_helper::DeletePtr< T >

Delete function for pointers w/ ownership . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

IO::EndOfDataException

EndOfDataException (p. 68) for signaling the end of a data stream . . . . . . . . . . . . . . 68

EVENT::EventException

EventException (p. 69) used for errors accessing the event data . . . . . . . . . . . . . . . . 69

EVENT::Exception

Base exception class for LCIO (p. 115) - all other exceptions extend this . . . . . . . . . . . . 69

lcrtrel_helper::FromRelation< U >

Helper class for relations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

HEPEVTIMPL::HEPEVT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 hepevt1_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 hepevt_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 IO::ILCFactory

Factory to create LCWriter (p. 166) and Reader object for the known data formats, e.g . . . . . 72

UTIL::ILDCellID0 Helper class to define constants for the canonical encoding of the CellID0 in tracking sub detectors in ILD (or ILD like detectors) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

UTIL::ILDCellIDEncoder< T > Special wrapper to the CellIDEncoder<T> that enforces cellID0 to use the encoder string defined in ILDCellID0::encoder_string (p. 73) . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

UTIL::ILDDetID

Define integer constants for identifying ILD sub detectors . . . . . . . . . . . . . . . . . . . . 75

UTIL::ILDTrkHitQualityBit Define integer constants for bits of the the lcio tracker hit quality used in the context of ILD reconstruction code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76

UTIL::ILDTrkHitTypeBit Define integer constants for bits of the the lcio tracker hit type used in the context of ILD reconstruction code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 SomeClass::Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 UTIL::IndexMap

Utility class to manage indexes according to Collection Parameters . . . . . . . . . . . . . . . 77

IO::IOException

IOException (p. 79) used for reading/writing errors . . . . . . . . . . . . . . . . . . . . . . . 79

lcrtrel::LC1To1Relation< U, From, To >

One to one relation between two objects of type From and To . . . . . . . . . . . . . . . . . . 80

lcrtrel::LC1ToNRelation< U, From, To >

One to many relation between one object of type From to many objects of type To . . . . . . . 80

lcrtrel_helper::LCBaseLinkContainerTraits< U, T, I, D, b >

Base class for all containers of extensions and relations, vectors, lists,.. . . . . . . . . . . . . 80

lcrtrel_helper::LCBaseLinkTraits< U, T, I, D, b >

Base class for all extensions and relations of single objects . . . . . . . . . . . . . . . . . . . 81

lcrtrel_helper::LCBaseTraits< U, T, I, D, b >

Vector of pointers to extension obbjects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82

EVENT::LCCollection

The generic collection used in LCIO (p. 115) . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

IOIMPL::LCCollectionIOVec

Adding stuff needed for io (friend declarations, etc.) . . . . . . . . . . . . . . . . . . . . . . . 86

IMPL::LCCollectionVec Implementation of the LCCollection using (inheriting from) an STL vector of LCObjects …. 87 EVENT::LCEvent

The main event interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

IMPL::LCEventImpl

Implementation of the main event class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

IOIMPL::LCEventIOImpl

Adding stuff needed for io (friend declarations, etc.) . . . . . . . . . . . . . . . . . . . . . . . 97

IO::LCEventListener

Listener for the occurence of LCEvents when reading a stream . . . . . . . . . . . . . . . . . 98

lcrtrel::LCExtension< U, T >

Simple Extension - pointer to an object of type T . . . . . . . . . . . . . . . . . . . . . . . . . 99

lcrtrel::LCExtensionList< U, T >

Extension list holding pointers to objects of type T - no ownership of the objects is taken …. 100

lcrtrel::LCExtensionVector< U, T >

Extension vector holding pointers to objects of type T - no ownership of the objects is taken . . 100

IOIMPL::LCFactory

Factory to create LCWriter and Reader object for the known data formats, e.g . . . . . . . . . 101

UTIL::LCFixedObject< NINT, NFLOAT, NDOUBLE >

Template class for fixed size LCGenericObject subclasses . . . . . . . . . . . . . . . . . . . 102

EVENT::LCFlag

Helper class to create and interpret the 32-bit flag word in LCCollections . . . . . . . . . . . . 104

IMPL::LCFlagImpl

Implementation of helper class to create and interpret the 32-bit flag word in LCCollections . . 104

lcrtrel::LCFloatExtension< U > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 EVENT::LCFloatVec

Float vector used for user extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106

UTIL::LCFourVector< TT >

Four vector used in LCIO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107

EVENT::LCGenericObject

Simple interface to store generic user data . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109

IMPL::LCGenericObjectImpl

Default LCIO implementation of the interface to store generic user data . . . . . . . . . . . . . 110

IOIMPL::LCGenericObjectIOImpl

Adding stuff needed for io (friend declarations, etc.) . . . . . . . . . . . . . . . . . . . . . . . 112

lcrtrel::LCIntExtension< U >

Special Extension that allows to write int extensions directly (not through a pointer !) . . . . . . 113

EVENT::LCIntVec

Int vector used for user extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114

EVENT::LCIO

Global constants used in LCIO (p. 115) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115

UTIL::LCIO_LONG< T > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 UTIL::lcio_short< T > . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 IMPL::LCIOExceptionHandler

Sets the default error handling for unexpected and uncaught exceptions . . . . . . . . . . . . 117

SIO::LCIORandomAccess

Implementation class for LCIORandomAccess (p. 118) records . . . . . . . . . . . . . . . . 118

SIO::LCIORandomAccessMgr

Manager class for LCIO direct access . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

UTIL::LCIterator< T >

Simple convenient iterator class for LCCollections that saves some boiler plate code . . . . . . 121

lcrtrel::LCNToNRelation< U, From, To >

Many to many relation between objects of type From to objects of type To . . . . . . . . . . . 122

EVENT::LCObject

The generic object that is held in an LCCollection (p. 83) . . . . . . . . . . . . . . . . . . . . 123

UTIL::LCObjectHandle< T >

Template handle class for LCObjects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124

lcrtrel::LCOwnedExtension< U, T > Simple Extension - pointer to an object of type T where the ownership is taken over by the object holding the extension, i.e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 lcrtrel::LCOwnedExtensionList< U, T >

Extension list holding pointers to objects of type T - ownership of the objects is taken, i.e … 125

lcrtrel::LCOwnedExtensionVector< U, T >

Extension vector holding pointers to objects of type T - ownership of the objects is taken, i.e . . 126

EVENT::LCParameters

Simple interface to store generic named parameters of type int, float and string . . . . . . . . 126

IMPL::LCParametersImpl

Implementation of Simple interface to store generic named parameters of type int, float and string 128

IO::LCReader

Interface for reading data from LCIO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130

EVENT::LCRelation

A single weighted relationship between two LCObjects . . . . . . . . . . . . . . . . . . . . . 136

IMPL::LCRelationImpl

Implementation of the LCRelation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137

IOIMPL::LCRelationIOImpl

Adding stuff needed for io (friend declarations, etc.) . . . . . . . . . . . . . . . . . . . . . . . 139

UTIL::LCRelationNavigator The LCRelationNavigator (p. 139) makes repeated lookup of relations more conveneient and efficient . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139

lcrtrel::LCRTRelations

Base class that provides run time (user) extensions and relation between objects . . . . . . . 142

EVENT::LCRunHeader

Interface for the run header . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145

IMPL::LCRunHeaderImpl

Implementation of LCRunHeader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146

IOIMPL::LCRunHeaderIOImpl

Adding stuff needed for io (friend declarations, etc.) . . . . . . . . . . . . . . . . . . . . . . . 148

IO::LCRunListener

Listener for the occurence of LCRunHeader when reading a stream . . . . . . . . . . . . . . 148

SIO::LCSIO

Collection of static helper functions for reading and writing data with SIO (p. 30) . . . . . . . . 149

UTIL::LCSplitWriter

LCWriter wrapper that automatically splits files if a given number of bytes is exceeded . . . . . 151

UTIL::LCStdHepRdr Basic utility for reading a binary stdhep file and filling a LCCollectionVec with MCParticles containing the stdhep file information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155

UTIL::LCStdHepRdrNew Basic utility for reading a binary stdhep file and filling a LCCollectionVec with MCParticles containing the stdhep file information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156

EVENT::LCStrVec

String vector used for user extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158

UTIL::LCTime Helper class that allows to convert time stamps as defined in LCEvent::getTimeStamp() ( ns since 1.1.1970 00:00:00 UTC) to ordinary date and time and vice versa . . . . . . . . . . . . 159

UTIL::LCTokenizer

Helper class for string tokenization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162

UTIL::LCTOOLS

Provides some basic functions for printing event data . . . . . . . . . . . . . . . . . . . . . . 162

UTIL::LCTypedVector< T >

Templated helper class that is an std::vector<T∗> of the elements in the LCCollection …. 164

UTIL::LCWarning

Utility class to show warnings in LCIO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165

IO::LCWriter

Interface for writing data with LCIO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166

Mass . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 EVENT::MCParticle

The LCIO (p. 115) Monte Carlo particle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169

IMPL::MCParticleCont . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 IMPL::MCParticleImpl

Implementation of MCParticle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173

IOIMPL::MCParticleIOImpl

Adding stuff needed for io (friend declarations, etc.) . . . . . . . . . . . . . . . . . . . . . . . 177

MCTruth2Rec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 MyUserExtension . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 lcrtrel_helper::NoDelete

Empty delete function for pointers w/o ownership . . . . . . . . . . . . . . . . . . . . . . . . 179

lcrtrel_helper::objorcont< is_container >

Helper functions that treat single objects and containers . . . . . . . . . . . . . . . . . . . . . 179

lcrtrel_helper::objorcont< false >

Helper functions specialization for single objects . . . . . . . . . . . . . . . . . . . . . . . . . 179

opts_t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 pagerInfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 ParentDaughter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 EVENT::ParticleID

Persistent interface for LCIO (p. 115) ParticleIDs (p. 185) . . . . . . . . . . . . . . . . . . . . 181

IMPL::ParticleIDImpl

Implementation of ParticleID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183

IOIMPL::ParticleIDIOImpl

Adding stuff needed for io (friend declarations, etc.) . . . . . . . . . . . . . . . . . . . . . . . 184

ParticleIDs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 UTIL::PIDHandler

Convenient class for setting and retrieving particle id information attached to Reconstructed-Particles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186

IMPL::PIDSort

Helper class to sort ParticleIDs (p. 185) wrt . . . . . . . . . . . . . . . . . . . . . . . . . . . 188

EVENT::RawCalorimeterHit

The generic calorimeter hit for real data (or simulation thereof) . . . . . . . . . . . . . . . . . 188

IMPL::RawCalorimeterHitImpl

Implementation of the real data RawCalorimeterHit . . . . . . . . . . . . . . . . . . . . . . . 190

IOIMPL::RawCalorimeterHitIOImpl

Adding stuff needed for io (friend declarations, etc.) . . . . . . . . . . . . . . . . . . . . . . . 191

EVENT::ReadOnlyException

EventException (p. 69) used for signaling a ’read only exception’ . . . . . . . . . . . . . . . . 192

EVENT::ReconstructedParticle

The LCIO (p. 115) reconstructedParticle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193

IMPL::ReconstructedParticleImpl

Implementation of ReconstructedParticle . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197

IOIMPL::ReconstructedParticleIOImpl

Adding stuff needed for io (friend declarations, etc.) . . . . . . . . . . . . . . . . . . . . . . . 200

lcrtrel_helper::RelationManySide< U, T >

Helper class for relations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201

lcrtrel_helper::RelationOneSide< U, T >

Helper class for relations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201

SIO::RunEvent

Helper struct that stores run and event positions in the file . . . . . . . . . . . . . . . . . . . . 202

SIO::RunEventMap

Map that holds positions of Run and Event records . . . . . . . . . . . . . . . . . . . . . . . 202

RunEventProcessor Little tool that copies LCIO files on an event by event and run by run basis, thus fixing files that have incomplete records at the end due to a program crash . . . . . . . . . . . . . . . . . . . 204

EVENT::SimCalorimeterHit

The generic simulated calorimeter hit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205

IMPL::SimCalorimeterHitImpl

Implementation of the generic SimCalorimeterHit . . . . . . . . . . . . . . . . . . . . . . . . 208

IOIMPL::SimCalorimeterHitIOImpl Adding stuff needed for io (friend declarations, etc.) . . . . . . . . . . . . . . . . . . . . . . . 212 lcrtrel_helper::SimplePtrInit

Simple init function for simple pointers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213

EVENT::SimTrackerHit

A generic simulated tracker hit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214

IMPL::SimTrackerHitImpl

Implementation of SimTrackerHit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216

IOIMPL::SimTrackerHitIOImpl

Adding stuff needed for io (friend declarations, etc.) . . . . . . . . . . . . . . . . . . . . . . . 219

SIO::SIOCalHitHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of CalorimeterHits . . . . . 220

SIO::SIOClusterHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of Clusters . . . . . . . . . 220

SIO::SIOCollectionHandler

Handler for LCCollection/LCCollectionIOVec objects for SIO (p. 30) . . . . . . . . . . . . . . . 221

SIO::SIOEventHandler

Handler for LCEvent/LCEventIOImpl objects . . . . . . . . . . . . . . . . . . . . . . . . . . . 222

SIO::SIOFloatVecHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of FloatVecs . . . . . . . . 223

SIO::SIOHandlerMgr

Singleton that holds instances of all types of SIOObjectHandlers . . . . . . . . . . . . . . . . 224

SIO::SIOIndexHandler

Handler for LCIOrandomAccess and LCIOIndex objects/blocks . . . . . . . . . . . . . . . . . 225

SIO::SIOIntVecHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of IntVecs . . . . . . . . . . 225

SIO::SIOLCGenericObjectHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of LCGenericObjects …. 226

SIO::SIOLCParameters

IO (p. 27) of LCParameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228

SIO::SIOLCRelationHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of CalorimeterHits . . . . . 228

SIO::SIOObjectHandler Interface for all lcio object SIO-handlers, has to be implemented for all event entities (hits, tracks, clusters,...) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229

SIO::SIOParticleHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of MCParticles . . . . . . . 231

SIO::SIORandomAccessHandler

Handler for LCIOrandomAccess and LCIOIndex objects/blocks . . . . . . . . . . . . . . . . . 232

SIO::SIORawCalHitHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of RawCalorimeterHits … 233

SIO::SIOReader

Concrete implementation of LCWriter using SIO (p. 30) . . . . . . . . . . . . . . . . . . . . . 234

SIO::SIOReconstructedParticleHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of ReconstructedParticles . 241

SIO::SIORecords

Manager class that holds instances of all known LCIO-SIO records . . . . . . . . . . . . . . . 241

SIO::SIORunHeaderHandler

Handler for LCRunHeader/LCRunHeaderImpl objects . . . . . . . . . . . . . . . . . . . . . . 242

SIO::SIOSimCalHitHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of SimCalorimeterHits … 243

SIO::SIOSimTrackHitHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of SimTrackerHits . . . . . . 244

SIO::SIOStrVecHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of StrVecs . . . . . . . . . 245

SIO::SIOTPCHitHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TPCHits . . . . . . . . . 246

SIO::SIOTrackerDataHandler Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TrackerDatas . . . . . . . 246 SIO::SIOTrackerHitHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of SimTrackerHits . . . . . . 247

SIO::SIOTrackerHitPlaneHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TrackerHitPlane . . . . . 248

SIO::SIOTrackerHitZCylinderHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TrackerHitZCylinder … 249

SIO::SIOTrackerPulseHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TrackerPulses . . . . . . 250

SIO::SIOTrackerRawDataHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TrackerRawDatas …. 250

SIO::SIOTrackHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of Tracks . . . . . . . . . . 251

SIO::SIOVertexHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of Vertexes . . . . . . . . . 252

SIO::SIOWriter

Concrete implementation of LCWriter using SIO (p. 30) . . . . . . . . . . . . . . . . . . . . . 253

SomeClass . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 TEST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258 lcrtrel_helper::ToRelation< U >

Helper class for relations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258

EVENT::TPCHit

Deprecated . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258

IMPL::TPCHitImpl

Implementation of the real data TPCHit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260

IOIMPL::TPCHitIOImpl

Adding stuff needed for io (friend declarations, etc.) . . . . . . . . . . . . . . . . . . . . . . . 262

EVENT::Track

The LCIO (p. 115) track class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 263

EVENT::TrackerData

TrackerData (p. 267) contains the corrected (calibrated) raw tracker data . . . . . . . . . . . . 267

IMPL::TrackerDataImpl

Default implementation of TrackerData . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269

IOIMPL::TrackerDataIOImpl

Adding stuff needed for io (friend declarations, etc.) . . . . . . . . . . . . . . . . . . . . . . . 270

EVENT::TrackerHit

A generic tracker hit to be used by pattern recognition . . . . . . . . . . . . . . . . . . . . . . 271

IMPL::TrackerHitImpl

Implementation of the generic tracker hit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 273

IOIMPL::TrackerHitIOImpl

Adding stuff needed for io (friend declarations, etc.) . . . . . . . . . . . . . . . . . . . . . . . 276

x,y,z, u(theta, phi), v(theta,phi), du, dv where (x,y,z) is the position in global cartesian coordinates, u,v are the the meassurement directions (unit vectors, spanning the plane) and du,dv are the measurement errors along these directions 277 IMPL::TrackerHitPlaneImpl

Implementation of the planar tracker hit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279

IOIMPL::TrackerHitPlaneIOImpl

Adding stuff needed for io (friend declarations, etc.) . . . . . . . . . . . . . . . . . . . . . . . 283

x,y,z, center(x,y), dRPhi, dz where (x,y,z) is the position in global cartesian coordinates, center(x,y) is the center of the cylinder and Rphi and z are the measurement errors 283 IMPL::TrackerHitZCylinderImpl

Implementation of the tracker hit on a cylindrical surface parallel to z . . . . . . . . . . . . . . 285

IOIMPL::TrackerHitZCylinderIOImpl

Adding stuff needed for io (friend declarations, etc.) . . . . . . . . . . . . . . . . . . . . . . . 288

EVENT::TrackerPulse Tracker pulses as computed from TrackerData (p. 267) objects or as directly measured by a specific subdetector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 289 IMPL::TrackerPulseImpl

Default implementation of TrackerPulse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291

IOIMPL::TrackerPulseIOImpl

Adding stuff needed for io (friend declarations, etc.) . . . . . . . . . . . . . . . . . . . . . . . 293

EVENT::TrackerRawData

Generic class for raw tracker data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294

IMPL::TrackerRawDataImpl

Default implementation of TrackerRawData . . . . . . . . . . . . . . . . . . . . . . . . . . . 295

IOIMPL::TrackerRawDataIOImpl

Adding stuff needed for io (friend declarations, etc.) . . . . . . . . . . . . . . . . . . . . . . . 297

IMPL::TrackImpl

Implementation of the LCIO track class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298

IOIMPL::TrackIOImpl

Adding stuff needed for io (friend declarations, etc.) . . . . . . . . . . . . . . . . . . . . . . . 304

EVENT::TrackState

The LCIO (p. 115) TrackState (p. 304) class . . . . . . . . . . . . . . . . . . . . . . . . . . . 304

IMPL::TrackStateImpl

Implementation of the LCIO TrackState class . . . . . . . . . . . . . . . . . . . . . . . . . . 307

IOIMPL::TrackStateIOImpl

Adding stuff needed for io (friend declarations, etc.) . . . . . . . . . . . . . . . . . . . . . . . 309

TrkCluLink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310 UTIL::UnknownAlgorithm

Exception for unknown algorithms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 311

SIO::SIORecords::Unpack

Helper class that sets the unpack flag of known SIO (p. 30) records as specified by the constructor 311

UserClass . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 UserStruct . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 EVENT::Vertex

For LCIO (p. 115) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312

IMPL::VertexImpl

Implementation of Vertex Class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 314

IOIMPL::VertexIOImpl Adding stuff needed for io (friend declarations, etc.) . . . . . . . . . . . . . . . . . . . . . . . 316

Namespace Documentation

6.1 DATA Namespace Reference

The namespace DATA (p. 21) has been removed from LCIO all data entities are now defined in EVENT (p. 21).

6.1.1 Detailed Description

The namespace DATA (p. 21) has been removed from LCIO all data entities are now defined in EVENT (p. 21).

6.2 EVENT Namespace Reference

The namespace EVENT (p. 21) holds all base interfaces of the event entities.

Classes

• class CalorimeterHit

The generic calorimeter hit - created from SimCalorimeterHit (p. 205) or RawCalorimeterHit (p. 188).

• class Cluster

The LCIO (p. 115) cluster.

• class LCCollection

The generic collection used in LCIO (p. 115).

• class LCEvent

The main event interface.

• class LCFlag

Helper class to create and interpret the 32-bit flag word in LCCollections.

• class LCFloatVec

Float vector used for user extensions.

• class LCGenericObject

Simple interface to store generic user data.

• class LCIntVec

Int vector used for user extensions.

• class LCIO

Global constants used in LCIO (p. 115).

• class LCObject

The generic object that is held in an LCCollection (p. 83).

• class LCParameters

22  Namespace Documentation

Simple interface to store generic named parameters of type int, float and string.

• class LCRelation

A single weighted relationship between two LCObjects.

• class LCRunHeader

Interface for the run header.

• class LCStrVec

String vector used for user extensions.

• class MCParticle

The LCIO (p. 115) Monte Carlo particle.

• class ParticleID

Persistent interface for LCIO (p. 115) ParticleIDs (p. 185).

• class RawCalorimeterHit

The generic calorimeter hit for real data (or simulation thereof).

• class ReconstructedParticle

The LCIO (p. 115) reconstructedParticle.

• class SimCalorimeterHit

The generic simulated calorimeter hit.

• class SimTrackerHit

A generic simulated tracker hit.

• class TPCHit

Deprecated.

• class Track

The LCIO (p. 115) track class.

• class TrackerData

TrackerData (p. 267) contains the corrected (calibrated) raw tracker data.

• class TrackerHit

A generic tracker hit to be used by pattern recognition.

• class TrackerHitPlane

A tracker hit on a planar surface - the hit is stored as: x,y,z, u(theta, phi), v(theta,phi), du, dv where (x,y,z) is the position in global cartesian coordinates, u,v are the the meassurement directions (unit vectors, spanning the plane) and du,dv are the measurement errors along these directions.

• class TrackerHitZCylinder

A tracker hit on a cylindrical surface that runs parallel to the z-axis - the hit is stored as: x,y,z, center(x,y), dRPhi, dz where (x,y,z) is the position in global cartesian coordinates, center(x,y) is the center of the cylinder and Rphi and z are the measurement errors.

• class TrackerPulse

Tracker pulses as computed from TrackerData (p. 267) objects or as directly measured by a specific subdetector.

• class TrackerRawData

Generic class for raw tracker data.

• class TrackState

The LCIO (p. 115) TrackState (p. 304) class.

• class Vertex

The Vertex (p. 312) class for LCIO (p. 115).

• class Exception

Base exception class for LCIO (p. 115) - all other exceptions extend this.

• class EventException

EventException (p. 69) used for errors accessing the event data.

• class DataNotAvailableException

EventException (p. 69) used for data not available.

• class ReadOnlyException

EventException (p. 69) used for signaling a ’read only exception’.

Typedefs

• typedef std::vector

< CalorimeterHit ∗ > CalorimeterHitVec

Vector of (pointers to) CalorimeterHits.

• typedef std::vector< Cluster ∗ > ClusterVec

Vector of (pointers to) Clusters.

• typedef std::vector< LCObject ∗ > LCObjectVec

Vector of (pointers to) LCObjects.

• typedef std::vector< MCParticle ∗ > MCParticleVec

Vector of (pointers to) MCParticles.

• typedef std::vector< ParticleID ∗ > ParticleIDVec

Vector of (pointers to) ParticleIDs (p. 185).

• typedef std::vector

< RawCalorimeterHit ∗ > RawCalorimeterHitVec

Vector of (pointers to) RawCalorimeterHits.

• typedef std::vector

< ReconstructedParticle ∗ > ReconstructedParticleVec

Vector of (pointers to) ReconstructedParticles.

• typedef std::vector< Track ∗ > TrackVec

Vector of (pointers to) Tracks.

• typedef std::vector< TrackerHit ∗ > TrackerHitVec

Vector of (pointers to) TrackerHits.

• typedef std::vector< TrackState ∗ > TrackStateVec

Vector of (pointers to) TrackStates.

• typedef std::vector< Vertex ∗ > VertexVec

Vector of (pointers to) Vertices.

• typedef std::vector< std::string > StringVec

Vector of strings.

• typedef std::vector< float > FloatVec

Vector of floats.

• typedef std::vector< double > DoubleVec

Vector of doubles.

• typedef std::vector< int > IntVec

Vector of ints.

• typedef std::vector< short > ShortVec

Vector of shorts.

• typedef long long long64

64 bit signed integer,e.g.to be used for timestamps

• typedef unsigned long long ulong64

64 bit unsigned integer,e.g.to be used for masks

6.2.1 Detailed Description

The namespace EVENT (p. 21) holds all base interfaces of the event entities. Typedefs used in LCIO (p. 115). Typedefs for std::vectors of basic types used in LCIO (p. 115).

6.2.2 Typedef Documentation

6.2.2.1 typedef std::vector<CalorimeterHit∗> EVENT::CalorimeterHitVec

Vector of (pointers to) CalorimeterHits.

24  Namespace Documentation

6.2.2.2 typedef std::vector<Cluster∗> EVENT::ClusterVec Vector of (pointers to) Clusters.

6.2.2.3 typedef std::vector< double > EVENT::DoubleVec Vector of doubles.

6.2.2.4 typedef std::vector< float > EVENT::FloatVec Vector of floats.

6.2.2.5 typedef std::vector< int > EVENT::IntVec Vector of ints.

6.2.2.6 typedef std::vector<LCObject∗> EVENT::LCObjectVec Vector of (pointers to) LCObjects.

6.2.2.7 typedef std::vector<MCParticle∗> EVENT::MCParticleVec Vector of (pointers to) MCParticles.

6.2.2.8 typedef std::vector<ParticleID∗> EVENT::ParticleIDVec Vector of (pointers to) ParticleIDs (p. 185).

6.2.2.9 typedef std::vector<RawCalorimeterHit∗> EVENT::RawCalorimeterHitVec Vector of (pointers to) RawCalorimeterHits.

6.2.2.10 typedef std::vector<ReconstructedParticle∗> EVENT::ReconstructedParticleVec Vector of (pointers to) ReconstructedParticles.

6.2.2.11 typedef std::vector< short > EVENT::ShortVec Vector of shorts.

6.2.2.12 typedef std::vector< std::string > EVENT::StringVec Vector of strings.

6.2.2.13 typedef std::vector<TrackerHit∗> EVENT::TrackerHitVec Vector of (pointers to) TrackerHits.

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

6.2.2.14 typedef std::vector<TrackState∗> EVENT::TrackStateVec

Vector of (pointers to) TrackStates.

6.2.2.15 typedef std::vector<Track∗> EVENT::TrackVec

Vector of (pointers to) Tracks.

6.2.2.16 typedef std::vector<Vertex∗> EVENT::VertexVec

Vector of (pointers to) Vertices.

6.3 IMPL Namespace Reference

The namespace IMPL (p. 25) holds the implementations of the base interfaces that are defined in EVENT (p. 21).

Classes

• class AccessChecked

Controls access to objects.

• class CalorimeterHitImpl

Implementation of the real data CalorimeterHit.

• class ClusterImpl

Implementation of Cluster.

• class LCCollectionVec

Implementation of the LCCollection using (inheriting from) an STL vector of LCObjects.

• class LCEventImpl

Implementation of the main event class.

• class LCFlagImpl

Implementation of helper class to create and interpret the 32-bit flag word in LCCollections.

• class LCGenericObjectImpl

Default LCIO implementation of the interface to store generic user data.

• class LCIOExceptionHandler

Sets the default error handling for unexpected and uncaught exceptions.

• class LCParametersImpl

Implementation of Simple interface to store generic named parameters of type int, float and string.

• class LCRelationImpl

Implementation of the LCRelation.

• class LCRunHeaderImpl

Implementation of LCRunHeader.

• class MCParticleImpl

Implementation of MCParticle.

• class PIDSort

Helper class to sort ParticleIDs (p. 185) wrt.

• class ParticleIDImpl

Implementation of ParticleID.

• class RawCalorimeterHitImpl

Implementation of the real data RawCalorimeterHit.

• class ReconstructedParticleImpl

Implementation of ReconstructedParticle.

• struct MCParticleCont

• class SimCalorimeterHitImpl

Implementation of the generic SimCalorimeterHit.

• class SimTrackerHitImpl

Implementation of SimTrackerHit.

• class TPCHitImpl

Implementation of the real data TPCHit.

• class TrackerDataImpl

Default implementation of TrackerData.

• class TrackerHitImpl

Implementation of the generic tracker hit.

• class TrackerHitPlaneImpl

Implementation of the planar tracker hit.

• class TrackerHitZCylinderImpl

Implementation of the tracker hit on a cylindrical surface parallel to z.

• class TrackerPulseImpl

Default implementation of TrackerPulse.

• class TrackerRawDataImpl

Default implementation of TrackerRawData.

• class TrackImpl

Implementation of the LCIO track class.

• class TrackStateImpl

Implementation of the LCIO TrackState class.

• class VertexImpl

Implementation of Vertex Class.

Typedefs

• typedef std::map< std::string, EVENT::LCCollection ∗ > LCCollectionMap

• typedef std::set < EVENT::LCCollection ∗ > LCCollectionSet

• typedef std::map< std::string, EVENT::IntVec > IntMap

• typedef std::map< std::string, EVENT::FloatVec > FloatMap

• typedef std::map< std::string, EVENT::StringVec > StringMap

• typedef std::vector < IMPL::MCParticleCont ∗ > MCParticleContVec

Functions

• void lcio_unexpected ()

Method catches any std::exception, prints the message to stdout and then exits the program.

6.3.1 Detailed Description

The namespace IMPL (p. 25) holds the implementations of the base interfaces that are defined in EVENT (p. 21).

6.4 IO Namespace Reference  27

6.4 IO Namespace Reference

The namespace IO (p. 27) holds the base interfaces for io of data.

Classes

• class ILCFactory

Factory to create LCWriter (p. 166) and Reader object for the known data formats, e.g.

• class LCEventListener

Listener for the occurence of LCEvents when reading a stream.

• class LCReader

Interface for reading data from LCIO.

• class LCRunListener

Listener for the occurence of LCRunHeader when reading a stream.

• class LCWriter

Interface for writing data with LCIO.

• class IOException

IOException (p. 79) used for reading/writing errors.

• class EndOfDataException

EndOfDataException (p. 68) for signaling the end of a data stream.

6.4.1 Detailed Description

The namespace IO (p. 27) holds the base interfaces for io of data.

6.5 IOIMPL Namespace Reference

The IOIMPL (p. 27) namespace holds extensions to the default implementations needed for IO (p. 27).

Classes

• class CalorimeterHitIOImpl

Adding stuff needed for io (friend declarations, etc.)

• class ClusterIOImpl

Adding stuff needed for io (friend declarations, etc.)

• class LCCollectionIOVec

Adding stuff needed for io (friend declarations, etc.)

• class LCEventIOImpl

Adding stuff needed for io (friend declarations, etc.)

• class LCFactory

Factory to create LCWriter and Reader object for the known data formats, e.g.

• class LCGenericObjectIOImpl

Adding stuff needed for io (friend declarations, etc.)

• class LCRelationIOImpl

Adding stuff needed for io (friend declarations, etc.)

• class LCRunHeaderIOImpl

Adding stuff needed for io (friend declarations, etc.)

• class MCParticleIOImpl

Adding stuff needed for io (friend declarations, etc.)

• class ParticleIDIOImpl

Adding stuff needed for io (friend declarations, etc.)

• class RawCalorimeterHitIOImpl

Adding stuff needed for io (friend declarations, etc.)

• class ReconstructedParticleIOImpl

Adding stuff needed for io (friend declarations, etc.)

• class SimCalorimeterHitIOImpl

Adding stuff needed for io (friend declarations, etc.)

• class SimTrackerHitIOImpl

Adding stuff needed for io (friend declarations, etc.)

• class TPCHitIOImpl

Adding stuff needed for io (friend declarations, etc.)

• class TrackerDataIOImpl

Adding stuff needed for io (friend declarations, etc.)

• class TrackerHitIOImpl

Adding stuff needed for io (friend declarations, etc.)

• class TrackerHitPlaneIOImpl

Adding stuff needed for io (friend declarations, etc.)

• class TrackerHitZCylinderIOImpl

Adding stuff needed for io (friend declarations, etc.)

• class TrackerPulseIOImpl

Adding stuff needed for io (friend declarations, etc.)

• class TrackerRawDataIOImpl

Adding stuff needed for io (friend declarations, etc.)

• class TrackIOImpl

Adding stuff needed for io (friend declarations, etc.)

• class TrackStateIOImpl

Adding stuff needed for io (friend declarations, etc.)

• class VertexIOImpl

Adding stuff needed for io (friend declarations, etc.)

6.5.1 Detailed Description

The IOIMPL (p. 27) namespace holds extensions to the default implementations needed for IO (p. 27). For internal use only - with the exception of IOIMPL::LCFactory (p. 101).

6.6 lcio Namespace Reference

The LCIO namespace combines EVENT (p. 21), IO (p. 27), IMPL (p. 25) and UTIL (p. 32) for user convenience.

6.6.1 Detailed Description

The LCIO namespace combines EVENT (p. 21), IO (p. 27), IMPL (p. 25) and UTIL (p. 32) for user convenience.

6.7 lcrtrel Namespace Reference

The namespace lcrtrel holds classes and functions for runtime extensions and relations see LCRTRelations (p. 142) for more detailed documentation.

6.7 lcrtrel Namespace Reference  29

Classes

• struct LCExtension

Simple Extension - pointer to an object of type T.

• class LCOwnedExtension

Simple Extension - pointer to an object of type T where the ownership is taken over by the object holding the extension, i.e.

• class LCExtensionVector

Extension vector holding pointers to objects of type T - no ownership of the objects is taken.

• class LCOwnedExtensionVector

Extension vector holding pointers to objects of type T - ownership of the objects is taken, i.e.

• class LCExtensionList

Extension list holding pointers to objects of type T - no ownership of the objects is taken.

• class LCOwnedExtensionList

Extension list holding pointers to objects of type T - ownership of the objects is taken, i.e.

• struct LC1To1Relation

One to one relation between two objects of type From and To.

• struct LC1ToNRelation

One to many relation between one object of type From to many objects of type To.

• struct LCNToNRelation

Many to many relation between objects of type From to objects of type To.

• struct LCIntExtension

Special Extension that allows to write int extensions directly (not through a pointer !).

• struct LCFloatExtension

• class LCRTRelations

Base class that provides run time (user) extensions and relation between objects.

Functions

• template<class R >

void set_relation (typename R::from::obj_ptr f, typename R::to::obj_ptr t)

Set the 1-to-1 relation between two objects - prexisting inconsistent relations involving the two objects are deleted to enforce a consistent set of from-to relations.

• template<class R >

void unset_relation (typename R::from::obj_ptr f)

Unset the 1-to-1 relation from f.

• template<class R >

void add_relation (typename R::from::obj_ptr f, typename R::to::obj_ptr t)

Add a link from f to t to an N-to-N relation ship.

• template<class R >

void remove_relation (typename R::from::obj_ptr f, typename R::to::obj_ptr t)

Remove the link from from f to t from the N-to-N relation ship.

• template<class R >

void remove_relations (typename R::from::obj_ptr f)

Removes all relations from the given object.

• template<class R >

void merge_relations (typename R::from::obj_ptr f1, typename R::from::obj_ptr f2)

Merge the relations from f2 to f1 - after this call f1 will hold all the relations and f2 will be empty.

6.7.1 Detailed Description

The namespace lcrtrel holds classes and functions for runtime extensions and relations see LCRTRelations (p. 142) for more detailed documentation.

6.7.2 Function Documentation

6.7.2.1 template<class R > void lcrtrel::set_relation ( typename R::from::obj_ptr f, typename R::to::obj_ptr t )

Set the 1-to-1 relation between two objects - prexisting inconsistent relations involving the two objects are deleted to enforce a consistent set of from-to relations.

6.8 SIO Namespace Reference

The SIO (p. 30) namespace holds the persistency implementation using SIO (p. 30).

Classes

• struct RunEvent

Helper struct that stores run and event positions in the file.

• class LCIORandomAccess

Implementation class for LCIORandomAccess (p. 118) records.

• class LCIORandomAccessMgr

Manager class for LCIO direct access.

• class SIORecords

Manager class that holds instances of all known LCIO-SIO records.

• class LCSIO

Collection of static helper functions for reading and writing data with SIO (p. 30).

• class RunEventMap

Map that holds positions of Run and Event records.

• class SIOCalHitHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of CalorimeterHits.

• class SIOClusterHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of Clusters.

• class SIOCollectionHandler

Handler for LCCollection/LCCollectionIOVec objects for SIO (p. 30).

• class SIOEventHandler

Handler for LCEvent/LCEventIOImpl objects.

• class SIOFloatVecHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of FloatVecs.

• class SIOHandlerMgr

Singleton that holds instances of all types of SIOObjectHandlers.

• class SIOIndexHandler

Handler for LCIOrandomAccess and LCIOIndex objects/blocks.

• class SIOIntVecHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of IntVecs.

• class SIOLCGenericObjectHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of LCGenericObjects.

• class SIOLCParameters

IO (p. 27) of LCParameters.

• class SIOLCRelationHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of CalorimeterHits.

• class SIOObjectHandler

Interface for all lcio object SIO-handlers, has to be implemented for all event entities (hits, tracks, clusters,...).

• class SIOParticleHandler

6.8 SIO Namespace Reference  31

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of MCParticles.

• class SIORandomAccessHandler

Handler for LCIOrandomAccess and LCIOIndex objects/blocks.

• class SIORawCalHitHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of RawCalorimeterHits.

• class SIOReader

Concrete implementation of LCWriter using SIO (p. 30).

• class SIOReconstructedParticleHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of ReconstructedParticles.

• class SIORunHeaderHandler

Handler for LCRunHeader/LCRunHeaderImpl objects.

• class SIOSimCalHitHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of SimCalorimeterHits.

• class SIOSimTrackHitHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of SimTrackerHits.

• class SIOStrVecHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of StrVecs.

• class SIOTPCHitHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TPCHits.

• class SIOTrackerDataHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TrackerDatas.

• class SIOTrackerHitHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of SimTrackerHits.

• class SIOTrackerHitPlaneHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TrackerHitPlane.

• class SIOTrackerHitZCylinderHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TrackerHitZCylinder.

• class SIOTrackerPulseHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TrackerPulses.

• class SIOTrackerRawDataHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TrackerRawDatas.

• class SIOTrackHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of Tracks.

• class SIOVertexHandler

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of Vertexes.

• class SIOWriter

Concrete implementation of LCWriter using SIO (p. 30).

Typedefs

• typedef std::map< std::string, SIOObjectHandler ∗ > SIOHandlerMap

Functions

• std::ostream & operator<< (std::ostream &os, const RunEvent &re)

• bool operator< (const RunEvent &r0, const RunEvent &other)

• std::ostream & operator<< (std::ostream &os, const LCIORandomAccess &ra)

• bool operator< (const LCIORandomAccess ra0, const LCIORandomAccess &other)

• std::ostream & operator<< (std::ostream &os, const LCIORandomAccessMgr &ra)

• std::ostream & operator<< (std::ostream &os, const RunEventMap &rm)

6.8.1 Detailed Description

The SIO (p. 30) namespace holds the persistency implementation using SIO (p. 30). Users should not use any of the classes defined here explicitly but through their base interfaces defined in IO (p. 27).

6.9 UTIL Namespace Reference

The namespace UTIL (p. 32) holds convenient and support classes and methods for the LCIO objects.

Classes

• class LCTokenizer

Helper class for string tokenization.

• class BitFieldValue

Helper class for BitField64 (p. 43) that corresponds to one field value.

• class BitField64

A bit field of 64bits that allows convenient declaration and manipulation of sub fields of various widths.

• class BitSet32

Convenient helper class for setting single bits in a 32bit-field, such as collection flags or hit types - extends std- ::bitset<32>.

• class CellIDDecoder

Convenient class for decoding cellIDs from collection parameter LCIO::CellIDEncoding.

• class CellIDEncoder

Specialization for SimTrackerHits that have only one cellID.

• class CollectionParameterMap

Helper class for setting and retrieving an std::map<string,int> as collection parameters.

• struct ILDCellID0

Helper class to define constants for the canonical encoding of the CellID0 in tracking sub detectors in ILD (or ILD like detectors).

• class ILDCellIDEncoder

Special wrapper to the CellIDEncoder<T> that enforces cellID0 to use the encoder string defined in ILDCellID0- ::encoder_string (p. 73).

• struct ILDDetID

Define integer constants for identifying ILD sub detectors.

• struct ILDTrkHitTypeBit

Define integer constants for bits of the the lcio tracker hit type used in the context of ILD reconstruction code.

• struct ILDTrkHitQualityBit

Define integer constants for bits of the the lcio tracker hit quality used in the context of ILD reconstruction code.

• class IndexMap

utility class to manage indexes according to Collection Parameters

• class LCFixedObject

Template class for fixed size LCGenericObject subclasses.

• class LCFourVector

Four vector used in LCIO.

• class LCIterator

Simple convenient iterator class for LCCollections that saves some boiler plate code.

• class LCObjectHandle

Template handle class for LCObjects.

• class LCRelationNavigator

The LCRelationNavigator (p. 139) makes repeated lookup of relations more conveneient and efficient.

6.9 UTIL Namespace Reference  33

• class LCSplitWriter

LCWriter wrapper that automatically splits files if a given number of bytes is exceeded.

• class LCStdHepRdr

Basic utility for reading a binary stdhep file and filling a LCCollectionVec with MCParticles containing the stdhep file information.

• class LCStdHepRdrNew

Basic utility for reading a binary stdhep file and filling a LCCollectionVec with MCParticles containing the stdhep file information.

• class LCTime

Helper class that allows to convert time stamps as defined in LCEvent::getTimeStamp() ( ns since 1.1.1970 00:00:00 UTC) to ordinary date and time and vice versa.

• class LCTOOLS

Provides some basic functions for printing event data.

• class LCTypedVector

Templated helper class that is an std::vector<T∗> of the elements in the LCCollection.

• class LCWarning

Utility class to show warnings in LCIO.

• class lcio_short

• class LCIO_LONG

• class PIDHandler

Convenient class for setting and retrieving particle id information attached to ReconstructedParticles.

• class UnknownAlgorithm

Exception for unknown algorithms.

Enumerations

• enum ObjectType { Cluster = 1, ReconstructedParticle }

Functions

• std::ostream & operator<< (std::ostream &os, const BitField64 &b)

Operator for dumping BitField64 (p. 43) to streams.

• BitSet32 make_bitset32 (int bit0)

Convenient helper that creates a BitSet32 (p. 47) with bit0 set.

• BitSet32 make_bitset32 (int bit0, int bit1)

Convenient helper that creates a BitSet32 (p. 47) with bit0 and bit1 set.

• BitSet32 make_bitset32 (int bit0, int bit1, int bit2)

Convenient helper that creates a BitSet32 (p. 47) with bit0 - bit2 set.

• BitSet32 make_bitset32 (int bit0, int bit1, int bit2, int bit3)

Convenient helper that creates a BitSet32 (p. 47) with bit0 - bit3 set.

• BitSet32 make_bitset32 (int bit0, int bit1, int bit2, int bit3, int bit4)

Convenient helper that creates a BitSet32 (p. 47) with bit0 - bit4 set.

• int set_bit (int flag, int bit)

Convenient helper to set a bit in a given int - for example: trkHit->setType( set_bit( trkHit->getType() , UTIL::ILDTrkHitTypeBit::ONE_DIMENSIONAL ) ) ;.

• int unset_bit (int flag, int bit)

Convenient helper to unset a bit in a given int - for example: trkHit->setType( set_bit( trkHit->getType() , UTIL::ILDTrkHitTypeBit::ONE_DIMENSIONAL ) ) ;.

• template<class T >

int CellIDEncoder_cellID1Bit ()

Helper function that returns the bit for cellid1 through template specialization or -1 if no cellid1 exists.

• template<>

int CellIDEncoder_cellID1Bit< EVENT::SimCalorimeterHit > ()

specialization that returns the proper bit for the second cellid

• template<>

int CellIDEncoder_cellID1Bit< EVENT::RawCalorimeterHit > ()

specialization that returns the proper bit for the second cellid

• template<>

int CellIDEncoder_cellID1Bit< EVENT::CalorimeterHit > ()

specialization that returns the proper bit for the second cellid

• template<>

int CellIDEncoder_cellID1Bit< EVENT::TrackerData > ()

specialization that returns the proper bit for the second cellid

• template<>

int CellIDEncoder_cellID1Bit< EVENT::TrackerPulse > ()

specialization that returns the proper bit for the second cellid

• template<>

int CellIDEncoder_cellID1Bit< EVENT::TrackerRawData > ()

specialization that returns the proper bit for the second cellid

• template<>

int CellIDEncoder_cellID1Bit< EVENT::TrackerHit > ()

specialization that returns the proper bit for the second cellid

• template<>

int CellIDEncoder_cellID1Bit< EVENT::TrackerHitPlane > ()

specialization that returns the proper bit for the second cellid

• template<>

int CellIDEncoder_cellID1Bit< EVENT::TrackerHitZCylinder > ()

specialization that returns the proper bit for the second cellid

• template<>

int CellIDEncoder_cellID1Bit< EVENT::SimTrackerHit > ()

specialization that returns the proper bit for the second cellid

• template<class T >

void CellIDEncoder_setCellID (T ∗hit, int low, int high)

Helper function that sets cellid1 and cellid2.

• template<class T >

const char ∗ lctypename ()

Template that returns the LCIO type name as used in the LCCollctions (and files), e.g.

• template<>

const char ∗ lctypename< EVENT::MCParticle > ()

• template<>

const char ∗ lctypename< EVENT::SimCalorimeterHit > ()

• template<>

const char ∗ lctypename< EVENT::CalorimeterHit > ()

• template<>

const char ∗ lctypename< EVENT::RawCalorimeterHit > ()

• template<>

const char ∗ lctypename< EVENT::SimTrackerHit > ()

• template<>

const char ∗ lctypename< EVENT::TrackerRawData > ()

• template<>

const char ∗ lctypename< EVENT::TrackerData > ()

• template<>

const char ∗ lctypename< EVENT::TrackerPulse > ()

• template<>

const char ∗ lctypename< EVENT::TrackerHit > ()

• template<>

const char ∗ lctypename< EVENT::TrackerHitPlane > ()

• template<>

const char ∗ lctypename< EVENT::TrackerHitZCylinder > ()

• template<>

const char ∗ lctypename< EVENT::LCStrVec > ()

• template<>

const char ∗ lctypename< EVENT::LCFloatVec > ()

• template<>

const char ∗ lctypename< EVENT::LCIntVec > ()

• template<>

const char ∗ lctypename< EVENT::Track > ()

• template<>

const char ∗ lctypename< EVENT::Cluster > ()

• template<>

const char ∗ lctypename< EVENT::ReconstructedParticle > ()

• template<>

const char ∗ lctypename< EVENT::LCRelation > ()

• template<>

const char ∗ lctypename< EVENT::LCGenericObject > ()

• template<>

const char ∗ lctypename< EVENT::Vertex > ()

• template<class T >

const char ∗ lctypename (const T ∗t)

Template that returns the LCIO type name as used in the LCCollctions (and files) based on the object’s type, e.g.

• template<>

const char ∗ lctypename< EVENT::LCObject > (const EVENT::LCObject ∗o)

Specialization for polymorphic LCObject pointers - slow.

• template<class T >

lcio_short< T > lcshort (const T ∗t, const EVENT::LCCollection ∗b=NULL)

• const std::string & header (const EVENT::Vertex ∗)

operator for detailed output of a vertex object (e.g.

• const std::string & tail (const EVENT::Vertex ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::Vertex > &sV)

• std::ostream & operator<< (std::ostream &out, const EVENT::Vertex &v)

• const std::string & header (const EVENT::MCParticle ∗)

• const std::string & tail (const EVENT::MCParticle ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::MCParticle > &sV)

• std::ostream & operator<< (std::ostream &out, const EVENT::MCParticle &)

• const std::string & header (const EVENT::TrackerHit ∗)

• const std::string & tail (const EVENT::TrackerHit ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::TrackerHit > &sV)

• std::ostream & operator<< (std::ostream &out, const EVENT::TrackerHit &)

• const std::string & header (const EVENT::TrackerHitPlane ∗)

• const std::string & tail (const EVENT::TrackerHitPlane ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::TrackerHitPlane > &sV)

• std::ostream & operator<< (std::ostream &out, const EVENT::TrackerHitPlane &)

• const std::string & header (const EVENT::TrackerHitZCylinder ∗)

• const std::string & tail (const EVENT::TrackerHitZCylinder ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::TrackerHitZCylinder > &sV)

• std::ostream & operator<< (std::ostream &out, const EVENT::TrackerHitZCylinder &)

• const std::string & header (const EVENT::SimTrackerHit ∗)

• const std::string & tail (const EVENT::SimTrackerHit ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::SimTrackerHit > &sV)

• std::ostream & operator<< (std::ostream &out, const EVENT::SimTrackerHit &)

• const std::string & header (const EVENT::CalorimeterHit ∗)

• const std::string & tail (const EVENT::CalorimeterHit ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::CalorimeterHit > &sV)

• std::ostream & operator<< (std::ostream &out, const EVENT::CalorimeterHit &)

• const std::string & header (const EVENT::SimCalorimeterHit ∗)

• const std::string & tail (const EVENT::SimCalorimeterHit ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::SimCalorimeterHit > &sV)

• std::ostream & operator<< (std::ostream &out, const EVENT::SimCalorimeterHit &)

• const std::string & header (const EVENT::ReconstructedParticle ∗)

• const std::string & tail (const EVENT::ReconstructedParticle ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::ReconstructedParticle > &)

• std::ostream & operator<< (std::ostream &out, const EVENT::ReconstructedParticle &)

• const std::string & header (const EVENT::Track ∗)

• const std::string & tail (const EVENT::Track ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::Track > &)

• std::ostream & operator<< (std::ostream &out, const EVENT::Track &)

• const std::string & header (const EVENT::TrackState ∗)

• const std::string & tail (const EVENT::TrackState ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::TrackState > &)

• std::ostream & operator<< (std::ostream &out, const EVENT::TrackState &)

• const std::string & header (const EVENT::Cluster ∗)

• const std::string & tail (const EVENT::Cluster ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::Cluster > &)

• std::ostream & operator<< (std::ostream &out, const EVENT::Cluster &)

• const std::string & header (const EVENT::LCRelation ∗)

• const std::string & tail (const EVENT::LCRelation ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::LCRelation > &)

• std::ostream & operator<< (std::ostream &out, const EVENT::LCRelation &)

• const std::string & header (const EVENT::LCFloatVec ∗)

• const std::string & tail (const EVENT::LCFloatVec ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::LCFloatVec > &)

• std::ostream & operator<< (std::ostream &out, const EVENT::LCFloatVec &)

• template<class T >

const std::string & header ()

• template<class T >

const std::string & tail ()

• template<class T >

LCIO_LONG< T > lcio_long (const T &o, const EVENT::LCCollection ∗c)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::CalorimeterHit > l)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::Cluster > l)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::Track > l)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::TrackState > l)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::ReconstructedParticle > ll)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::SimCalorimeterHit > ll)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::TrackerHit > ll)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::TrackerHitPlane > ll)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::TrackerHitZCylinder > ll)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::SimTrackerHit > ll)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::MCParticle > ll)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::Vertex > ll)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::LCRelation > ll)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::LCFloatVec > ll)

• const std::string & header (const EVENT::LCCollection ∗)

• const std::string & tail (const EVENT::LCCollection ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::LCCollection > &)

• std::ostream & operator<< (std::ostream &out, const EVENT::LCCollection &)

• const std::string & header (const EVENT::LCEvent ∗)

• const std::string & tail (const EVENT::LCEvent ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::LCEvent > &)

• std::ostream & operator<< (std::ostream &out, const EVENT::LCEvent &)

• const std::string & header (const EVENT::LCFlag ∗)

• const std::string & tail (const EVENT::LCFlag ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::LCFlag > &)

• std::ostream & operator<< (std::ostream &out, const EVENT::LCFlag &)

• const std::string & header (const EVENT::LCGenericObject ∗, const EVENT::LCCollection ∗v=NULL)

• const std::string & tail (const EVENT::LCGenericObject ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::LCGenericObject > &)

• std::ostream & operator<< (std::ostream &out, const EVENT::LCGenericObject &)

• const std::string & header (const EVENT::LCIntVec ∗)

• const std::string & tail (const EVENT::LCIntVec ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::LCIntVec > &)

• std::ostream & operator<< (std::ostream &out, const EVENT::LCIntVec &)

• const std::string & header (const EVENT::LCObject ∗)

• const std::string & tail (const EVENT::LCObject ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::LCObject > &)

• std::ostream & operator<< (std::ostream &out, const EVENT::LCObject &)

• const std::string & header (const EVENT::LCParameters ∗)

• const std::string & tail (const EVENT::LCParameters ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::LCParameters > &)

• std::ostream & operator<< (std::ostream &out, const EVENT::LCParameters &)

• const std::string & header (const EVENT::LCRunHeader ∗)

• const std::string & tail (const EVENT::LCRunHeader ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::LCRunHeader > &)

• std::ostream & operator<< (std::ostream &out, const EVENT::LCRunHeader &)

• const std::string & header (const EVENT::ParticleID ∗)

• const std::string & tail (const EVENT::ParticleID ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::ParticleID > &)

• std::ostream & operator<< (std::ostream &out, const EVENT::ParticleID &)

• const std::string & header (const EVENT::RawCalorimeterHit ∗)

• const std::string & tail (const EVENT::RawCalorimeterHit ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::RawCalorimeterHit > &)

• std::ostream & operator<< (std::ostream &out, const EVENT::RawCalorimeterHit &)

• const std::string & header (const EVENT::TrackerData ∗)

• const std::string & tail (const EVENT::TrackerData ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::TrackerData > &)

• std::ostream & operator<< (std::ostream &out, const EVENT::TrackerData &)

• const std::string & header (const EVENT::TrackerPulse ∗)

• const std::string & tail (const EVENT::TrackerPulse ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::TrackerPulse > &)

• std::ostream & operator<< (std::ostream &out, const EVENT::TrackerPulse &)

• const std::string & header (const EVENT::TrackerRawData ∗)

• const std::string & tail (const EVENT::TrackerRawData ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::TrackerRawData > &)

• std::ostream & operator<< (std::ostream &out, const EVENT::TrackerRawData &)

• const std::string & header (const EVENT::LCIO ∗)

• const std::string & tail (const EVENT::LCIO ∗)

• std::ostream & operator<< (std::ostream &out, const UTIL::lcio_short< EVENT::LCIO > &)

• std::ostream & operator<< (std::ostream &out, const EVENT::LCIO &)

• template<class T >

std::string toString (const T ∗obj)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::LCCollection > l)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::LCEvent > l)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::LCFlag > l)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::LCGenericObject > l)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::LCIntVec > l)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::LCObject > l)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::LCParameters > l)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::LCRunHeader > l)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::LCIO > l)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::ParticleID > l)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::RawCalorimeterHit > l)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::TrackerData > l)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::TrackerPulse > l)

• std::ostream & operator<< (std::ostream &out, const LCIO_LONG< EVENT::TrackerRawData > l)

Variables

• static int MAX_HITS = 1000

6.9.1 Detailed Description

The namespace UTIL (p. 32) holds convenient and support classes and methods for the LCIO objects.

6.9.2 Function Documentation

6.9.2.1 template<> int UTIL::CellIDEncoder_cellID1Bit< EVENT::SimCalorimeterHit > ( )

specialization that returns the proper bit for the second cellid Helper function that returns the bit for cellid1 through template specialization or -1 if no cellid1 exists. specialization that returns the proper bit for the second cellid

6.9.2.2 const std::string & UTIL::header ( const EVENT::Vertex ∗ )

operator for detailed output of a vertex object (e.g. cout << vertexObj << endl; ) for printing the header of a vertex object (e.g. cout << header(vertexObj) << endl; ) for printing the tail of a vertex object (e.g. cout << tail(vertexObj) << endl; ) Referenced by UTIL::LCTOOLS::printCalorimeterHits(), UTIL::LCTOOLS::printClusters(), UTIL::LCTOOLS::print-LCFloatVecs(), UTIL::LCTOOLS::printLCGenericObjects(), UTIL::LCTOOLS::printLCIntVecs(), UTIL::LCTOOLS- ::printRawCalorimeterHits(), UTIL::LCTOOLS::printReconstructedParticles(), UTIL::LCTOOLS::printRelation(), U-TIL::LCTOOLS::printSimCalorimeterHits(), UTIL::LCTOOLS::printSimTrackerHits(), UTIL::LCTOOLS::printTracker-Data(), UTIL::LCTOOLS::printTrackerHitPlane(), UTIL::LCTOOLS::printTrackerHits(), UTIL::LCTOOLS::print-TrackerHitZCylinder(), UTIL::LCTOOLS::printTrackerPulse(), UTIL::LCTOOLS::printTrackerRawData(), UTIL::L-CTOOLS::printTracks(), and UTIL::LCTOOLS::printVertices().

6.9.2.3 template<class T > const char ∗ UTIL::lctypename ( )

Template that returns the LCIO type name as used in the LCCollctions (and files), e.g. Specialization for implemention classes. std::cout << lctypename<MCParticle>() (p. 39) << std::endl ; This can be used in user templates to create a new collection of the template type.

6.9.2.4 template<class T > const char∗ UTIL::lctypename ( const T ∗ t )

Template that returns the LCIO type name as used in the LCCollctions (and files) based on the object’s type, e.g. MCParticle∗ mcp ; std::cout << lctypename( mcp ) << std::endl ;

6.9.2.5 template<> const char∗ UTIL::lctypename< EVENT::LCObject > ( const EVENT::LCObject ∗ o )

Specialization for polymorphic LCObject pointers - slow. E.g.: LCObject∗ obj = new McParticle ; std::cout << lctypename( obj ) << std::endl ;

Class Documentation

7.1 IMPL::AccessChecked Class Reference

Controls access to objects.

#include <IMPL/AccessChecked.h>

Inheritance diagram for IMPL::AccessChecked:

Public Member Functions

• virtual int simpleUID () const

Protected Member Functions

• virtual void setReadOnly (bool readOnly)

• void checkAccess () throw ( EVENT::ReadOnlyException )

• void checkAccess (const char ∗what) throw ( EVENT::ReadOnlyException )

Protected Attributes

• bool _readOnly

• int _id

Friends

• class LCEventImpl

• class LCCollectionVec

7.1.1 Detailed Description

Controls access to objects.

Author gaede

Version Sep 4, 2003

The documentation for this class was generated from the following files:

• AccessChecked.h

• AccessChecked.cc

7.2 lcrtrel_helper::BiDirectional< From, To > Struct Template Reference

Helper class for biderectional relations provides the to and from type.

#include <LCRTRelations.h>

Public Types

• typedef From from

• typedef To to

7.2.1 Detailed Description

template<class From, class To>struct lcrtrel_helper::BiDirectional< From, To >

Helper class for biderectional relations provides the to and from type.

The documentation for this struct was generated from the following file:

• LCRTRelations.h

7.3 UTIL::BitField64 Class Reference

A bit field of 64bits that allows convenient declaration and manipulation of sub fields of various widths.

#include <UTIL/BitField64.h>

Inheritance diagram for UTIL::BitField64:

Public Types

• typedef std::map< std::string, unsigned int > IndexMap

Public Member Functions

• BitField64 (const std::string &initString)

The c’tor takes an initialization string of the form:

<fieldDesc>[,<fieldDesc>...]

fieldDesc = name:[start]:[-]length

where:

name: The name of the field start: The start bit of the field.

• lcio::long64 getValue () const

Returns the current 64bit value.

• void setValue (lcio::long64 value)

Set a new 64bit value.

• void reset ()

Reset - same as setValue(0) - useful if the same encoder is used for many objects.

• BitFieldValue & operator[] (size_t theIndex)

Acces to field through index.

• const BitFieldValue & operator[] (size_t theIndex) const

Const acces to field through index.

• unsigned highestBit () const

Highest bit used in fields [0-63].

• size_t size () const

Number of values.

• size_t index (const std::string &name) const

Index (p. 77) for field named ’name’.

• BitFieldValue & operator[] (const std::string &name)

Access to field through name .

• const BitFieldValue & operator[] (const std::string &name) const

Const Access to field through name .

• unsigned lowWord () const

The low word, bits 0-31.

• unsigned highWord () const

The high word, bits 32-63.

• std::string fieldDescription () const

Return a valid description string of all fields.

• std::string valueString () const

Return a string with a comma separated list of the current sub field values.

Protected Member Functions

• void addField (const std::string &name, unsigned offset, int width)

Add an additional field to the list.

• void init (const std::string &initString)

Decode the initialization string as described in the constructor.

• BitField64 ()

No default c’tor.

Protected Attributes

• std::vector< BitFieldValue ∗ > _fields

• lcio::long64 _value

• IndexMap _map

• lcio::long64 _joined

Friends

• std::ostream & operator<< (std::ostream &os, const BitField64 &b)

Operator for dumping BitField64 (p. 43) to streams.

7.3.1 Detailed Description

A bit field of 64bits that allows convenient declaration and manipulation of sub fields of various widths. Example: BitField64 (p. 43) b("layer:7,system:-3,barrel:3,theta:32:11,phi:11" ) ; b[ "layer" ] = 123 ; b[ "system" ] = -4 ; b[ "barrel" ] = 7 ; b[ "theta" ] = 180 ; b[ "phi" ] = 270 ; ... int theta = b["theta"] ; ... unsigned phiIndex = b.index("phi) ; int phi = b[ phiIndex ] ;

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.3.2 Constructor & Destructor Documentation

7.3.2.1 UTIL::BitField64::BitField64 ( const std::string & initString ) [inline]

The c’tor takes an initialization string of the form: <fieldDesc>[,<fieldDesc>...] fieldDesc = name:[start]:[-]length where: name: The name of the field start: The start bit of the field. If omitted assumed to start immediately following previous field, or at the least significant bit if the first field. length: The number of bits in the field. If preceeded by ’-’ the field is signed, otherwise unsigned. Bit numbering is from the least significant bit (bit 0) to the most significant (bit 63). Example: "layer:7,system:-3,barrel:3,theta:32:11,phi:11" References init().

7.3.3 Member Function Documentation

7.3.3.1 void UTIL::BitField64::init ( const std::string & initString ) [protected]

Decode the initialization string as described in the constructor.

See Also BitField64( const std::string& initString ) (p. 45)

References addField(). Referenced by BitField64(). The documentation for this class was generated from the following files:

• BitField64.h

• BitField64.cc

7.4 UTIL::BitFieldValue Class Reference

Helper class for BitField64 (p. 43) that corresponds to one field value.

#include <UTIL/BitField64.h>

Public Member Functions

• BitFieldValue (lcio::long64 &bitfield, const std::string &name, unsigned offset, int signedWidth)

The default c’tor.

• lcio::long64 value () const

Returns the current field value.

• BitFieldValue & operator= (lcio::long64 in)

Assignment operator for user convenience.

• operator lcio::long64 () const

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Conversion operator for lcio::long64 - allows to write:

lcio::long64 index = myBitFieldValue ;.

• const std::string & name () const

fg: removed because it causes ambiguities with operator lcio::long64().

• unsigned offset () const

The field’s offset.

• unsigned width () const

The field’s width.

• bool isSigned () const

True if field is interpreted as signed.

• lcio::ulong64 mask () const

The field’s mask.

Protected Attributes

• lcio::long64 & _b

• lcio::ulong64 _mask

• std::string _name

• unsigned _offset

• unsigned _width

• int _minVal

• int _maxVal

• bool _isSigned

7.4.1 Detailed Description

Helper class for BitField64 (p. 43) that corresponds to one field value.

7.4.2 Constructor & Destructor Documentation

7.4.2.1 UTIL::BitFieldValue::BitFieldValue ( lcio::long64 & bitfield, const std::string & name, unsigned offset, int signedWidth )

The default c’tor. Parameters

bitfield

reference to the 64bit bitfield

offset offset of field

signedWidth

width of field, negative if field is signed

References offset().

7.4.3 Member Function Documentation

7.4.3.1 const std::string& UTIL::BitFieldValue::name ( ) const [inline]

fg: removed because it causes ambiguities with operator lcio::long64().

Conversion operator for int - allows to write:

int index = myBitFieldValue ;The field’s name

The documentation for this class was generated from the following files:

• BitField64.h

• BitField64.cc

7.5 UTIL::BitSet32 Class Reference

Convenient helper class for setting single bits in a 32bit-field, such as collection flags or hit types - extends std- ::bitset<32>.

#include <UTIL/BitSet32.h>

Inheritance diagram for UTIL::BitSet32:

Public Member Functions

• BitSet32 (int val)

Allow to construct this from an int.

• BitSet32 & operator= (int val)

Allow assignment from an int.

• operator int ()

Direct conversion to int.

7.5.1 Detailed Description

Convenient helper class for setting single bits in a 32bit-field, such as collection flags or hit types - extends std- ::bitset<32>.

Use convenient methods make_bitset32(...) to create a bit set with some bits set.

Examples:

  // set collection flags :
  LCCollectionVec* calVec = new LCCollectionVec( LCIO::SIMCALORIMETERHIT )  ;
  calVec->setFlag( UTIL::make_bitset32(  LCIO::CHBIT_LONG, LCIO::CHBIT_STEP ) );

  // check TrackerHit type:
  BitSet32 (p. 47) type( hit->getType() ) ;
  bool isComposite = type.test( ILDTrkHitTypeBit::COMPOSITE_SPACEPOINT ) ; // test w/ range check
  // or shorter:
  if( BitSet32 (p. 47)( hit->getType() )[ ILDTrkHitTypeBit::COMPOSITE_SPACEPOINT ] ) {...}

The documentation for this class was generated from the following file:

• BitSet32.h

7.6 UTIL::LCTime::CalendarTime Struct Reference

Helper struct that holds the calendar time.

#include <UTIL/LCTime.h>

Public Attributes

• short year

• short month

• short day

• short hour

• short min

• short sec

• int ns

7.6.1 Detailed Description

Helper struct that holds the calendar time. The documentation for this struct was generated from the following file:

• LCTime.h

7.7 CalibrationConstant Class Reference

Example for a simple calibration class based on the LCFixedObject template.

#include </opt/ilcsoft/HEAD-2016-12-02/lcio/HEAD/src/cpp/src/EXAMPLE/Calibration-
Constant.h>

Inheritance diagram for CalibrationConstant:

Public Member Functions

• CalibrationConstant (int cellID, float offset, float gain)

Convenient c’tor.

• CalibrationConstant (LCObject ∗obj)

’Copy constructor’ needed to interpret LCCollection read from file/database.

• virtual ∼CalibrationConstant ()

Important for memory handling.

• int getCellID ()

• float getOffset ()

• float getGain ()

• void print (std::ostream &os)

• const std::string getTypeName () const

• const std::string getDataDescription () const

7.7.1 Detailed Description

Example for a simple calibration class based on the LCFixedObject template. LCFixedObject uses an instance of LCGenericObjectImpl that holds the data, thus there is no overhead when the data is read from a database or file for copying it to some local structure (Decorator pattern). The documentation for this class was generated from the following file:

• CalibrationConstant.h

7.8 EVENT::CalorimeterHit Class Reference

The generic calorimeter hit - created from SimCalorimeterHit (p. 205) or RawCalorimeterHit (p. 188).

#include <pre-generated/EVENT/CalorimeterHit.h>

Inheritance diagram for EVENT::CalorimeterHit:

Public Types

• typedef CalorimeterHit lcobject_type

Useful typedef for template programming with LCIO (p. 115).

Public Member Functions

• virtual ∼CalorimeterHit ()

Destructor.

• virtual int getCellID0 () const =0

Returns the detector specific (geometrical) cell id.

• virtual int getCellID1 () const =0

Returns the second detector specific (geometrical) cell id.

• virtual float getEnergy () const =0

Returns the energy of the hit in [GeV].

• virtual float getEnergyError () const =0

Returns the error of the hit energy in [GeV].

• virtual float getTime () const =0

Returns the time of the hit in [ns].

• virtual const float ∗ getPosition () const =0

Returns the position of the hit in world coordinates.

• virtual int getType () const =0

Type of hit.

• virtual LCObject ∗ getRawHit () const =0

The RawCalorimeterHit (p. 188).

Additional Inherited Members

7.8.1 Detailed Description

The generic calorimeter hit - created from SimCalorimeterHit (p. 205) or RawCalorimeterHit (p. 188). The difference to RawCalorimeterHit (p. 188) is that amplitude ( energy) and time are now float and an optional position (x,y,z) can be stored with the hit. In order to suppress a (32-bit) pointer tag stored with every hit set flag(LCIO::RCHBIT_NO_PTR)==1.

Author gaede

Version

Id: CalorimeterHit.aid,v 1.16 2008-06-04 17:37:15 engels Exp

7.8.2 Member Function Documentation

7.8.2.1 virtual int EVENT::CalorimeterHit::getCellID1 ( ) const [pure virtual]

Returns the second detector specific (geometrical) cell id. Optional, check/set flag(LCIO::RCHBIT_ID1)==1. Implemented in IMPL::CalorimeterHitImpl (p. 53). Referenced by SIO::SIOCalHitHandler::write().

7.8.2.2 virtual float EVENT::CalorimeterHit::getEnergyError ( ) const [pure virtual]

Returns the error of the hit energy in [GeV]. Optional, check/set flag(LCIO::RCHBIT_ENERGY_ERROR)==1. Implemented in IMPL::CalorimeterHitImpl (p. 53). Referenced by SIO::SIOCalHitHandler::write().

7.8.2.3 virtual const float∗ EVENT::CalorimeterHit::getPosition ( ) const [pure virtual]

Returns the position of the hit in world coordinates. NULL if information is not stored - check flag(LCIO::RCHBIT_LONG)==1. Implemented in IMPL::CalorimeterHitImpl (p. 53). Referenced by SIO::SIOCalHitHandler::write().

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.8.2.4 virtual LCObject∗ EVENT::CalorimeterHit::getRawHit ( ) const [pure virtual]

The RawCalorimeterHit (p. 188). As in the future there might be other types of raw calorimeter hits the return type is LCObject (p. 123). Implemented in IMPL::CalorimeterHitImpl (p. 53). Referenced by SIO::SIOCalHitHandler::write().

7.8.2.5 virtual float EVENT::CalorimeterHit::getTime ( ) const [pure virtual]

Returns the time of the hit in [ns]. Optional, check/set flag(LCIO::RCHBIT_TIME)==1. Implemented in IMPL::CalorimeterHitImpl (p. 54). Referenced by SIO::SIOCalHitHandler::write().

7.8.2.6 virtual int EVENT::CalorimeterHit::getType ( ) const [pure virtual]

Type of hit. Mapping of integer types to type names through collection parameters "CalorimeterHitTypeNames" and "- CalorimeterHitTypeValues". Implemented in IMPL::CalorimeterHitImpl (p. 54). Referenced by SIO::SIOCalHitHandler::write(). The documentation for this class was generated from the following file:

• CalorimeterHit.h

7.9 IMPL::CalorimeterHitImpl Class Reference

Implementation of the real data CalorimeterHit.

#include <IMPL/CalorimeterHitImpl.h>

Inheritance diagram for IMPL::CalorimeterHitImpl:

Public Member Functions

• CalorimeterHitImpl ()

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Default constructor, initializes values to 0.

• virtual ∼CalorimeterHitImpl ()

Copy constructor.

• virtual int id () const

Returns an object id for internal (debugging) use in LCIO.

• virtual int getCellID0 () const

Returns the detector specific (geometrical) cell id.

• virtual int getCellID1 () const

Returns the second detector specific (geometrical) cell id.

• virtual float getEnergy () const

Returns the energy of the hit.

• virtual float getEnergyError () const

Returns the error of the hit energy.

• virtual float getTime () const

Returns the time of the hit in [ns].

• virtual const float ∗ getPosition () const

Returns the position of the hit in world coordinates.

• virtual int getType () const

Type of hit.

• virtual EVENT::LCObject ∗ getRawHit () const

The RawCalorimeterHit.

• void setCellID0 (int id0)

Sets the first cell id;.

• void setCellID1 (int id1)

Sets the second cell id; Only store if the flag word (bit RCHBIT_ID1) of the collection is set.

• void setEnergy (float en)

Sets the energy/amplitude.

• void setEnergyError (float enerr)

Sets the error of the hit energy.

• void setTime (float t)

Sets the time.

• void setPosition (const float pos[3])

Sets the position.

• void setType (int type)

• void setRawHit (EVENT::LCObject ∗rawHit)

Protected Attributes

• int _cellID0

• int _cellID1

• float _energy

• float _energyError

• float _time

• float _position [3]

• int _type

• EVENT::LCObject ∗ _rawHit

Additional Inherited Members

7.9.1 Detailed Description

Implementation of the real data CalorimeterHit.

Author gaede

Version

Id: CalorimeterHitImpl.h (p. ?? ),v 1.16 2008-06-04 17:37:16 engels Exp

7.9.2 Constructor & Destructor Documentation

7.9.2.1 IMPL::CalorimeterHitImpl::∼CalorimeterHitImpl ( ) [virtual]

Copy constructor. Not yet - needs pointer chasing ... Destructor.

7.9.3 Member Function Documentation

7.9.3.1 int IMPL::CalorimeterHitImpl::getCellID1 ( ) const [virtual]

Returns the second detector specific (geometrical) cell id. 0 if information is not stored - check the flag word (bit RCHBIT_ID1) of the collection. Default is to store only cellid0. Implements EVENT::CalorimeterHit (p. 50).

7.9.3.2 float IMPL::CalorimeterHitImpl::getEnergyError ( ) const [virtual]

Returns the error of the hit energy. Optional, check/set flag(LCIO::RCHBIT_ENERGY_ERROR)==1. Implements EVENT::CalorimeterHit (p. 50).

7.9.3.3 const float ∗ IMPL::CalorimeterHitImpl::getPosition ( ) const [virtual]

Returns the position of the hit in world coordinates. NULL if information is not stored. Ask collection for flag, only available if bit LCIO.RCHBIT_LONG is set. Implements EVENT::CalorimeterHit (p. 50).

7.9.3.4 EVENT::LCObject ∗ IMPL::CalorimeterHitImpl::getRawHit ( ) const [virtual]

The RawCalorimeterHit. As in the future there might be other types of raw calorimeter hits the return type is LCObject. Implements EVENT::CalorimeterHit (p. 51).

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.9.3.5 float IMPL::CalorimeterHitImpl::getTime ( ) const [virtual]

Returns the time of the hit in [ns]. Optional, check/set flag(LCIO::RCHBIT_TIME)==1. Implements EVENT::CalorimeterHit (p. 51).

7.9.3.6 int IMPL::CalorimeterHitImpl::getType ( ) const [virtual]

Type of hit. Mapping of integer types to type names through collection parameters "CalorimeterHitTypeNames" and "- CalorimeterHitTypeValues". Implements EVENT::CalorimeterHit (p. 51).

7.9.3.7 void IMPL::CalorimeterHitImpl::setCellID1 ( int id1 )

Sets the second cell id; Only store if the flag word (bit RCHBIT_ID1) of the collection is set. Default is to store only cellid0.

7.9.3.8 void IMPL::CalorimeterHitImpl::setEnergyError ( float enerr )

Sets the error of the hit energy. Not stored if flag RCHBIT_ENERGY_ERROR isn’t set

7.9.3.9 void IMPL::CalorimeterHitImpl::setPosition ( const float pos[3] )

Sets the position. Not stored if LCIO.RCHBIT_LONG isn’t set. The documentation for this class was generated from the following files:

• CalorimeterHitImpl.h

• CalorimeterHitImpl.cc

7.10 IOIMPL::CalorimeterHitIOImpl Class Reference

Adding stuff needed for io (friend declarations, etc.)

#include <IOIMPL/CalorimeterHitIOImpl.h>

Inheritance diagram for IOIMPL::CalorimeterHitIOImpl:

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.11 lcrtrel_helper::can_call_ext< B > Struct Template Reference  55

Friends

• class SIO::SIOCalHitHandler

Additional Inherited Members 7.10.1 Detailed Description

Adding stuff needed for io (friend declarations, etc.)

Author gaede

Version Aug 8, 2003 The documentation for this class was generated from the following file:

• CalorimeterHitIOImpl.h

7.11 lcrtrel_helper::can_call_ext< B > Struct Template Reference

The documentation for this struct was generated from the following file:

• LCRTRelations.h

7.12 lcrtrel_helper::can_call_ext< true > Struct Template Reference

Static Public Member Functions

• static void check () The documentation for this struct was generated from the following file:

• LCRTRelations.h

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

56  Class Documentation

7.13 UTIL::CellIDDecoder< T > Class Template Reference

Convenient class for decoding cellIDs from collection parameter LCIO::CellIDEncoding.

#include <UTIL/CellIDDecoder.h>

Public Member Functions

• CellIDDecoder (const std::string &encoder_str)

Constructor takes encoding string as argument.

• CellIDDecoder (const EVENT::LCCollection ∗col)

Constructor reads encoding string from collection parameter LCIO::CellIDEncoding.

• const BitField64 & operator() (const T ∗hit)

Provides access to the bit fields, e.g.

Static Public Member Functions

• static void setDefaultEncoding (const std::string &defaultEncoding)

This can be used to set the default encoding that is used if no CellIDEncoding parameter is set in the collection, e.g.

Protected Attributes

• BitField64 ∗ _b

• const T ∗ _oldHit

Static Protected Attributes

• static std::string ∗ _defaultEncoding = new std::string("byte0:8,byte1:8,byte2:8,byte3:8,byte4:8,byte5-

:8,byte6:8,byte7:8")

7.13.1 Detailed Description

template<class T>class UTIL::CellIDDecoder< T > Convenient class for decoding cellIDs from collection parameter LCIO::CellIDEncoding. See UTIL::BitField64 (p. 43) for a description of the encoding string. See Also BitField64 (p. 43) Version

Id: CellIDDecoder.h (p. ?? ),v 1.9.16.1 2011-03-04 14:09:07 engels Exp

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.14 UTIL::CellIDEncoder< T > Class Template Reference  57

7.13.2 Member Function Documentation

7.13.2.1 template<class T> const BitField64& UTIL::CellIDDecoder< T >::operator() ( const T ∗ hit ) [inline] Provides access to the bit fields, e.g. int layer = myCellIDEncoding( hit )[ "layer" ] ;

7.13.2.2 template<class T> static void UTIL::CellIDDecoder< T >::setDefaultEncoding ( const std::string &

defaultEncoding ) [inline], [static] This can be used to set the default encoding that is used if no CellIDEncoding parameter is set in the collection, e.g. in older lcio files. The documentation for this class was generated from the following file:

• CellIDDecoder.h

7.14 UTIL::CellIDEncoder< T > Class Template Reference

Specialization for SimTrackerHits that have only one cellID.

#include <UTIL/CellIDEncoder.h>

Inheritance diagram for UTIL::CellIDEncoder< T >:

Public Member Functions

• CellIDEncoder (const std::string &cellIDEncoding, EVENT::LCCollection ∗col)

Constructor, sets collection parameter LCIO::CellIDEncoding to the given encoding string.

• void setCellID (T ∗hit)

• void setCellIDFlag ()

Helper method that sets/unsets the proper bit for storing a second cellid word.

Protected Attributes

• EVENT::LCCollection ∗ _col

Additional Inherited Members 7.14.1 Detailed Description

template<class T>class UTIL::CellIDEncoder< T > Specialization for SimTrackerHits that have only one cellID.

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Convenient class for encoding cellIDs for various hit objects. It sets the proper collection parameter LCIO::CellID-Encoding and sets the proper flag bit for storing a second cellid if necessary. See UTIL::BitField64 (p. 43) for a description of the encoding string. Example: CellIDEncoder<SimCalorimeterHitImpl> cd( "i:20,j:20,k:20" ,calVec ) ; for(int j=0;j<NHITS;j++){ SimCalorimeterHitImpl∗ hit = new SimCalorimeterHitImpl ; cd["i"] = j ; cd["j"] = j + 100 ; cd["k"] = j + 200 ; cd.setCellID( hit ) ; }

See Also BitField64 (p. 43)

Version

Id: CellIDEncoder.h (p. ?? ),v 1.6 2008-01-22 14:38:55 gaede Exp

The documentation for this class was generated from the following file:

• CellIDEncoder.h

7.15 EVENT::Cluster Class Reference

The LCIO (p. 115) cluster.

#include <pre-generated/EVENT/Cluster.h>

Inheritance diagram for EVENT::Cluster:

Public Types

• typedef Cluster lcobject_type

Useful typedef for template programming with LCIO (p. 115).

Public Member Functions

• virtual ∼Cluster ()

Destructor.

• virtual int getType () const =0

Flagword that defines the type of cluster.

• virtual float getEnergy () const =0

Energy of the cluster.

• virtual float getEnergyError () const =0

Returns the error on the energy of the cluster.

• virtual const float ∗ getPosition () const =0

Position of the cluster.

• virtual const FloatVec & getPositionError () const =0

Covariance matrix of the position (6 Parameters)

• virtual float getIPhi () const =0

Intrinsic direction of cluster at position: Phi.

• virtual float getITheta () const =0

Intrinsic direction of cluster at position: Theta.

• virtual const FloatVec & getDirectionError () const =0

Covariance matrix of the direction (3 Parameters).

• virtual const FloatVec & getShape () const =0

Shape parameters - check/set collection parameter ClusterShapeParameters for size and names of parameters.

• virtual const ParticleIDVec & getParticleIDs () const =0

The particle Id’s sorted by their likelihood.

• virtual const ClusterVec & getClusters () const =0

The clusters that have been combined to this cluster.

• virtual const CalorimeterHitVec & getCalorimeterHits () const =0

The hits that have been combined to this cluster.

• virtual const FloatVec & getHitContributions () const =0

Returns the energy contribution of the hits Runs parallel to the CalorimeterHitVec from getCalorimeterHits() (p. 60)

• virtual const FloatVec & getSubdetectorEnergies () const =0

A vector that holds the energy observed in a particular subdetectors.

Additional Inherited Members

7.15.1 Detailed Description

The LCIO (p. 115) cluster.

Author gaede

Version

Id: Cluster.aid,v 1.14 2006-08-03 16:53:34 gaede Exp

7.15.2 Member Function Documentation

7.15.2.1 virtual const CalorimeterHitVec& EVENT::Cluster::getCalorimeterHits ( ) const [pure virtual]

The hits that have been combined to this cluster. Only available if collection flag bit LCIO::CLBIT_HITS==1 and if the CalorimeterHit (p. 49) objects have not been saved with LCIO::RCHBIT_NO_PTR==1.

See Also CalorimeterHit (p. 49)

Implemented in IMPL::ClusterImpl (p. 62). Referenced by SIO::SIOClusterHandler::write().

7.15.2.2 virtual const FloatVec& EVENT::Cluster::getDirectionError ( ) const [pure virtual]

Covariance matrix of the direction (3 Parameters). Stored as lower triangle matrix where the order of parameters is: iPhi, iTheta. So we have cov(iPhi,iPhi), cov( iPhi, iTheta), cov( iTheta, iTheta). Implemented in IMPL::ClusterImpl (p. 62). Referenced by SIO::SIOClusterHandler::write().

7.15.2.3 virtual float EVENT::Cluster::getIPhi ( ) const [pure virtual]

Intrinsic direction of cluster at position: Phi. Not to be confused with direction cluster is seen from IP. Implemented in IMPL::ClusterImpl (p. 62). Referenced by SIO::SIOClusterHandler::write().

7.15.2.4 virtual float EVENT::Cluster::getITheta ( ) const [pure virtual]

Intrinsic direction of cluster at position: Theta. Not to be confused with direction cluster is seen from IP. Implemented in IMPL::ClusterImpl (p. 62). Referenced by SIO::SIOClusterHandler::write().

7.15.2.5 virtual const ParticleIDVec& EVENT::Cluster::getParticleIDs ( ) const [pure virtual]

The particle Id’s sorted by their likelihood.

See Also ParticleID (p. 181)

Implemented in IMPL::ClusterImpl (p. 63). Referenced by UTIL::LCTOOLS::printClusters(), and SIO::SIOClusterHandler::write().

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.15.2.6 virtual const FloatVec& EVENT::Cluster::getSubdetectorEnergies ( ) const [pure virtual]

A vector that holds the energy observed in a particular subdetectors. The mapping of indices to subdetectors is implementation dependent. To be used as convenient information or if hits are not stored in the data set, e.g. DST or FastMC. Check/set collection parameter ClusterSubdetectorNames for decoding the indices of the array. Implemented in IMPL::ClusterImpl (p. 64). Referenced by SIO::SIOClusterHandler::write().

7.15.2.7 virtual int EVENT::Cluster::getType ( ) const [pure virtual]

Flagword that defines the type of cluster. Bits 0-15 can be used to denote the subdetectors that have contributed hits to the cluster. For the definition of the bits check/Set the collection variables ClusterTypeBitNames and ClusterTypeBitIndices. Bits 16-31 are used internally. Implemented in IMPL::ClusterImpl (p. 64). Referenced by SIO::SIOClusterHandler::write(). The documentation for this class was generated from the following file:

• Cluster.h

7.16 IMPL::ClusterImpl Class Reference

Implementation of Cluster.

#include <IMPL/ClusterImpl.h>

Inheritance diagram for IMPL::ClusterImpl:

Public Member Functions

• ClusterImpl ()

Default constructor, initializes values to 0.

• virtual ∼ClusterImpl ()

Destructor.

• virtual int id () const

Returns an object id for internal (debugging) use in LCIO.

• virtual int getType () const

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Flagword that defines the type of cluster.

• virtual float getEnergy () const

Energy of the cluster.

• virtual float getEnergyError () const

Returns the error on the energy of the cluster.

• virtual const float ∗ getPosition () const

Position of the cluster.

• virtual const EVENT::FloatVec & getPositionError () const

Covariance matrix of the position (6 Parameters)

• virtual float getITheta () const

Intrinsic direction of cluster at position: Theta.

• virtual float getIPhi () const

Intrinsic direction of cluster at position: Phi.

• virtual const EVENT::FloatVec & getDirectionError () const

Covariance matrix of the direction (3 Parameters)

• virtual const EVENT::FloatVec & getShape () const

Shape parameters - check/set collection parameter ClusterShapeParameters for size and names of parameters.

• virtual const

EVENT::ParticleIDVec & getParticleIDs () const

Type hypotheses: 3 Parameters: compatible with EM, HAD, muon cluster.

• virtual const EVENT::ClusterVec & getClusters () const

The clusters that have been combined to this cluster.

• virtual const

EVENT::CalorimeterHitVec & getCalorimeterHits () const

The hits that have been combined to this cluster.

• virtual const EVENT::FloatVec & getHitContributions () const

Returns the energy contribution of the hits Runs parallel to the CalorimeterHitVec from getCalorimeterHits() (p. 62)

• virtual const EVENT::FloatVec & getSubdetectorEnergies () const

A vector that holds the energy observed in a particular subdetectors.

• void setTypeBit (int index, bool val=true)

• void setEnergy (float energy)

• void setEnergyError (float energyError)

• void setPosition (const float ∗position)

• void setPositionError (const EVENT::FloatVec &errpos)

• void setPositionError (const float ∗errpos)

• void setITheta (float theta)

• void setIPhi (float phi)

• void setDirectionError (const EVENT::FloatVec &errdir)

• void setDirectionError (const float ∗errdir)

• void setShape (const EVENT::FloatVec &shape)

• void addParticleID (EVENT::ParticleID ∗pid)

• void addCluster (EVENT::Cluster ∗cluster)

• void addHit (EVENT::CalorimeterHit ∗hit, float contribution)

• EVENT::FloatVec & subdetectorEnergies ()

To be used for modifying the subdetector energies, e.g.

Protected Member Functions

• void setType (int type)

Protected Attributes

• std::bitset< 32 > _type

• float _energy

• float _energyError

• float _position [3]

• EVENT::FloatVec _errpos

• float _theta

• float _phi

• EVENT::FloatVec _errdir

• EVENT::FloatVec _shape

• EVENT::ParticleIDVec _pid

• EVENT::ClusterVec _clusters

• EVENT::CalorimeterHitVec _hits

• EVENT::FloatVec _weights

• EVENT::FloatVec _subdetectorEnergies

Additional Inherited Members

7.16.1 Detailed Description

Implementation of Cluster.

See Also Cluster

Author gaede

Version

Id: ClusterImpl.h (p. ?? ),v 1.13 2006-08-04 16:52:46 gaede Exp

7.16.2 Member Function Documentation

7.16.2.1 const EVENT::ParticleIDVec & IMPL::ClusterImpl::getParticleIDs ( ) const [virtual]

Type hypotheses: 3 Parameters: compatible with EM, HAD, muon cluster. The particle Id’s sorted by their probability.

See Also ParticleID

Implements EVENT::Cluster (p. 60).

7.16.2.2 const EVENT::FloatVec & IMPL::ClusterImpl::getSubdetectorEnergies ( ) const [virtual]

A vector that holds the energy observed in a particular subdetectors. The mapping of indices to subdetectors is implementation dependent. To be used as convenient information or if hits are not stored in the data set, e.g. DST or FastMC. Check/set collection parameter ClusterSubdetectorNames for decoding the indices of the array. Implements EVENT::Cluster (p. 61).

7.16.2.3 int IMPL::ClusterImpl::getType ( ) const [virtual]

Flagword that defines the type of cluster. Bits 0-15 can be used to denote the subdetectors that have contributed hits to the cluster. The definition of the bits has to be done elsewhere, e.g. in the run header. Bits 16-31 are used internally. Implements EVENT::Cluster (p. 61).

7.16.2.4 EVENT::FloatVec & IMPL::ClusterImpl::subdetectorEnergies ( )

To be used for modifying the subdetector energies, e.g. clu->subdetectorEnergies() (p. 64).resize(3) ; clu->subdetectorEnergies() (p. 64)[0] = 123.4567 ; clu- >subdetectorEnergies() (p. 64)[1] = 12.34567 ; clu->subdetectorEnergies() (p. 64)[2] = 1.234567 ; The documentation for this class was generated from the following files:

• ClusterImpl.h

• ClusterImpl.cc

7.17 IOIMPL::ClusterIOImpl Class Reference

Adding stuff needed for io (friend declarations, etc.)

#include <IOIMPL/ClusterIOImpl.h>

Inheritance diagram for IOIMPL::ClusterIOImpl:

Friends

• class SIO::SIOClusterHandler

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Additional Inherited Members

7.17.1 Detailed Description

Adding stuff needed for io (friend declarations, etc.)

Author gaede

Version Mar 15, 2004

The documentation for this class was generated from the following file:

• ClusterIOImpl.h

7.18 UTIL::CollectionParameterMap Class Reference

Helper class for setting and retrieving an std::map<string,int> as collection parameters.

#include <UTIL/CollectionParameterMap.h>

Public Types

• typedef std::map< std::string, int > map_type

Public Member Functions

• CollectionParameterMap (const std::string &keyName, const std::string &valueName, LCCollection ∗col)

Create CollectionParameterMap (p. 65) for given collection - read the collection parameters keyName and value-Name if they exist.

• CollectionParameterMap (const std::string &keyName, const std::string &valueName, const LCCollection

∗col)

Create CollectionParameterMap (p. 65) for given collection - read the collection parameters keyName and value-Name if they exist.

• ∼CollectionParameterMap ()

Update the collection parameters keyName and valueName when going out of scope.

• map_type & map ()

The std::map< std::string, int >.

Protected Member Functions

• void init (const LCCollection ∗col)

Protected Attributes

• std::string _keyName

• std::string _valueName

• LCCollection ∗ _col

• map_type _map

7.18.1 Detailed Description

Helper class for setting and retrieving an std::map<string,int> as collection parameters. Depending on which constructor was used the collection parameters will be updated if an object of this class goes out of scope. The corresponding collection parameters will hold the current state of the map. If the a const LC-Collection∗ was specified (read use case) no parameters will be updated.

See Also LCCollection::getParameters()

LCParameters

Author F.Gaede, DESY

Version

Id: CollectionParameterMap.h (p. ?? ),v 1.1 2008-05-30 13:26:16 gaede Exp

The documentation for this class was generated from the following files:

• CollectionParameterMap.h

• CollectionParameterMap.cc

7.19 lcrtrel_helper::CreationPtrInit< T > Struct Template Reference

Factory for objects of type T.

#include <LCRTRelations.h>

Static Public Member Functions

• static void ∗ init ()

7.19.1 Detailed Description

template<class T>struct lcrtrel_helper::CreationPtrInit< T >

Factory for objects of type T. The documentation for this struct was generated from the following file:

• LCRTRelations.h

7.20 EVENT::DataNotAvailableException Class Reference

EventException (p. 69) used for data not available.

#include <Exceptions.h>

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.21 lcrtrel_helper::DeleteElements< T > Struct Template Reference  67

Inheritance diagram for EVENT::DataNotAvailableException:

Public Member Functions

• DataNotAvailableException (std::string text)

Additional Inherited Members 7.20.1 Detailed Description

EventException (p. 69) used for data not available. Author gaede

Version Jun 5, 2003 The documentation for this class was generated from the following file:

• Exceptions.h

7.21 lcrtrel_helper::DeleteElements< T > Struct Template Reference

Delete function for containers of owned objects.

#include <LCRTRelations.h>

Static Public Member Functions

• static void clean (void ∗v)

7.21.1 Detailed Description

template<class T>struct lcrtrel_helper::DeleteElements< T > Delete function for containers of owned objects. The documentation for this struct was generated from the following file:

• LCRTRelations.h

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.22 lcrtrel_helper::DeletePtr< T > Struct Template Reference

Delete function for pointers w/ ownership.

#include <LCRTRelations.h>

Static Public Member Functions

• static void clean (void ∗v)

7.22.1 Detailed Description

template<class T>struct lcrtrel_helper::DeletePtr< T >

Delete function for pointers w/ ownership. The documentation for this struct was generated from the following file:

• LCRTRelations.h

7.23 IO::EndOfDataException Class Reference

EndOfDataException (p. 68) for signaling the end of a data stream.

#include <Exceptions.h>

Inheritance diagram for IO::EndOfDataException:

Public Member Functions

• EndOfDataException (std::string text)

Additional Inherited Members

7.23.1 Detailed Description

EndOfDataException (p. 68) for signaling the end of a data stream.

Author gaede Version Jun 5, 2003

The documentation for this class was generated from the following file:

• Exceptions.h

7.24 EVENT::EventException Class Reference

EventException (p. 69) used for errors accessing the event data.

#include <Exceptions.h>

Inheritance diagram for EVENT::EventException:

Public Member Functions

• EventException (std::string text)

Additional Inherited Members

7.24.1 Detailed Description

EventException (p. 69) used for errors accessing the event data.

Author gaede

Version Apr 30, 2003

The documentation for this class was generated from the following file:

• Exceptions.h

7.25 EVENT::Exception Class Reference

Base exception class for LCIO (p. 115) - all other exceptions extend this.

#include <Exceptions.h>

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Inheritance diagram for EVENT::Exception:

Public Member Functions

• Exception (const std::string &text)

• virtual const char ∗ what () const throw ()

Protected Attributes

• std::string message

7.25.1 Detailed Description

Base exception class for LCIO (p. 115) - all other exceptions extend this.

Author gaede

Version Apr 30, 2003 The documentation for this class was generated from the following file:

• Exceptions.h

7.26 lcrtrel_helper::FromRelation< U > Struct Template Reference

Helper class for relations.

#include <LCRTRelations.h>

7.26.1 Detailed Description

template<class U>struct lcrtrel_helper::FromRelation< U > Helper class for relations. The documentation for this struct was generated from the following file:

• LCRTRelations.h

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.27 HEPEVTIMPL::HEPEVT Class Reference

Static Public Member Functions

• static void fromHepEvt (EVENT::LCEvent ∗evt, const char ∗mcColName=0)

StdHep fortran interface to read from or write to the COMMON /HEPEVT/.

• static void toHepEvt (const EVENT::LCEvent ∗evt, const char ∗mcColName=0)

7.27.1 Member Function Documentation

7.27.1.1 void HEPEVTIMPL::HEPEVT::fromHepEvt ( EVENT::LCEvent ∗ evt, const char ∗ mcColName = 0 ) [static]

StdHep fortran interface to read from or write to the COMMON /HEPEVT/.

fromHepEvt reads from COMMON /HEPEVT/. toHepEvt fills COMMON /HEPEVT/.

References IMPL::MCParticleImpl::addParent(), IMPL::LCEventImpl::getCollection(), EVENT::LCCollection::get-ElementAt(), IMPL::LCCollectionVec::getElementAt(), EVENT::LCCollection::getNumberOfElements(), IMPL::MC-ParticleImpl::setCharge(), IMPL::MCParticleImpl::setGeneratorStatus(), IMPL::MCParticleImpl::setMass(), IMPL- ::MCParticleImpl::setMomentum(), IMPL::MCParticleImpl::setPDG(), IMPL::MCParticleImpl::setSimulatorStatus(), and IMPL::MCParticleImpl::setVertex().

The documentation for this class was generated from the following files:

• HEPEVT.h

• HEPEVT.cc

7.28 hepevt1_t Struct Reference

Public Attributes

• PTRTYPE mcpointerv [NMXHEP]

• float mcchargev [NMXHEP]

The documentation for this struct was generated from the following file:

• hepevt1.h

7.29 hepevt_t Struct Reference

Public Attributes

• int nevhep

• int nhep

• int isthep [NMXHEP]

• int idhep [NMXHEP]

• int jmohep [NMXHEP][2]

• int jdahep [NMXHEP][2]

• double phep [NMXHEP][5]

• double vhep [NMXHEP][4]

The documentation for this struct was generated from the following file:

• hepevt0.h

7.30 IO::ILCFactory Class Reference

Factory to create LCWriter (p. 166) and Reader object for the known data formats, e.g.

#include <pre-generated/IO/ILCFactory.h>

Inheritance diagram for IO::ILCFactory:

Public Member Functions

• virtual ∼ILCFactory ()

Destructor.

• virtual LCWriter ∗ createLCWriter ()=0

Returns the instance of the factory.

• virtual LCReader ∗ createLCReader (int readerFlag=0)=0

Creates an LCReader (p. 130) object for the current persistency type.

7.30.1 Detailed Description

Factory to create LCWriter (p. 166) and Reader object for the known data formats, e.g. SIO (p. 30). The implementation (LCFactory) will be a singleton class that knows all concrete reader/writer implementations.

Author gaede

Version Mar 7, 2003

7.30.2 Member Function Documentation

7.30.2.1 virtual LCWriter∗ IO::ILCFactory::createLCWriter ( ) [pure virtual]

Returns the instance of the factory. In the future we need to specify the data format here... This doesn’t work for java as an interface can’t have static members ...Creates an LCWriter (p. 166) object for the current persistency type. Implemented in IOIMPL::LCFactory (p. 101). The documentation for this class was generated from the following file:

• ILCFactory.h

7.31 UTIL::ILDCellID0 Struct Reference

Helper class to define constants for the canonical encoding of the CellID0 in tracking sub detectors in ILD (or ILD like detectors).

#include <UTIL/ILDConf.h>

Static Public Attributes

• static std::string encoder_string = "subdet:5,side:-2,layer:9,module:8,sensor:8"

The canonical encoding string to be used for writing the CellID0 with the CellIDEncoder<T>:

• static const unsigned subdet = 0

Index (p. 77) of the field subdet in encoder_string - sub detector Id as defined in DetID, e.g.

• static const unsigned side = 1

Index (p. 77) of the field side in encoder_string - one of DetID::bwd, DetID::barrel, DetID::fwd.

• static const unsigned layer = 2

Index (p. 77) of the field layer in encoder_string - local sub detector layer ID starting from 0 going away from the origin (IP).

• static const unsigned module = 3

Index (p. 77) of the field module in encoder_string - module ID as defined for the given subdetetor.

• static const unsigned sensor = 4

Index (p. 77) of the field sensor in encoder_string - sensor ID as defined for the given subdetetor module.

7.31.1 Detailed Description

Helper class to define constants for the canonical encoding of the CellID0 in tracking sub detectors in ILD (or ILD like detectors). F.Gaede, DESY, 2011

Version $Id:$

7.31.2 Member Data Documentation

7.31.2.1 std::string UTIL::ILDCellID0::encoder_string = "subdet:5,side:-2,layer:9,module:8,sensor:8" [static]

The canonical encoding string to be used for writing the CellID0 with the CellIDEncoder<T>:

"subdet:5,side:-2,layer:9,module:8:sensor:8"

Use the CellIdDecoder to access the elements:

     CellIDDecoder<TrackerHit> idDec( trkHitCol ) ;
      // ...
      TrackerHit* hit = (TrackerHit*) trkHitCol.getElementAt( i ) ;
      int layer  = idDec( hit )[ ILDCellID0::layer (p. 73) ] ;

7.31.2.2 const unsigned UTIL::ILDCellID0::module = 3 [static]

Index (p. 77) of the field module in encoder_string - module ID as defined for the given subdetetor.

74  Class Documentation

7.31.2.3 const unsigned UTIL::ILDCellID0::sensor = 4 [static] Index (p. 77) of the field sensor in encoder_string - sensor ID as defined for the given subdetetor module.

7.31.2.4 const unsigned UTIL::ILDCellID0::subdet = 0 [static] Index (p. 77) of the field subdet in encoder_string - sub detector Id as defined in DetID, e.g. DetID::VXD or DetID::FTD. The documentation for this struct was generated from the following files:

• ILDConf.h

• ILDConf.cc

7.32 UTIL::ILDCellIDEncoder< T > Class Template Reference

Special wrapper to the CellIDEncoder<T> that enforces cellID0 to use the encoder string defined in ILDCellID0- ::encoder_string (p. 73).

#include <UTIL/ILDConf.h>

Inheritance diagram for UTIL::ILDCellIDEncoder< T >:

Public Member Functions

• ILDCellIDEncoder (EVENT::LCCollection ∗col)

Constructor for using the canonical cellID0 as defined in ILDCellID0::encoder_string (p. 73) - cellID1 will not be used.

• ILDCellIDEncoder (const std::string &cellIDEncoding, EVENT::LCCollection ∗col)

Constructor that prepends given ILDCellID0::encoder_string (p. 73) to the given encoding string.

Additional Inherited Members 7.32.1 Detailed Description

template<class T>class UTIL::ILDCellIDEncoder< T > Special wrapper to the CellIDEncoder<T> that enforces cellID0 to use the encoder string defined in ILDCellID0- ::encoder_string (p. 73).

7.32.2 Constructor & Destructor Documentation

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.32.2.1 template<class T> UTIL::ILDCellIDEncoder< T >::ILDCellIDEncoder ( const std::string & cellIDEncoding, EVENT::LCCollection ∗ col ) [inline]

Constructor that prepends given ILDCellID0::encoder_string (p. 73) to the given encoding string.

Throws an Exception if called with an empty string.

The documentation for this class was generated from the following file:

• ILDConf.h

7.33 UTIL::ILDDetID Struct Reference

Define integer constants for identifying ILD sub detectors.

#include <UTIL/ILDConf.h>

Static Public Attributes

• static const int NOTUSED = 0

• static const int VXD = 1

• static const int SIT = 2

• static const int FTD = 3

• static const int TPC = 4

• static const int SET = 5

• static const int ETD = 6

• static const int ECAL = 20

• static const int ECAL_PLUG = 21

• static const int HCAL = 22

• static const int HCAL_RING = 23

• static const int LCAL = 24

• static const int BCAL = 25

• static const int LHCAL = 26

• static const int YOKE = 27

• static const int COIL = 28

• static const int ECAL_ENDCAP = 29

• static const int HCAL_ENDCAP = 30

• static const int YOKE_ENDCAP = 31

• static const int bwd = - 1

• static const int barrel = 0

• static const int fwd = 1

7.33.1 Detailed Description

Define integer constants for identifying ILD sub detectors.

Typically used in CellIDs, for example together with ILD::CellID0::encoder_string.

F.Gaede, DESY, 2011

Version $Id:$

The documentation for this struct was generated from the following files:

• ILDConf.h

• ILDConf.cc

7.34 UTIL::ILDTrkHitQualityBit Struct Reference

Define integer constants for bits of the the lcio tracker hit quality used in the context of ILD reconstruction code.

#include <UTIL/ILDConf.h>

Static Public Attributes

• static const int USED_IN_FIT = 30

• static const int USED_IN_TRACK = 29

• static const int DOUBLE_HIT_CANDIDATE = 28

• static const int GOOD = 27

7.34.1 Detailed Description

Define integer constants for bits of the the lcio tracker hit quality used in the context of ILD reconstruction code.

F.Gaede, DESY, 2012

Version $Id:$

The documentation for this struct was generated from the following files:

• ILDConf.h

• ILDConf.cc

7.35 UTIL::ILDTrkHitTypeBit Struct Reference

Define integer constants for bits of the the lcio tracker hit type used in the context of ILD reconstruction code.

#include <UTIL/ILDConf.h>

Static Public Attributes

• static const int COMPOSITE_SPACEPOINT = 30

• static const int ONE_DIMENSIONAL = 29

7.35.1 Detailed Description

Define integer constants for bits of the the lcio tracker hit type used in the context of ILD reconstruction code.

S.J. Aplin, DESY, 2012

Version $Id:$

The documentation for this struct was generated from the following files:

• ILDConf.h

• ILDConf.cc

7.36 Index Struct Reference

Inheritance diagram for Index:

The documentation for this struct was generated from the following file:

• lcrtrelation.cc

7.37 SomeClass::Index Struct Reference

Inheritance diagram for SomeClass::Index:

The documentation for this struct was generated from the following file:

• test_RTextension.cc

7.38 UTIL::IndexMap Class Reference

utility class to manage indexes according to Collection Parameters

#include <UTIL/IndexMap.h>

Public Member Functions

• IndexMap (const EVENT::LCCollection ∗col, const std::string &key1, const std::string &key2)

Constructor: takes as arguments the collection where the parameters are stored and the parameter names to manage Before you loop over a collection create an IndexMap (p. 77) object with the keys to manage.

• int encode (const std::string name)

encode string to index (e.g.

• const std::string & decode (int type) const

decode int to string (e.g.

Protected Attributes

• const EVENT::LCCollection ∗ _col

• const std::string _key1

• const std::string _key2

• EVENT::StringVec _strvec

• EVENT::IntVec _intvec

7.38.1 Detailed Description

utility class to manage indexes according to Collection Parameters EXP: UNDER DEVELOPMENT!!! - Don’t use this class yet

Author engels

Version

Id: IndexMap.h (p. ?? ),v 1.2 2006-09-21 06:10:38 gaede Exp

7.38.2 Constructor & Destructor Documentation

7.38.2.1 UTIL::IndexMap::IndexMap ( const EVENT::LCCollection ∗ col, const std::string & key1, const std::string & key2 )

Constructor: takes as arguments the collection where the parameters are stored and the parameter names to manage Before you loop over a collection create an IndexMap (p. 77) object with the keys to manage. Example: LCCollection∗ col = evt->getCollection( "Vertices" ) ; IndexMap (p. 77) imvtx( col, "AlgorithmNames", "AlgorithmTypes" ) ; Loop over Collection.... Vertex∗ v = dynamic_cast<Vertex∗>( col->getElementAt( i ) ) ; cout << "vertex " << i << has been created with " << imvtx.decode( v->getAlgorithmType() ) << endl; End Loop For using the encoding use similar process: IndexMap (p. 77) imvtx( col, "AlgorithmNames", "AlgorithmTypes" ) ; Loop for creating vertices... VertexImpl∗ v = new VertexImpl ; v->setAlgorithmType( imvtx.encode( "ZvTop" ) ); // If the parameter "ZvTop" or even the collection parameters "AlgorithmNames" and "AlgorithmTypes" do not exist // they are automatically added and ZvTop gets a corresponding new AlgorithmType which is also automatically added End Loop References EVENT::LCParameters::getIntVals(), EVENT::LCCollection::getParameters(), and EVENT::LC-Parameters::getStringVals().

7.38.3 Member Function Documentation

7.38.3.1 const std::string & UTIL::IndexMap::decode ( int type ) const

decode int to string (e.g. cout << "vertex has been created with " << imvtx.decode( vtx->getAlgorithmType() ) << endl;

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.38.3.2 int UTIL::IndexMap::encode ( const std::string name )

encode string to index (e.g. vertex->setAlgorithmType( imvtx.enconde( "ZvTop" ) ); The documentation for this class was generated from the following files:

• IndexMap.h

• IndexMap.cc

7.39 IO::IOException Class Reference

IOException (p. 79) used for reading/writing errors.

#include <Exceptions.h>

Inheritance diagram for IO::IOException:

Public Member Functions

• IOException (std::string text)

Additional Inherited Members

7.39.1 Detailed Description

IOException (p. 79) used for reading/writing errors.

Author gaede

Version Apr 30, 2003

The documentation for this class was generated from the following file:

• Exceptions.h

7.40 lcrtrel::LC1To1Relation< U, From, To > Struct Template Reference

One to one relation between two objects of type From and To.

#include <LCRTRelations.h>

Inheritance diagram for lcrtrel::LC1To1Relation< U, From, To >:

Additional Inherited Members

7.40.1 Detailed Description

template<class U, class From, class To>struct lcrtrel::LC1To1Relation< U, From, To >

One to one relation between two objects of type From and To. The documentation for this struct was generated from the following file:

• LCRTRelations.h

7.41 lcrtrel::LC1ToNRelation< U, From, To > Struct Template Reference

One to many relation between one object of type From to many objects of type To.

#include <LCRTRelations.h>

Inheritance diagram for lcrtrel::LC1ToNRelation< U, From, To >:

Additional Inherited Members

7.41.1 Detailed Description

template<class U, class From, class To>struct lcrtrel::LC1ToNRelation< U, From, To >

One to many relation between one object of type From to many objects of type To. The documentation for this struct was generated from the following file:

• LCRTRelations.h

7.42 lcrtrel_helper::LCBaseLinkContainerTraits< U, T, I, D, b > Struct Template Reference

Base class for all containers of extensions and relations, vectors, lists,...

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.43 lcrtrel_helper::LCBaseLinkTraits< U, T, I, D, b > Struct Template Reference  81

#include <LCRTRelations.h>

Inheritance diagram for lcrtrel_helper::LCBaseLinkContainerTraits< U, T, I, D, b >:

Public Types

• typedef LCBaseTraits< U, T, I, D, b > base

• typedef T ∗ ext_type

• typedef const T ∗ rel_type

• typedef T::value_type obj_ptr

• typedef T::iterator iterator

• typedef T::const_iterator const_iterator

Static Public Attributes

• static const bool is_container =true

Additional Inherited Members

7.42.1 Detailed Description

template<class U, class T, class I = CreationPtrInit<T>, class D = DeletePtr<T>, bool b = 1>struct lcrtrel_helper::LCBaseLink-ContainerTraits< U, T, I, D, b >

Base class for all containers of extensions and relations, vectors, lists,... The documentation for this struct was generated from the following file:

• LCRTRelations.h

7.43 lcrtrel_helper::LCBaseLinkTraits< U, T, I, D, b > Struct Template Reference

Base class for all extensions and relations of single objects.

#include <LCRTRelations.h>

Inheritance diagram for lcrtrel_helper::LCBaseLinkTraits< U, T, I, D, b >:

Public Types

• typedef LCBaseTraits< U, T, I, D, b > base

• typedef T ∗& ext_type

• typedef T ∗ rel_type

• typedef base::ptr obj_ptr

Static Public Attributes

• static const bool is_container =false

Additional Inherited Members

7.43.1 Detailed Description

template<class U, class T, class I = SimplePtrInit, class D = NoDelete, bool b = 1>struct lcrtrel_helper::LCBaseLinkTraits< U, T, I, D, b >

Base class for all extensions and relations of single objects.

The documentation for this struct was generated from the following file:

• LCRTRelations.h

7.44 lcrtrel_helper::LCBaseTraits< U, T, I, D, b > Struct Template Reference

Vector of pointers to extension obbjects.

#include <LCRTRelations.h>

Inheritance diagram for lcrtrel_helper::LCBaseTraits< U, T, I, D, b >:

Public Types

• typedef T ∗ ptr

base pointer type

• typedef U tag

Static Public Member Functions

• static void clean (void ∗v)

• static ptr init ()

• static DeleteFPtr deletePtr ()

Static Public Attributes

• static const int allowed_to_call_ext = b

7.44.1 Detailed Description

template<class U, class T, class I, class D, bool b>struct lcrtrel_helper::LCBaseTraits< U, T, I, D, b >

Vector of pointers to extension obbjects. Base class for all extensions and relations The documentation for this struct was generated from the following file:

• LCRTRelations.h

7.45 EVENT::LCCollection Class Reference

The generic collection used in LCIO (p. 115).

#include <pre-generated/EVENT/LCCollection.h>

Inheritance diagram for EVENT::LCCollection:

Public Member Functions

• virtual ∼LCCollection ()

Destructor.

• virtual int getNumberOfElements () const =0

Returns the number of elements in the collection.

• virtual const std::string & getTypeName () const =0

Returns the type name of the collection - valid names are defined in LCIO (p. 115).

• virtual LCObject ∗ getElementAt (int index) const =0

Returns pointer to element at index - no range check, use getNumberOfEntries().

• virtual int getFlag () const =0

Returns flag word for collection.

• virtual bool isTransient () const =0

True if collection is transient, i.e.

• virtual bool isDefault () const =0

True if collection is the default collection for the given type.

• virtual bool isSubset () const =0

True if the collection holds a subset of objects from other collections.

• virtual void addElement (LCObject ∗obj)=0 throw (ReadOnlyException, std::exception )

Adds the given element to (end of) the collection.

• virtual void removeElementAt (int i)=0 throw (ReadOnlyException, std::exception )

Removes the i-th element from the collection.

• virtual void setFlag (int flag)=0

Set the flag word.

• virtual const LCParameters & getParameters () const =0

84  Class Documentation

Parameters defined for this collection.

• virtual LCParameters & parameters ()=0

Parameters defined for this collection.

Static Public Attributes

• static const int BITTransient = 16

Transient bit in flag word.

• static const int BITDefault = 17

• static const int BITSubset = 18

7.45.1 Detailed Description

The generic collection used in LCIO (p. 115). Holds arbitrary objects of type LCObject (p. 123).

Author gaede

Version

Id: LCCollection.aid,v 1.11 2005-03-02 16:22:58 gaede Exp

See Also LCObject (p. 123)

LCIO (p. 115)

7.45.2 Member Function Documentation

7.45.2.1 virtual void EVENT::LCCollection::addElement ( LCObject ∗ obj ) throw ReadOnlyException, std::exception)

       [pure virtual]

Adds the given element to (end of) the collection. Throws an exception if the collection (event) is ’read only’. Exceptions

ReadOnlyException

(p. 192)

Implemented in IMPL::LCCollectionVec (p. 89).

7.45.2.2 virtual int EVENT::LCCollection::getFlag ( ) const [pure virtual] Returns flag word for collection. Bits 16-31 are reserved for LCIO (p. 115) Depending on the object type stored they have a special meaning, e.g. for SimCalorimeterHits: CHBIT_LONG = 31 - store position

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

CHBIT_BARREL = 30 - endcap or barrel

CHBIT_ID1 = 29 - cellid1 is sored

CHBIT_PDG = 28 - store pdg of secondaries

Bit 16 is used to flag collection as transient

Bit 17 is used to flag collection as default

Bit 18 is used to flag collection as subset

Bits 0-15 are subdetector/user specific.

See Also isTransient() (p. 86)

Implemented in IMPL::LCCollectionVec (p. 89).

Referenced by SIO::SIOVertexHandler::init(), SIO::SIOLCGenericObjectHandler::init(), SIO::SIOObjectHandler- ::init(), UTIL::LCTOOLS::printCalorimeterHits(), UTIL::LCTOOLS::printClusters(), UTIL::LCTOOLS::printLCFloat-Vecs(), UTIL::LCTOOLS::printLCGenericObjects(), UTIL::LCTOOLS::printLCIntVecs(), UTIL::LCTOOLS::printL-CStrVecs(), UTIL::LCTOOLS::printMCParticles(), UTIL::LCTOOLS::printRawCalorimeterHits(), UTIL::LCTOOL-S::printReconstructedParticles(), UTIL::LCTOOLS::printRelation(), UTIL::LCTOOLS::printSimCalorimeterHits(), UTIL::LCTOOLS::printSimTrackerHits(), UTIL::LCTOOLS::printTPCHits(), UTIL::LCTOOLS::printTrackerData(), UTIL::LCTOOLS::printTrackerHitPlane(), UTIL::LCTOOLS::printTrackerHits(), UTIL::LCTOOLS::printTrackerHitZ-Cylinder(), UTIL::LCTOOLS::printTrackerPulse(), UTIL::LCTOOLS::printTrackerRawData(), UTIL::LCTOOLS::print-Tracks(), and UTIL::CellIDEncoder< T >::setCellIDFlag().

7.45.2.3 virtual int EVENT::LCCollection::getNumberOfElements ( ) const [pure virtual]

Returns the number of elements in the collection.

Same as size().

Implemented in IMPL::LCCollectionVec (p. 88).

Referenced by UTIL::LCTOOLS::dumpEvent(), HEPEVTIMPL::HEPEVT::fromHepEvt(), SIO::SIOVertexHandler- ::init(), SIO::SIOLCGenericObjectHandler::init(), UTIL::LCIterator< T >::LCIterator(), UTIL::LCTOOLS::print-CalorimeterHits(), UTIL::LCTOOLS::printClusters(), UTIL::LCTOOLS::printLCFloatVecs(), UTIL::LCTOOLS::printL-CGenericObjects(), UTIL::LCTOOLS::printLCIntVecs(), UTIL::LCTOOLS::printLCStrVecs(), UTIL::LCTOOLS::print-MCParticles(), UTIL::LCTOOLS::printRawCalorimeterHits(), UTIL::LCTOOLS::printReconstructedParticles(), UTI-L::LCTOOLS::printRelation(), UTIL::LCTOOLS::printSimCalorimeterHits(), UTIL::LCTOOLS::printSimTrackerHits(), UTIL::LCTOOLS::printTPCHits(), UTIL::LCTOOLS::printTrackerData(), UTIL::LCTOOLS::printTrackerHitPlane(), UTIL::LCTOOLS::printTrackerHits(), UTIL::LCTOOLS::printTrackerHitZCylinder(), UTIL::LCTOOLS::printTracker-Pulse(), UTIL::LCTOOLS::printTrackerRawData(), UTIL::LCTOOLS::printTracks(), and UTIL::LCTOOLS::print-Vertices().

7.45.2.4 virtual bool EVENT::LCCollection::isDefault ( ) const [pure virtual]

True if collection is the default collection for the given type.

This implies that the collection is complete and unambigous. Convenient method that checks bit BITDefault of the flag word.

Implemented in IMPL::LCCollectionVec (p. 90).

7.45.2.5 virtual bool EVENT::LCCollection::isSubset ( ) const [pure virtual]

True if the collection holds a subset of objects from other collections.

If the collection is not transient only the pointers/references to the original objects will be stored. Convenient method that checks bit BITSubset of the flag word. Implemented in IMPL::LCCollectionVec (p. 90).

7.45.2.6 virtual bool EVENT::LCCollection::isTransient ( ) const [pure virtual]

True if collection is transient, i.e. will not be written to any LCIO (p. 115) file. Convenient method that checks bit BITTransient of the flag word. Implemented in IMPL::LCCollectionVec (p. 90). Referenced by SIO::SIOWriter::setUpHandlers().

7.45.2.7 virtual LCParameters& EVENT::LCCollection::parameters ( ) [pure virtual]

Parameters defined for this collection. Can be used to modify the collection paramters. Implemented in IMPL::LCCollectionVec (p. 88). Referenced by UTIL::CellIDEncoder< T >::CellIDEncoder(), SIO::SIOVertexHandler::init(), SIO::SIOLCGeneric-ObjectHandler::init(), SIO::SIOObjectHandler::init(), UTIL::CollectionParameterMap::∼CollectionParameterMap(), and UTIL::PIDHandler::∼PIDHandler().

7.45.2.8 virtual void EVENT::LCCollection::removeElementAt ( int i ) throw ReadOnlyException, std::exception) [pure

       virtual]

Removes the i-th element from the collection. Throws an exception if the collection (event) is ’read only’. Exceptions

ReadOnlyException (p. 192)

Implemented in IMPL::LCCollectionVec (p. 90).

7.45.2.9 virtual void EVENT::LCCollection::setFlag ( int flag ) [pure virtual]

Set the flag word. This is allowed in ’read only’ mode. Implemented in IMPL::LCCollectionVec (p. 88). Referenced by SIO::SIOVertexHandler::init(), SIO::SIOLCGenericObjectHandler::init(), SIO::SIOObjectHandler- ::init(), and UTIL::CellIDEncoder< T >::setCellIDFlag(). The documentation for this class was generated from the following file:

• LCCollection.h

7.46 IOIMPL::LCCollectionIOVec Class Reference

Adding stuff needed for io (friend declarations, etc.)

#include <IOIMPL/LCCollectionIOVec.h>

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Inheritance diagram for IOIMPL::LCCollectionIOVec:

Public Member Functions

• LCCollectionIOVec (const std::string &type)

Friends

• class SIO::SIOCollectionHandler

Additional Inherited Members 7.46.1 Detailed Description

Adding stuff needed for io (friend declarations, etc.)

Author gaede

Version Mar 6, 2003 The documentation for this class was generated from the following file:

• LCCollectionIOVec.h

7.47 IMPL::LCCollectionVec Class Reference

Implementation of the LCCollection using (inheriting from) an STL vector of LCObjects.

#include <IMPL/LCCollectionVec.h>

Inheritance diagram for IMPL::LCCollectionVec:

Public Member Functions

• LCCollectionVec ()

Default Constructor should be protected - every LCCollection needs to know the type of its elements.

• LCCollectionVec (const std::string &type)

The public default constructur that takes the name of the type of the elements.

• virtual ∼LCCollectionVec ()

Destructor.

• virtual int getNumberOfElements () const

Returns the number of entries in the collection.

• virtual const std::string & getTypeName () const

Returns the type name of the collection - valid names are defined in LCIO.

• virtual EVENT::LCObject ∗ getElementAt (int index) const

Returns pointer to element at index - no range check !.

• virtual int getFlag () const

Returns flag word for collection.

• virtual bool isTransient () const

True if collection is transient, i.e.

• void setTransient (bool val=true)

Sets the transient flag for this collection.

• virtual bool isDefault () const

True if collection is the default collection for the given type.

• void setDefault (bool val=true)

Sets the default flag for this collection.

• virtual bool isSubset () const

True if the collection holds a subset of objects from other collections.

• void setSubset (bool val=true)

Sets the subset flag for this collection.

• void setFlag (int flag)

Sets the flag word for this collection.

• virtual void addElement (EVENT::LCObject ∗obj) throw (EVENT::ReadOnlyException)

Adds the given element to (end of) the collection.

• virtual void removeElementAt (int i) throw (EVENT::ReadOnlyException)

Removes the i-th element from the collection.

• virtual const EVENT::LCParameters & getParameters () const

Parameters defined for this run.

• virtual EVENT::LCParameters & parameters ()

Parameters defined for this run.

Protected Member Functions

• void setReadOnly (bool readOnly)

Protected Attributes

• std::string _typeName

• int _flag

• LCParametersImpl _params

Additional Inherited Members

7.47.1 Detailed Description

Implementation of the LCCollection using (inheriting from) an STL vector of LCObjects.

Author gaede

Version

Id: LCCollectionVec.h (p. ?? ),v 1.12 2010-01-21 16:27:36 gaede Exp

See Also LCObject LCCollection

7.47.2 Constructor & Destructor Documentation

7.47.2.1 IMPL::LCCollectionVec::LCCollectionVec ( const std::string & type )

The public default constructur that takes the name of the type of the elements.

7.47.3 Member Function Documentation

7.47.3.1 void IMPL::LCCollectionVec::addElement ( EVENT::LCObject ∗ obj ) throw EVENT::ReadOnlyException)

       [virtual]

Adds the given element to (end of) the collection. Throws an exception if the collection (event) is ’read only’. Exceptions

ReadOnlyException

Implements EVENT::LCCollection (p. 84). Referenced by UTIL::LCRelationNavigator::createLCCollection().

7.47.3.2 int IMPL::LCCollectionVec::getFlag ( ) const [virtual]

Returns flag word for collection. Bits 16-31 are reserved for LCIO Depending on the object type stored they have a special meaning, e.g. for Sim-CalorimeterHits: CHBIT_LONG = 31 - store position CHBIT_BARREL = 30 - endcap or barrel CHBIT_ID1 = 29 - cellid1 is sored CHBIT_PDG = 28 - store pdg of secondaries

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Bit 16 is used to flag collection as transient Bit 17 is used to flag collection as default Bit 18 is used to flag collection as subset Bits 0-15 are subdetector/user specific.

See Also isTransient() (p. 90)

Implements EVENT::LCCollection (p. 84).

7.47.3.3 bool IMPL::LCCollectionVec::isDefault ( ) const [virtual]

True if collection is the default collection for the given type. This implies that the collection is complete and unambigous. Convenient method that checks bit BITDefault of the flag word. Implements EVENT::LCCollection (p. 85).

7.47.3.4 bool IMPL::LCCollectionVec::isSubset ( ) const [virtual]

True if the collection holds a subset of objects from other collections. If the collection is not transient only the pointers/references to the original objects will be stored. Convenient method that checks bit BITSubset of the flag word. Implements EVENT::LCCollection (p. 85). Referenced by ∼LCCollectionVec().

7.47.3.5 bool IMPL::LCCollectionVec::isTransient ( ) const [virtual]

True if collection is transient, i.e. will not be written to any LCIO file. Convenient method that checks bit 16 of the flag word. Implements EVENT::LCCollection (p. 86). References EVENT::LCCollection::BITTransient.

7.47.3.6 void IMPL::LCCollectionVec::removeElementAt ( int i ) throw EVENT::ReadOnlyException) [virtual]

Removes the i-th element from the collection. Throws an exception if the collection (event) is ’read only’. Exceptions

ReadOnlyException

Implements EVENT::LCCollection (p. 86).

7.47.3.7 void IMPL::LCCollectionVec::setDefault ( bool val = true )

Sets the default flag for this collection. User need to make sure this flag is unique for a type.

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.47.3.8 void IMPL::LCCollectionVec::setSubset ( bool val = true )

Sets the subset flag for this collection. Collections of subsets are only written as pointers to LCIO files, i.e. the objects have to exist in another non-transient collections. NB: Do not add any ’real’ objects to a collection flagged as subset, as this will lead to memory leaks !

7.47.3.9 void IMPL::LCCollectionVec::setTransient ( bool val = true )

Sets the transient flag for this collection. Transient collections are not written to LCIO files. References EVENT::LCCollection::BITTransient. Referenced by IMPL::LCEventImpl::takeCollection(). The documentation for this class was generated from the following files:

• LCCollectionVec.h

• LCCollectionVec.cc

7.48 EVENT::LCEvent Class Reference

The main event interface.

#include <pre-generated/EVENT/LCEvent.h>

Inheritance diagram for EVENT::LCEvent:

Public Member Functions

• virtual ∼LCEvent ()

Destructor.

• virtual int getRunNumber () const =0

Return the run number off this event.

• virtual int getEventNumber () const =0

Returns this event’s number .

• virtual const std::string & getDetectorName () const =0

Returns the name of the detector setup used in the simulation.

• virtual long64 getTimeStamp () const =0

Returns the time stamp of the event.

• virtual double getWeight () const =0

Returns the event weight.

• virtual const std::vector < std::string > ∗ getCollectionNames () const =0

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Returns the names of the collections in the event.

• virtual LCCollection ∗ getCollection (const std::string &name) const =0 throw (DataNotAvailableException,

std::exception )

Returns the collection for the given name.

• virtual LCCollection ∗ takeCollection (const std::string &name) const =0 throw (DataNotAvailableException,

std::exception )

Returns the collection for the given name and transfers the ownership of the collection to the caller.

• virtual void addCollection (LCCollection ∗col, const std::string &name)=0 throw (EventException, std-

::exception )

Adds a collection with the given name (has to be a valid C/C++ variable name).

• virtual void removeCollection (const std::string &name)=0 throw (ReadOnlyException, std::exception )

Removes (and deletes) the collection with name (if it exists in the event).

• virtual const LCParameters & getParameters () const =0

Parameters defined for this event.

• virtual LCParameters & parameters ()=0

Parameters defined for this event.

7.48.1 Detailed Description

The main event interface.

Holds named collections with elements of type LCObject (p. 123) (’untyped’). Users have to cast to the corresponding types.

Author gaede

Version

Id: LCEvent.aid,v 1.14 2007-11-08 19:35:37 gaede Exp

See Also LCCollection (p. 83)

7.48.2 Member Function Documentation

7.48.2.1 virtual void EVENT::LCEvent::addCollection ( LCCollection ∗ col, const std::string & name ) throw EventException, std::exception) [pure virtual]

Adds a collection with the given name (has to be a valid C/C++ variable name).

Throws an exception if the name already exists in the event. NB: Adding collections is allowed even when the event is ’read only’.

See Also validateCollectionName Exceptions

EventException (p. 69)

Implemented in IMPL::LCEventImpl (p. 96).

7.48.2.2 virtual LCCollection∗ EVENT::LCEvent::getCollection ( const std::string & name ) const throw DataNotAvailableException, std::exception) [pure virtual]

Returns the collection for the given name. Exceptions

DataNotAvailable-

Exception (p. 66)

Implemented in IMPL::LCEventImpl (p. 96). Referenced by UTIL::LCTOOLS::dumpEvent(), UTIL::LCTOOLS::dumpEventDetailed(), and SIO::SIOWriter::set-UpHandlers().

7.48.2.3 virtual long64 EVENT::LCEvent::getTimeStamp ( ) const [pure virtual]

Returns the time stamp of the event. Timestamp is a unique 64 bit integer. Definition: ns since January 1, 1970, 00:00:00 GMT Implemented in IMPL::LCEventImpl (p. 94). Referenced by UTIL::LCTOOLS::dumpEventDetailed().

7.48.2.4 virtual void EVENT::LCEvent::removeCollection ( const std::string & name ) throw ReadOnlyException, std::exception) [pure virtual]

Removes (and deletes) the collection with name (if it exists in the event). Throws an exception if the event is ’read only’ as defined by the read mode in LCReader. Exceptions

ReadOnlyException (p. 192)

Implemented in IMPL::LCEventImpl (p. 97).

7.48.2.5 virtual LCCollection∗ EVENT::LCEvent::takeCollection ( const std::string & name ) const throw DataNotAvailableException, std::exception) [pure virtual]

Returns the collection for the given name and transfers the ownership of the collection to the caller. The caller is responsible for deleting the collection after the Event is deleted. The collection is still in the event but flagged as transient after the call returns. This is usefull when you want to keep the collection for the next events. Use with care!

Exceptions

DataNotAvailable-

Exception (p. 66)

Implemented in IMPL::LCEventImpl (p. 97). The documentation for this class was generated from the following file:

• LCEvent.h

7.49 IMPL::LCEventImpl Class Reference

Implementation of the main event class.

#include <IMPL/LCEventImpl.h>

Inheritance diagram for IMPL::LCEventImpl:

Public Member Functions

• virtual ∼LCEventImpl ()

Copy contructor, creates a deep copy of the event.

• virtual int getRunNumber () const

Return the run number off this event.

• virtual int getEventNumber () const

Returns this event’s number .

• virtual const std::string & getDetectorName () const

Returns the name of the detector setup used in the simulation.

• virtual EVENT::long64 getTimeStamp () const

Returns the time stamp of the event.

• virtual double getWeight () const

Returns the event weight.

• virtual const std::vector

< std::string > ∗ getCollectionNames () const

Returns the names of the collections in the event.

• virtual EVENT::LCCollection ∗ getCollection (const std::string &name) const throw (EVENT::DataNot-

AvailableException, std::exception)

Returns the collection for the given name.

• virtual EVENT::LCCollection ∗ takeCollection (const std::string &name) const throw (EVENT::DataNot-

AvailableException, std::exception )

Returns the collection for the given name and transfers the ownership of the collection to the caller.

• virtual void addCollection (EVENT::LCCollection ∗col, const std::string &name) throw (EVENT::Event-

Exception, std::exception)

Adds a collection with the given name (has to be a valid C/C++ variable name).

• virtual void removeCollection (const std::string &name) throw (EVENT::ReadOnlyException, std::exception)

Removes (and deletes) the collection with name (if it exists in the event).

• virtual const EVENT::LCParameters & getParameters () const

Returns the names of the relations in the event.

• virtual EVENT::LCParameters & parameters ()

Parameters defined for this run.

• void setRunNumber (int rn)

Sets the run number.

• void setEventNumber (int en)

Sets the event number.

• void setDetectorName (const std::string &dn)

Sets the detector name.

• void setTimeStamp (EVENT::long64 ts)

Sets the event time stamp.

• void setWeight (double w)

Set the event weight.

Protected Member Functions

• void setAccessMode (int accessMode)

• bool validateCollectionName (const char ∗name)

Tests the validity of a collection name.

Protected Attributes

• int _runNumber

• int _eventNumber

• EVENT::long64 _timeStamp

• std::string _detectorName

• LCCollectionMap _colMap

• std::vector< std::string > _colNames

• LCParametersImpl _params

• LCCollectionSet _notOwned

7.49.1 Detailed Description

Implementation of the main event class.

This is used by LCIO for reading events from file.

Author gaede

Version Mar 5, 2003

See Also LCEvent LCCollection

96  Class Documentation

7.49.2 Constructor & Destructor Documentation

7.49.2.1 IMPL::LCEventImpl::∼LCEventImpl ( ) [virtual] Copy contructor, creates a deep copy of the event. Not yet - needs pointer chasing ...Destructor.

7.49.3 Member Function Documentation

7.49.3.1 void IMPL::LCEventImpl::addCollection ( EVENT::LCCollection ∗ col, const std::string & name ) throw

EVENT::EventException, std::exception) [virtual] Adds a collection with the given name (has to be a valid C/C++ variable name). Throws an exception if the name already exists in the event. NB: Adding collections is allowed even when the event is ’read only’.

See Also validateCollectionName (p. 97) Exceptions

EventException

Implements EVENT::LCEvent (p. 92). Referenced by UTIL::LCStdHepRdr::updateNextEvent(), and UTIL::LCStdHepRdrNew::updateNextEvent().

7.49.3.2 LCCollection ∗ IMPL::LCEventImpl::getCollection ( const std::string & name ) const throw

EVENT::DataNotAvailableException, std::exception) [virtual] Returns the collection for the given name. Exceptions

DataNotAvailableException Implements EVENT::LCEvent (p. 93). Referenced by HEPEVTIMPL::HEPEVT::fromHepEvt().

7.49.3.3 virtual const EVENT::LCParameters& IMPL::LCEventImpl::getParameters ( ) const [inline],

       [virtual]

Returns the names of the relations in the event. Returns the relation for the given name. Exceptions

DataNotAvailableException Adds a relation with the given name. Throws an exception if the name already exists in the event. NB: Adding relations is allowed even when the event is ’read only’.

EventException Removes (and deletes) the relation with name (if it exists in the event). Throws

an exception if the event is ’read only’ as defined by the read mode in LCReader.

ReadOnlyException-Parameters

defined for this event.

Implements EVENT::LCEvent (p. 92).

7.49.3.4 void IMPL::LCEventImpl::removeCollection ( const std::string & name ) throw EVENT::ReadOnlyException, std::exception) [virtual]

Removes (and deletes) the collection with name (if it exists in the event). Throws an exception if the event is ’read only’ as defined by the read mode in LCReader. Exceptions

ReadOnlyException

Implements EVENT::LCEvent (p. 93).

7.49.3.5 LCCollection ∗ IMPL::LCEventImpl::takeCollection ( const std::string & name ) const throw EVENT::DataNotAvailableException, std::exception) [virtual]

Returns the collection for the given name and transfers the ownership of the collection to the caller. The caller is responsible for deleting the collection after the Event is deleted. The collection is still in the event but flagged as transient after the call returns. This is usefull when you want to keep the collection for the next events. Use with care! Exceptions

DataNotAvailableException

Implements EVENT::LCEvent (p. 93). References IMPL::LCCollectionVec::setTransient().

7.49.3.6 bool IMPL::LCEventImpl::validateCollectionName ( const char ∗ name ) [protected]

Tests the validity of a collection name. LCIO only accepts collection names starting with (regular expression) [A-Za-z_] and continuing with A-Za-z0-9-

_.

SIO (p. 30) only accepts names starting with (regular expression) [A-Za-z_] and continuing with A-Za-z0-9_. The documentation for this class was generated from the following files:

• LCEventImpl.h

• LCEventImpl.cc

7.50 IOIMPL::LCEventIOImpl Class Reference

Adding stuff needed for io (friend declarations, etc.)

#include <IOIMPL/LCEventIOImpl.h>

Inheritance diagram for IOIMPL::LCEventIOImpl:

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Friends

• class SIO::SIOReader

• class SIO::SIOEventHeaderHandler

• class SIO::SIOEventHandler

Additional Inherited Members

7.50.1 Detailed Description

Adding stuff needed for io (friend declarations, etc.)

Author gaede

Version Mar 6, 2003

The documentation for this class was generated from the following file:

• LCEventIOImpl.h

7.51 IO::LCEventListener Class Reference

Listener for the occurence of LCEvents when reading a stream.

#include <pre-generated/IO/LCEventListener.h>

Public Member Functions

• virtual ∼LCEventListener ()

Destructor.

• virtual void processEvent (EVENT::LCEvent ∗evt)=0

Call back for analyzing an LCEvent.

• virtual void modifyEvent (EVENT::LCEvent ∗evt)=0

Call back for updating an LCEvent.

7.52 lcrtrel::LCExtension< U, T > Struct Template Reference  99

7.51.1 Detailed Description

Listener for the occurence of LCEvents when reading a stream. User classes should implement this interface for their analysis code and then register the listener with LCReader (p. 130).

Author gaede

Version Mar 11, 2003

See Also LCEvent LCReader (p. 130)

7.51.2 Member Function Documentation

7.51.2.1 virtual void IO::LCEventListener::modifyEvent ( EVENT::LCEvent ∗ evt ) [pure virtual]

Call back for updating an LCEvent. Here all modifications are allowed. Only use this method if you need to remove a collection from the event or you need to modify (fix) exisiting data.

7.51.2.2 virtual void IO::LCEventListener::processEvent ( EVENT::LCEvent ∗ evt ) [pure virtual]

Call back for analyzing an LCEvent. The access to the event is ’read only’. Adding new collections to the event is allowed wheras existing data collections and elements can not be modified. The documentation for this class was generated from the following file:

• LCEventListener.h

7.52 lcrtrel::LCExtension< U, T > Struct Template Reference

Simple Extension - pointer to an object of type T.

#include <LCRTRelations.h>

Inheritance diagram for lcrtrel::LCExtension< U, T >:

Additional Inherited Members

7.52.1 Detailed Description

template<class U, typename T>struct lcrtrel::LCExtension< U, T >

Simple Extension - pointer to an object of type T. The class U needs to be the subclass type, e.g. struct MyAttributes : public LCExtension<MyAttributes,Attributes> {} ; The documentation for this struct was generated from the following file:

• LCRTRelations.h

7.53 lcrtrel::LCExtensionList< U, T > Class Template Reference

Extension list holding pointers to objects of type T - no ownership of the objects is taken.

#include <LCRTRelations.h>

Inheritance diagram for lcrtrel::LCExtensionList< U, T >:

Additional Inherited Members

7.53.1 Detailed Description

template<class U, class T>class lcrtrel::LCExtensionList< U, T >

Extension list holding pointers to objects of type T - no ownership of the objects is taken. The documentation for this class was generated from the following file:

• LCRTRelations.h

7.54 lcrtrel::LCExtensionVector< U, T > Class Template Reference

Extension vector holding pointers to objects of type T - no ownership of the objects is taken.

#include <LCRTRelations.h>

Inheritance diagram for lcrtrel::LCExtensionVector< U, T >:

Additional Inherited Members

7.54.1 Detailed Description

template<class U, class T>class lcrtrel::LCExtensionVector< U, T >

Extension vector holding pointers to objects of type T - no ownership of the objects is taken. The documentation for this class was generated from the following file:

• LCRTRelations.h

7.55 IOIMPL::LCFactory Class Reference

Factory to create LCWriter and Reader object for the known data formats, e.g.

#include <IOIMPL/LCFactory.h>

Inheritance diagram for IOIMPL::LCFactory:

Public Member Functions

• virtual ∼LCFactory ()

Destructor.

• virtual IO::LCWriter ∗ createLCWriter ()

Creates an LCWriter object for the current persistency type.

• virtual IO::LCReader ∗ createLCReader (int lcReaderFlag=0)

Creates an LCReader object for the current persistency type.

Static Public Member Functions

• static LCFactory ∗ getInstance ()

Returns the instance of the factory.

7.55.1 Detailed Description

Factory to create LCWriter and Reader object for the known data formats, e.g. SIO (p. 30). Singleton class that knows all concrete reader/writer implementations.

Author gaede

Version Mar 10, 2003

7.55.2 Member Function Documentation

7.55.2.1 LCReader ∗ IOIMPL::LCFactory::createLCReader ( int lcReaderFlag = 0 ) [virtual]

Creates an LCReader object for the current persistency type.

lcReaderFlag: configuration options for the LCReader object - combine multible options with ’|’. So far only LC-Reader::directAccess.

Implements IO::ILCFactory (p. 72).

7.55.2.2 LCFactory ∗ IOIMPL::LCFactory::getInstance ( ) [static]

Returns the instance of the factory.

In the future we need to specify the data format here... This doesn’t work for java as an interface can’t have static members ...

The documentation for this class was generated from the following files:

• LCFactory.h

• LCFactory.cc

7.56 UTIL::LCFixedObject< NINT, NFLOAT, NDOUBLE > Class Template Reference

Template class for fixed size LCGenericObject subclasses.

#include <UTIL/LCFixedObject.h>

Inheritance diagram for UTIL::LCFixedObject< NINT, NFLOAT, NDOUBLE >:

Public Member Functions

• LCFixedObject ()

Default c’tor.

• LCFixedObject (LCObject ∗obj)

C’tor to be used for elements of LCGenericObjects read from an LCIO file or the database.

• LCGenericObjectImpl ∗ obj ()

The LCGenericObjectImpl .

• virtual ∼LCFixedObject ()

Clean up if we created a new LCGenericObjectImpl.

• virtual int id () const

Return the id of the underlying LCGenericObjectImpl.

• int getNInt () const

• int getNFloat () const

• int getNDouble () const

• int getIntVal (int index) const

• float getFloatVal (int index) const

• double getDoubleVal (int index) const

• bool isFixedSize () const

7.56 UTIL::LCFixedObject< NINT, NFLOAT, NDOUBLE > Class Template Reference  103

Protected Attributes

• LCGenericObjectImpl ∗ _obj

• bool _createdObject

7.56.1 Detailed Description

template<int NINT, int NFLOAT, int NDOUBLE>class UTIL::LCFixedObject< NINT, NFLOAT, NDOUBLE >

Template class for fixed size LCGenericObject subclasses. By inheriting from this class users can very easily implement their own objects. It uses an instance of LCGenericObjectImpl that holds the data, thus there is no overhead when the data is read from a database or file for copying it to some local structure (Decorator pattern). This is still an abstract class: subclasses have to implement LCGenericObject::getTypeName() and LCGeneric-Object::getDataDescription().

Author gaede

Version

Id: LCFixedObject.h (p. ?? ),v 1.3 2011-03-13 12:41:49 gaede Exp

7.56.2 Constructor & Destructor Documentation

7.56.2.1 template<int NINT, int NFLOAT, int NDOUBLE> UTIL::LCFixedObject< NINT, NFLOAT, NDOUBLE >::LCFixedObject ( LCObject ∗ obj ) [inline]

C’tor to be used for elements of LCGenericObjects read from an LCIO file or the database. Subclasses should ’override’ this, e.g.: Myclass(LCObject∗ obj) : LCFixedObject(obj) {} References UTIL::LCFixedObject< NINT, NFLOAT, NDOUBLE >::obj().

7.56.3 Member Function Documentation

7.56.3.1 template<int NINT, int NFLOAT, int NDOUBLE> LCGenericObjectImpl∗ UTIL::LCFixedObject< NINT, NFLOAT, NDOUBLE >::obj ( ) [inline]

The LCGenericObjectImpl . Sublcasses use this to access their data. Referenced by UTIL::LCFixedObject< NINT, NFLOAT, NDOUBLE >::LCFixedObject(). The documentation for this class was generated from the following file:

• LCFixedObject.h

7.57 EVENT::LCFlag Class Reference

Helper class to create and interpret the 32-bit flag word in LCCollections.

#include <pre-generated/EVENT/LCFlag.h>

Inheritance diagram for EVENT::LCFlag:

Public Member Functions

• virtual ∼LCFlag ()

Destructor.

• virtual bool bitSet (int index) const =0

Returns true if bit at given index is set.

• virtual int getFlag () const =0

Returns the flag word.

• virtual void setBit (int bit)=0

Sets bit to 1.

• virtual void unsetBit (int bit)=0

Sets bit to 0.

7.57.1 Detailed Description

Helper class to create and interpret the 32-bit flag word in LCCollections. Bits are defined in class LCIO (p. 115).

See Also LCIO (p. 115)

Author gaede

Version May 6, 2003

The documentation for this class was generated from the following file:

• LCFlag.h

7.58 IMPL::LCFlagImpl Class Reference

Implementation of helper class to create and interpret the 32-bit flag word in LCCollections.

#include <IMPL/LCFlagImpl.h>

7.59 lcrtrel::LCFloatExtension< U > Struct Template Reference  105

Inheritance diagram for IMPL::LCFlagImpl:

Public Member Functions

• LCFlagImpl ()

Constructor initializing flag with 0.

• LCFlagImpl (int flag)

Constructor for evaluating an existing flag word.

• virtual ∼LCFlagImpl ()

Destructor.

• virtual bool bitSet (int index) const

Returns true if bit at given index is set.

• virtual int getFlag () const

Returns the flag word.

• virtual void setBit (int bit)

Sets bit to 1.

• virtual void unsetBit (int bit)

Sets bit to 0.

7.58.1 Detailed Description

Implementation of helper class to create and interpret the 32-bit flag word in LCCollections.

Bits are defined in class LCIO.

See Also LCIO

Author gaede

Version May 6, 2003

The documentation for this class was generated from the following files:

• LCFlagImpl.h

• LCFlagImpl.cc

7.59 lcrtrel::LCFloatExtension< U > Struct Template Reference

Public Types

• typedef double ptr

• typedef double & ext_type

• typedef U tag

106  Class Documentation

Static Public Member Functions

• static void clean (void ∗)

• static ptr init ()

• static DeleteFPtr deletePtr ()

Static Public Attributes

• static const int allowed_to_call_ext = 1 The documentation for this struct was generated from the following file:

• LCRTRelations.h

7.60 EVENT::LCFloatVec Class Reference

Float vector used for user extensions.

#include <pre-generated/EVENT/LCFloatVec.h>

Inheritance diagram for EVENT::LCFloatVec:

Public Types

• typedef LCFloatVec lcobject_type

Public Member Functions

• virtual ∼LCFloatVec ()

Destructor.

• int id () const

Returns an object id for internal (debugging) use in LCIO (p. 115).

Protected Attributes

• IMPL::AccessChecked _acc

Additional Inherited Members 7.60.1 Detailed Description

Float vector used for user extensions. The cpp implementation is a std::vector<float>

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.61 UTIL::LCFourVector< TT > Class Template Reference  107

Author gaede

Version Mar 12, 2003

See Also LCObject (p. 123)

The documentation for this class was generated from the following file:

• pre-generated/EVENT/LCFloatVec.h

7.61 UTIL::LCFourVector< TT > Class Template Reference

Four vector used in LCIO.

#include <UTIL/LCFourVector.h>

Inheritance diagram for UTIL::LCFourVector< TT >:

Public Member Functions

• LCFourVector (const TT ∗lcObj)

Constructor for templated type,e.g.

• LCFourVector (const EVENT::LCObject ∗lcObj)

Constructor for LCObject.

• const TT ∗ operator-> () const

Instances of LCFourVector (p. 107) serve as a handle to the templated type.

• const TT ∗ lcObj () const

Pointer to the LCObject that has been used to create the four vector.

Protected Attributes

• const TT ∗ _lcObj

7.61.1 Detailed Description

template<class TT>class UTIL::LCFourVector< TT >

Four vector used in LCIO. The cpp implementation is a CLHEP::HepLorentzVector. Instances of this class can be used as CLHEP::Hep-LorentzVector AND pointers to the corresponding LCIO class. Currently available are:

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

• LCFourVector<MCParticle>/MCParticle4V

• LCFourVector<ReconstructedParticle>/ReconstructedParticle4V

Author gaede

Version Mar 12, 2004

7.61.2 Constructor & Destructor Documentation

7.61.2.1 template<class TT> UTIL::LCFourVector< TT >::LCFourVector ( const TT ∗ lcObj ) [inline]

Constructor for templated type,e.g. LCFourVector<MCParticle>( myMCParticle ).

7.61.2.2 template<class TT> UTIL::LCFourVector< TT >::LCFourVector ( const EVENT::LCObject ∗ lcObj )

       [inline]

Constructor for LCObject. Can be used to save the dynamic_cast to the explicit type from an LCCollection element, e.g. LCFourVector<MC-Particle>( col->getElement(i) ).

7.61.3 Member Function Documentation

7.61.3.1 template<class TT> const TT∗ UTIL::LCFourVector< TT >::lcObj ( ) const [inline]

Pointer to the LCObject that has been used to create the four vector. To be used when the object is needed as a function argument or optionally to call methods of the object, e.g. particle4V.lcObj()->getMomentum() ;

7.61.3.2 template<class TT> const TT∗ UTIL::LCFourVector< TT >::operator-> ( ) const [inline]

Instances of LCFourVector (p. 107) serve as a handle to the templated type. Methods of the original type, e.g. MCParticle can be accessed via the ’->’ operator. Methods of the four vector implementation (HepLorentzvector) are accessed via the ’.’ operator. For Example: LCFourVector<MCParticle> particle4V( col->getElementAt(i) ) particle4V->getMomentum() // LCIO object particle4V.m() // mass from HepLorentzVector The documentation for this class was generated from the following file:

• LCFourVector.h

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.62 EVENT::LCGenericObject Class Reference  109

7.62 EVENT::LCGenericObject Class Reference

Simple interface to store generic user data.

#include <pre-generated/EVENT/LCGenericObject.h>

Inheritance diagram for EVENT::LCGenericObject:

Public Types

• typedef LCGenericObject lcobject_type

Useful typedef for template programming with LCIO (p. 115).

Public Member Functions

• virtual ∼LCGenericObject ()

Destructor.

• virtual int getNInt () const =0

Number of integer values stored in this object.

• virtual int getNFloat () const =0

Number of float values stored in this object.

• virtual int getNDouble () const =0

Number of double values stored in this object.

• virtual int getIntVal (int index) const =0

Returns the integer value for the given index.

• virtual float getFloatVal (int index) const =0

Returns the float value for the given index.

• virtual double getDoubleVal (int index) const =0

Returns the double value for the given index.

• virtual bool isFixedSize () const =0

True if objects of the implementation class have a fixed size, i.e getNInt, getNFloat and getNDouble will return values that are constant during the lifetime of the object.

• virtual const std::string getTypeName () const =0

The type name of the user class (typically the class name) This type name is stored as a collection parameter "Type-Name" with every collection of LCGenericObject (p. 109) subclasses.

• virtual const std::string getDataDescription () const =0

The description string.

• virtual int id () const

Provide default implementation for id.

Additional Inherited Members

7.62.1 Detailed Description

Simple interface to store generic user data. To store your own classes they have to implement this interface. 2005-02-11, gaede: Changed the return value of getTypeName() (p. 110) and getDataDescription() (p. 110) from const std::string& to const std::string to make the implementation easier

Author gaede

Version

Id: LCGenericObject.aid,v 1.7 2006-08-04 16:52:44 gaede Exp

7.62.2 Member Function Documentation

7.62.2.1 virtual const std::string EVENT::LCGenericObject::getDataDescription ( ) const [pure virtual]

The description string. A comma separated list of pairs of type identifier, one of ’i’,’f’,’d’ followed by ’:’ and an attribute name, e.g. "i:cellId,f- :offset,f:gain". Implemented in IMPL::LCGenericObjectImpl (p. 112).

7.62.2.2 virtual const std::string EVENT::LCGenericObject::getTypeName ( ) const [pure virtual]

The type name of the user class (typically the class name) This type name is stored as a collection parameter "TypeName" with every collection of LCGenericObject (p. 109) subclasses. Implemented in IMPL::LCGenericObjectImpl (p. 111). The documentation for this class was generated from the following file:

• LCGenericObject.h

7.63 IMPL::LCGenericObjectImpl Class Reference

Default LCIO implementation of the interface to store generic user data.

#include <IMPL/LCGenericObjectImpl.h>

Inheritance diagram for IMPL::LCGenericObjectImpl:

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Public Member Functions

• LCGenericObjectImpl ()

Variable size c’tor.

• LCGenericObjectImpl (int nInt, int nFloat, int nDouble)

Fixed size c’tor.

• virtual ∼LCGenericObjectImpl ()

Destructor.

• virtual int id () const

Provide default implementation for id.

• virtual int getNInt () const

Number of integer values stored in this object.

• virtual int getNFloat () const

Number of float values stored in this object.

• virtual int getNDouble () const

Number of double values stored in this object.

• virtual int getIntVal (int index) const

Returns the integer value for the given index.

• virtual float getFloatVal (int index) const

Returns the float value for the given index.

• virtual double getDoubleVal (int index) const

Returns the double value for the given index.

• virtual void setIntVal (unsigned index, int value)

Sets the integer value at the given index.

• virtual void setFloatVal (unsigned index, float value)

Sets the float value at the given index.

• virtual void setDoubleVal (unsigned index, double value)

Sets the double value at the given index.

• virtual bool isFixedSize () const

True if objects of the implementation class have a fixed size, i.e getNInt, getNFloat and getNDouble will return values that are constant during the lifetime of the object.

• virtual const std::string getTypeName () const

The type name of the user class (typically the class name)

• virtual const std::string getDataDescription () const

The description string.

112  Class Documentation

Protected Attributes

• std::vector< int > _intVec

• std::vector< float > _floatVec

• std::vector< double > _doubleVec

• bool _isFixedSize

Static Protected Attributes

• static std::string _typeName = "LCGenericObject"

• static std::string _dataDescription = ""

Additional Inherited Members 7.63.1 Detailed Description

Default LCIO implementation of the interface to store generic user data.

Author gaede

Version

Id: LCGenericObjectImpl.h (p. ?? ),v 1.5 2006-08-04 16:52:46 gaede Exp

7.63.2 Member Function Documentation

7.63.2.1 const std::string IMPL::LCGenericObjectImpl::getDataDescription ( ) const [virtual] The description string. A comma separated list of pairs of type identifier, one of ’i’,’f’,’d’ followed by ’:’ and an attribute name, e.g. "i:cellId,f- :offset,f:gain". Implements EVENT::LCGenericObject (p. 110). The documentation for this class was generated from the following files:

• LCGenericObjectImpl.h

• LCGenericObjectImpl.cc

7.64 IOIMPL::LCGenericObjectIOImpl Class Reference

Adding stuff needed for io (friend declarations, etc.)

#include <IOIMPL/LCGenericObjectIOImpl.h>

Inheritance diagram for IOIMPL::LCGenericObjectIOImpl:

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.65 lcrtrel::LCIntExtension< U > Struct Template Reference  113

Public Member Functions

• LCGenericObjectIOImpl (int nInt, int nFloat, int nDouble)

Friends

• class SIO::SIOLCGenericObjectHandler

Additional Inherited Members 7.64.1 Detailed Description

Adding stuff needed for io (friend declarations, etc.)

Author gaede

Version

Id: LCGenericObjectIOImpl.h (p. ?? ),v 1.2 2005-04-15 08:37:40 gaede Exp The documentation for this class was generated from the following file:

• LCGenericObjectIOImpl.h

7.65 lcrtrel::LCIntExtension< U > Struct Template Reference

Special Extension that allows to write int extensions directly (not through a pointer !).

#include <LCRTRelations.h>

Public Types

• typedef long ptr

• typedef long & ext_type

• typedef U tag

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

114  Class Documentation

Static Public Member Functions

• static void clean (void ∗)

• static ptr init ()

• static DeleteFPtr deletePtr ()

Static Public Attributes

• static const int allowed_to_call_ext = 1

7.65.1 Detailed Description

template<class U>struct lcrtrel::LCIntExtension< U > Special Extension that allows to write int extensions directly (not through a pointer !). The documentation for this struct was generated from the following file:

• LCRTRelations.h

7.66 EVENT::LCIntVec Class Reference

Int vector used for user extensions.

#include <pre-generated/EVENT/LCIntVec.h>

Inheritance diagram for EVENT::LCIntVec:

Public Types

• typedef LCIntVec lcobject_type

Public Member Functions

• virtual ∼LCIntVec ()

Destructor.

• int id () const

Returns an object id for internal (debugging) use in LCIO (p. 115).

Protected Attributes

• IMPL::AccessChecked _acc

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Additional Inherited Members

7.66.1 Detailed Description

Int vector used for user extensions.

The cpp implementation is a std::vector<int>

Author gaede

Version Aug 8, 2003

See Also LCObject (p. 123)

The documentation for this class was generated from the following file:

• pre-generated/EVENT/LCIntVec.h

7.67 EVENT::LCIO Class Reference

Global constants used in LCIO (p. 115).

#include <pre-generated/EVENT/LCIO.h>

Public Member Functions

• virtual ∼LCIO ()

Destructor.

Static Public Attributes

• static const int MAJORVERSION = LCIO_MAJOR_VERSION

• static const int MINORVERSION = LCIO_MINOR_VERSION

• static const int CHBIT_LONG = 31

• static const int CHBIT_BARREL = 30

• static const int CHBIT_ID1 = 29

• static const int CHBIT_PDG = 28

• static const int CHBIT_STEP = 28

• static const int RCHBIT_LONG = 31

• static const int RCHBIT_BARREL = 30

• static const int RCHBIT_ID1 = 29

• static const int RCHBIT_NO_PTR = 28

• static const int RCHBIT_TIME = 27

• static const int RCHBIT_ENERGY_ERROR = 26

• static const int TRAWBIT_ID1 = 31

• static const int TRAWBIT_CM = 30

• static const int RTHBIT_ID1 = 31

• static const int RTHPBIT_ID1 = 31 • static const int RTHZBIT_ID1 = 31

• static const int THBIT_BARREL = 31

• static const int THBIT_MOMENTUM = 30

• static const int THBIT_ID1 = 29

• static const int TRBIT_HITS = 31

• static const int CLBIT_HITS = 31

• static const int TPCBIT_RAW = 31

• static const int TPCBIT_NO_PTR = 30

• static const int LCREL_WEIGHTED = 31

• static const int GOBIT_FIXED = 31

• static const int READ_ONLY = 0

• static const int UPDATE = 1

• static const int ERROR = 0

• static const int SUCCESS = 1

• static const int WRITE_NEW = 0

• static const int WRITE_APPEND = 1

• static const char ∗ LCEVENT = "LCEvent"

• static const char ∗ LCRUNHEADER = "LCRunHeader"

• static const char ∗ LCCOLLECTION = "LCCollection"

• static const char ∗ MCPARTICLE = "MCParticle"

• static const char ∗ SIMCALORIMETERHIT = "SimCalorimeterHit"

• static const char ∗ RAWCALORIMETERHIT = "RawCalorimeterHit"

• static const char ∗ CALORIMETERHIT = "CalorimeterHit"

• static const char ∗ SIMTRACKERHIT = "SimTrackerHit"

• static const char ∗ TPCHIT = "TPCHit"

• static const char ∗ TRACKERRAWDATA = "TrackerRawData"

• static const char ∗ TRACKERDATA = "TrackerData"

• static const char ∗ TRACKERPULSE = "TrackerPulse"

• static const char ∗ TRACKERHIT = "TrackerHit"

• static const char ∗ TRACKERHITPLANE = "TrackerHitPlane"

• static const char ∗ TRACKERHITZCYLINDER = "TrackerHitZCylinder"

• static const char ∗ LCSTRVEC = "LCStrVec"

• static const char ∗ LCFLOATVEC = "LCFloatVec"

• static const char ∗ LCINTVEC = "LCIntVec"

• static const char ∗ TRACK = "Track"

• static const char ∗ TRACKSTATE = "TrackState"

• static const char ∗ CLUSTER = "Cluster"

• static const char ∗ RECONSTRUCTEDPARTICLE = "ReconstructedParticle"

• static const char ∗ LCRELATION = "LCRelation"

• static const char ∗ LCGENERICOBJECT = "LCGenericObject"

• static const char ∗ PARTICLEID = "ParticleID"

• static const char ∗ VERTEX = "Vertex"

• static const char ∗ CellIDEncoding = "CellIDEncoding"

7.67.1 Detailed Description

Global constants used in LCIO (p. 115).

Author gaede

7.68 UTIL::LCIO_LONG< T > Class Template Reference  117

Version

Id: LCIO.aid,v 1.77 2010-10-29 10:26:09 gaede Exp

See Also LCObject (p. 123)

LCIO (p. 115) The documentation for this class was generated from the following files:

• LCIO.h

• LCIO.cc

7.68 UTIL::LCIO_LONG< T > Class Template Reference

Public Member Functions

• LCIO_LONG (const T &o, const EVENT::LCCollection ∗c)

• const T ∗ object (void) const

• const EVENT::LCCollection ∗ collection (void) const The documentation for this class was generated from the following file:

• Operators.h

7.69 UTIL::lcio_short< T > Class Template Reference

Public Member Functions

• lcio_short (const T ∗t, const EVENT::LCCollection ∗b=NULL)

Public Attributes

• const T ∗ obj

• const EVENT::LCCollection ∗ col The documentation for this class was generated from the following file:

• Operators.h

7.70 IMPL::LCIOExceptionHandler Class Reference

Sets the default error handling for unexpected and uncaught exceptions.

#include <IMPL/LCIOExceptionHandler.h>

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Static Public Member Functions

• static LCIOExceptionHandler ∗ createInstance ()

7.70.1 Detailed Description

Sets the default error handling for unexpected and uncaught exceptions. The documentation for this class was generated from the following files:

• LCIOExceptionHandler.h

• LCIOExceptionHandler.cc

7.71 SIO::LCIORandomAccess Class Reference

Implementation class for LCIORandomAccess (p. 118) records.

#include <SIO/LCIORandomAccess.h>

Public Member Functions

• long64 getIndexLocation () const

• long64 getPrevLocation () const

• long64 getFirstRecordLocation () const

• void setIndexLocation (long64 il)

• void setPreviousLocation (long64 pl)

• void setFirstRecordLocation (long64 fl)

Protected Attributes

• RunEvent _minRunEvt

• RunEvent _maxRunEvt

• int _nRunHeaders

• int _nEvents

• int _recordsAreInOrder

• long64 _indexLocation

• long64 _prevLocation

• long64 _nextLocation

• long64 _firstRecordLocation

Friends

• class SIORandomAccessHandler

• class LCIORandomAccessMgr

• std::ostream & operator<< (std::ostream &os, const LCIORandomAccess &ra)

• bool operator< (const LCIORandomAccess ra0, const LCIORandomAccess &other)

7.72 SIO::LCIORandomAccessMgr Class Reference  119

7.71.1 Detailed Description

Implementation class for LCIORandomAccess (p. 118) records.

Author gaede

Version

Id: LCIORandomAccess.h (p. ?? ),v 1.2 2010-06-22 13:49:54 gaede Exp

The documentation for this class was generated from the following file:

• LCIORandomAccess.h

7.72 SIO::LCIORandomAccessMgr Class Reference

Manager class for LCIO direct access.

#include <SIO/LCIORandomAccessMgr.h>

Public Member Functions

• long64 getPosition (const RunEvent &re)

Return the position of the specified Event record or Run record respectively (if EventNum == -1 ).

• void add (const RunEvent &re, long64 pos)

Add a new entry to the event map - if the RunEvent (p. 202) already exists the new position will be stored.

• bool createEventMap (SIO_stream ∗s)

Get the run and event header map from the stream - either by reading the random access records or by recreating it for old files.

• const RunEventMap & getEventMap ()

Return the event map - it will be empty, if not yet created.

• bool initAppend (SIO_stream ∗s)

Initialize random access for append mode: read last LCIORandomAccess (p. 118) record if it exists - recreate the RunEvent (p. 202) map from the file if not (old files).

• void writeRandomAccessRecords (SIO_stream ∗stream)

Write the current random access records LCIOIndex and LCIORandomAccess (p. 118) to the stream.

• void clear ()

Clear all lists and maps before closing a file.

Protected Member Functions

• LCIORandomAccess ∗ createFromEventMap ()

Prepare an LCIORandomAccess (p. 118) object from the current contents of RunEventMap (p. 202) (all file locations set to 0).

• void addLCIORandomAccess (LCIORandomAccess ∗ra)

Add a new LCIORandomAccess (p. 118) object to the list.

• bool readLCIORandomAccessAt (SIO_stream ∗stream, long64 pos)

Read the LCIORandomAccess (p. 118) record at the specified position.

• bool readLCIOIndexAt (SIO_stream ∗stream, long64 pos)

Read the LCIOIndex record at the specified position.

• bool recreateEventMap (SIO_stream ∗)

Fill the RunEventMap (p. 202) from the event and run header records in the file.

• bool readLCIORandomAccess (SIO_stream ∗stream)

Helper for reading the next LCIORandomAccess (p. 118) record (need preceeding call to LCSIO::seek() )

• bool readLCIOIndex (SIO_stream ∗stream)

Helper for reading the next LCIOIndex record (need preceeding call to LCSIO::seek() )

• void createFileRecord ()

Create file record from all LCIORandomAccess (p. 118) records.

• const LCIORandomAccess ∗ lastLCIORandomAccess ()

Pointer to the last LCIORandomAccess (p. 118) in the list.

Protected Attributes

• RunEventMap _runEvtMap

• std::list< LCIORandomAccess ∗ > _list

• LCIORandomAccess ∗ _fileRecord

Friends

• class SIOIndexHandler

• class SIORandomAccessHandler

• std::ostream & operator<< (std::ostream &os, const LCIORandomAccessMgr &ra)

7.72.1 Detailed Description

Manager class for LCIO direct access.

Provides the functionality for reading and writing the LCIORandomAccess (p. 118) and LCIOIndex records as needed by SIOReader (p. 234) and SIOWriter (p. 253). Direct Access is implemented through appending LCIO-Index and LCIORandomAccess (p. 118) and records to the end of the LCIO file. The LCIOIndex records hold the locations of the RunHeader and EventHeader in the file and are referenced by corresponding LCIORandomAccess (p. 118) records. The last record in the file will always be an LCIORandomAccess (p. 118) record - and previous records can be found through pointers to the previous record (simply linked list). LCIORandomAccess (p. 118) are always stored uncompressed. When reading a file with direct access mode a RunEvent (p. 202) map is created by reading all LCIORandomAccess (p. 118) and LCIOIndex records or - for old files - is recreated from all RunHeader and EventHeader records in the file. If an old file is opened in APPEND mode the corresponding records are created and written at the end on close()

See Also writeRandomAccessRecords() (p. 119). : currently the last LCIORandomAccess (p. 118) record is found by seeking a fixed size from the end - need to store the actual length as last word in the file for future changes of the records length

Author gaede

Version

Id: LCIORandomAccessMgr.h (p. ?? ),v 1.3 2011-03-03 16:00:12 gaede Exp

7.73 UTIL::LCIterator< T > Class Template Reference  121

7.72.2 Member Function Documentation

7.72.2.1 LCIORandomAccess ∗ SIO::LCIORandomAccessMgr::createFromEventMap ( ) [protected]

Prepare an LCIORandomAccess (p. 118) object from the current contents of RunEventMap (p. 202) (all file locations set to 0). Referenced by writeRandomAccessRecords().

7.72.2.2 long64 SIO::LCIORandomAccessMgr::getPosition ( const RunEvent & re ) [inline]

Return the position of the specified Event record or Run record respectively (if EventNum == -1 ). Returns RunEventMap::NPos if no record found. References SIO::RunEventMap::getPosition(). The documentation for this class was generated from the following files:

• LCIORandomAccessMgr.h

• LCIORandomAccessMgr.cc

7.73 UTIL::LCIterator< T > Class Template Reference

Simple convenient iterator class for LCCollections that saves some boiler plate code.

#include <UTIL/LCIterator.h>

Public Member Functions

• LCIterator (EVENT::LCEvent ∗evt, const std::string &name)

Constructor that reads the collection with the given name from the event - note: no exception is thrown if the collection is not in the event, this will behave the same as an empty collection - use operator() to test, if the collection exists.

• LCIterator (EVENT::LCCollection ∗col)

Constructor for the given collection.

• T ∗ next ()

Returns the next element as long as there is one, otherwise 0 is returned.

• int size ()

Size of the collection.

• EVENT::LCCollection ∗ operator-> ()

Serves as a handle to the LCCollection itself, to provide access to the collection parameters etc.

• EVENT::LCCollection ∗ operator() ()

Return pointer to LCCollection, e.g.

7.73.1 Detailed Description

template<class T>class UTIL::LCIterator< T >

Simple convenient iterator class for LCCollections that saves some boiler plate code. LCIterators can be constructed either form an LCCollection or from the LCEvent and and a collection name. Examples:

 // ------ use with while():
 LCIterator<TrackerHit> it( evt, "VXDCollection" ) ;
 while( TrackerHit* hit = it.next()  ){

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

     hitMap[ hit->getCellID0() ] = hit ;
 }
 // --- with for() :
 for(  LCIterator<Track> it( col ) ;  Track* trk = it.next()  ; ) {
     std::cout << trk->getTrackState( TrackState::AtIP ) << std::endl  ;
 }

Author F.Gaede, DESY

Version $Id:$

7.73.2 Member Function Documentation

7.73.2.1 template<class T> EVENT::LCCollection∗ UTIL::LCIterator< T >::operator() ( ) [inline]

Return pointer to LCCollection, e.g. for testing whether the collections was in the event. The documentation for this class was generated from the following file:

• LCIterator.h

7.74 lcrtrel::LCNToNRelation< U, From, To > Struct Template Reference

Many to many relation between objects of type From to objects of type To.

#include <LCRTRelations.h>

Inheritance diagram for lcrtrel::LCNToNRelation< U, From, To >:

Additional Inherited Members

7.74.1 Detailed Description

template<class U, class From, class To>struct lcrtrel::LCNToNRelation< U, From, To >

Many to many relation between objects of type From to objects of type To. The documentation for this struct was generated from the following file:

• LCRTRelations.h

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.75 EVENT::LCObject Class Reference

The generic object that is held in an LCCollection (p. 83).

#include <pre-generated/EVENT/LCObject.h>

Inheritance diagram for EVENT::LCObject:

Public Member Functions

• virtual ∼LCObject ()

Destructor.

• virtual int id () const =0

Returns an object id for internal (debugging) use in LCIO (p. 115).

• virtual LCObject ∗ clone () const

Returns an object id for internal (debugging) use in LCIO (p. 115).

Additional Inherited Members

7.75.1 Detailed Description

The generic object that is held in an LCCollection (p. 83). Every object stored in an LCCollection (p. 83), i.e. in an LCEvent (p. 91) has to implement this interface.

Author gaede

Version Jun 7, 2003 fg 20040405 moved here from hep.lcio.data

7.75.2 Member Function Documentation

7.75.2.1 virtual LCObject∗ EVENT::LCObject::clone ( ) const [inline], [virtual] Returns an object id for internal (debugging) use in LCIO (p. 115). Deep copy of the object. C++ implementations should return the concrete type (virtual constructor). No reasonable implementation yet. Do not use this method. The documentation for this class was generated from the following file:

• LCObject.h

7.76 UTIL::LCObjectHandle< T > Class Template Reference

Template handle class for LCObjects.

#include <UTIL/LCObjectHandle.h>

Public Member Functions

• LCObjectHandle (T ∗lcObj)

• LCObjectHandle (EVENT::LCObject ∗lcObj)

• T ∗ operator-> ()

• T ∗ lcObj ()

Protected Attributes

• T ∗ _lcObj

7.76.1 Detailed Description

template<class T>class UTIL::LCObjectHandle< T > Template handle class for LCObjects.

Author gaede Version

Id: LCObjectHandle.h,v 1.4 2005-02-03 14:11:00 gaede Exp The documentation for this class was generated from the following file:

• UTIL/LCObjectHandle.h

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.77 lcrtrel::LCOwnedExtension< U, T > Class Template Reference  125

7.77 lcrtrel::LCOwnedExtension< U, T > Class Template Reference

Simple Extension - pointer to an object of type T where the ownership is taken over by the object holding the extension, i.e.

#include <LCRTRelations.h>

Inheritance diagram for lcrtrel::LCOwnedExtension< U, T >:

Additional Inherited Members

7.77.1 Detailed Description

template<class U, typename T>class lcrtrel::LCOwnedExtension< U, T >

Simple Extension - pointer to an object of type T where the ownership is taken over by the object holding the extension, i.e. it deletes the object when itself is deleted. The class U needs to be the subclass type, e.g. struct MyAttributes : public LCExtension<MyAttributes,Attributes> {} ; The documentation for this class was generated from the following file:

• LCRTRelations.h

7.78 lcrtrel::LCOwnedExtensionList< U, T > Class Template Reference

Extension list holding pointers to objects of type T - ownership of the objects is taken, i.e.

#include <LCRTRelations.h>

Inheritance diagram for lcrtrel::LCOwnedExtensionList< U, T >:

Additional Inherited Members

7.78.1 Detailed Description

template<class U, class T>class lcrtrel::LCOwnedExtensionList< U, T >

Extension list holding pointers to objects of type T - ownership of the objects is taken, i.e.

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

all objects pointed to in the vector are deleted when the object itself is deleted. The documentation for this class was generated from the following file:

• LCRTRelations.h

7.79 lcrtrel::LCOwnedExtensionVector< U, T > Class Template Reference

Extension vector holding pointers to objects of type T - ownership of the objects is taken, i.e.

#include <LCRTRelations.h>

Inheritance diagram for lcrtrel::LCOwnedExtensionVector< U, T >:

Additional Inherited Members

7.79.1 Detailed Description

template<class U, class T>class lcrtrel::LCOwnedExtensionVector< U, T >

Extension vector holding pointers to objects of type T - ownership of the objects is taken, i.e. all objects pointed to in the vector are deleted when the object itself is deleted. The documentation for this class was generated from the following file:

• LCRTRelations.h

7.80 EVENT::LCParameters Class Reference

Simple interface to store generic named parameters of type int, float and string.

#include <pre-generated/EVENT/LCParameters.h>

Inheritance diagram for EVENT::LCParameters:

Public Member Functions

• virtual ∼LCParameters ()

Destructor.

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

• virtual int getIntVal (const std::string &key) const =0

Returns the first integer value for the given key.

• virtual float getFloatVal (const std::string &key) const =0

Returns the first float value for the given key.

• virtual const std::string & getStringVal (const std::string &key) const =0

Returns the first string value for the given key.

• virtual IntVec & getIntVals (const std::string &key, IntVec &values) const =0

Adds all integer values for the given key to values.

• virtual FloatVec & getFloatVals (const std::string &key, FloatVec &values) const =0

Adds all float values for the given key to values.

• virtual StringVec & getStringVals (const std::string &key, StringVec &values) const =0

Adds all float values for the given key to values.

• virtual const StringVec & getIntKeys (StringVec &keys) const =0

Returns a list of all keys of integer parameters.

• virtual const StringVec & getFloatKeys (StringVec &keys) const =0

Returns a list of all keys of float parameters.

• virtual const StringVec & getStringKeys (StringVec &keys) const =0

Returns a list of all keys of string parameters.

• virtual int getNInt (const std::string &key) const =0

The number of integer values stored for this key.

• virtual int getNFloat (const std::string &key) const =0

The number of float values stored for this key.

• virtual int getNString (const std::string &key) const =0

The number of string values stored for this key.

• virtual void setValue (const std::string &key, int value)=0

Set integer value for the given key.

• virtual void setValue (const std::string &key, float value)=0

Set float value for the given key.

• virtual void setValue (const std::string &key, const std::string &value)=0

Set string value for the given key.

• virtual void setValues (const std::string &key, const IntVec &values)=0

Set integer values for the given key.

• virtual void setValues (const std::string &key, const FloatVec &values)=0

Set float values for the given key.

• virtual void setValues (const std::string &key, const StringVec &values)=0

Set string values for the given key.

7.80.1 Detailed Description

Simple interface to store generic named parameters of type int, float and string. It can be used to store (user) meta data that is run, event or collection dependent.

Author gaede

Version

Id: LCParameters.aid,v 1.3 2009-11-20 11:20:20 gaede Exp See Also LCRunHeader.parameters() (p. 145)

LCEvent.parameters() (p. 92) LCCollection.parameters() (p. 86)

7.80.2 Member Function Documentation

7.80.2.1 virtual FloatVec& EVENT::LCParameters::getFloatVals ( const std::string & key, FloatVec & values ) const

       [pure virtual]

Adds all float values for the given key to values. Returns a reference to values for convenience. Referenced by UTIL::LCTOOLS::printParameters(), and SIO::SIOLCParameters::write().

7.80.2.2 virtual IntVec& EVENT::LCParameters::getIntVals ( const std::string & key, IntVec & values ) const [pure

       virtual]

Adds all integer values for the given key to values. Returns a reference to values for convenience. Referenced by UTIL::IndexMap::IndexMap(), UTIL::LCTOOLS::printParameters(), and SIO::SIOLCParameters- ::write().

7.80.2.3 virtual StringVec& EVENT::LCParameters::getStringVals ( const std::string & key, StringVec & values ) const

       [pure virtual]

Adds all float values for the given key to values. Returns a reference to values for convenience. Referenced by UTIL::IndexMap::IndexMap(), SIO::SIOVertexHandler::init(), UTIL::LCTOOLS::printParameters(), and SIO::SIOLCParameters::write(). The documentation for this class was generated from the following file:

• LCParameters.h

7.81 IMPL::LCParametersImpl Class Reference

Implementation of Simple interface to store generic named parameters of type int, float and string.

#include <IMPL/LCParametersImpl.h>

Inheritance diagram for IMPL::LCParametersImpl:

Public Member Functions

• virtual ∼LCParametersImpl ()

Destructor.

• virtual int getIntVal (const std::string &key) const

Returns the first integer value for the given key.

• virtual float getFloatVal (const std::string &key) const

Returns the first float value for the given key.

• virtual const std::string & getStringVal (const std::string &key) const

Returns the first string value for the given key.

• virtual EVENT::IntVec & getIntVals (const std::string &key, EVENT::IntVec &values) const

Adds all integer values for the given key to values.

• virtual EVENT::FloatVec & getFloatVals (const std::string &key, EVENT::FloatVec &values) const

Adds all float values for the given key to values.

• virtual EVENT::StringVec & getStringVals (const std::string &key, EVENT::StringVec &values) const

Adds all float values for the given key to values.

• virtual const EVENT::StringVec & getIntKeys (EVENT::StringVec &keys) const

Returns a list of all keys of integer parameters.

• virtual const EVENT::StringVec & getFloatKeys (EVENT::StringVec &keys) const

Returns a list of all keys of float parameters.

• virtual const EVENT::StringVec & getStringKeys (EVENT::StringVec &keys) const

Returns a list of all keys of string parameters.

• virtual int getNInt (const std::string &key) const

The number of integer values stored for this key.

• virtual int getNFloat (const std::string &key) const

The number of float values stored for this key.

• virtual int getNString (const std::string &key) const

The number of string values stored for this key.

• virtual void setValue (const std::string &key, int value)

Set integer value for the given key.

• virtual void setValue (const std::string &key, float value)

Set float value for the given key.

• virtual void setValue (const std::string &key, const std::string &value)

Set string value for the given key.

• virtual void setValues (const std::string &key, const EVENT::IntVec &values)

Set integer values for the given key.

• virtual void setValues (const std::string &key, const EVENT::FloatVec &values)

Set float values for the given key.

• virtual void setValues (const std::string &key, const EVENT::StringVec &values)

Set string values for the given key.

Protected Attributes

• IntMap _intMap

• FloatMap _floatMap

• StringMap _stringMap

Friends

• class LCRunHeaderImpl

• class LCEventImpl

• class LCCollectionVec

Additional Inherited Members

7.81.1 Detailed Description

Implementation of Simple interface to store generic named parameters of type int, float and string. It can be used to store (user) meta data that is run, event or collection dependent.

Author gaede

Version Jun 23, 2004

See Also LCRunHeader.parameters()

LCEvent.parameters() LCCollection.parameters()

7.81.2 Member Function Documentation

7.81.2.1 FloatVec & IMPL::LCParametersImpl::getFloatVals ( const std::string & key, EVENT::FloatVec & values ) const

       [virtual]

Adds all float values for the given key to values. Returns a reference to values for convenience.

7.81.2.2 IntVec & IMPL::LCParametersImpl::getIntVals ( const std::string & key, EVENT::IntVec & values ) const

       [virtual]

Adds all integer values for the given key to values. Returns a reference to values for convenience.

7.81.2.3 StringVec & IMPL::LCParametersImpl::getStringVals ( const std::string & key, EVENT::StringVec & values )

       const [virtual]

Adds all float values for the given key to values. Returns a reference to values for convenience. The documentation for this class was generated from the following files:

• LCParametersImpl.h

• LCParametersImpl.cc

7.82 IO::LCReader Class Reference

Interface for reading data from LCIO.

#include <pre-generated/IO/LCReader.h>

Inheritance diagram for IO::LCReader:

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Public Member Functions

• virtual ∼LCReader ()

Destructor.

• virtual void open (const std::string &filename)=0 throw (IOException, std::exception )

Opens a file for reading (read-only).

• virtual void open (const std::vector< std::string > &filenames)=0 throw (IOException, std::exception )

Opens a list of files for reading (read-only).

• virtual EVENT::LCRunHeader ∗ readNextRunHeader ()=0 throw (IOException, std::exception )

Reads the next run header from the file.

• virtual EVENT::LCRunHeader ∗ readNextRunHeader (int accessMode)=0 throw (IOException, std-

::exception )

Same as readNextRunHeader() (p. 134) but allows to set the access mode LCIO::READ_ONLY (default) or LCIO::- Update.

• virtual EVENT::LCEvent ∗ readNextEvent ()=0 throw (IOException, std::exception )

Reads the next event from the file.

• virtual EVENT::LCEvent ∗ readNextEvent (int accessMode)=0 throw (IOException, std::exception )

Same as readNextEvent() (p. 134) but allows to set the access mode LCIO::READ_ONLY (default) or LCIO::Update.

• virtual int getNumberOfEvents ()=0 throw (IOException, std::exception )

Return the number of events in the file - the file has to be open.

• virtual int getNumberOfRuns ()=0 throw (IOException, std::exception )

Return the number of runs (run headers) in the file - the file has to be open.

• virtual void getRuns (EVENT::IntVec &runs)=0

Return the run numbers of the runs (run headers) in the file - the file has to be open.

• virtual void getEvents (EVENT::IntVec &events)=0

Return the run and event numbers of the events in the file - the file has to be open.

• virtual void setReadCollectionNames (const std::vector< std::string > &colnames)=0

Limit the collection names that are going to be read to the subset given in the vector - all other collection will be ignored.

• virtual void skipNEvents (int n)=0 throw (IOException, std::exception )

Skips the next n events from the current position.

• virtual EVENT::LCRunHeader ∗ readRunHeader (int runNumber)=0 throw (IOException, std::exception )

Reads the specified runHeader from file.

• virtual EVENT::LCRunHeader ∗ readRunHeader (int runNumber, int accessMode)=0 throw (IOException,

std::exception )

Same as LCEvent∗ readRunHeader(int runNumber) (p. 135) allowing to set the access mode LCIO::READ_ONLY (default) or LCIO::Update.

• virtual EVENT::LCEvent ∗ readEvent (int runNumber, int evtNumber)=0 throw (IOException, std::exception

)

Reads the specified event from file.

• virtual EVENT::LCEvent ∗ readEvent (int runNumber, int evtNumber, int accessMode)=0 throw (IO-

Exception, std::exception )

Same as LCEvent∗ readEvent(int runNumber, int evtNumber) (p. 134) allowing to set the access mode LCIO::RE-AD_ONLY (default) or LCIO::Update.

• virtual void close ()=0 throw (IOException, std::exception )

Closes the output file/stream etc.

• virtual void registerLCEventListener (LCEventListener ∗ls)=0

Registers a listener for reading LCEvents from a stream.

• virtual void removeLCEventListener (LCEventListener ∗ls)=0

Remove a listener for reading LCEvents from a stream.

• virtual void registerLCRunListener (LCRunListener ∗ls)=0

Registers a listener for reading LCEventsLCRunHeaders from a stream.

• virtual void removeLCRunListener (LCRunListener ∗ls)=0

Remove a listener for reading LCRunHeaders from a stream.

• virtual void readStream ()=0 throw (IOException, std::exception )

Reads the input stream and notifies registered listeners according to the object type found in the stream.

• virtual void readStream (int maxRecord)=0 throw (IOException, std::exception )

Reads maxRecord from the input stream and notifies registered listeners according to the object type found in the stream.

Static Public Attributes

• static const int directAccess = 0x00000001 << 0

Configuration flags for the LCReader (p. 130) instance - to be used with LCFactory::createLCReader().

7.82.1 Detailed Description

Interface for reading data from LCIO. Use LCFactory to instantiate a corresponding LCWriter (p. 166) object for the output format at hand (SIO (p. 30) only, so far).

See Also LCFactory

Author gaede

Version Mar 4, 2003

7.82.2 Member Function Documentation

7.82.2.1 virtual void IO::LCReader::close ( ) throw IOException, std::exception) [pure virtual]

Closes the output file/stream etc. Exceptions

IOException (p. 79)

Implemented in SIO::SIOReader (p. 236).

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.82.2.2 virtual void IO::LCReader::getEvents ( EVENT::IntVec & events ) [pure virtual]

Return the run and event numbers of the events in the file - the file has to be open. In case several input files are specified in the open() (p. 133) method - the run and event numbers of the events in the file that is currently open are returned. The size of the vector events will be twice the number of events, where i-th run number will be in events[2∗i] and the i-th event number in events[2∗i+]. Implemented in SIO::SIOReader (p. 236).

7.82.2.3 virtual int IO::LCReader::getNumberOfEvents ( ) throw IOException, std::exception) [pure virtual]

Return the number of events in the file - the file has to be open. In case several input files are specified in the open() (p. 133) method - the number of events in the file that is currently open is returned. Implemented in SIO::SIOReader (p. 236).

7.82.2.4 virtual int IO::LCReader::getNumberOfRuns ( ) throw IOException, std::exception) [pure virtual]

Return the number of runs (run headers) in the file - the file has to be open. In case several input files are specified in the open() (p. 133) method - the number of runs (run headers) in the file that is currently open is returned. Implemented in SIO::SIOReader (p. 236).

7.82.2.5 virtual void IO::LCReader::getRuns ( EVENT::IntVec & runs ) [pure virtual]

Return the run numbers of the runs (run headers) in the file - the file has to be open. In case several input files are specified in the open() (p. 133) method - the run numbers of the runs (run headers) in the file that is currently open is returned. Implemented in SIO::SIOReader (p. 237).

7.82.2.6 virtual void IO::LCReader::open ( const std::string & filename ) throw IOException, std::exception) [pure

       virtual]

Opens a file for reading (read-only). Exceptions

IOException (p. 79)

Implemented in SIO::SIOReader (p. 237).

7.82.2.7 virtual void IO::LCReader::open ( const std::vector< std::string > & filenames ) throw IOException, std::exception)

       [pure virtual]

Opens a list of files for reading (read-only). All subsequent read operations will operate on the list, i.e. if an EOF is encountered the next file in the list will be opened and read transparently to the user.

Exceptions IOException (p. 79)

Implemented in SIO::SIOReader (p. 237).

7.82.2.8 virtual EVENT::LCEvent∗ IO::LCReader::readEvent ( int runNumber, int evtNumber ) throw IOException, std::exception) [pure virtual]

Reads the specified event from file. Returns NULL if the specified event hasn’t been found in the file. Exceptions

IOException (p. 79)

Implemented in SIO::SIOReader (p. 237).

7.82.2.9 virtual EVENT::LCEvent∗ IO::LCReader::readEvent ( int runNumber, int evtNumber, int accessMode ) throw IOException, std::exception) [pure virtual]

Same as LCEvent∗ readEvent(int runNumber, int evtNumber) (p. 134) allowing to set the access mode LCIO::R-EAD_ONLY (default) or LCIO::Update. Exceptions

IOException (p. 79)

Implemented in SIO::SIOReader (p. 238).

7.82.2.10 virtual EVENT::LCEvent∗ IO::LCReader::readNextEvent ( ) throw IOException, std::exception) [pure

        virtual]

Reads the next event from the file. Returns NULL if ’EOF’ read. Exceptions

IOException (p. 79)

Implemented in SIO::SIOReader (p. 239).

7.82.2.11 virtual EVENT::LCEvent∗ IO::LCReader::readNextEvent ( int accessMode ) throw IOException, std::exception)

        [pure virtual]

Same as readNextEvent() (p. 134) but allows to set the access mode LCIO::READ_ONLY (default) or LCIO::- Update. Exceptions

IOException (p. 79)

Implemented in SIO::SIOReader (p. 239).

7.82.2.12 virtual EVENT::LCRunHeader∗ IO::LCReader::readNextRunHeader ( ) throw IOException, std::exception)

        [pure virtual]

Reads the next run header from the file. Returns NULL if ’EOF’ read.

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Exceptions IOException (p. 79)

Implemented in SIO::SIOReader (p. 239).

7.82.2.13 virtual EVENT::LCRunHeader∗ IO::LCReader::readNextRunHeader ( int accessMode ) throw IOException, std::exception) [pure virtual]

Same as readNextRunHeader() (p. 134) but allows to set the access mode LCIO::READ_ONLY (default) or LCIO- ::Update. Exceptions

IOException (p. 79)

Implemented in SIO::SIOReader (p. 239).

7.82.2.14 virtual EVENT::LCRunHeader∗ IO::LCReader::readRunHeader ( int runNumber ) throw IOException, std::exception) [pure virtual]

Reads the specified runHeader from file. Returns NULL if the specified runHeader hasn’t been found in the file. Exceptions

IOException (p. 79)

Implemented in SIO::SIOReader (p. 239).

7.82.2.15 virtual EVENT::LCRunHeader∗ IO::LCReader::readRunHeader ( int runNumber, int accessMode ) throw IOException, std::exception) [pure virtual]

Same as LCEvent∗ readRunHeader(int runNumber) (p. 135) allowing to set the access mode LCIO::READ_ONLY (default) or LCIO::Update. Exceptions

IOException (p. 79)

Implemented in SIO::SIOReader (p. 240).

7.82.2.16 virtual void IO::LCReader::readStream ( ) throw IOException, std::exception) [pure virtual]

Reads the input stream and notifies registered listeners according to the object type found in the stream. Exceptions

IOException (p. 79)

Implemented in SIO::SIOReader (p. 240).

7.82.2.17 virtual void IO::LCReader::readStream ( int maxRecord ) throw IOException, std::exception) [pure

        virtual]

Reads maxRecord from the input stream and notifies registered listeners according to the object type found in the stream. An exception is thrown if less than maxRecord where read from the file.

Exceptions IOException (p. 79)

Implemented in SIO::SIOReader (p. 240).

7.82.2.18 virtual void IO::LCReader::setReadCollectionNames ( const std::vector< std::string > & colnames ) [pure

        virtual]

Limit the collection names that are going to be read to the subset given in the vector - all other collection will be ignored. This might improve the reading performance considerably in cases where only a small subset of the collections in the event is needed. Implemented in SIO::SIOReader (p. 240).

7.82.2.19 virtual void IO::LCReader::skipNEvents ( int n ) throw IOException, std::exception) [pure virtual]

Skips the next n events from the current position. Exceptions

IOException (p. 79)

Implemented in SIO::SIOReader (p. 240).

7.82.3 Member Data Documentation

7.82.3.1 const int IO::LCReader::directAccess = 0x00000001 << 0 [static]

Configuration flags for the LCReader (p. 130) instance - to be used with LCFactory::createLCReader(). The documentation for this class was generated from the following file:

• LCReader.h

7.83 EVENT::LCRelation Class Reference

A single weighted relationship between two LCObjects.

#include <pre-generated/EVENT/LCRelation.h>

Inheritance diagram for EVENT::LCRelation:

Public Types

• typedef LCRelation lcobject_type

Useful typedef for template programming with LCIO (p. 115).

Public Member Functions

• virtual ∼LCRelation ()

Destructor.

• virtual LCObject ∗ getFrom () const =0

The ’from’ object of the given relation.

• virtual LCObject ∗ getTo () const =0

The ’to’ object of the given relation.

• virtual float getWeight () const =0

The weight of the given relation - only if collection flag bit LCIO::LCREL_WEIGHTED is set.

Additional Inherited Members

7.83.1 Detailed Description

A single weighted relationship between two LCObjects. Stored in an LCCollection (p. 83) like any other LCObject (p. 123). Use inplementations of LCRelationNavigator to navigate the relation efficiently. the types of the objects is decoded in the collection paramters ’RelationFromType’ and ’RelationToType’. In order to store weights != 1. with the relation set the collection flag bit LCIO::LCREL_WEI-GHTED, this is done automatically in UTIL::LCRelationNavigator (p. 139) if a weight != 1. exists.

Author gaede

Version

Id: LCRelation.aid,v 1.7 2006-09-14 10:16:11 gaede Exp

See Also LCRelationNavigator

The documentation for this class was generated from the following file:

• LCRelation.h

7.84 IMPL::LCRelationImpl Class Reference

Implementation of the LCRelation.

#include <IMPL/LCRelationImpl.h>

Inheritance diagram for IMPL::LCRelationImpl:

Public Member Functions

• LCRelationImpl (EVENT::LCObject ∗from, EVENT::LCObject ∗to, float weight=1.0f)

• int id () const

Returns an object id for internal (debugging) use in LCIO.

• EVENT::LCObject ∗ getFrom () const

The ’from’ object of the given relation.

• EVENT::LCObject ∗ getTo () const

The ’to’ object of the given relation.

• float getWeight () const

The weight of the given relation - only if collection flag bit LCIO::LCREL_WEIGHTED is set.

• void setFrom (EVENT::LCObject ∗from)

• void setTo (EVENT::LCObject ∗to)

• void setWeight (float weight)

Protected Attributes

• EVENT::LCObject ∗ _from

• EVENT::LCObject ∗ _to

• float _weight

Additional Inherited Members

7.84.1 Detailed Description

Implementation of the LCRelation.

Author gaede

Version

id:

The documentation for this class was generated from the following file:

• LCRelationImpl.h

7.85 IOIMPL::LCRelationIOImpl Class Reference

Adding stuff needed for io (friend declarations, etc.)

#include <IOIMPL/LCRelationIOImpl.h>

Inheritance diagram for IOIMPL::LCRelationIOImpl:

Public Member Functions

• LCRelationIOImpl (EVENT::LCObject ∗from, EVENT::LCObject ∗to, float weight=1.0f)

Friends

• class SIO::SIOLCRelationHandler

Additional Inherited Members

7.85.1 Detailed Description

Adding stuff needed for io (friend declarations, etc.)

Author gaede

Version

Id: LCRelationIOImpl.h (p. ?? ),v 1.4 2005-04-15 08:37:40 gaede Exp

The documentation for this class was generated from the following file:

• LCRelationIOImpl.h

7.86 UTIL::LCRelationNavigator Class Reference

The LCRelationNavigator (p. 139) makes repeated lookup of relations more conveneient and efficient.

#include <UTIL/LCRelationNavigator.h>

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Public Member Functions

• LCRelationNavigator (const std::string &fromType, const std::string &toType)

Default constructor.

• LCRelationNavigator (const EVENT::LCCollection ∗col)

Create the navigator object from an existing collection of relations.

• virtual ∼LCRelationNavigator ()

Destructor.

• virtual const std::string & getFromType () const

The type of the ’from’ objects in this relation.

• virtual const std::string & getToType () const

The type of the ’to’ objects in this relation.

• virtual const EVENT::LCObjectVec & getRelatedToObjects (EVENT::LCObject ∗from) const

All objects that the given from-object is related to.

• virtual const EVENT::LCObjectVec & getRelatedFromObjects (EVENT::LCObject ∗to) const

All from-objects related to the given object ( the inverse relationship).

• virtual const EVENT::FloatVec & getRelatedToWeights (EVENT::LCObject ∗from) const

The weights of the relations returned by a call to getRelatedToObjects(from).

• virtual const EVENT::FloatVec & getRelatedFromWeights (EVENT::LCObject ∗to) const

The weights of the relations returned by a call to getRelatedFromObjects(to).

• virtual void addRelation (EVENT::LCObject ∗from, EVENT::LCObject ∗to, float weight=1.0)

Adds a relation.

• virtual void removeRelation (EVENT::LCObject ∗from, EVENT::LCObject ∗to)

Remove a given relation.

• virtual EVENT::LCCollection ∗ createLCCollection ()

Remove a given relation.

Protected Member Functions

• virtual void initialize (const EVENT::LCCollection ∗col)

• void removeRelation (EVENT::LCObject ∗from, EVENT::LCObject ∗to, RelMap &map)

• void addRelation (EVENT::LCObject ∗from, EVENT::LCObject ∗to, float weight, RelMap &map)

Protected Attributes

• RelMap _map

• RelMap _rMap

• std::string _from

• std::string _to

7.86.1 Detailed Description

The LCRelationNavigator (p. 139) makes repeated lookup of relations more conveneient and efficient.

The relations are treated symmetrical, i.e. lookup of relations is equally efficient and fast for either direction (from-to and to-from) at the price of a slower (by a factor of ∼2) modification speed.

Author gaede Version

Id: LCRelationNavigator.h (p. ?? ),v 1.2 2004-09-06 14:35:51 gaede Exp

7.86.2 Member Function Documentation

7.86.2.1 void UTIL::LCRelationNavigator::addRelation ( EVENT::LCObject ∗ from, EVENT::LCObject ∗ to, float weight =

       1.0 ) [virtual]

Adds a relation.

If there is already an existing relation between the two given objects the weight (or default weight 1.0) is added to that relationship’s weight.

References addRelation().

Referenced by addRelation().

7.86.2.2 EVENT::LCCollection ∗ UTIL::LCRelationNavigator::createLCCollection ( ) [virtual]

Remove a given relation.

To reduce the weight of the relationship, call addRelation( from, to, weight ) with weight<0.

References IMPL::LCCollectionVec::addElement(), IMPL::LCFlagImpl::getFlag(), getFromType(), getToType(), IMP-L::LCCollectionVec::parameters(), IMPL::LCFlagImpl::setBit(), IMPL::LCCollectionVec::setFlag(), and EVENT::LC-Parameters::setValue().

7.86.2.3 const EVENT::LCObjectVec & UTIL::LCRelationNavigator::getRelatedFromObjects ( EVENT::LCObject ∗ to )

       const [virtual]

All from-objects related to the given object ( the inverse relationship).

LCObjects are of type getFromType() (p. 140).

7.86.2.4 const EVENT::FloatVec & UTIL::LCRelationNavigator::getRelatedFromWeights ( EVENT::LCObject ∗ to ) const

       [virtual]

The weights of the relations returned by a call to getRelatedFromObjects(to).

See Also getRelatedFromObjects (p. 141)

7.86.2.5 const EVENT::LCObjectVec & UTIL::LCRelationNavigator::getRelatedToObjects ( EVENT::LCObject ∗ from )

       const [virtual]

All objects that the given from-object is related to.

LCObjects are of type getToType() (p. 140).

7.86.2.6 const EVENT::FloatVec & UTIL::LCRelationNavigator::getRelatedToWeights ( EVENT::LCObject ∗ from ) const

       [virtual]

The weights of the relations returned by a call to getRelatedToObjects(from).

See Also getRelatedToObjects (p. 141)

The documentation for this class was generated from the following files:

• LCRelationNavigator.h

• LCRelationNavigator.cc

7.87 lcrtrel::LCRTRelations Class Reference

Base class that provides run time (user) extensions and relation between objects.

#include <LCRTRelations.h>

Inheritance diagram for lcrtrel::LCRTRelations:

Public Member Functions

• template<class V >

V::ext_type ext ()

Provides access to an extension object - the type and ownership is defined by the class V which should be a subtype of LCExtension (p. 99), LCOwnedExtension (p. 125), LCExtensionVector (p. 100), LCExtensionList (p. 100),...

• template<class V >

const V::ext_type ext () const

• template<class V >

V::rel_type rel ()

Provides read access to relations - the object types and their connectivity are defined by the class V which has to be a subtype of either LC1To1Relation (p. 80), LC1ToNRelation (p. 80) or LCNToNRelation (p. 122).

Protected Member Functions

• template<class V >

V::ptr & ptr () const

Returns the reference to the pointer to the extension/relation object.

Friends

• template<class R >

void set_relation (typename R::from::obj_ptr f, typename R::to::obj_ptr t)

Set the 1-to-1 relation between two objects - prexisting inconsistent relations involving the two objects are deleted to enforce a consistent set of from-to relations.

• template<class R >

void unset_relation (typename R::from::obj_ptr f)

Unset the 1-to-1 relation from f.

• template<class R >

void add_relation (typename R::from::obj_ptr f, typename R::to::obj_ptr t)

Add a link from f to t to an N-to-N relation ship.

• template<class R >

void remove_relation (typename R::from::obj_ptr f, typename R::to::obj_ptr t)

Remove the link from from f to t from the N-to-N relation ship.

• template<class R >

void remove_relations (typename R::from::obj_ptr f)

Removes all relations from the given object.

• template<class R >

void merge_relations (typename R::from::obj_ptr f1, typename R::from::obj_ptr f2)

Merge the relations from f2 to f1 - after this call f1 will hold all the relations and f2 will be empty.

7.87.1 Detailed Description

Base class that provides run time (user) extensions and relation between objects.

Every subclass object of LCRTRelations (p. 142) (and thus LCObbject) will automatically have the following functionality:

• extension of the object with arbitrary (even non-LCObject) classes

• extension of single objects or vectors, lists of objects

• optionally ownership is taken for extension objects (memory management)

• relations to other subclasses of LCRTRelations (p. 142):

– one to one

– one to many

– many to many

The described functionality is provided through the two templated member functions:

ext<class V>() (p. 142)

rel<class V>() (p. 142) the class V is a user defined so called traits class, that uniquely tags the extension/relationship and defines the types of the objects involved. For extensions users have to subclass one of the following classes: LCExtension (p. 99), LCOwnedExtension (p. 125), LCIntExtension (p. 113), LCFloatExtension (p. 105), LCExtensionVector (p. 100), LCExtensionList (p. 100), LCOwnedExtensionVector (p. 126), LCOwned-ExtensionList (p. 125). For example the following defines a user extension of a vector of strings that are owned by the object (i.e. deleted when the object is deleted): struct ParticleIDs (p. 185) : LCOwnedExtensionVector<ParticleIDs, std::string> {};

(note: the first template parameter has to be the class itself !) This extension can then be used anywhere in the following code for all LCObjects, e.g.: MCParticle∗ mcp = dynamic_cast<MCParticle∗>( mcpcol->getElementAt(i) ) ; mcp->ext<ParticleIDs>() (p. 142)->push_back( new std::string("charged") ) ; mcp->ext<ParticleIDs>() (p. 142)->push_back( new std::string("hadron") ) ; mcp->ext<ParticleIDs>() (p. 142)->push_back( new std::string("pion") ) ; and be read out again: ParticleIDs::ext_type pidv = mcp->ext<ParticleIDs>() (p. 142) ; for( ParticleIDs::const_iterator ipid = pidv->begin() ; ipid != pidv->end(); ++ipid){

std::cout << ∗∗ipid << ", " ; } Similarily the following defines a one to many relationship between Tracks and Clusters: struct TrkCluLink (p. 310) : LC1ToNRelation<TrkCluLink,Track,Cluster> {} ; Relations are allways biderectional, i.e. there is a from and a to side. They can then be set and modified with the following functions: set_relation() (p. 145), unset_relation() (p. 143), add_relation() (p. 143), remove_relation() (p. 143), remove_relations() (p. 143), merge_relations() (p. 143). For example: Track∗ trk = dynamic_cast<Track∗> ( trkcol->getElementAt(j) ) ; //... Cluster∗ clu = dynamic_cast<Cluster∗> ( clucol->getElementAt(k) ) ; add_relation<TrkCluLink>( trk ,clu ); The many side can then be read out with a const_iterator and the one side with a normal pointer: Track∗ trk = clu->rel<TrkCluLink::from>() (p. 142) ; //... TrkCluLink::to::rel_type clulist = trk->rel<TrkCluLink::to>() (p. 142) ; for( TrkCluLink::to::const_iterator iclu = clulist->begin() ; iclu != clulist->end() ; ++iclu ){ Cluster∗ clu = ∗iclu ; // iterator is of type pointer to container element std::cout << " assigned cluster with E = " << clu->getEnergy() << std::endl ; } More examples can be found in $LCIO/src/cpp/src/EXAMPLES/lcrtrelations.cc.

7.87.2 Friends And Related Function Documentation

7.87.2.1 template<class R > void set_relation ( typename R::from::obj_ptr f, typename R::to::obj_ptr t ) [friend]

Set the 1-to-1 relation between two objects - prexisting inconsistent relations involving the two objects are deleted to enforce a consistent set of from-to relations. The documentation for this class was generated from the following file:

• LCRTRelations.h

7.88 EVENT::LCRunHeader Class Reference

Interface for the run header.

#include <pre-generated/EVENT/LCRunHeader.h>

Inheritance diagram for EVENT::LCRunHeader:

Public Types

• typedef LCRunHeader lcobject_type

Useful typedef for template programming with LCIO (p. 115).

Public Member Functions

• virtual ∼LCRunHeader ()

Destructor.

• virtual int getRunNumber () const =0

Returns the run number.

• virtual const std::string & getDetectorName () const =0

Returns the name of the detector setup used in the simulation.

• virtual const std::string & getDescription () const =0

Description of the simulation, physics channels etc.

• virtual const std::vector

< std::string > ∗ getActiveSubdetectors () const =0

Returns the names of the active subdetectors used in the simulation.

• virtual const LCParameters & getParameters () const =0

Parameters defined for this run.

• virtual LCParameters & parameters ()=0

Parameters defined for this run.

Additional Inherited Members

7.88.1 Detailed Description

Interface for the run header.

Author gaede

Version

Id: LCRunHeader.aid,v 1.8 2006-03-24 13:25:52 gaede Exp

The documentation for this class was generated from the following file:

• LCRunHeader.h

7.89 IMPL::LCRunHeaderImpl Class Reference

Implementation of LCRunHeader.

#include <IMPL/LCRunHeaderImpl.h>

Inheritance diagram for IMPL::LCRunHeaderImpl:

Public Member Functions

• virtual ∼LCRunHeaderImpl ()

Destructor.

• virtual int id () const

Returns an object id for internal (debugging) use in LCIO.

• virtual int getRunNumber () const

Returns the run number.

• virtual const std::string & getDetectorName () const

Returns the name of the detector setup used in the simulation.

• virtual const std::string & getDescription () const

Description of the simulation conditions, e.g.

• virtual const std::vector

< std::string > ∗ getActiveSubdetectors () const

Returns the names of the active subdetectors used in the simulation.

• virtual const EVENT::LCParameters & getParameters () const

Parameters defined for this run.

• virtual EVENT::LCParameters & parameters ()

Parameters defined for this run.

• void setRunNumber (int rn)

Sets the run number.

• void setDetectorName (const std::string &dn)

Sets the detector name.

• void setDescription (const std::string &dsc)

Sets the description of the run.

• void addActiveSubdetector (const std::string &adn)

Adds an active sub detector name.

Protected Member Functions

• virtual void setReadOnly (bool readOnly)

Protected Attributes

• int _runNumber

• std::string _detectorName

• std::string _description

• std::vector< std::string > _activeSubdetectors

• LCParametersImpl _params

Additional Inherited Members

7.89.1 Detailed Description

Implementation of LCRunHeader.

Author gaede

Version Mar 11, 2003

7.89.2 Member Function Documentation

7.89.2.1 const std::string & IMPL::LCRunHeaderImpl::getDescription ( ) const [virtual]

Description of the simulation conditions, e.g.

physics channels.

Implements EVENT::LCRunHeader (p. 145).

The documentation for this class was generated from the following files:

• LCRunHeaderImpl.h

• LCRunHeaderImpl.cc

7.90 IOIMPL::LCRunHeaderIOImpl Class Reference

Adding stuff needed for io (friend declarations, etc.)

#include <IOIMPL/LCRunHeaderIOImpl.h>

Inheritance diagram for IOIMPL::LCRunHeaderIOImpl:

Friends

• class SIO::SIOReader

Additional Inherited Members

7.90.1 Detailed Description

Adding stuff needed for io (friend declarations, etc.)

Author gaede

Version Sep 9, 2003

The documentation for this class was generated from the following file:

• LCRunHeaderIOImpl.h

7.91 IO::LCRunListener Class Reference

Listener for the occurence of LCRunHeader when reading a stream.

#include <pre-generated/IO/LCRunListener.h>

Public Member Functions

• virtual ∼LCRunListener ()

Destructor.

• virtual void processRunHeader (EVENT::LCRunHeader ∗rh)=0

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Call back for processing an LCRunHeader.

• virtual void modifyRunHeader (EVENT::LCRunHeader ∗rh)=0

Call back for modifying an existing LCRunHeader.

7.91.1 Detailed Description

Listener for the occurence of LCRunHeader when reading a stream.

User classes should implement this interface for their analysis code and then register the listener with LCReader (p. 130).

Author gaede

Version Mar 11, 2003

See Also LCRunHeader LCReader (p. 130)

7.91.2 Member Function Documentation

7.91.2.1 virtual void IO::LCRunListener::processRunHeader ( EVENT::LCRunHeader ∗ rh ) [pure virtual]

Call back for processing an LCRunHeader.

Access is ’read only’.

The documentation for this class was generated from the following file:

• LCRunListener.h

7.92 SIO::LCSIO Class Reference

Collection of static helper functions for reading and writing data with SIO (p. 30).

#include <SIO/LCSIO.h>

Static Public Member Functions

• static SIORecords & records ()

• static void checkVersion (int versionID)

Checks the version of the file - oldefile (version < v00-08) are no longer supported.

• static unsigned int read (SIO_stream ∗stream, char ∗∗c, int ∗len=0)

Read a string from the stream into a dummy buffer.

• static unsigned int write (SIO_stream ∗stream, int i)

This version checks the versionId to be able to read ’old’ files with trailing ’\00’ (version <= 00-02).

• static unsigned int write (SIO_stream ∗stream, unsigned int i)

Write an unsigned int to the stream.

• static unsigned int write (SIO_stream ∗stream, EVENT::long64 i)

Write a vector size to the stream (needed for AMD64 architectures)

• static unsigned int write (SIO_stream ∗stream, float f)

Write a float to the stream.

• static unsigned int write (SIO_stream ∗stream, const std::string &s)

Write a string to the stream.

• static std::string getValidSIOName (const std::string &aName)

Creates a valid SIO (p. 30) name (basically equivalent to a valid C++ name) by replacing every [.

• static void seekStream (SIO_stream ∗stream, EVENT::long64 pos)

Seek stream to the given absolute position - if pos<0 from end of file.

Static Public Attributes

• static const char ∗ FILE_EXTENSION =".slcio"

• static const bool COMPRESSION = true

the compression mode for SIO (p. 30)

7.92.1 Detailed Description

Collection of static helper functions for reading and writing data with SIO (p. 30).

Could go to the SIO_functions class. Also defines some constant names for lcio-SIO.

Author gaede

Version Mar 6, 2003

7.92.2 Member Function Documentation

7.92.2.1 std::string SIO::LCSIO::getValidSIOName ( const std::string & aName ) [static]

Creates a valid SIO (p. 30) name (basically equivalent to a valid C++ name) by replacing every [.

], [] and [/] with underscores and removing all nonalphanumeric characters and adding an ’A’ to the beginning, if necessary.

Referenced by SIO::SIOReader::open(), and SIO::SIOWriter::open().

7.92.2.2 unsigned int SIO::LCSIO::read ( SIO_stream ∗ stream, char ∗∗ c, int ∗ len = 0 ) [static]

Read a string from the stream into a dummy buffer.

Warning the same buffer is used for each call. So the return value needs to be copied to its final memory destination. len holds the length of the string (can contain ’\0’).

7.92.2.3 void SIO::LCSIO::seekStream ( SIO_stream ∗ stream, EVENT::long64 pos ) [static]

Seek stream to the given absolute position - if pos<0 from end of file.

Exceptions IOException

Referenced by SIO::LCIORandomAccessMgr::createEventMap(), SIO::SIOWriter::open(), SIO::LCIORandom-AccessMgr::readLCIOIndexAt(), SIO::LCIORandomAccessMgr::readLCIORandomAccessAt(), and SIO::LCIO-RandomAccessMgr::recreateEventMap().

7.92.2.4 unsigned int SIO::LCSIO::write ( SIO_stream ∗ stream, int i ) [static]

This version checks the versionId to be able to read ’old’ files with trailing ’\00’ (version <= 00-02). Remove this method after a reasonable transition period...Write an int to the stream.

7.92.2.5 unsigned int SIO::LCSIO::write ( SIO_stream ∗ stream, EVENT::long64 i ) [static]

Write a vector size to the stream (needed for AMD64 architectures) Write a 64bit integer to the stream

7.92.2.6 unsigned int SIO::LCSIO::write ( SIO_stream ∗ stream, const std::string & s ) [static]

Write a string to the stream. Writes the length of the string first and adds ’\0’ to the end. The documentation for this class was generated from the following files:

• LCSIO.h

• LCSIO.cc

7.93 UTIL::LCSplitWriter Class Reference

LCWriter wrapper that automatically splits files if a given number of bytes is exceeded.

#include <UTIL/LCSplitWriter.h>

Inheritance diagram for UTIL::LCSplitWriter:

Public Member Functions

• LCSplitWriter (IO::LCWriter ∗wrt, EVENT::long64 maxBytes)

The constructor.

• virtual ∼LCSplitWriter ()

Destructor.

• virtual void open (const std::string &filename) throw (IO::IOException, std::exception )

Opens a file for writing where the filename has to include the extension but not the number, e.g.

• virtual void open (const std::string &filename, int writeMode) throw (IO::IOException, std::exception )

Not implemented - will throw an Exception if called.

• virtual void writeRunHeader (const EVENT::LCRunHeader ∗hdr) throw (IO::IOException, std::exception )

Writes the given run header to file.

• virtual void setCompressionLevel (int level)

Set CompressionLevel.

• virtual void writeEvent (const EVENT::LCEvent ∗evt) throw (IO::IOException, std::exception )

Writes the given file to file.

• virtual void close () throw (IO::IOException, std::exception )

Closes the output file/stream.

• virtual void flush () throw (IO::IOException, std::exception )

Flushes the output file/stream.

• EVENT::long64 fileSize ()

Return the current file size in bytes.

Protected Member Functions

• EVENT::long64 file_size (const char ∗fname)

Helper function that returns the file size in bytes - o if the file does not exist.

• const std::string & getFilename ()

Get the complete filename w/o extensiomn,e.g.

• void setBaseFilename (const std::string &filename)

Helper function to determine base file name, i.e.

• std::string getCountingString (unsigned count)

Returns the string representation of the file number, e.g.

Protected Attributes

• IO::LCWriter ∗ _wrt

• EVENT::long64 _maxBytes

• std::string _baseFilename

• std::string _filename

• std::string _extension

• unsigned _count

• unsigned _lastCount

7.93.1 Detailed Description

LCWriter wrapper that automatically splits files if a given number of bytes is exceeded. File names are extended by a file number of the form .000 - .999 ...

Author gaede

Version

Id: LCSplitWriter.h (p. ?? ),v 1.2 2007-11-09 21:00:55 gaede Exp

7.93 UTIL::LCSplitWriter Class Reference  153

7.93.2 Constructor & Destructor Documentation

7.93.2.1 UTIL::LCSplitWriter::LCSplitWriter ( IO::LCWriter ∗ wrt, EVENT::long64 maxBytes ) [inline] The constructor. Takes a pointer to a valid LCWriter instance and the file size in bytes that causes a new file to be opened if exceeded.example: // replace LCWriter∗ lcWrt = LCFactory::getInstance()->createLCWriter() ; // with LCSplitWriter∗ lcWrt = new LCSplitWriter (p. 151)( LCFactory::getInstance()->createLCWriter() , 2040109465 ) ; to automatically split files after 1.9 GByte file size has been exceeded.

7.93.3 Member Function Documentation

7.93.3.1 void UTIL::LCSplitWriter::close ( ) throw IO::IOException, std::exception) [virtual] Closes the output file/stream. Exceptions

IO::IOException (p. 79) Implements IO::LCWriter (p. 167).

7.93.3.2 void UTIL::LCSplitWriter::flush ( ) throw IO::IOException, std::exception) [virtual] Flushes the output file/stream. Exceptions

IO::IOException (p. 79) Implements IO::LCWriter (p. 167).

7.93.3.3 std::string UTIL::LCSplitWriter::getCountingString ( unsigned count ) [protected] Returns the string representation of the file number, e.g. "007".

7.93.3.4 const std::string & UTIL::LCSplitWriter::getFilename ( ) [protected] Get the complete filename w/o extensiomn,e.g. MyFilename.007

7.93.3.5 void UTIL::LCSplitWriter::open ( const std::string & filename ) throw IO::IOException, std::exception)

       [virtual]

Opens a file for writing where the filename has to include the extension but not the number, e.g. myfile.slcio. Note that this is different from the LCWriter specification. \underline{Exceptions}

154  Class Documentation

IO::IOException (p. 79) Implements IO::LCWriter (p. 167). 7.93.3.6 void UTIL::LCSplitWriter::open ( const std::string & filename, int writeMode ) throw IO::IOException,

std::exception) [virtual] Not implemented - will throw an Exception if called. Overwriting of or appending to split files is not straight forward. Pleas use the default write mode and remove exisiting files. Exceptions

IO::IOException (p. 79) Implements IO::LCWriter (p. 168). 7.93.3.7 void UTIL::LCSplitWriter::setBaseFilename ( const std::string & filename ) [protected] Helper function to determine base file name, i.e. w/o extension. 7.93.3.8 virtual void UTIL::LCSplitWriter::setCompressionLevel ( int level ) [inline], [virtual] Set CompressionLevel.

See Also LCWriter::setCompressionLevel() Implements IO::LCWriter (p. 168). References IO::LCWriter::setCompressionLevel(). 7.93.3.9 void UTIL::LCSplitWriter::writeEvent ( const EVENT::LCEvent ∗ evt ) throw IO::IOException, std::exception)

       [virtual]

Writes the given file to file. Opens a new file if the given file size is already exceeded before the execution of the write access. Exceptions

IO::IOException (p. 79) Implements IO::LCWriter (p. 168). 7.93.3.10 void UTIL::LCSplitWriter::writeRunHeader ( const EVENT::LCRunHeader ∗ hdr ) throw IO::IOException,

std::exception) [virtual]

Writes the given run header to file. Opens a new file if the given file size is already exceeded before the execution of the write access.

Exceptions IO::IOException (p. 79)

Implements IO::LCWriter (p. 168). The documentation for this class was generated from the following files:

• LCSplitWriter.h

• LCSplitWriter.cc

7.94 UTIL::LCStdHepRdr Class Reference

Basic utility for reading a binary stdhep file and filling a LCCollectionVec with MCParticles containing the stdhep file information.

#include <UTIL/LCStdHepRdr.h>

Public Member Functions

• LCStdHepRdr (const char ∗evfile)

Open the stdhep input file in the constructer.

• ∼LCStdHepRdr ()

noop

• long getNumberOfEvents () const

Get number of events in the stdhep file.

• long getNumberOfTotalEventsExpected () const

Get total number of expected events in the whole set of stdhep files from which this stdhep file belongs to.

• IMPL::LCCollectionVec ∗ readEvent ()

Read an event and return an LCCollectionVec of MCParticles.

• void updateNextEvent (IMPL::LCEventImpl ∗evt, const char ∗colName=EVENT::LCIO::MCPARTICLE)

Reads the next stdhep event and adds a new MCParticle collection to the the event with default name ’MCParticle’.

• void printHeader (std::ostream &os=std::cout)

Print the file header to the given ostream.

• int threeCharge (int pdgID) const

Return the charge of the particle times 3 - code copied from HepPDT package.

7.94.1 Detailed Description

Basic utility for reading a binary stdhep file and filling a LCCollectionVec with MCParticles containing the stdhep file information.

Author cassell

Version

Id: LCStdHepRdr.h (p. ?? ),v 1.4 2007-11-12 16:39:04 gaede Exp

7.94.2 Member Function Documentation

7.94.2.1 long UTIL::LCStdHepRdr::getNumberOfEvents ( ) const [inline]

Get number of events in the stdhep file.

This number is read from the file header (no guarantee that it is correct)

7.94.2.2 long UTIL::LCStdHepRdr::getNumberOfTotalEventsExpected ( ) const [inline]

Get total number of expected events in the whole set of stdhep files from which this stdhep file belongs to.

This number is read from the file header (no guarantee that it is correct)

7.94.2.3 IMPL::LCCollectionVec ∗ UTIL::LCStdHepRdr::readEvent ( )

Read an event and return an LCCollectionVec of MCParticles.

Deprecated please use updateEvent()

References IMPL::MCParticleImpl::addParent(), IMPL::MCParticleImpl::getDaughters(), IMPL::LCCollectionVec- ::getElementAt(), IMPL::MCParticleImpl::getGeneratorStatus(), IMPL::MCParticleImpl::getParents(), IMPL::MC-ParticleImpl::getPDG(), IMPL::LCCollectionVec::parameters(), IMPL::MCParticleImpl::setCharge(), IMPL::MC-ParticleImpl::setColorFlow(), IMPL::MCParticleImpl::setGeneratorStatus(), IMPL::MCParticleImpl::setMass(), IMP-L::MCParticleImpl::setMomentum(), IMPL::MCParticleImpl::setPDG(), IMPL::MCParticleImpl::setSimulatorStatus(), IMPL::MCParticleImpl::setSpin(), IMPL::MCParticleImpl::setTime(), EVENT::LCParameters::setValue(), and IMPL- ::MCParticleImpl::setVertex().

7.94.2.4 int UTIL::LCStdHepRdr::threeCharge ( int pdgID ) const

Return the charge of the particle times 3 - code copied from HepPDT package.

PID digits (base 10) are: n nr nl nq1 nq2 nq3 nj The location enum provides a convenient index into the PID.

7.94.2.5 void UTIL::LCStdHepRdr::updateNextEvent ( IMPL::LCEventImpl ∗ evt, const char ∗ colName =

       EVENT::LCIO::MCPARTICLE )

Reads the next stdhep event and adds a new MCParticle collection to the the event with default name ’MCParticle’.

Exceptions

IO::EndOfDataException (p. 68)

if no event in stdhep file

References IMPL::LCEventImpl::addCollection(), EVENT::LCParameters::getFloatVal(), EVENT::LCParameters- ::getIntVal(), IMPL::LCCollectionVec::getParameters(), IMPL::LCEventImpl::parameters(), EVENT::LCParameters- ::setValue(), and IMPL::LCEventImpl::setWeight().

The documentation for this class was generated from the following files:

• LCStdHepRdr.h

• LCStdHepRdr.cc

7.95 UTIL::LCStdHepRdrNew Class Reference

Basic utility for reading a binary stdhep file and filling a LCCollectionVec with MCParticles containing the stdhep file information.

#include <UTIL/LCStdHepRdrNew.h>

Public Member Functions

• LCStdHepRdrNew (const char ∗evfile)

Open the stdhep input file in the constructer.

• ∼LCStdHepRdrNew ()

noop

• long getNumberOfEvents () const

Get number of events in the stdhep file.

• long getNumberOfTotalEventsExpected () const

Get total number of expected events in the whole set of stdhep files from which this stdhep file belongs to.

• IMPL::LCCollectionVec ∗ readEvent ()

Read an event and return an LCCollectionVec of MCParticles.

• void updateNextEvent (IMPL::LCEventImpl ∗evt, const char ∗colName=EVENT::LCIO::MCPARTICLE)

Reads the next stdhep event and adds a new MCParticle collection to the the event with default name ’MCParticle’.

• void printHeader (std::ostream &os=std::cout)

Print the file header to the given ostream.

• int threeCharge (int pdgID) const

Return the charge of the particle times 3 - code copied from HepPDT package.

• void setWriteEventNumber (bool writeEventNumber)

7.95.1 Detailed Description

Basic utility for reading a binary stdhep file and filling a LCCollectionVec with MCParticles containing the stdhep file information.

Author cassell

Version

Id: LCStdHepRdr.h (p. ?? ),v 1.4 2007-11-12 16:39:04 gaede Exp

7.95.2 Member Function Documentation

7.95.2.1 long UTIL::LCStdHepRdrNew::getNumberOfEvents ( ) const [inline]

Get number of events in the stdhep file. This number is read from the file header (no guarantee that it is correct)

7.95.2.2 long UTIL::LCStdHepRdrNew::getNumberOfTotalEventsExpected ( ) const [inline]

Get total number of expected events in the whole set of stdhep files from which this stdhep file belongs to. This number is read from the file header (no guarantee that it is correct)

7.95.2.3 IMPL::LCCollectionVec ∗ UTIL::LCStdHepRdrNew::readEvent ( )

Read an event and return an LCCollectionVec of MCParticles.

Deprecated please use updateEvent()

References IMPL::MCParticleImpl::addParent(), IMPL::MCParticleImpl::getDaughters(), IMPL::LCCollectionVec- ::getElementAt(), IMPL::MCParticleImpl::getGeneratorStatus(), IMPL::MCParticleImpl::getParents(), IMPL::MC-ParticleImpl::getPDG(), IMPL::LCCollectionVec::parameters(), IMPL::MCParticleImpl::setCharge(), IMPL::MC-ParticleImpl::setColorFlow(), IMPL::MCParticleImpl::setGeneratorStatus(), IMPL::MCParticleImpl::setMass(), IMP-L::MCParticleImpl::setMomentum(), IMPL::MCParticleImpl::setPDG(), IMPL::MCParticleImpl::setSimulatorStatus(), IMPL::MCParticleImpl::setSpin(), IMPL::MCParticleImpl::setTime(), EVENT::LCParameters::setValue(), and IMPL- ::MCParticleImpl::setVertex().

7.95.2.4 int UTIL::LCStdHepRdrNew::threeCharge ( int pdgID ) const

Return the charge of the particle times 3 - code copied from HepPDT package.

PID digits (base 10) are: n nr nl nq1 nq2 nq3 nj The location enum provides a convenient index into the PID.

7.95.2.5 void UTIL::LCStdHepRdrNew::updateNextEvent ( IMPL::LCEventImpl ∗ evt, const char ∗ colName =

       EVENT::LCIO::MCPARTICLE )

Reads the next stdhep event and adds a new MCParticle collection to the the event with default name ’MCParticle’.

Exceptions

IO::EndOfDataException (p. 68)

if no event in stdhep file

References IMPL::LCEventImpl::addCollection(), EVENT::LCParameters::getFloatVal(), EVENT::LCParameters- ::getIntVal(), IMPL::LCCollectionVec::getParameters(), IMPL::LCEventImpl::parameters(), IMPL::LCEventImpl::set-EventNumber(), EVENT::LCParameters::setValue(), and IMPL::LCEventImpl::setWeight().

The documentation for this class was generated from the following files:

• LCStdHepRdrNew.h

• LCStdHepRdrNew.cc

7.96 EVENT::LCStrVec Class Reference

String vector used for user extensions.

#include <pre-generated/EVENT/LCStrVec.h>

Inheritance diagram for EVENT::LCStrVec:

Public Types

• typedef LCStrVec lcobject_type

Public Member Functions

• virtual ∼LCStrVec ()

Destructor.

• int id () const

Returns an object id for internal (debugging) use in LCIO (p. 115).

Protected Attributes

• IMPL::AccessChecked _acc

Additional Inherited Members 7.96.1 Detailed Description

String vector used for user extensions. The cpp implementation is a std::vector<string>

Author vogt

Version Jun 18, 2004

See Also LCObject (p. 123) The documentation for this class was generated from the following file:

• pre-generated/EVENT/LCStrVec.h

7.97 UTIL::LCTime Class Reference

Helper class that allows to convert time stamps as defined in LCEvent::getTimeStamp() ( ns since 1.1.1970 00:00:00 UTC) to ordinary date and time and vice versa.

#include <UTIL/LCTime.h>

Classes

• struct CalendarTime

Helper struct that holds the calendar time.

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

160  Class Documentation

Public Member Functions

• LCTime ()

Init with current time ( accuracy depends on OS/machine,etc.) in UTC/GMT.

• LCTime (EVENT::long64 time)

Init from 64 bit time stamp ( ns since 1.1.1970 00:00:00 ).

• LCTime (int year, int month, int day, int hour=0, int min=0, int s=0)

Init from date and time.

• LCTime (int unixTime)

Init from unix time ( s since 1.1.1970 00:00:00 ).

• int unixTime () const

Unix time ( s since 1.1.1970 00:00:00 ).

• EVENT::long64 timeStamp () const

64bit time stamp as used in LCEvent::getTimestamp().

• bool isLeapYear (int year) const

True if the year is a leap year in the Gregorian calendar: year is multiple of 4 and not multiple of 100 or year is multiple of 400.

• int daysInYear (int year) const

The number if days in the given year in the Gregorian calendar.

• int daysInMonth (int month, int year) const

The number if days in the given month and year in the Gregorian calendar.

• std::string getDateString () const

Date in human readable format, e.g.

• const CalendarTime & calendarTime () const

Calendar time: year,month,day hour,min,sec and ns.

• int year () const

Calendar time: year.

• int month () const

Calendar time: month.

• int day () const

Calendar time: day.

• int hour () const

Calendar time: hour.

• int min () const

Calendar time: min.

• int sec () const

Calendar time: sec.

• int ns () const

Calendar time: ns.

• void operator+= (EVENT::long64 t)

Add the specified number of ns to the time.

Static Public Member Functions

• static bool test (int nDates)

Tests the LCTime (p. 159) class with nDates random dates Returns true if successful - throws Exception in case of error.

Protected Member Functions

• void convertToCalTime ()

• void convertFromCalTime ()

Protected Attributes

• EVENT::long64 _t

• CalendarTime _d

Static Protected Attributes

• static int dpm [13] = { 0,31,28,31,30,31,30,31,31,30,31,30,31 }

7.97.1 Detailed Description

Helper class that allows to convert time stamps as defined in LCEvent::getTimeStamp() ( ns since 1.1.1970 00:00:00 UTC) to ordinary date and time and vice versa. Dates and times are in UTC/GMT by default. Negative time stamps are not supported and mapped to 0LL and 1.1.1970 00:00:00 UTC respectively.

Author gaede

Version

Id: LCTime.h (p. ?? ),v 1.4 2005-04-22 15:39:01 gaede Exp

7.97.2 Member Function Documentation

7.97.2.1 int UTIL::LCTime::daysInMonth ( int month, int year ) const

The number if days in the given month and year in the Gregorian calendar. References isLeapYear(), and month(). Referenced by LCTime(), and test().

7.97.2.2 int UTIL::LCTime::daysInYear ( int year ) const

The number if days in the given year in the Gregorian calendar. References isLeapYear().

7.97.2.3 std::string UTIL::LCTime::getDateString ( ) const

Date in human readable format, e.g. 10.02.2005 10:54:29.123456789 : dd.mm.yyyy hh:mm:ss._ms_us_ns

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Referenced by UTIL::LCTOOLS::dumpEventDetailed(), LCTime(), and test(). The documentation for this class was generated from the following files:

• LCTime.h

• LCTime.cc

7.98 UTIL::LCTokenizer Class Reference

Helper class for string tokenization.

#include <UTIL/BitField64.h>

Public Member Functions

• LCTokenizer (std::vector< std::string > &tokens, char del)

Only c’tor, give (empty) token vector and delimeter character.

• void operator() (const char &c)

Operator for use with algorithms, e.g.

7.98.1 Detailed Description

Helper class for string tokenization. Usage: std::vector<std::string> tokens ; LCTokenizer (p. 162) t( tokens ,’,’) ; std::for_each( aString.begin(), aString.end(), t ) ;

7.98.2 Member Function Documentation

7.98.2.1 void UTIL::LCTokenizer::operator() ( const char & c ) [inline]

Operator for use with algorithms, e.g. for_each The documentation for this class was generated from the following file:

• BitField64.h

7.99 UTIL::LCTOOLS Class Reference

Provides some basic functions for printing event data.

#include <UTIL/LCTOOLS.h>

Static Public Member Functions

• static void dumpEvent (const EVENT::LCEvent ∗evt)

Simple function to dump event contents to the screen.

• static void dumpEventDetailed (const EVENT::LCEvent ∗evt)

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Dumps the whole event to the screen, using the dedicated print functions for the different collection types.

• static void dumpRunHeader (const EVENT::LCRunHeader ∗run)

Simple function to dump the run header to the screen.

• static void printMCParticles (const EVENT::LCCollection ∗col)

Complete printout of all MCParticles in the collection.

• static void printSimTrackerHits (const EVENT::LCCollection ∗col)

Complete printout of all SimTrackerHits in the collection.

• static void printTPCHits (const EVENT::LCCollection ∗col)

Complete printout of all TPCHits in the collection.

• static void printTrackerRawData (const EVENT::LCCollection ∗col)

Complete printout of all TrackerRawData objects in the collection.

• static void printTrackerData (const EVENT::LCCollection ∗col)

Complete printout of all TrackerData objects in the collection.

• static void printTrackerPulse (const EVENT::LCCollection ∗col)

Complete printout of all TrackerPulse objects in the collection.

• static void printTrackerHits (const EVENT::LCCollection ∗col)

Complete printout of all TrackerHits in the collection.

• static void printTrackerHitPlane (const EVENT::LCCollection ∗col)

Complete printout of all TrackerHits in the collection.

• static void printTrackerHitZCylinder (const EVENT::LCCollection ∗col)

Complete printout of all TrackerHits in the collection.

• static void printSimCalorimeterHits (const EVENT::LCCollection ∗col)

Complete printout of all SimCalorimeterHits in the collection.

• static void printCalorimeterHits (const EVENT::LCCollection ∗col)

Complete printout of all CalorimeterHits in the collection.

• static void printRawCalorimeterHits (const EVENT::LCCollection ∗col)

Complete printout of all RawCalorimeterHits in the collection.

• static void printLCFloatVecs (const EVENT::LCCollection ∗col)

Print LCFloatVec user extension vectors in the collection.

• static void printLCIntVecs (const EVENT::LCCollection ∗col)

Print LCIntVecs user extension vectors in the collection.

• static void printLCStrVecs (const EVENT::LCCollection ∗col)

Print LCStrVecs user extension vectors in the collection.

• static void printTracks (const EVENT::LCCollection ∗col)

Complete printout of all Tracks in the collection.

• static void printClusters (const EVENT::LCCollection ∗col)

Complete printout of all Clusters in the collection.

• static void printVertices (const EVENT::LCCollection ∗col)

Complete printout of all Vertices in the collection.

• static void printReconstructedParticles (const EVENT::LCCollection ∗col)

Complete printout of all ReconstructedParticles in the collection.

• static void printLCGenericObjects (const EVENT::LCCollection ∗col)

Complete printout of all LCGenericObjects in the collection.

• static void printRelation (const EVENT::LCCollection ∗col)

Print the relation.

• static void printParameters (const EVENT::LCParameters &params)

Print the parameters.

• static std::string getSimulatorStatusString (const EVENT::MCParticle ∗mcp=0)

Returns a string describing the simulator flag bits set for this particle.

• static int printDaughterParticles (const EVENT::MCParticle ∗part, int index)

7.99.1 Detailed Description

Provides some basic functions for printing event data.

Author gaede

Version Mar 10, 2003

See Also LCEvent

7.99.2 Member Function Documentation

7.99.2.1 void UTIL::LCTOOLS::dumpEvent ( const EVENT::LCEvent ∗ evt ) [static]

Simple function to dump event contents to the screen. Also demonstrates the access to event data through the EVENT (p. 21) interface. This is only a short format - one line per collection. References EVENT::LCEvent::getCollection(), EVENT::LCEvent::getCollectionNames(), EVENT::LCEvent::get-DetectorName(), EVENT::LCEvent::getEventNumber(), EVENT::LCCollection::getNumberOfElements(), EVENT::- LCEvent::getRunNumber(), and EVENT::LCCollection::getTypeName().

7.99.2.2 std::string UTIL::LCTOOLS::getSimulatorStatusString ( const EVENT::MCParticle ∗ mcp = 0 ) [static]

Returns a string describing the simulator flag bits set for this particle. If no MCParticle given a string describing the bits is returned: References EVENT::MCParticle::getSimulatorStatus(), EVENT::MCParticle::hasLeftDetector(), EVENT::M-CParticle::isBackscatter(), EVENT::MCParticle::isCreatedInSimulation(), EVENT::MCParticle::isDecayedIn-Calorimeter(), EVENT::MCParticle::isDecayedInTracker(), EVENT::MCParticle::isOverlay(), EVENT::MCParticle- ::isStopped(), and EVENT::MCParticle::vertexIsNotEndpointOfParent(). The documentation for this class was generated from the following files:

• UTIL/LCTOOLS.h

• LCTOOLS.cc

7.100 UTIL::LCTypedVector< T > Class Template Reference

Templated helper class that is an std::vector<T∗> of the elements in the LCCollection.

#include <UTIL/LCTypedVector.h>

Inheritance diagram for UTIL::LCTypedVector< T >:

Public Member Functions

• LCTypedVector (EVENT::LCCollection ∗col)

• EVENT::LCCollection ∗ col ()

Protected Attributes

• EVENT::LCCollection ∗ _col

7.100.1 Detailed Description

template<class T>class UTIL::LCTypedVector< T >

Templated helper class that is an std::vector<T∗> of the elements in the LCCollection. This can be used, e.g. if the objects are to be passed to an algorithm in terms of iterators. For example: LCTypedVector<SimTrackerHits> sim( evt->getCollection("tpc") ) ; LCCollectionVec∗ trackerHits = new LCCollectionVec( LCIO::TRACKERHIT ) ; std::transform( sim.begin(), sim.end(), std::back_inserter( ∗trackerHits ) , digitizer ) ; Note that instantiating an LCTypedVector (p. 164) involves copying and casting of the pointers in the collection. In a future release of LCIO this might not be neccessary any more.

Author gaede

Version

Id: LCTypedVector.h (p. ?? ),v 1.3 2007-02-19 17:38:46 gaede Exp

The documentation for this class was generated from the following file:

• LCTypedVector.h

7.101 UTIL::LCWarning Class Reference

Utility class to show warnings in LCIO.

#include <UTIL/LCWarning.h>

Public Member Functions

• void registerWarning (const std::string &id, const std::string &txt, int max=10)

register a new warning

• void printWarning (const std::string &id)

print the warning text associated to the given id

Static Public Member Functions

• static LCWarning & getInstance ()

7.101.1 Detailed Description

Utility class to show warnings in LCIO. The maximum amount of warnings printed is configurable and per default set to 10. Warnings are shown one last time when application ends.

Author engels

Version

7.101.2 Member Function Documentation

7.101.2.1 void UTIL::LCWarning::registerWarning ( const std::string & id, const std::string & txt, int max = 10 )

register a new warning id = "a_unique_string_to_identify_the_warning" txt = "the world will end tomorrow, leave the room immediately and go get some fun ;)" max = the maximum amount of times to display the warning The documentation for this class was generated from the following files:

• LCWarning.h

• LCWarning.cc

7.102 IO::LCWriter Class Reference

Interface for writing data with LCIO.

#include <pre-generated/IO/LCWriter.h>

Inheritance diagram for IO::LCWriter:

Public Member Functions

• virtual ∼LCWriter ()

Destructor.

• virtual void open (const std::string &filename)=0 throw (IOException, std::exception )

Opens a file for writing.

• virtual void open (const std::string &filename, int writeMode)=0 throw (IOException, std::exception )

Opens a file for writing.

• virtual void setCompressionLevel (int level)=0

Set the compression level - needs to be called before open() (p. 167) otherwise call will have no effect.

• virtual void writeRunHeader (const EVENT::LCRunHeader ∗hdr)=0 throw (IOException, std::exception )

Writes the given run header to file.

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.102 IO::LCWriter Class Reference  167

• virtual void writeEvent (const EVENT::LCEvent ∗evt)=0 throw (IOException, std::exception )

Writes the given event to file.

• virtual void close ()=0 throw (IOException, std::exception )

Closes the output file/stream.

• virtual void flush ()=0 throw (IOException, std::exception )

Flushes the output file/stream.

7.102.1 Detailed Description

Interface for writing data with LCIO. Uses interfaces from EVENT/hep.lcio.event. Use LCFactory to instantiate a corresponding LCWriter (p. 166) object for the output format at hand (SIO (p. 30) only, so far).

See Also LCFactory

Author gaede

Version Mar 4, 2003

7.102.2 Member Function Documentation

7.102.2.1 virtual void IO::LCWriter::close ( ) throw IOException, std::exception) [pure virtual] Closes the output file/stream. Exceptions

IOException (p. 79)

Implemented in SIO::SIOWriter (p. 255), and UTIL::LCSplitWriter (p. 153).

7.102.2.2 virtual void IO::LCWriter::flush ( ) throw IOException, std::exception) [pure virtual] Flushes the output file/stream. Exceptions

IOException (p. 79)

Implemented in SIO::SIOWriter (p. 256), and UTIL::LCSplitWriter (p. 153).

7.102.2.3 virtual void IO::LCWriter::open ( const std::string & filename ) throw IOException, std::exception) [pure

        virtual]

Opens a file for writing. If file with given name exists, an exception is thrown. Use append or new mode instead.

Exceptions IOException (p. 79)

Implemented in SIO::SIOWriter (p. 256), and UTIL::LCSplitWriter (p. 153).

7.102.2.4 virtual void IO::LCWriter::open ( const std::string & filename, int writeMode ) throw IOException, std::exception)

        [pure virtual]

Opens a file for writing. Possible write modes are: LCIO::WRITE_NEW (existing files are replaced) and LCIO::WRITE_APPEND. Exceptions

IOException (p. 79)

Implemented in SIO::SIOWriter (p. 256), and UTIL::LCSplitWriter (p. 154).

7.102.2.5 virtual void IO::LCWriter::setCompressionLevel ( int level ) [pure virtual]

Set the compression level - needs to be called before open() (p. 167) otherwise call will have no effect. If not called the Writer will use default compression. Valid compression levels are:

• level < 0 : default compression

• level == 0 : no compression

• level > 0 : compression level (typically 1 (fastest) - 9 (best compression))

Status: (v01-09) C++: experimental code - don’t use for production Java: not implemented Parameters

level

compression level

Implemented in SIO::SIOWriter (p. 256), and UTIL::LCSplitWriter (p. 154). Referenced by UTIL::LCSplitWriter::setCompressionLevel().

7.102.2.6 virtual void IO::LCWriter::writeEvent ( const EVENT::LCEvent ∗ evt ) throw IOException, std::exception)

        [pure virtual]

Writes the given event to file. Exceptions

IOException (p. 79)

Implemented in SIO::SIOWriter (p. 257), and UTIL::LCSplitWriter (p. 154).

7.102.2.7 virtual void IO::LCWriter::writeRunHeader ( const EVENT::LCRunHeader ∗ hdr ) throw IOException, std::exception) [pure virtual]

Writes the given run header to file.

Exceptions IOException (p. 79)

Implemented in SIO::SIOWriter (p. 257), and UTIL::LCSplitWriter (p. 154). The documentation for this class was generated from the following file:

• LCWriter.h

7.103 Mass Struct Reference

Inheritance diagram for Mass:

The documentation for this struct was generated from the following file:

• lcrtrelation.cc

7.104 EVENT::MCParticle Class Reference

The LCIO (p. 115) Monte Carlo particle.

#include <pre-generated/EVENT/MCParticle.h>

Inheritance diagram for EVENT::MCParticle:

Public Types

• typedef MCParticle lcobject_type

Useful typedef for template programming with LCIO (p. 115).

Public Member Functions

• virtual ∼MCParticle ()

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Destructor.

• virtual double getEnergy () const =0

Returns the energy of the particle (at the vertex) in [GeV] computed from the particle’s momentum and mass - only float used in files.

• virtual const float ∗ getSpin () const =0

Returns the spin (helicity) vector of the particle.

• virtual const int ∗ getColorFlow () const =0

Returns the color flow as defined by the generator.

• virtual const MCParticleVec & getParents () const =0

Returns the parents of this particle.

• virtual const MCParticleVec & getDaughters () const =0

Returns the daughters of this particle.

• virtual int getPDG () const =0

Returns the PDG code of the particle.

• virtual int getGeneratorStatus () const =0

Returns the status for particles as defined by the generator, typically 0 empty line 1 undecayed particle, stable in the generator 2 particle decayed in the generator 3 documentation line.

• virtual int getSimulatorStatus () const =0

Returns the status for particles from the simulation, e.g.

• virtual bool isCreatedInSimulation () const =0

True if the particle has been created by the simulation program (rather than the generator).

• virtual bool isBackscatter () const =0

True if the particle was created by the simulator as a result of an interaction or decay in non-tracking region, e.g.

• virtual bool vertexIsNotEndpointOfParent () const =0

True if the particle was created as a result of a continuous process where the parent particle continues, i.e.

• virtual bool isDecayedInTracker () const =0

True if the particle decayed or interacted in a tracking region.

• virtual bool isDecayedInCalorimeter () const =0

True if the particle decayed or interacted (non-continuous interaction, particle terminated) in non-tracking region.

• virtual bool hasLeftDetector () const =0

True if the particle left the world volume undecayed.

• virtual bool isStopped () const =0

True if the particle lost all kinetic energy inside the world volume and did not decay.

• virtual bool isOverlay () const =0

True if the particle has been overlayed by the simulation (or digitization) program.

• virtual const double ∗ getVertex () const =0

Returns the production vertex of the particle in [mm].

• virtual float getTime () const =0

The creation time of the particle in [ns] wrt.

• virtual const double ∗ getEndpoint () const =0

Returns the endpoint of the particle in [mm] if the endpoint has been set explicetly.

• virtual const double ∗ getMomentum () const =0

Returns the particle’s 3-momentum at the production vertex in [GeV].

• virtual const double ∗ getMomentumAtEndpoint () const =0

Returns the particle’s 3-momentum at the endpoint in [GeV] Only float used in files.

• virtual double getMass () const =0

Returns the mass of the particle in [GeV] - only float used in files.

• virtual float getCharge () const =0

Returns the particle’s charge.

Static Public Attributes

• static const int BITEndpoint = 31

• static const int BITCreatedInSimulation = 30

• static const int BITBackscatter = 29

• static const int BITVertexIsNotEndpointOfParent = 28

• static const int BITDecayedInTracker = 27

• static const int BITDecayedInCalorimeter = 26

• static const int BITLeftDetector = 25

• static const int BITStopped = 24

• static const int BITOverlay = 23

Additional Inherited Members

7.104.1 Detailed Description

The LCIO (p. 115) Monte Carlo particle.

Author gaede

Version

Id: MCParticle.aid,v 1.26 2006-08-03 16:53:34 gaede Exp

7.104.2 Member Function Documentation

7.104.2.1 virtual const double∗ EVENT::MCParticle::getEndpoint ( ) const [pure virtual]

Returns the endpoint of the particle in [mm] if the endpoint has been set explicetly. If not, the vertex of the first daughter that has vertexIsNotEndpointOfParent==false is returned. If no such daughter exists (0,0,0) is returned. Implemented in IMPL::MCParticleImpl (p. 176). Referenced by UTIL::LCTOOLS::printMCParticles(), and SIO::SIOParticleHandler::write().

7.104.2.2 virtual const double∗ EVENT::MCParticle::getMomentum ( ) const [pure virtual]

Returns the particle’s 3-momentum at the production vertex in [GeV].

• only float used in files.

Implemented in IMPL::MCParticleImpl (p. 174). Referenced by UTIL::LCTOOLS::printMCParticles(), and SIO::SIOParticleHandler::write().

7.104.2.3 virtual int EVENT::MCParticle::getSimulatorStatus ( ) const [pure virtual]

Returns the status for particles from the simulation, e.g. decayed in flight. Bits 31-16 are used to decode the information. Use the followng boolean functions to determine the proper simulator status:

See Also isCreatedInSimulation() (p. 170)

isBackscatter() (p. 172)

vertexIsNotEndpointOfParent() (p. 172) isDecayedInTracker() (p. 170) isDecayedInCalorimeter() (p. 170) hasLeftDetector() (p. 170) isStopped() (p. 170) isOverlay() (p. 170)

Implemented in IMPL::MCParticleImpl (p. 176). Referenced by UTIL::LCTOOLS::getSimulatorStatusString(), and SIO::SIOParticleHandler::write().

7.104.2.4 virtual float EVENT::MCParticle::getTime ( ) const [pure virtual]

The creation time of the particle in [ns] wrt. the event, e.g. for preassigned decays or decays in flight from the simulator. Implemented in IMPL::MCParticleImpl (p. 177). Referenced by SIO::SIOParticleHandler::write().

7.104.2.5 virtual bool EVENT::MCParticle::isBackscatter ( ) const [pure virtual]

True if the particle was created by the simulator as a result of an interaction or decay in non-tracking region, e.g. a calorimeter. By convention, such particles are not saved. However, if this particle creates a tracker hit, the particle is added to the MCParticle (p. 169) list with this flag set, and the parent set to the particle that initially decayed or interacted in a non-tracking region. Implemented in IMPL::MCParticleImpl (p. 174). Referenced by UTIL::LCTOOLS::getSimulatorStatusString().

7.104.2.6 virtual bool EVENT::MCParticle::vertexIsNotEndpointOfParent ( ) const [pure virtual]

True if the particle was created as a result of a continuous process where the parent particle continues, i.e. hard ionization, Bremsstrahlung, elastic interactions, etc. Implemented in IMPL::MCParticleImpl (p. 174). Referenced by UTIL::LCTOOLS::getSimulatorStatusString(). The documentation for this class was generated from the following files:

• MCParticle.h

• MCParticleImpl.cc

7.105 IMPL::MCParticleCont Struct Reference

Public Member Functions

• MCParticleCont (EVENT::MCParticle ∗part, float e, float t, int pdg, const float ∗step)

Public Attributes

• EVENT::MCParticle ∗ Particle

• float Energy

• float Time

• int PDG

• float StepPosition [3]

The documentation for this struct was generated from the following file:

• SimCalorimeterHitImpl.h

7.106 IMPL::MCParticleImpl Class Reference

Implementation of MCParticle.

#include <IMPL/MCParticleImpl.h>

Inheritance diagram for IMPL::MCParticleImpl:

Public Member Functions

• virtual ∼MCParticleImpl ()

Copy c’tor.

• virtual int id () const

Returns an object id for internal (debugging) use in LCIO.

• virtual const

EVENT::MCParticleVec & getParents () const

Returns the parents of this particle.

• virtual const

EVENT::MCParticleVec & getDaughters () const

Returns the daughters of this particle.

• virtual const double ∗ getEndpoint () const

Returns the i-th daughter of this particle.

• virtual int getPDG () const

Returns the number of daughters of this particle.

• virtual int getGeneratorStatus () const

Returns the status for particles from the generator 0 empty line 1 undecayed particle, stable in the generator 2 particle decayed in the generator 3 documentation line.

• virtual int getSimulatorStatus () const

Returns the status for particles from the simulation, e.g.

• virtual bool isCreatedInSimulation () const

True if the particle has been created by the simulation program (rather than the generator).

• virtual bool isBackscatter () const

True if the particle is the result of a backscatter from a calorimeter shower.

• virtual bool vertexIsNotEndpointOfParent () const

True if the particle’s vertex is not the endpoint of the parent particle.

• virtual bool isDecayedInTracker () const

True if the particle has interacted in a tracking region.

• virtual bool isDecayedInCalorimeter () const

True if the particle has interacted in a calorimeter region.

• virtual bool hasLeftDetector () const

True if the particle has left the world volume undecayed.

• virtual bool isStopped () const

True if the particle has been stopped by the simulation program.

• virtual bool isOverlay () const

True if the particle has been overlayed by the simulation (or digitization) program.

• virtual const double ∗ getVertex () const

Returns the production vertex of the particle.

• virtual float getTime () const

The creation time of the particle in [ns] wrt.

• virtual const double ∗ getMomentum () const

Returns the particle momentum at the production vertex.

• virtual const double ∗ getMomentumAtEndpoint () const

Returns the particle momentum at the endpoint.

• virtual double getMass () const

Returns the mass of the particle in [GeV].

• virtual float getCharge () const

Returns the particle’s charge.

• virtual double getEnergy () const

Returns the energy of the particle (at the vertex) in [GeV] computed from the particle’s momentum and mass.

• virtual const float ∗ getSpin () const

Returns the spin.

• virtual const int ∗ getColorFlow () const

Returns the color flow.

• void addParent (EVENT::MCParticle ∗mom)

Adds a parent particle.

• void setPDG (int pdg)

Sets the parent.

• void setGeneratorStatus (int status)

Sets the Generator status.

• void setSimulatorStatus (int status)

Sets the Simulator status.

• void setVertex (const double vtx[3])

Sets the production vertex.

• void setTime (float time)

Sets the createion time.

• void setEndpoint (const double pnt[3])

Sets the particle endpoint.

• void setMomentum (const float p[3])

Sets the momentum.

• void setMomentum (const double p[3])

Sets the momentum.

• void setMomentumAtEndpoint (const float p[3])

Sets the momentum at the endpoint.

• void setMomentumAtEndpoint (const double p[3])

Sets the momentum at the endpoint.

• void setMass (float m)

Sets the mass.

• void setCharge (float c)

Sets the charge.

• void setSpin (const float spin[3])

Sets the spin.

• void setColorFlow (const int cflow[2])

Sets the color flow.

• virtual void setCreatedInSimulation (bool val)

• virtual void setBackscatter (bool val)

• virtual void setVertexIsNotEndpointOfParent (bool val)

• virtual void setDecayedInTracker (bool val)

• virtual void setDecayedInCalorimeter (bool val)

• virtual void setHasLeftDetector (bool val)

• virtual void setStopped (bool val)

• virtual void setOverlay (bool val)

Protected Member Functions

• void addDaughter (EVENT::MCParticle ∗daughter)

Adds a daughter particle - only called from addParent() (p. 174).

Protected Attributes

• int _pdg

• int _genstatus

• std::bitset< 32 > _simstatus

• double _vertex [3]

• double _endpoint [3]

• double _pEndpoint [3]

• double _p [3]

• double _mass

• float _charge

• float _time

• EVENT::MCParticleVec _parents

• EVENT::MCParticleVec _daughters

• bool _endpointSet

• float _spin [3]

• int _colorFlow [2]

176  Class Documentation

Additional Inherited Members 7.106.1 Detailed Description

Implementation of MCParticle.

Author gaede

Version

Id: MCParticleImpl.h (p. ?? ),v 1.24 2010-01-21 16:28:25 gaede Exp

7.106.2 Constructor & Destructor Documentation

7.106.2.1 IMPL::MCParticleImpl::∼MCParticleImpl ( ) [virtual] Copy c’tor. Destructor.

7.106.3 Member Function Documentation

7.106.3.1 const double ∗ IMPL::MCParticleImpl::getEndpoint ( ) const [virtual] Returns the i-th daughter of this particle. Same as getDaughter() except for return type.

See Also getNumberOfDaughtersReturns the number of parents of this particle - 0 if mother.Returns the i-th parent of this particle.Returns the endpoint of the particle in [mm] - returns NULL if undefined (not set). Definition of the enpoint depends on the application that created the particle, e.g. the start point of the shower in a calorimeter.

Implements EVENT::MCParticle (p. 171).

7.106.3.2 int IMPL::MCParticleImpl::getPDG ( ) const [virtual] Returns the number of daughters of this particle. Returns the PDG code of the particle. Implements EVENT::MCParticle (p. 170). Referenced by UTIL::LCStdHepRdr::readEvent(), and UTIL::LCStdHepRdrNew::readEvent().

7.106.3.3 int IMPL::MCParticleImpl::getSimulatorStatus ( ) const [virtual] Returns the status for particles from the simulation, e.g. decayed in flight. Bits 31-16 are used to decode the information. Use the followng boolean functions to determine the proper simulator status:

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

See Also isCreatedInSimulation() (p. 174)

isBackscatter() (p. 174)

vertexIsNotEndpointOfParent() (p. 174) isDecayedInTracker() (p. 174) isDecayedInCalorimeter() (p. 174) hasLeftDetector() (p. 174) isStopped() (p. 174) isOverlay() (p. 174)

Implements EVENT::MCParticle (p. 172).

7.106.3.4 float IMPL::MCParticleImpl::getTime ( ) const [virtual]

The creation time of the particle in [ns] wrt.

the event, e.g. for preassigned decays or decays in flight from the simulator.

Implements EVENT::MCParticle (p. 172).

7.106.3.5 void IMPL::MCParticleImpl::setPDG ( int pdg )

Sets the parent.

Sets a second parent.Sets the PDG code.

Referenced by HEPEVTIMPL::HEPEVT::fromHepEvt(), UTIL::LCStdHepRdrNew::readEvent(), and UTIL::LCStd-HepRdr::readEvent().

The documentation for this class was generated from the following files:

• MCParticleImpl.h

• MCParticleImpl.cc

7.107 IOIMPL::MCParticleIOImpl Class Reference

Adding stuff needed for io (friend declarations, etc.)

#include <IOIMPL/MCParticleIOImpl.h>

Inheritance diagram for IOIMPL::MCParticleIOImpl:

178  Class Documentation

Friends

• class SIO::SIOParticleHandler

Additional Inherited Members 7.107.1 Detailed Description

Adding stuff needed for io (friend declarations, etc.)

Author gaede

Version Mar 7, 2003 The documentation for this class was generated from the following file:

• MCParticleIOImpl.h

7.108 MCTruth2Rec Struct Reference

Inheritance diagram for MCTruth2Rec:

The documentation for this struct was generated from the following file:

• test_RTextension.cc

7.109 MyUserExtension Struct Reference

Inheritance diagram for MyUserExtension:

The documentation for this struct was generated from the following file:

• lcrtrelation.cc

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.110 lcrtrel_helper::NoDelete Struct Reference

Empty delete function for pointers w/o ownership.

#include <LCRTRelations.h>

Static Public Member Functions

• static void clean (void ∗)

7.110.1 Detailed Description

Empty delete function for pointers w/o ownership. The documentation for this struct was generated from the following file:

• LCRTRelations.h

7.111 lcrtrel_helper::objorcont< is_container > Struct Template Reference

Helper functions that treat single objects and containers.

#include <LCRTRelations.h>

Static Public Member Functions

• template<class T , class S >

static void add (T t, S s)

• template<class T , class S >

static void remove (T t, S s)

7.111.1 Detailed Description

template<bool is_container>struct lcrtrel_helper::objorcont< is_container >

Helper functions that treat single objects and containers. The documentation for this struct was generated from the following file:

• LCRTRelations.h

7.112 lcrtrel_helper::objorcont< false > Struct Template Reference

Helper functions specialization for single objects.

#include <LCRTRelations.h>

Static Public Member Functions

• template<class T , class S >

static void add (T &t, S s)

• template<class T , class S >

static void remove (T &t, S)

7.112.1 Detailed Description

template<>struct lcrtrel_helper::objorcont< false >

Helper functions specialization for single objects. The documentation for this struct was generated from the following file:

• LCRTRelations.h

7.113 opts_t Struct Reference

Public Attributes

• const char ∗ colName

• char ∗∗ inputFiles

• int inputFilesNum

• int verbosity

• bool pedantic

• bool average

• int startevent

• int maxevents

• int expelements

• int minelements

• int maxelements

• int abselementerror

• int abseventerror

• float relelementerror

• float releventerror

• int elementsTotal

• int eventsTotal

• int eventsFailed

• int eventsSkipped

The documentation for this struct was generated from the following file:

• lcio_check_col_elements.cc

7.114 pagerInfo Struct Reference

Public Attributes

• char ∗ filename

• int fd_temp

• int fd_old

• bool save

The documentation for this struct was generated from the following file:

• lsh.cc

7.115 ParentDaughter Struct Reference

Inheritance diagram for ParentDaughter:

The documentation for this struct was generated from the following file:

• lcrtrelation.cc

7.116 EVENT::ParticleID Class Reference

Persistent interface for LCIO (p. 115) ParticleIDs (p. 185).

#include <pre-generated/EVENT/ParticleID.h>

Inheritance diagram for EVENT::ParticleID:

Public Types

• typedef ParticleID lcobject_type

Useful typedef for template programming with LCIO (p. 115).

Public Member Functions

• virtual ∼ParticleID ()

Destructor.

• virtual int getType () const =0

Type - userdefined.

• virtual int getPDG () const =0

The PDG code of this id - UnknownPDG ( 999999 ) if unknown.

• virtual float getLikelihood () const =0

The likelihood of this hypothesis - in a user defined normalization.

• virtual int getAlgorithmType () const =0

Type of the algorithm/module that created this hypothesis - NOTE: must be unique within one collection.

• virtual const FloatVec & getParameters () const =0

Parameters associated with this hypothesis.

Static Public Attributes

• static const int UnknownPDG = 999999

Constant to be used if the PDG code is not known or undefined.

Additional Inherited Members

7.116.1 Detailed Description

Persistent interface for LCIO (p. 115) ParticleIDs (p. 185). Used by ReconstructedParticle (p. 193) and Cluster (p. 58) for different hypotheses on the particle type.

Author gaede

Version

Id: ParticleID.aid,v 1.11 2008-05-30 13:23:27 gaede Exp

See Also ReconstructedParticle.getParticleIDs() (p. 196)

Cluster.getParticleIDs() (p. 60)

7.116.2 Member Function Documentation

7.116.2.1 virtual int EVENT::ParticleID::getAlgorithmType ( ) const [pure virtual]

Type of the algorithm/module that created this hypothesis - NOTE: must be unique within one collection. Check/set collection parameters PIDAlgorithmTypeName and PIDAlgorithmTypeID. Implemented in IMPL::ParticleIDImpl (p. 184). Referenced by UTIL::LCTOOLS::printClusters(), UTIL::LCTOOLS::printReconstructedParticles(), SIO::SIOCluster-Handler::write(), and SIO::SIOReconstructedParticleHandler::write().

7.116.2.2 virtual const FloatVec& EVENT::ParticleID::getParameters ( ) const [pure virtual]

Parameters associated with this hypothesis. Check/set collection parameters ParameterNames_PIDAlgorithmTypeName for decoding the indices. Implemented in IMPL::ParticleIDImpl (p. 184). Referenced by UTIL::LCTOOLS::printClusters(), UTIL::LCTOOLS::printReconstructedParticles(), SIO::SIOCluster-Handler::write(), and SIO::SIOReconstructedParticleHandler::write(). The documentation for this class was generated from the following file:

• ParticleID.h

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.117 IMPL::ParticleIDImpl Class Reference

Implementation of ParticleID.

#include <IMPL/ParticleIDImpl.h>

Inheritance diagram for IMPL::ParticleIDImpl:

Public Member Functions

• ParticleIDImpl ()

Default constructor, initializes values to 0.

• virtual ∼ParticleIDImpl ()

Destructor.

• virtual int id () const

Returns an object id for internal (debugging) use in LCIO.

• virtual int getType () const

Type - userdefined.

• virtual int getPDG () const

The PDG code of this id - UnknownPDG ( 999999 ) if unknown.

• virtual float getLikelihood () const

The likelihood of this hypothesis - in a user defined normalization.

• virtual int getAlgorithmType () const

Type of the algorithm/module that created this hypothesis - NOTE: must be unique within one collection.

• virtual const EVENT::FloatVec & getParameters () const

Parameters associated with this hypothesis.

• virtual EVENT::FloatVec & parameters ()

Access to parameters associated with this hypothesis.

• void setType (int type)

• void setPDG (int pdg)

• void setLikelihood (float logL)

• void setAlgorithmType (int algorithmType)

• void addParameter (float p)

Protected Attributes

• int _type

• int _pdg

• float _likelihood

• int _algorithmType

• EVENT::FloatVec _parameters

Additional Inherited Members

7.117.1 Detailed Description

Implementation of ParticleID.

See Also ParticleID

Author gaede

Version

Id: ParticleIDImpl.h (p. ?? ),v 1.11 2008-05-30 13:23:27 gaede Exp

7.117.2 Member Function Documentation

7.117.2.1 int IMPL::ParticleIDImpl::getAlgorithmType ( ) const [virtual]

Type of the algorithm/module that created this hypothesis - NOTE: must be unique within one collection. Check/set collection parameters PIDAlgorithmTypeName and PIDAlgorithmTypeID. Implements EVENT::ParticleID (p. 182).

7.117.2.2 const EVENT::FloatVec & IMPL::ParticleIDImpl::getParameters ( ) const [virtual]

Parameters associated with this hypothesis. Check/set collection parameter ParameterNames_PIDAlgorithmTypeName for decoding the indices. Implements EVENT::ParticleID (p. 182). The documentation for this class was generated from the following files:

• ParticleIDImpl.h

• ParticleIDImpl.cc

7.118 IOIMPL::ParticleIDIOImpl Class Reference

Adding stuff needed for io (friend declarations, etc.)

#include <IOIMPL/ParticleIDIOImpl.h>

Inheritance diagram for IOIMPL::ParticleIDIOImpl:

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Friends

• class SIO::SIOReconstructedParticleHandler

• class SIO::SIOClusterHandler

Additional Inherited Members 7.118.1 Detailed Description

Adding stuff needed for io (friend declarations, etc.)

Author gaede

Version Mar 31, 2004 The documentation for this class was generated from the following file:

• ParticleIDIOImpl.h

7.119 ParticleIDs Struct Reference

Inheritance diagram for ParticleIDs:

The documentation for this struct was generated from the following file:

• lcrtrelation.cc

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.120 UTIL::PIDHandler Class Reference

Convenient class for setting and retrieving particle id information attached to ReconstructedParticles.

#include <UTIL/PIDHandler.h>

Public Member Functions

• PIDHandler (const LCCollection ∗col)

Create PIDHandler (p. 186) for reading from the given collection - read the collection parameters PIDAlgorithmType-Name, PIDAlgorithmTypeID and the corresponding ParameterNames_PIDAlgorithmTypeName if they exist.

• PIDHandler (LCCollection ∗col)

Create PIDHandler (p. 186) for setting PID information in the given collection.

• ∼PIDHandler ()

Update the collection parameter when going out of scope.

• int addAlgorithm (const std::string &algoName, const StringVec &parameterNames)

Add a new algorithm and return the corresponding algorithm id.

• int getAlgorithmID (const std::string &algoName)

Return the typeID of algorithm algoName - throws UnknownAlgorithm (p. 311).

• const std::string & getAlgorithmName (int algoID)

Return the name of the algorithm with id - throws UnknownAlgorithm (p. 311).

• int getParameterIndex (int algorithmID, const std::string &pName)

The index of parameter pName for the algorithm with algorithmID - throws UnknownAlgoritm.

• const ParticleID & getParticleID (LCObject ∗particle, int algorithmID)

Return the (first) ParticleID object for the given algorithm and particle (or cluster) - throws UnknownAlgorithm (p. 311).

• ParticleIDVec getParticleIDs (LCObject ∗p, int id)

Return all PID objects for a given algorithm - ordered with decreasing likelihood - throws UnknownAlgorithm (p. 311).

• void setParticleIDUsed (IMPL::ReconstructedParticleImpl ∗particle, int algorithmID)

Set the particleID algorithm that is used for this particle’s kinematic variables.

• const StringVec & getParameterNames (int algorithmID)

The names of parameters for the algorithm with algorithmID - throws UnknownAlgoritm.

• const IntVec & getAlgorithmIDs ()

Return the IDs of all known Algorithms.

• void setParticleID (LCObject ∗p, int userType, int PDG, float likelihood, int algorithmID, const FloatVec

&params)

Set the particleID information for this particle (or cluster) - throws UnknownAlgorithm (p. 311).

Protected Member Functions

• int nextID ()

• void init (const LCCollection ∗col)

Protected Attributes

• LCCollection ∗ _col

• CPM _cpm

• int _type

• int _maxID

• PNM _pNames

• CPMINV _cpmInv

• IntVec _ids

7.120.1 Detailed Description

Convenient class for setting and retrieving particle id information attached to ReconstructedParticles.

Use one instance of PIDHandler (p. 186) for one LCCollection only. When used for setting particle id information the scope of the PIDHandler (p. 186) must match that of the collection, as the collection parameters are updated automatically when the PIDHandler (p. 186) object goes out of scope. Reads (and updates) the collection parameters PIDAlgorithmTypeName, PIDAlgorithmTypeID and the corresponding ParameterNames_PIDAlgorithmTypeName for every algorithm used.

See Also ReconstructedParticle ParticleID

Author F.Gaede, DESY

Version

Id: PIDHandler.h (p. ?? ),v 1.2 2010-01-21 16:28:25 gaede Exp

7.120.2 Constructor & Destructor Documentation

7.120.2.1 UTIL::PIDHandler::PIDHandler ( LCCollection ∗ col )

Create PIDHandler (p. 186) for setting PID information in the given collection.

• read and update the collection parameters PIDAlgorithmTypeName, PIDAlgorithmTypeID and the corresponding ParameterNames_PIDAlgorithmTypeName if they exist.

7.120.3 Member Function Documentation

7.120.3.1 const ParticleID & UTIL::PIDHandler::getParticleID ( LCObject ∗ particle, int algorithmID )

Return the (first) ParticleID object for the given algorithm and particle (or cluster) - throws UnknownAlgorithm (p. 311).

Only use if you know there is only one PID object for the algorithms or if you simply want the most likely PID for this algorithm.

References getParticleIDs().

7.120.3.2 void UTIL::PIDHandler::setParticleIDUsed ( IMPL::ReconstructedParticleImpl ∗ particle, int algorithmID )

Set the particleID algorithm that is used for this particle’s kinematic variables.

• throws UnknownAlgorithm (p. 311).

References IMPL::ReconstructedParticleImpl::getParticleIDs().

The documentation for this class was generated from the following files:

• PIDHandler.h

• PIDHandler.cc

7.121 IMPL::PIDSort Class Reference

Helper class to sort ParticleIDs (p. 185) wrt.

#include <IMPL/ParticleIDImpl.h>

Inheritance diagram for IMPL::PIDSort:

Public Member Functions

• bool operator() (const EVENT::ParticleID ∗p1, const EVENT::ParticleID ∗p2)

7.121.1 Detailed Description

Helper class to sort ParticleIDs (p. 185) wrt. their likelihood. The documentation for this class was generated from the following file:

• ParticleIDImpl.h

7.122 EVENT::RawCalorimeterHit Class Reference

The generic calorimeter hit for real data (or simulation thereof).

#include <pre-generated/EVENT/RawCalorimeterHit.h>

Inheritance diagram for EVENT::RawCalorimeterHit:

Public Types

• typedef RawCalorimeterHit lcobject_type

Useful typedef for template programming with LCIO (p. 115).

Public Member Functions

• virtual ∼RawCalorimeterHit ()

Destructor.

• virtual int getCellID0 () const =0

Returns the detector specific (geometrical) cell id.

• virtual int getCellID1 () const =0

Returns the second detector specific (geometrical) cell id.

• virtual int getAmplitude () const =0

Returns the amplitude of the hit in ADC counts.

• virtual int getTimeStamp () const =0

Returns a time stamp for the hit.

Additional Inherited Members

7.122.1 Detailed Description

The generic calorimeter hit for real data (or simulation thereof). In order to reduce the file size you can suppress a (32-bit) pointer tag stored with every hit by setting the flag(LCIO- ::RCHBIT_NO_PTR)==1. NB: If you apply this flag to reduce the file size you won’t be able to point/refer to the RawCalorimeterHits, e.g. from an LCRelation (p. 136) object.

Author gaede

Version

Id: RawCalorimeterHit.aid,v 1.4 2006-08-03 16:53:35 gaede Exp

7.122.2 Member Function Documentation

7.122.2.1 virtual int EVENT::RawCalorimeterHit::getCellID1 ( ) const [pure virtual]

Returns the second detector specific (geometrical) cell id. Optional, check/set flag(LCIO::RCHBIT_ID1)==1. Implemented in IMPL::RawCalorimeterHitImpl (p. 191). Referenced by SIO::SIORawCalHitHandler::write().

7.122.2.2 virtual int EVENT::RawCalorimeterHit::getTimeStamp ( ) const [pure virtual]

Returns a time stamp for the hit. Optional, check/set flag(LCIO::RCHBIT_TIME)==1. Implemented in IMPL::RawCalorimeterHitImpl (p. 191). Referenced by SIO::SIORawCalHitHandler::write(). The documentation for this class was generated from the following file:

• RawCalorimeterHit.h

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.123 IMPL::RawCalorimeterHitImpl Class Reference

Implementation of the real data RawCalorimeterHit.

#include <IMPL/RawCalorimeterHitImpl.h>

Inheritance diagram for IMPL::RawCalorimeterHitImpl:

Public Member Functions

• RawCalorimeterHitImpl ()

Default constructor, initializes values to 0.

• virtual ∼RawCalorimeterHitImpl ()

Destructor.

• virtual int id () const

Returns an object id for internal (debugging) use in LCIO.

• virtual int getCellID0 () const

Returns the detector specific (geometrical) cell id.

• virtual int getCellID1 () const

Returns the second detector specific (geometrical) cell id.

• virtual int getAmplitude () const

Returns the amplitude of the hit in ADC counts.

• virtual int getTimeStamp () const

Returns a time stamp for the hit.

• void setCellID0 (int id0)

• void setCellID1 (int id1)

Sets the second cell id; Only store if the flag word (bit CHBIT_ID1) of the collection is set.

• void setAmplitude (int amplitude)

Sets the amplitude.

• void setTimeStamp (int time)

Set the time stamp.

Protected Attributes

• int _cellID0

• int _cellID1

• int _amplitude

• int _timeStamp

Additional Inherited Members

7.123.1 Detailed Description

Implementation of the real data RawCalorimeterHit.

Author gaede

Version

Id: RawCalorimeterHitImpl.h (p. ?? ),v 1.3 2006-08-04 16:52:46 gaede Exp

7.123.2 Member Function Documentation

7.123.2.1 int IMPL::RawCalorimeterHitImpl::getCellID1 ( ) const [virtual]

Returns the second detector specific (geometrical) cell id. Optional, check/set flag(LCIO::RCHBIT_ID1)==1. Implements EVENT::RawCalorimeterHit (p. 189).

7.123.2.2 int IMPL::RawCalorimeterHitImpl::getTimeStamp ( ) const [virtual]

Returns a time stamp for the hit. Optional, check/set flag(LCIO::RCHBIT_TIME)==1. Implements EVENT::RawCalorimeterHit (p. 189).

7.123.2.3 void IMPL::RawCalorimeterHitImpl::setCellID1 ( int id1 )

Sets the second cell id; Only store if the flag word (bit CHBIT_ID1) of the collection is set. Default is to store only cellid0. The documentation for this class was generated from the following files:

• RawCalorimeterHitImpl.h

• RawCalorimeterHitImpl.cc

7.124 IOIMPL::RawCalorimeterHitIOImpl Class Reference

Adding stuff needed for io (friend declarations, etc.)

#include <IOIMPL/RawCalorimeterHitIOImpl.h>

Inheritance diagram for IOIMPL::RawCalorimeterHitIOImpl:

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Friends

• class SIO::SIORawCalHitHandler

Additional Inherited Members

7.124.1 Detailed Description

Adding stuff needed for io (friend declarations, etc.)

Author gaede

Version

Id: RawCalorimeterHitIOImpl.h (p. ?? ),v 1.2 2005-04-15 08:37:40 gaede Exp

The documentation for this class was generated from the following file:

• RawCalorimeterHitIOImpl.h

7.125 EVENT::ReadOnlyException Class Reference

EventException (p. 69) used for signaling a ’read only exception’.

#include <Exceptions.h>

Inheritance diagram for EVENT::ReadOnlyException:

Public Member Functions

• ReadOnlyException (std::string text)

Additional Inherited Members 7.125.1 Detailed Description

EventException (p. 69) used for signaling a ’read only exception’.

Author gaede

Version Jun 5, 2003 The documentation for this class was generated from the following file:

• Exceptions.h

7.126 EVENT::ReconstructedParticle Class Reference

The LCIO (p. 115) reconstructedParticle.

#include <pre-generated/EVENT/ReconstructedParticle.h>

Inheritance diagram for EVENT::ReconstructedParticle:

Public Types

• typedef ReconstructedParticle lcobject_type

Useful typedef for template programming with LCIO (p. 115).

Public Member Functions

• virtual ∼ReconstructedParticle ()

Destructor.

• virtual int getType () const =0

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Type of reconstructed particle.

• virtual bool isCompound () const =0

Return particles primary flag.

• virtual const double ∗ getMomentum () const =0

The reconstructed particle’s 3-momentum - only float used in files,.

• virtual double getEnergy () const =0

Energy of the reconstructed particle - only float used in files.

• virtual const FloatVec & getCovMatrix () const =0

Covariance matrix of the reconstructed particle’s 4vector (10 parameters).

• virtual double getMass () const =0

Mass (p. 169) of the reconstructed particle, set independently from four vector quantities.

• virtual float getCharge () const =0

Charge of the reconstructed particle.

• virtual const float ∗ getReferencePoint () const =0

Reference point of the reconstructedParticle parameters.

• virtual const ParticleIDVec & getParticleIDs () const =0

The particle Id’s sorted by their likelihood.

• virtual ParticleID ∗ getParticleIDUsed () const =0

The particle Id used for the kinematics of this particle.

• virtual float getGoodnessOfPID () const =0

The overall goodness of the PID on a scale of [0;1].

• virtual const

ReconstructedParticleVec & getParticles () const =0

The reconstructed particles that have been combined to this particle.

• virtual const ClusterVec & getClusters () const =0

The weights of the reconstructed particles combined to this particle.

• virtual const TrackVec & getTracks () const =0

The weights of cluster contributions to this particle.

• virtual Vertex ∗ getStartVertex () const =0

The start vertex associated to this particle.

• virtual Vertex ∗ getEndVertex () const =0

The vertex where the particle decays This method actually returns the start vertex from the first daughter particle found.

• virtual void addParticleID (ParticleID ∗pid)=0

Set the type of reconstructed particle, one of: ReconstructedParticle::SINGLE, ReconstructedParticle::V0, ReconstructedParticle::COMPOUND, ReconstructedParticle::JET

• virtual void addParticle (ReconstructedParticle ∗particle)=0

Add a particle that has been used to create this particle.

• virtual void addCluster (Cluster ∗cluster)=0

Add a cluster that has been used to create this particle.

• virtual void addTrack (Track ∗track)=0

Add a track that has been used to create this particle.

Additional Inherited Members

7.126.1 Detailed Description

The LCIO (p. 115) reconstructedParticle.

Author gaede

Version

Id: ReconstructedParticle.aid,v 1.18 2006-09-21 06:10:35 gaede Exp

7.126.2 Member Function Documentation

7.126.2.1 virtual void EVENT::ReconstructedParticle::addParticleID ( ParticleID ∗ pid ) [pure virtual]

Set the type of reconstructed particle, one of: ReconstructedParticle::SINGLE, ReconstructedParticle::V0, ReconstructedParticle::COMPOUND, ReconstructedParticle::JET Set primary flag.Set particle momentum (px,py,pz).Set energy.Set the values of the covariance matrix of the particles four vector (px,py,pz,E).Set the values of the covariance matrix of the particles four vector (px,py,pz,E).Set mass - independently from four vector. Set the charge. Set the reference point, i.e. the point where the four vector is defined.Add a ParticleID (p. 181) object.

See Also ParticleID (p. 181)

Implemented in IMPL::ReconstructedParticleImpl (p. 199).

7.126.2.2 virtual const ClusterVec& EVENT::ReconstructedParticle::getClusters ( ) const [pure virtual]

The weights of the reconstructed particles combined to this particle. The clusters that have been used for this particle. Implemented in IMPL::ReconstructedParticleImpl (p. 198). Referenced by SIO::SIOReconstructedParticleHandler::write().

7.126.2.3 virtual const FloatVec& EVENT::ReconstructedParticle::getCovMatrix ( ) const [pure virtual]

Covariance matrix of the reconstructed particle’s 4vector (10 parameters). Stored as lower triangle matrix of the four momentum (px,py,pz,E), i.e. cov(px,px), cov(py,px), cov( py,py ) , .... Implemented in IMPL::ReconstructedParticleImpl (p. 197). Referenced by SIO::SIOReconstructedParticleHandler::write().

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.126.2.4 virtual double EVENT::ReconstructedParticle::getMass ( ) const [pure virtual]

Mass (p. 169) of the reconstructed particle, set independently from four vector quantities.

• only float used in files.

Implemented in IMPL::ReconstructedParticleImpl (p. 197). Referenced by SIO::SIOReconstructedParticleHandler::write().

7.126.2.5 virtual const ParticleIDVec& EVENT::ReconstructedParticle::getParticleIDs ( ) const [pure virtual]

The particle Id’s sorted by their likelihood.

See Also ParticleID (p. 181)

Implemented in IMPL::ReconstructedParticleImpl (p. 199). Referenced by UTIL::LCTOOLS::printReconstructedParticles(), and SIO::SIOReconstructedParticleHandler- ::write().

7.126.2.6 virtual ParticleID∗ EVENT::ReconstructedParticle::getParticleIDUsed ( ) const [pure virtual]

The particle Id used for the kinematics of this particle.

See Also ParticleID (p. 181)

Implemented in IMPL::ReconstructedParticleImpl (p. 200). Referenced by SIO::SIOReconstructedParticleHandler::write().

7.126.2.7 virtual const TrackVec& EVENT::ReconstructedParticle::getTracks ( ) const [pure virtual]

The weights of cluster contributions to this particle. The tracks that have been used for this particle. Implemented in IMPL::ReconstructedParticleImpl (p. 198). Referenced by SIO::SIOReconstructedParticleHandler::write().

7.126.2.8 virtual int EVENT::ReconstructedParticle::getType ( ) const [pure virtual]

Type of reconstructed particle. Check/set collection parameters ReconstructedParticleTypeNames and ReconstructedParticleTypeValues. Implemented in IMPL::ReconstructedParticleImpl (p. 200). Referenced by SIO::SIOReconstructedParticleHandler::write().

7.126.2.9 virtual bool EVENT::ReconstructedParticle::isCompound ( ) const [pure virtual]

Return particles primary flag.

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

All particles in the ReconstructedParticle (p. 193) collection should by definition return true. Compound partciles will return false.Convenient method - same as (getParticles() (p. 194).size() > 0 ). Implemented in IMPL::ReconstructedParticleImpl (p. 197). The documentation for this class was generated from the following file:

• ReconstructedParticle.h

7.127 IMPL::ReconstructedParticleImpl Class Reference

Implementation of ReconstructedParticle.

#include <IMPL/ReconstructedParticleImpl.h>

Inheritance diagram for IMPL::ReconstructedParticleImpl:

Public Member Functions

• ReconstructedParticleImpl ()

Default constructor, initializes values to 0.

• virtual ∼ReconstructedParticleImpl ()

Destructor.

• virtual int id () const

Returns an object id for internal (debugging) use in LCIO.

• virtual int getType () const

Type of reconstructed particle.

• virtual bool isCompound () const

Convenient method - same as (getParticles() (p. 198).size() > 0 ).

• virtual const double ∗ getMomentum () const

The magnitude of the reconstructed particle’s momentum,.

• virtual double getEnergy () const

Energy of the reconstructed particle.

• virtual const EVENT::FloatVec & getCovMatrix () const

Covariance matrix of the reconstructed particle’s 4vector (10 parameters).

• virtual double getMass () const

Mass (p. 169) of the reconstructed particle, set independently from four vector quantities.

• virtual float getCharge () const

Charge of the reconstructed particle.

• virtual const float ∗ getReferencePoint () const

Reference point of the reconstructedParticle parameters.

• virtual const

EVENT::ParticleIDVec & getParticleIDs () const

The particle Id’s sorted by their likelihood.

• virtual EVENT::ParticleID ∗ getParticleIDUsed () const

The particle Id used for the kinematics of this particle.

• virtual float getGoodnessOfPID () const

The overall goodness of the PID on a scale of [0;1].

• virtual const

EVENT::ReconstructedParticleVec & getParticles () const

The reconstructed particles that have been combined to this particle.

• virtual const EVENT::ClusterVec & getClusters () const

The clusters that have been used for this particle.

• virtual const EVENT::TrackVec & getTracks () const

The tracks that have been used for this particle.

• virtual EVENT::Vertex ∗ getStartVertex () const

The start vertex associated to this particle.

• virtual EVENT::Vertex ∗ getEndVertex () const

The vertex where the particle decays This method actually returns the start vertex from the first daughter particle found.

• void setType (int type)

• void setMomentum (const float ∗momentum)

• void setMomentum (const double ∗momentum)

• void setEnergy (float energy)

• void setCovMatrix (const float ∗cov)

• void setCovMatrix (const EVENT::FloatVec &)

• void setMass (float mass)

• void setCharge (float charge)

• void setReferencePoint (const float ∗reference)

• void addParticleID (EVENT::ParticleID ∗pid)

Set the type of reconstructed particle, one of:

ReconstructedParticle::SINGLE,

ReconstructedParticle::V0,

ReconstructedParticle::COMPOUND,

ReconstructedParticle::JET

• void setParticleIDUsed (EVENT::ParticleID ∗pid)

• void setGoodnessOfPID (float goodness)

• void addParticle (EVENT::ReconstructedParticle ∗particle)

Add a particle that has been used to create this particle.

• void addCluster (EVENT::Cluster ∗cluster)

Add a cluster that has been used to create this particle.

• void addTrack (EVENT::Track ∗track)

Add a track that has been used to create this particle.

• void setStartVertex (EVENT::Vertex ∗sv)

Protected Attributes

• int _type

• double _momentum [3]

• double _energy

• EVENT::FloatVec _cov

• double _mass

• float _charge

• float _reference [3] • EVENT::ParticleID ∗ _pidUsed

• float _goodnessOfPID

• EVENT::ParticleIDVec _pid

• EVENT::ReconstructedParticleVec _particles

• EVENT::ClusterVec _clusters

• EVENT::TrackVec _tracks

• EVENT::Vertex ∗ _sv

Additional Inherited Members

7.127.1 Detailed Description

Implementation of ReconstructedParticle.

See Also ReconstructedParticle

Author gaede

Version Mar 30, 2003

7.127.2 Member Function Documentation

7.127.2.1 void IMPL::ReconstructedParticleImpl::addParticleID ( EVENT::ParticleID ∗ pid ) [virtual]

Set the type of reconstructed particle, one of:

ReconstructedParticle::SINGLE,

ReconstructedParticle::V0,

ReconstructedParticle::COMPOUND,

ReconstructedParticle::JET

Set primary flag.Set particle momentum (px,py,pz).Set energy.Set the values of the covariance matrix of the particles four vector (px,py,pz,E).Set the values of the covariance matrix of the particles four vector (px,py,pz,E).Set mass - independently from four vector. Set the charge. Set the reference point, i.e. the point where the four vector is defined.Add a ParticleID object.

See Also ParticleID

Implements EVENT::ReconstructedParticle (p. 195).

Referenced by SIO::SIOReconstructedParticleHandler::read().

7.127.2.2 const EVENT::ParticleIDVec & IMPL::ReconstructedParticleImpl::getParticleIDs ( ) const [virtual]

The particle Id’s sorted by their likelihood.

See Also ParticleID

Implements EVENT::ReconstructedParticle (p. 196). Referenced by UTIL::PIDHandler::setParticleIDUsed().

7.127.2.3 EVENT::ParticleID ∗ IMPL::ReconstructedParticleImpl::getParticleIDUsed ( ) const [virtual]

The particle Id used for the kinematics of this particle.

See Also ParticleID

Implements EVENT::ReconstructedParticle (p. 196).

7.127.2.4 int IMPL::ReconstructedParticleImpl::getType ( ) const [virtual]

Type of reconstructed particle. Check/set collection parameterrs ReconstructedParticleTypeNames and ReconstructedParticleTypeValues. Implements EVENT::ReconstructedParticle (p. 196). The documentation for this class was generated from the following files:

• ReconstructedParticleImpl.h

• ReconstructedParticleImpl.cc

7.128 IOIMPL::ReconstructedParticleIOImpl Class Reference

Adding stuff needed for io (friend declarations, etc.)

#include <IOIMPL/ReconstructedParticleIOImpl.h>

Inheritance diagram for IOIMPL::ReconstructedParticleIOImpl:

Friends

• class SIO::SIOReconstructedParticleHandler

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.129 lcrtrel_helper::RelationManySide< U, T > Struct Template Reference  201

Additional Inherited Members

7.128.1 Detailed Description

Adding stuff needed for io (friend declarations, etc.)

Author gaede

Version Mar 31, 2004

The documentation for this class was generated from the following file:

• ReconstructedParticleIOImpl.h

7.129 lcrtrel_helper::RelationManySide< U, T > Struct Template Reference

Helper class for relations.

#include <LCRTRelations.h>

Inheritance diagram for lcrtrel_helper::RelationManySide< U, T >:

Additional Inherited Members

7.129.1 Detailed Description

template<class U, class T>struct lcrtrel_helper::RelationManySide< U, T >

Helper class for relations. The documentation for this struct was generated from the following file:

• LCRTRelations.h

7.130 lcrtrel_helper::RelationOneSide< U, T > Struct Template Reference

Helper class for relations.

#include <LCRTRelations.h>

Inheritance diagram for lcrtrel_helper::RelationOneSide< U, T >:

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Additional Inherited Members

7.130.1 Detailed Description

template<class U, class T>struct lcrtrel_helper::RelationOneSide< U, T >

Helper class for relations. The documentation for this struct was generated from the following file:

• LCRTRelations.h

7.131 SIO::RunEvent Struct Reference

Helper struct that stores run and event positions in the file.

#include <SIO/LCIORandomAccess.h>

Public Member Functions

• RunEvent (int run, int evt)

• RunEvent (long64 runEvt)

• operator long64 () const

Public Attributes

• int RunNum

• int EvtNum

7.131.1 Detailed Description

Helper struct that stores run and event positions in the file. The operator<() defines a lexicographical ordering in RunNum and EvtNum where all run records (EvtNum=-1) are orderd first. The documentation for this struct was generated from the following file:

• LCIORandomAccess.h

7.132 SIO::RunEventMap Class Reference

Map that holds positions of Run and Event records.

#include <SIO/RunEventMap.h>

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Public Types

• typedef std::map< RunEvent, long64 > MapType

• typedef MapType::iterator Map_IT

• typedef MapType::const_iterator Map_cIT

Public Member Functions

• void add (const RunEvent &re, long64 pos)

Add a new entry to the event map - if the RunEvent (p. 202) already exists the new position will be stored.

• long64 getPosition (const RunEvent &re)

Return the position of the specified Event record or Run record respectively (if EventNum == -1 ).

• long64 getPosition (long64 re)

Same as getPosition(const RunEvent& re ) (p. 204).

• RunEvent minRunEvent ()

• RunEvent maxRunEvent ()

• int getNumberOfEventRecords () const

• int getNumberOfRunRecords () const

• size_t size ()

• Map_IT begin ()

• Map_IT end ()

• Map_cIT begin () const

• Map_cIT end () const

• bool empty () const

• template<class InputIterator >

void insert (InputIterator b, InputIterator e)

• void clear ()

Static Public Attributes

• static const int NPos = -1

Protected Attributes

• MapType _map

• int _nRun

• int _nEvt

Friends

• std::ostream & operator<< (std::ostream &os, const RunEventMap &rm)

7.132.1 Detailed Description

Map that holds positions of Run and Event records.

Author gaede Version

Id: RunEventMap.h (p. ?? ),v 1.3 2011-03-03 16:00:12 gaede Exp

7.132.2 Member Function Documentation

7.132.2.1 long64 SIO::RunEventMap::getPosition ( const RunEvent & re ) [inline]

Return the position of the specified Event record or Run record respectively (if EventNum == -1 ). Returns RunEventMap::NPos if no record found. Referenced by SIO::LCIORandomAccessMgr::getPosition(). The documentation for this class was generated from the following files:

• RunEventMap.h

• RunEventMap.cc

7.133 RunEventProcessor Class Reference

Little tool that copies LCIO files on an event by event and run by run basis, thus fixing files that have incomplete records at the end due to a program crash. Inheritance diagram for RunEventProcessor:

Public Member Functions

• RunEventProcessor (const char ∗outFileName)

• void modifyEvent (LCEvent ∗evt)

• void processEvent (LCEvent ∗evt)

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

• void modifyRunHeader (LCRunHeader ∗run)

• void processRunHeader (LCRunHeader ∗run)

• RunEventProcessor (const char ∗outFileName, unsigned int nFiles)

• void modifyEvent (LCEvent ∗evt)

• void processEvent (LCEvent ∗evt)

• void modifyRunHeader (LCRunHeader ∗run)

• void processRunHeader (LCRunHeader ∗run)

• RunEventProcessor (const char ∗outFileName, int splitSize)

• void modifyEvent (LCEvent ∗evt)

• void processEvent (LCEvent ∗evt)

• void modifyRunHeader (LCRunHeader ∗run)

• void processRunHeader (LCRunHeader ∗run)

• void processEvent (LCEvent ∗evt)

• void modifyEvent (LCEvent ∗evt)

• void processRunHeader (LCRunHeader ∗run)

• void modifyRunHeader (LCRunHeader ∗run)

Protected Attributes

• LCWriter ∗ lcWrt

• int nEvent

• unsigned int _nFiles

7.133.1 Detailed Description

Little tool that copies LCIO files on an event by event and run by run basis, thus fixing files that have incomplete records at the end due to a program crash. Example of reading events from the file, add sth. Little tool that splits large LCIO files. lcio tool for merging files on an event by event and run by run basis to the event and write it to a new file. This uses the listener mechanism to read the input file (run headers and events) record by record. The RunEventProcessor (p. 204) class is defined for processing run and event records. This is our analysis module. For simplicity it is defined in the same file - in a real world application it should of course be defined in separate header and source files. The documentation for this class was generated from the following files:

• copyfix.cc

• lcio_merge_files.cc

• lcio_split_file.cc

• recjob.cc

7.134 EVENT::SimCalorimeterHit Class Reference

The generic simulated calorimeter hit.

#include <pre-generated/EVENT/SimCalorimeterHit.h>

Inheritance diagram for EVENT::SimCalorimeterHit:

Public Types

• typedef SimCalorimeterHit lcobject_type

Useful typedef for template programming with LCIO (p. 115).

Public Member Functions

• virtual ∼SimCalorimeterHit ()

Destructor.

• virtual int getCellID0 () const =0

Returns the detector specific (geometrical) cell id.

• virtual int getCellID1 () const =0

Returns the second detector specific (geometrical) cell id.

• virtual float getEnergy () const =0

Returns the energy of the hit in [GeV].

• virtual const float ∗ getPosition () const =0

Returns the position of the hit in world coordinates.

• virtual const float ∗ getStepPosition (int i) const =0

Returns the position where the energy deposited (step) occurred optional, only if bit LCIO::CHBIT_STEP is set.

• virtual int getNMCParticles () const =0

Returns the number of MC contributions to the hit.

• virtual int getNMCContributions () const =0

Returns the number of MC contributions to the hit.

• virtual float getEnergyCont (int i) const =0

Returns the energy in [GeV] of the i-th contribution to the hit.

• virtual float getTimeCont (int i) const =0

Returns the time of the i-th in [ns] contribution to the hit.

• virtual int getPDGCont (int i) const =0

Returns the PDG code of the shower particle that caused this contribution.

• virtual MCParticle ∗ getParticleCont (int i) const =0

Returns the MCParticle (p. 169) that caused the shower responsible for this contribution to the hit.

Additional Inherited Members

7.134.1 Detailed Description

The generic simulated calorimeter hit.

Author gaede

Version

Id: SimCalorimeterHit.aid,v 1.8 2006-03-24 13:25:52 gaede Exp

7.134.2 Member Function Documentation

7.134.2.1 virtual int EVENT::SimCalorimeterHit::getCellID1 ( ) const [pure virtual]

Returns the second detector specific (geometrical) cell id. 0 if information is not stored - check the flag word (bit CHBIT_ID1) of the collection. Default is to store only cellid0. Implemented in IMPL::SimCalorimeterHitImpl (p. 211). Referenced by SIO::SIOSimCalHitHandler::write().

7.134.2.2 virtual float EVENT::SimCalorimeterHit::getEnergyCont ( int i ) const [pure virtual]

Returns the energy in [GeV] of the i-th contribution to the hit.

See Also getNMCContributions() (p. 207)

Implemented in IMPL::SimCalorimeterHitImpl (p. 211). Referenced by SIO::SIOSimCalHitHandler::write().

7.134.2.3 virtual int EVENT::SimCalorimeterHit::getNMCContributions ( ) const [pure virtual]

Returns the number of MC contributions to the hit. 0 if information is not stored. There are two levels of detail: if collection flag bit LCIO.CHBIT_STEP==1 then all simulator steps’ contributions to the hit are stored, otherwise there is only one contribution for every particle entering the calorimeter. Implemented in IMPL::SimCalorimeterHitImpl (p. 211). Referenced by SIO::SIOSimCalHitHandler::write().

7.134.2.4 virtual int EVENT::SimCalorimeterHit::getNMCParticles ( ) const [pure virtual]

Returns the number of MC contributions to the hit. 0 if information is not stored. Renamed to getNMCContributions.

Deprecated See Also getNMCContributions() (p. 207)

Implemented in IMPL::SimCalorimeterHitImpl (p. 211).

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.134.2.5 virtual MCParticle∗ EVENT::SimCalorimeterHit::getParticleCont ( int i ) const [pure virtual]

Returns the MCParticle (p. 169) that caused the shower responsible for this contribution to the hit. This is the particle that flew into the calorimeter and not the shower particle that made the energy deposition.

See Also getNMCContributions() (p. 207)

Implemented in IMPL::SimCalorimeterHitImpl (p. 211). Referenced by SIO::SIOSimCalHitHandler::write().

7.134.2.6 virtual int EVENT::SimCalorimeterHit::getPDGCont ( int i ) const [pure virtual]

Returns the PDG code of the shower particle that caused this contribution. Check the flag word bit LCIO.CHBIT_STEP of the collection whether this information is available.

See Also getNMCContributions() (p. 207)

Implemented in IMPL::SimCalorimeterHitImpl (p. 212). Referenced by SIO::SIOSimCalHitHandler::write().

7.134.2.7 virtual const float∗ EVENT::SimCalorimeterHit::getPosition ( ) const [pure virtual]

Returns the position of the hit in world coordinates. NULL if information is not stored - check the flag word (bit CHBIT_LONG) of the collection. Implemented in IMPL::SimCalorimeterHitImpl (p. 212). Referenced by SIO::SIOSimCalHitHandler::write().

7.134.2.8 virtual float EVENT::SimCalorimeterHit::getTimeCont ( int i ) const [pure virtual]

Returns the time of the i-th in [ns] contribution to the hit.

See Also getNMCContributions() (p. 207)

Implemented in IMPL::SimCalorimeterHitImpl (p. 212). Referenced by SIO::SIOSimCalHitHandler::write(). The documentation for this class was generated from the following file:

• SimCalorimeterHit.h

7.135 IMPL::SimCalorimeterHitImpl Class Reference

Implementation of the generic SimCalorimeterHit.

#include <IMPL/SimCalorimeterHitImpl.h>

Inheritance diagram for IMPL::SimCalorimeterHitImpl:

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.135 IMPL::SimCalorimeterHitImpl Class Reference  209

Public Member Functions

• SimCalorimeterHitImpl ()

Default constructor, initializes values to 0.

• SimCalorimeterHitImpl (const EVENT::SimCalorimeterHit &hit)

Copy constructor - shallow copy.

• virtual ∼SimCalorimeterHitImpl ()

Destructor.

• virtual int id () const

Returns an object id for internal (debugging) use in LCIO.

• virtual int getCellID0 () const

Create a deep copy of itself.

• virtual int getCellID1 () const

Returns the second detector specific (geometrical) cell id.

• virtual float getEnergy () const

Returns the energy of the hit.

• virtual const float ∗ getPosition () const

Returns the position of the hit in world coordinates.

• virtual const float ∗ getStepPosition (int i) const

Returns the position where the energy deposited (step) occurred optional, only if bit LCIO::CHBIT_STEP is set.

• virtual int getNMCParticles () const

Returns the number of MC contributions to the hit.

• virtual int getNMCContributions () const

Returns the number of MC contributions to the hit.

• virtual float getEnergyCont (int i) const

Returns the energy in [GeV] of the i-th contribution to the hit.

• virtual float getTimeCont (int i) const

Returns the time of the i-th in [ns] contribution to the hit.

• virtual int getPDGCont (int i) const

Returns the PDG code of the shower particle that caused this contribution.

• virtual EVENT::MCParticle ∗ getParticleCont (int i) const

Returns the MCParticle that caused the shower responsible for this contribution to the hit.

• void setCellID0 (int id0)

Sets the first cell id;.

• void setCellID1 (int id1)

Sets the second cell id; Only store if the flag word (bit CHBIT_ID1) of the collection is set.

• void setEnergy (float en)

Sets the energy.

• void setPosition (const float pos[3])

Sets the position.

• void addMCParticleContribution (EVENT::MCParticle ∗p, float en, float t)

Adds an MCParticle contribution to the hit - contributions for the same MCParticle are combined into one, where the energy is accumulated and the time is the energy weighted mean time.

• void addMCParticleContribution (EVENT::MCParticle ∗p, float en, float t, int pdg, float ∗stepPos=0)

Adds a detailed MCParticle contribution to the hit.

Protected Attributes

• int _cellID0

• int _cellID1

• float _energy

• float _position [3]

• MCParticleContVec _vec

Additional Inherited Members

7.135.1 Detailed Description

Implementation of the generic SimCalorimeterHit.

Author gaede

Version Mar 5, 2003

7.135.2 Member Function Documentation

7.135.2.1 void IMPL::SimCalorimeterHitImpl::addMCParticleContribution ( EVENT::MCParticle ∗ p, float en, float t )

Adds an MCParticle contribution to the hit - contributions for the same MCParticle are combined into one, where the energy is accumulated and the time is the energy weighted mean time. Use this method for the standard LCIO mode, where one energy contribution per primary particle entering the calorimeter is stored.

7.135.2.2 void IMPL::SimCalorimeterHitImpl::addMCParticleContribution ( EVENT::MCParticle ∗ p, float en, float t, int pdg, float ∗ stepPos = 0 )

Adds a detailed MCParticle contribution to the hit. This method should be used for the detailed mode, where one MCParticleContribution is stored for every simulator step. If stepPos==0, (0,0,0) will be stored. NB: The flag word bit LCIO::CHBIT_STEP (or LCIO::CHBIT_PDG) has to be set, in order for the PDG and step position to be stored.

7.135.2.3 int IMPL::SimCalorimeterHitImpl::getCellID0 ( ) const [virtual]

Create a deep copy of itself. Overwrites return type in LCObject::clone(). Not yet - as AID genereates virtual inheritance and g++ doesn’t support virtual constructors with virtual inheritance TODO - change org.freehep.aid.CPPHeaderGenerator !Returns the detector specific (geometrical) cell id. Implements EVENT::SimCalorimeterHit (p. 206).

7.135.2.4 int IMPL::SimCalorimeterHitImpl::getCellID1 ( ) const [virtual]

Returns the second detector specific (geometrical) cell id. 0 if information is not stored - check the flag word (bit CHBIT_ID1) of the collection. Default is to store only cellid0. Implements EVENT::SimCalorimeterHit (p. 207).

7.135.2.5 float IMPL::SimCalorimeterHitImpl::getEnergyCont ( int i ) const [virtual]

Returns the energy in [GeV] of the i-th contribution to the hit.

See Also getNMCContributions() (p. 211)

Implements EVENT::SimCalorimeterHit (p. 207).

7.135.2.6 int IMPL::SimCalorimeterHitImpl::getNMCContributions ( ) const [virtual]

Returns the number of MC contributions to the hit. 0 if information is not stored. There are two levels of detail: if collection flag bit LCIO.CHBIT_STEP==1 then all simulator steps’ contributions to the hit are stored, otherwise there is only one contribution for every particle entering the calorimeter. Implements EVENT::SimCalorimeterHit (p. 207). Referenced by getNMCParticles().

7.135.2.7 int IMPL::SimCalorimeterHitImpl::getNMCParticles ( ) const [virtual]

Returns the number of MC contributions to the hit. 0 if information is not stored. Renamed to getNMCContributions.

Deprecated See Also getNMCContributions() (p. 211)

Implements EVENT::SimCalorimeterHit (p. 207). References getNMCContributions(), and UTIL::LCWarning::printWarning().

7.135.2.8 EVENT::MCParticle ∗ IMPL::SimCalorimeterHitImpl::getParticleCont ( int i ) const [virtual]

Returns the MCParticle that caused the shower responsible for this contribution to the hit. This is the particle that flew into the calorimeter and not the shower particle that made the energy deposition.

See Also getNMCContributions() (p. 211)

Implements EVENT::SimCalorimeterHit (p. 208).

7.135.2.9 int IMPL::SimCalorimeterHitImpl::getPDGCont ( int i ) const [virtual]

Returns the PDG code of the shower particle that caused this contribution. Check the flag word bit LCIO.CHBIT_STEP of the collection whether this information is available.

See Also getNMCContributions() (p. 211)

Implements EVENT::SimCalorimeterHit (p. 208).

7.135.2.10 const float ∗ IMPL::SimCalorimeterHitImpl::getPosition ( ) const [virtual]

Returns the position of the hit in world coordinates. NULL if information is not stored. Ask collection for flag, only available if bit LCIO.CHBIT_LONG is set. Implements EVENT::SimCalorimeterHit (p. 208).

7.135.2.11 float IMPL::SimCalorimeterHitImpl::getTimeCont ( int i ) const [virtual]

Returns the time of the i-th in [ns] contribution to the hit.

See Also getNMCContributions() (p. 211)

Implements EVENT::SimCalorimeterHit (p. 208).

7.135.2.12 void IMPL::SimCalorimeterHitImpl::setCellID1 ( int id1 )

Sets the second cell id; Only store if the flag word (bit CHBIT_ID1) of the collection is set. Default is to store only cellid0.

7.135.2.13 void IMPL::SimCalorimeterHitImpl::setPosition ( const float pos[3] )

Sets the position. Not stored if LCIO.CHBIT_LONG isn’t set. The documentation for this class was generated from the following files:

• SimCalorimeterHitImpl.h

• SimCalorimeterHitImpl.cc

7.136 IOIMPL::SimCalorimeterHitIOImpl Class Reference

Adding stuff needed for io (friend declarations, etc.)

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

#include <IOIMPL/SimCalorimeterHitIOImpl.h>

Inheritance diagram for IOIMPL::SimCalorimeterHitIOImpl:

Friends

• class SIO::SIOSimCalHitHandler

Additional Inherited Members 7.136.1 Detailed Description

Adding stuff needed for io (friend declarations, etc.)

Author gaede

Version Mar 7, 2003 The documentation for this class was generated from the following file:

• SimCalorimeterHitIOImpl.h

7.137 lcrtrel_helper::SimplePtrInit Struct Reference

Simple init function for simple pointers.

#include <LCRTRelations.h>

Static Public Member Functions

• static void ∗ init ()

7.137.1 Detailed Description

Simple init function for simple pointers. The documentation for this struct was generated from the following file:

• LCRTRelations.h

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.138 EVENT::SimTrackerHit Class Reference

A generic simulated tracker hit.

#include <pre-generated/EVENT/SimTrackerHit.h>

Inheritance diagram for EVENT::SimTrackerHit:

Public Types

• typedef SimTrackerHit lcobject_type

Useful typedef for template programming with LCIO (p. 115).

Public Member Functions

• virtual ∼SimTrackerHit ()

Destructor.

• virtual int getCellID0 () const =0

Same name as in CalorimeterHit (p. 49), even though there are no ’cells’ in this case.

• virtual int getCellID1 () const =0

Same name as in CalorimeterHit (p. 49), even though there are no ’cells’ in this case Optional, check/set flag(LCIO- ::THBIT_ID1)==1.

• virtual int getCellID () const =0

DEPRECATED: use ()

• virtual const double ∗ getPosition () const =0

Returns the hit position in [mm].

• virtual float getdEdx () const =0

Returns the dE/dx of the hit in [GeV].

• virtual float getEDep () const =0

The energy deposited on the hit [GeV].

• virtual float getTime () const =0

Returns the time of the hit in [ns].

• virtual MCParticle ∗ getMCParticle () const =0

Returns the MC particle that caused the hit.

• virtual const float ∗ getMomentum () const =0

Returns the 3-momentum of the particle at the hits position in [GeV] - optional, only if bit LCIO::THBIT_MOMENTUM is set.

• virtual float getPathLength () const =0

The path length of the particle in the sensitive material that resulted in this hit.

7.138 EVENT::SimTrackerHit Class Reference  215

Additional Inherited Members 7.138.1 Detailed Description

A generic simulated tracker hit.

Author gaede

Version

Id: SimTrackerHit.aid,v 1.11 2010-06-17 12:28:59 engels Exp

7.138.2 Member Function Documentation

7.138.2.1 virtual float EVENT::SimTrackerHit::getdEdx ( ) const [pure virtual] Returns the dE/dx of the hit in [GeV].

Deprecated See Also getEDep() (p. 214) Implemented in IMPL::SimTrackerHitImpl (p. 218).

7.138.2.2 virtual MCParticle∗ EVENT::SimTrackerHit::getMCParticle ( ) const [pure virtual] Returns the MC particle that caused the hit.

See Also MCParticle (p. 169) Implemented in IMPL::SimTrackerHitImpl (p. 218). Referenced by SIO::SIOSimTrackHitHandler::write().

7.138.2.3 virtual float EVENT::SimTrackerHit::getPathLength ( ) const [pure virtual] The path length of the particle in the sensitive material that resulted in this hit. This is only stored together with momentum, i.e. if LCIO::THBIT_MOMENTUM is set. Implemented in IMPL::SimTrackerHitImpl (p. 218). Referenced by SIO::SIOSimTrackHitHandler::write().

7.138.2.4 virtual float EVENT::SimTrackerHit::getTime ( ) const [pure virtual] Returns the time of the hit in [ns]. TO DO needs definition. Implemented in IMPL::SimTrackerHitImpl (p. 218).

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Referenced by SIO::SIOSimTrackHitHandler::write(). The documentation for this class was generated from the following file:

• SimTrackerHit.h

7.139 IMPL::SimTrackerHitImpl Class Reference

Implementation of SimTrackerHit.

#include <IMPL/SimTrackerHitImpl.h>

Inheritance diagram for IMPL::SimTrackerHitImpl:

Public Member Functions

• SimTrackerHitImpl ()

Default constructor, initializes values to 0.

• virtual ∼SimTrackerHitImpl ()

Destructor.

• virtual int id () const

Returns an object id for internal (debugging) use in LCIO.

• virtual int getCellID () const

DEPRECATED: use ()

• virtual int getCellID0 () const

Same name as in CalorimeterHit, even though there are no ’cells’ in this case.

• virtual int getCellID1 () const

Same name as in CalorimeterHit, even though there are no ’cells’ in this case 0 if information is not stored - check the flag word (bit THBIT_ID1) of the collection.

• virtual const double ∗ getPosition () const

Returns the hit position.

• virtual float getdEdx () const

Returns the dE/dx of the hit.

• virtual float getEDep () const

Returns the deposited energy of the hit [GeV].

• virtual float getTime () const

Returns the time of the hit.

• virtual EVENT::MCParticle ∗ getMCParticle () const

Returns the MC particle that caused the hit.

• virtual const float ∗ getMomentum () const

Returns the 3-momentum of the particle at the hits position in [GeV] - optional, only if bit LCIO::THBIT_MOMENTUM is set.

• virtual float getPathLength () const

The path length of the particle in the sensitive material that resulted in this hit.

• void setCellID (int id)

DEPRECATED: use .

• void setCellID0 (int id0)

Sets the first cell id;.

• void setCellID1 (int id1)

Sets the second cell id; Only store if the flag word (bit THBIT_ID1) of the collection is set.

• void setPosition (const double pos[3])

Sets the position.

• void setdEdx (float dEdX)

Sets dE/dx.

• void setEDep (float e)

Sets EDep.

• void setTime (float t)

Sets the time.

• void setMCParticle (EVENT::MCParticle ∗particle)

Sets the MCParticle that caused the hit.

• void setMomentum (const float p[3])

Sets the momentum of the particle at the hit’s position.

• void setMomentum (float px, float py, float pz)

Sets the momentum of the particle at the hit’s position.

• void setPathLength (float pathLength)

Protected Attributes

• int _cellID0

• int _cellID1

• double _pos [3]

• float _EDep

• float _time

• EVENT::MCParticle ∗ _particle

• float _p [3]

<– fixme: ROOT cannot handle this pointer !!????

• float _pathLength

Additional Inherited Members

7.139.1 Detailed Description

Implementation of SimTrackerHit.

See Also SimTrackerHit

Author gaede

Version Mar 6, 2003

218  Class Documentation

7.139.2 Member Function Documentation

7.139.2.1 int IMPL::SimTrackerHitImpl::getCellID1 ( ) const [virtual] Same name as in CalorimeterHit, even though there are no ’cells’ in this case 0 if information is not stored - check the flag word (bit THBIT_ID1) of the collection. Default is to store only cellid0. Implements EVENT::SimTrackerHit (p. 214).

7.139.2.2 float IMPL::SimTrackerHitImpl::getdEdx ( ) const [virtual] Returns the dE/dx of the hit.

Deprecated See Also getEDep() (p. 216) Implements EVENT::SimTrackerHit (p. 215). References getEDep(), and UTIL::LCWarning::printWarning().

7.139.2.3 EVENT::MCParticle ∗ IMPL::SimTrackerHitImpl::getMCParticle ( ) const [virtual] Returns the MC particle that caused the hit.

See Also MCParticle Implements EVENT::SimTrackerHit (p. 215).

7.139.2.4 float IMPL::SimTrackerHitImpl::getPathLength ( ) const [virtual] The path length of the particle in the sensitive material that resulted in this hit. This is only stored together with momentum, i.e. if LCIO::THBIT_MOMENTUM is set. Implements EVENT::SimTrackerHit (p. 215).

7.139.2.5 float IMPL::SimTrackerHitImpl::getTime ( ) const [virtual] Returns the time of the hit. TO DO needs definition. Implements EVENT::SimTrackerHit (p. 215).

7.139.2.6 void IMPL::SimTrackerHitImpl::setCellID1 ( int id1 ) Sets the second cell id; Only store if the flag word (bit THBIT_ID1) of the collection is set. Default is to store only cellid0.

7.139.2.7 void IMPL::SimTrackerHitImpl::setdEdx ( float dEdX ) Sets dE/dx.

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Deprecated See Also setEDep() (p. 217)

References UTIL::LCWarning::printWarning(), and setEDep(). The documentation for this class was generated from the following files:

• SimTrackerHitImpl.h

• SimTrackerHitImpl.cc

7.140 IOIMPL::SimTrackerHitIOImpl Class Reference

Adding stuff needed for io (friend declarations, etc.)

#include <IOIMPL/SimTrackerHitIOImpl.h>

Inheritance diagram for IOIMPL::SimTrackerHitIOImpl:

Friends

• class SIO::SIOSimTrackHitHandler

Additional Inherited Members

7.140.1 Detailed Description

Adding stuff needed for io (friend declarations, etc.)

Author gaede

Version Mar 12, 2003

The documentation for this class was generated from the following file:

• SimTrackerHitIOImpl.h

7.141 SIO::SIOCalHitHandler Class Reference

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of CalorimeterHits.

#include <SIO/SIOCalHitHandler.h>

Inheritance diagram for SIO::SIOCalHitHandler:

Public Member Functions

• virtual unsigned int read (SIO_stream ∗stream, EVENT::LCObject ∗∗objP)

Reads lcio calorimeter hit objects from an SIO (p. 30) stream.

• virtual unsigned int write (SIO_stream ∗stream, const EVENT::LCObject ∗obj)

Writes lcio calorimeter hit objects to an SIO (p. 30) stream.

Additional Inherited Members

7.141.1 Detailed Description

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of CalorimeterHits.

Author gaede

Version

Id: SIOCalHitHandler.h (p. ?? ),v 1.7 2005-04-15 08:37:42 gaede Exp

The documentation for this class was generated from the following files:

• SIOCalHitHandler.h

• SIOCalHitHandler.cc

7.142 SIO::SIOClusterHandler Class Reference

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of Clusters.

#include <SIO/SIOClusterHandler.h>

Inheritance diagram for SIO::SIOClusterHandler:

Public Member Functions

• virtual unsigned int read (SIO_stream ∗stream, EVENT::LCObject ∗∗objP)

Reads lcio clusterer hit objects from an SIO (p. 30) stream.

• virtual unsigned int write (SIO_stream ∗stream, const EVENT::LCObject ∗obj)

Writes lcio objects to an SIO (p. 30) stream.

Additional Inherited Members

7.142.1 Detailed Description

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of Clusters.

Author gaede

Version

Id: SIOClusterHandler.h (p. ?? ),v 1.4 2005-04-15 08:37:42 gaede Exp

The documentation for this class was generated from the following files:

• SIOClusterHandler.h

• SIOClusterHandler.cc

7.143 SIO::SIOCollectionHandler Class Reference

Handler for LCCollection/LCCollectionIOVec objects for SIO (p. 30).

#include <SIO/SIOCollectionHandler.h>

Inheritance diagram for SIO::SIOCollectionHandler:

Public Member Functions

• SIOCollectionHandler (const std::string &name, const std::string &type, IOIMPL::LCEventIOImpl ∗∗anEvt-

P=0) throw (EVENT::Exception)

The default constructor needs the name, the type and a pointer the address of the collection.

• const std::string & getTypeName () const

• virtual unsigned int xfer (SIO_stream ∗, SIO_operation, unsigned int)

• virtual unsigned int version ()

• void setCollection (const EVENT::LCCollection ∗col)

• void setEvent (IOIMPL::LCEventIOImpl ∗∗anEvtP)

7.143.1 Detailed Description

Handler for LCCollection/LCCollectionIOVec objects for SIO (p. 30).

Author gaede

Version

Id: SIOCollectionHandler.h (p. ?? ),v 1.8 2005-04-15 08:37:42 gaede Exp

7.143.2 Constructor & Destructor Documentation

7.143.2.1 SIO::SIOCollectionHandler::SIOCollectionHandler ( const std::string & name, const std::string & type, IOIMPL::LCEventIOImpl ∗∗ anEvtP = 0 ) throw EVENT::Exception)

The default constructor needs the name, the type and a pointer the address of the collection.

Throws an exception if handler for given type doesn’t exist.

Exceptions Exception

The documentation for this class was generated from the following files:

• SIOCollectionHandler.h

• SIOCollectionHandler.cc

7.144 SIO::SIOEventHandler Class Reference

Handler for LCEvent/LCEventIOImpl objects.

#include <SIO/SIOEventHandler.h>

Inheritance diagram for SIO::SIOEventHandler:

Public Member Functions

• SIOEventHandler (const std::string &name)

• SIOEventHandler (const std::string &name, IOIMPL::LCEventIOImpl ∗∗evtP)

• virtual unsigned int xfer (SIO_stream ∗, SIO_operation, unsigned int)

• virtual unsigned int version ()

• void setEvent (const EVENT::LCEvent ∗evt)

• void setEventPtr (IOIMPL::LCEventIOImpl ∗∗evtP)

• void setReadCollectionNames (const std::vector< std::string > &colnames)

7.144.1 Detailed Description

Handler for LCEvent/LCEventIOImpl objects.

Author gaede

Version

Id: SIOEventHandler.h (p. ?? ),v 1.8 2005-04-15 08:37:42 gaede Exp The documentation for this class was generated from the following files:

• SIOEventHandler.h

• SIOEventHandler.cc

7.145 SIO::SIOFloatVecHandler Class Reference

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of FloatVecs.

#include <SIO/SIOFloatVecHandler.h>

Inheritance diagram for SIO::SIOFloatVecHandler:

Public Member Functions

• virtual unsigned int read (SIO_stream ∗stream, EVENT::LCObject ∗∗objP)

Reads lcio FloatVecs objects from an SIO (p. 30) stream.

• virtual unsigned int write (SIO_stream ∗stream, const EVENT::LCObject ∗obj)

Writes lcio FloatVecs to an SIO (p. 30) stream.

Additional Inherited Members 7.145.1 Detailed Description

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of FloatVecs.

Author gaede Version

Id: SIOFloatVecHandler.h (p. ?? ),v 1.6 2005-04-15 08:37:42 gaede Exp

The documentation for this class was generated from the following files:

• SIOFloatVecHandler.h

• SIOFloatVecHandler.cc

7.146 SIO::SIOHandlerMgr Class Reference

Singleton that holds instances of all types of SIOObjectHandlers.

#include <SIO/SIOHandlerMgr.h>

Public Member Functions

• SIOObjectHandler ∗ getHandler (const std::string &type)

Returns the instance of the handler for the given type - null if it doesn’t exist.

• bool registerHandler (const std::string &type, SIOObjectHandler ∗handler)

Registers an instance of a handler for the given type.

Static Public Member Functions

• static SIOHandlerMgr ∗ instance ()

Returns the instance of this class.

7.146.1 Detailed Description

Singleton that holds instances of all types of SIOObjectHandlers. New types need to have their handler classes registered here.

Author gaede

Version

Id: SIOHandlerMgr.h (p. ?? ),v 1.5 2005-04-15 08:37:42 gaede Exp

7.146.2 Member Function Documentation

7.146.2.1 bool SIO::SIOHandlerMgr::registerHandler ( const std::string & type, SIOObjectHandler ∗ handler )

Registers an instance of a handler for the given type. Fails if type name allready exists. The documentation for this class was generated from the following files:

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

• SIOHandlerMgr.h

• SIOHandlerMgr.cc

7.147 SIO::SIOIndexHandler Class Reference

Handler for LCIOrandomAccess and LCIOIndex objects/blocks.

#include <SIO/SIOIndexHandler.h>

Inheritance diagram for SIO::SIOIndexHandler:

Public Member Functions

• SIOIndexHandler (const std::string &name, LCIORandomAccessMgr ∗raMgr)

C’tor.

• virtual unsigned int xfer (SIO_stream ∗, SIO_operation, unsigned int)

• virtual unsigned int version ()

7.147.1 Detailed Description

Handler for LCIOrandomAccess and LCIOIndex objects/blocks.

Author gaede

Version

Id: SIOIndexHandler.h (p. ?? ),v 1.2 2010-06-22 13:49:54 gaede Exp

The documentation for this class was generated from the following files:

• SIOIndexHandler.h

• SIOIndexHandler.cc

7.148 SIO::SIOIntVecHandler Class Reference

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of IntVecs.

#include <SIO/SIOIntVecHandler.h>

Inheritance diagram for SIO::SIOIntVecHandler:

Public Member Functions

• virtual unsigned int read (SIO_stream ∗stream, EVENT::LCObject ∗∗objP)

Reads IntVec objects from an SIO (p. 30) stream.

• virtual unsigned int write (SIO_stream ∗stream, const EVENT::LCObject ∗obj)

Writes lcio IntVec objects to an SIO (p. 30) stream.

Additional Inherited Members

7.148.1 Detailed Description

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of IntVecs.

Author gaede

Version

Id: SIOIntVecHandler.h (p. ?? ),v 1.4 2005-04-15 08:37:42 gaede Exp

The documentation for this class was generated from the following files:

• SIOIntVecHandler.h

• SIOIntVecHandler.cc

7.149 SIO::SIOLCGenericObjectHandler Class Reference

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of LCGenericObjects.

#include <SIO/SIOLCGenericObjectHandler.h>

Inheritance diagram for SIO::SIOLCGenericObjectHandler:

Public Member Functions

• virtual unsigned int read (SIO_stream ∗stream, EVENT::LCObject ∗∗objP)

Reads lcio LCGenericObject objects from an SIO (p. 30) stream.

• virtual unsigned int write (SIO_stream ∗stream, const EVENT::LCObject ∗obj)

Writes lcio LCGenericObject objects to an SIO (p. 30) stream.

• virtual unsigned int init (SIO_stream ∗stream, SIO_operation op, EVENT::LCCollection ∗col, unsigned int

vers)

Initialize the handler and/or the collection.

Protected Attributes

• int _nInt

• int _nFloat

• int _nDouble

• bool _isFixedSize

7.149.1 Detailed Description

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of LCGenericObjects.

Author gaede

Version

Id: SIOLCGenericObjectHandler.h (p. ?? ),v 1.3 2005-04-15 08:37:42 gaede Exp

7.149.2 Member Function Documentation

7.149.2.1 unsigned int SIO::SIOLCGenericObjectHandler::init ( SIO_stream ∗ stream, SIO_operation op, EVENT::LCCollection ∗ col, unsigned int vers ) [virtual]

Initialize the handler and/or the collection. Read/write the flag and the parameters. Overwrite for classes that need specific processing. Reimplemented from SIO::SIOObjectHandler (p. 231). References IMPL::LCFlagImpl::bitSet(), EVENT::LCCollection::getElementAt(), EVENT::LCCollection::getFlag(), EVENT::LCCollection::getNumberOfElements(), EVENT::LCCollection::getParameters(), EVENT::LCParameters- ::getStringVal(), EVENT::LCGenericObject::isFixedSize(), EVENT::LCCollection::parameters(), EVENT::LC-Collection::setFlag(), and EVENT::LCParameters::setValue(). The documentation for this class was generated from the following files:

• SIOLCGenericObjectHandler.h

• SIOLCGenericObjectHandler.cc

7.150 SIO::SIOLCParameters Class Reference

IO (p. 27) of LCParameters.

#include <SIO/SIOLCParameters.h>

Inheritance diagram for SIO::SIOLCParameters:

Static Public Member Functions

• static unsigned int read (SIO_stream ∗stream, LCParameters &params, unsigned int vers)

Reads objects from an SIO (p. 30) stream.

• static unsigned int write (SIO_stream ∗stream, const LCParameters &params)

Writes lcio objects to an SIO (p. 30) stream.

Additional Inherited Members

7.150.1 Detailed Description

IO (p. 27) of LCParameters.

Author gaede

Version

Id: SIOLCParameters.h (p. ?? ),v 1.4 2005-04-15 08:37:42 gaede Exp

The documentation for this class was generated from the following files:

• SIOLCParameters.h

• SIOLCParameters.cc

7.151 SIO::SIOLCRelationHandler Class Reference

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of CalorimeterHits.

#include <SIO/SIOLCRelationHandler.h>

Inheritance diagram for SIO::SIOLCRelationHandler:

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Public Member Functions

• virtual unsigned int read (SIO_stream ∗stream, EVENT::LCObject ∗∗objP)

Reads lcio relation objects from an SIO (p. 30) stream.

• virtual unsigned int write (SIO_stream ∗stream, const EVENT::LCObject ∗obj)

Writes lcio objects to an SIO (p. 30) stream.

Additional Inherited Members

7.151.1 Detailed Description

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of CalorimeterHits.

Author gaede

Version

Id: SIOLCRelationHandler.h (p. ?? ),v 1.2 2005-04-15 08:37:42 gaede Exp

The documentation for this class was generated from the following files:

• SIOLCRelationHandler.h

• SIOLCRelationHandler.cc

7.152 SIO::SIOObjectHandler Class Reference

Interface for all lcio object SIO-handlers, has to be implemented for all event entities (hits, tracks, clusters,...).

#include <SIO/SIOObjectHandler.h>

Inheritance diagram for SIO::SIOObjectHandler:

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Public Member Functions

• virtual unsigned int read (SIO_stream ∗stream, EVENT::LCObject ∗∗objP)=0

Reads lcio objects from an SIO (p. 30) stream.

• virtual unsigned int write (SIO_stream ∗stream, const EVENT::LCObject ∗obj)=0

Writes lcio objects to an SIO (p. 30) stream.

• virtual unsigned int init (SIO_stream ∗stream, SIO_operation op, EVENT::LCCollection ∗col, unsigned int

vers)

Initialize the handler and/or the collection.

• virtual unsigned int readBase (SIO_stream ∗stream, EVENT::LCObject ∗∗objP)

Calls read() (p. 230) if the collection is not a subset otherwise only reads the pointers.

• virtual unsigned int writeBase (SIO_stream ∗stream, const EVENT::LCObject ∗obj)

Calls write() (p. 230) if the collection is not a subset otherwise only writes the pointers.

Protected Attributes

• unsigned int _flag

• unsigned int _vers

7.152.1 Detailed Description

Interface for all lcio object SIO-handlers, has to be implemented for all event entities (hits, tracks, clusters,...).

Author gaede

Version

Id: SIOObjectHandler.h (p. ?? ),v 1.10 2008-11-27 16:53:35 engels Exp

7.152.2 Member Function Documentation

7.152.2.1 unsigned int SIO::SIOObjectHandler::init ( SIO_stream ∗ stream, SIO_operation op, EVENT::LCCollection ∗ col, unsigned int vers ) [virtual]

Initialize the handler and/or the collection. Read/write the flag and the parameters. Overwrite for classes that need specific processing. Reimplemented in SIO::SIOLCGenericObjectHandler (p. 227), and SIO::SIOVertexHandler (p. 253). References EVENT::LCCollection::getFlag(), EVENT::LCCollection::getParameters(), EVENT::LCCollection- ::parameters(), and EVENT::LCCollection::setFlag().

7.152.2.2 unsigned int SIO::SIOObjectHandler::readBase ( SIO_stream ∗ stream, EVENT::LCObject ∗∗ objP )

        [virtual]

Calls read() (p. 230) if the collection is not a subset otherwise only reads the pointers.

7.152.2.3 unsigned int SIO::SIOObjectHandler::writeBase ( SIO_stream ∗ stream, const EVENT::LCObject ∗ obj )

        [virtual]

Calls write() (p. 230) if the collection is not a subset otherwise only writes the pointers. The documentation for this class was generated from the following files:

• SIOObjectHandler.h

• SIOObjectHandler.cc

7.153 SIO::SIOParticleHandler Class Reference

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of MCParticles.

#include <SIO/SIOParticleHandler.h>

Inheritance diagram for SIO::SIOParticleHandler:

Public Member Functions

• virtual unsigned int read (SIO_stream ∗stream, EVENT::LCObject ∗∗objP)

Reads lcio MCParticle objects from an SIO (p. 30) stream.

• virtual unsigned int write (SIO_stream ∗stream, const EVENT::LCObject ∗obj)

Writes lcio MCParticle objects to an SIO (p. 30) stream.

Static Public Member Functions

• static void restoreParentDaughterRelations (EVENT::LCEvent ∗evt)

Additional Inherited Members

7.153.1 Detailed Description

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of MCParticles.

Author gaede

Version

Id: SIOParticleHandler.h (p. ?? ),v 1.8 2005-04-15 08:37:42 gaede Exp

The documentation for this class was generated from the following files:

• SIOParticleHandler.h

• SIOParticleHandler.cc

7.154 SIO::SIORandomAccessHandler Class Reference

Handler for LCIOrandomAccess and LCIOIndex objects/blocks.

#include <SIO/SIORandomAccessHandler.h>

Inheritance diagram for SIO::SIORandomAccessHandler:

Public Member Functions

• SIORandomAccessHandler (const std::string &name, LCIORandomAccessMgr ∗raMgr)

C’tor.

• virtual unsigned int xfer (SIO_stream ∗, SIO_operation, unsigned int)

• virtual unsigned int version ()

7.154.1 Detailed Description

Handler for LCIOrandomAccess and LCIOIndex objects/blocks.

Author gaede

Version

Id: SIORandomAccessHandler.h (p. ?? ),v 1.2 2010-06-22 13:49:54 gaede Exp The documentation for this class was generated from the following files:

• SIORandomAccessHandler.h

• SIORandomAccessHandler.cc

7.155 SIO::SIORawCalHitHandler Class Reference

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of RawCalorimeterHits.

#include <SIO/SIORawCalHitHandler.h>

Inheritance diagram for SIO::SIORawCalHitHandler:

Public Member Functions

• virtual unsigned int read (SIO_stream ∗stream, EVENT::LCObject ∗∗objP)

Reads lcio raw calorimeter hit objects from an SIO (p. 30) stream.

• virtual unsigned int write (SIO_stream ∗stream, const EVENT::LCObject ∗obj)

Writes lcio raw calorimeter hit objects to an SIO (p. 30) stream.

Additional Inherited Members 7.155.1 Detailed Description

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of RawCalorimeterHits.

Author gaede Version

Id: SIORawCalHitHandler.h (p. ?? ),v 1.2 2005-04-15 08:37:42 gaede Exp

The documentation for this class was generated from the following files:

• SIORawCalHitHandler.h

• SIORawCalHitHandler.cc

7.156 SIO::SIOReader Class Reference

Concrete implementation of LCWriter using SIO (p. 30).

#include <SIO/SIOReader.h>

Inheritance diagram for SIO::SIOReader:

Public Member Functions

• SIOReader (int lcReaderFlag=0)

Default constructor.

• virtual void open (const std::vector< std::string > &filenames) throw (IO::IOException, std::exception)

Opens a list of files for reading (read-only).

• virtual void open (const std::string &filename) throw (IO::IOException, std::exception)

Opens a file for reading (read-only).

• virtual EVENT::LCRunHeader ∗ readNextRunHeader () throw (IO::IOException, std::exception)

Reads the next run header from the file.

• virtual EVENT::LCRunHeader ∗ readNextRunHeader (int accessMode) throw (IO::IOException, std-

::exception)

Same as readNextRunHeader() (p. 239) but allows to set the access mode LCIO::READ_ONLY (default) or LCIO::- Update.

• virtual EVENT::LCEvent ∗ readNextEvent () throw (IO::IOException, std::exception)

Reads the next event from the file.

• virtual EVENT::LCEvent ∗ readNextEvent (int accessMode) throw (IO::IOException, std::exception)

Same as readNextRunHeader() (p. 239) but allows to set the access mode LCIO::READ_ONLY (default) or LCIO::- Update.

• virtual int getNumberOfEvents () throw (IO::IOException, std::exception )

Return the number of events in the file - the file has to be open.

• virtual int getNumberOfRuns () throw (IO::IOException, std::exception )

Return the number of runs (run headers) in the file - the file has to be open.

• virtual void getRuns (EVENT::IntVec &runs)

Return the run numbers of the runs (run headers) in the file - the file has to be open.

• virtual void getEvents (EVENT::IntVec &events)

Return the run and event numbers of the events in the file - the file has to be open.

• virtual void skipNEvents (int n) throw (IO::IOException, std::exception )

Skips the next n events from the current position.

• virtual void setReadCollectionNames (const std::vector< std::string > &colnames)

Limit the collection names that are going to be read to the subset given in the vector - all other collection will be ignored.

• virtual EVENT::LCRunHeader ∗ readRunHeader (int runNumber) throw (IO::IOException , std::exception)

Reads the specified runHeader from file.

• virtual EVENT::LCRunHeader ∗ readRunHeader (int runNumber, int accessMode) throw (IO::IOException ,

std::exception)

Same as LCEvent∗ readRunHeader(int runNumber) (p. 239) allowing to set the access mode LCIO::READ_ONLY (default) or LCIO::Update.

• virtual EVENT::LCEvent ∗ readEvent (int runNumber, int evtNumber) throw (IO::IOException, std::exception

)

Reads the specified event from file.

• virtual EVENT::LCEvent ∗ readEvent (int runNumber, int evtNumber, int accessMode) throw (IO::IO-

Exception, std::exception )

Same as LCEvent∗ readEvent(int runNumber, int evtNumber allowing to set the access mode LCIO::READ_ONLY (default) or LCIO::Update.

• virtual void close () throw (IO::IOException, std::exception)

Closes the output file/stream etc.

• virtual void registerLCEventListener (IO::LCEventListener ∗ls)

Registers a listener for reading LCEvents from a stream.

• virtual void removeLCEventListener (IO::LCEventListener ∗ls)

Remove a listener for reading LCEvents from a stream.

• virtual void registerLCRunListener (IO::LCRunListener ∗ls)

Registers a listener for reading LCEventsLCRunHeaders from a stream.

• virtual void removeLCRunListener (IO::LCRunListener ∗ls)

Remove a listener for reading LCRunHeaders from a stream.

• virtual void readStream () throw (IO::IOException, std::exception)

Reads the input stream and notifies registered listeners according to the object type found in the stream.

• virtual void readStream (int maxRecord) throw (IO::IOException, std::exception)

Reads maxRecord from the input stream and notifies registered listeners according to the object type found in the stream.

Protected Member Functions

• void setUpHandlers ()

• void readRecord () throw (IO::IOException , IO::EndOfDataException , std::exception)

• void postProcessEvent ()

• void getEventMap ()

• void recreateEventMap ()

Protected Attributes

• SIO_record ∗ _dummyRecord

• SIO_stream ∗ _stream

• SIORunHeaderHandler ∗ _runHandler

• SIOEventHandler ∗ _evtHandler

Additional Inherited Members

7.156.1 Detailed Description

Concrete implementation of LCWriter using SIO (p. 30).

Author gaede

Version

Id: SIOReader.h (p. ?? ),v 1.29 2011-03-03 16:00:12 gaede Exp

7.156.2 Member Function Documentation

7.156.2.1 void SIO::SIOReader::close ( ) throw IO::IOException, std::exception) [virtual]

Closes the output file/stream etc.

Exceptions IOException

Implements IO::LCReader (p. 132). References SIO::LCIORandomAccessMgr::clear().

7.156.2.2 void SIO::SIOReader::getEvents ( EVENT::IntVec & events ) [virtual]

Return the run and event numbers of the events in the file - the file has to be open. In case several input files are specified in the open() (p. 237) method - the run and event numbers of the events in the file that is currently open are returned. The size of the vector events will be twice the number of events, where i-th run number will be in events[2∗i] and the i-th event number in events[2∗i+]. Implements IO::LCReader (p. 133). References SIO::LCIORandomAccessMgr::getEventMap(), getNumberOfEvents(), and getNumberOfRuns().

7.156.2.3 int SIO::SIOReader::getNumberOfEvents ( ) throw IO::IOException, std::exception) [virtual]

Return the number of events in the file - the file has to be open. In case several input files are specified in the open() (p. 237) method - the number of events in the file that is currently open is returned. Implements IO::LCReader (p. 133). References SIO::LCIORandomAccessMgr::getEventMap(). Referenced by getEvents().

7.156.2.4 int SIO::SIOReader::getNumberOfRuns ( ) throw IO::IOException, std::exception) [virtual]

Return the number of runs (run headers) in the file - the file has to be open.

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

In case several input files are specified in the open() (p. 237) method - the number of runs (run headers) in the file that is currently open is returned. Implements IO::LCReader (p. 133). References SIO::LCIORandomAccessMgr::getEventMap(). Referenced by getEvents(), and getRuns().

7.156.2.5 void SIO::SIOReader::getRuns ( EVENT::IntVec & runs ) [virtual]

Return the run numbers of the runs (run headers) in the file - the file has to be open. In case several input files are specified in the open() (p. 237) method - the run numbers of the runs (run headers) in the file that is currently open is returned. Implements IO::LCReader (p. 133). References SIO::LCIORandomAccessMgr::getEventMap(), and getNumberOfRuns().

7.156.2.6 void SIO::SIOReader::open ( const std::vector< std::string > & filenames ) throw IO::IOException, std::exception) [virtual]

Opens a list of files for reading (read-only). All subsequent read operations will operate on the list, i.e. if an EOF is encountered the next file in the list will be opened and read transparently to the user.

Exceptions IOException

Implements IO::LCReader (p. 133).

7.156.2.7 void SIO::SIOReader::open ( const std::string & filename ) throw IO::IOException, std::exception)

        [virtual]

Opens a file for reading (read-only).

Exceptions IOException

Implements IO::LCReader (p. 133). References SIO::LCSIO::getValidSIOName().

7.156.2.8 EVENT::LCEvent ∗ SIO::SIOReader::readEvent ( int runNumber, int evtNumber ) throw IO::IOException, std::exception) [virtual]

Reads the specified event from file. Returns NULL if the specified event hasn’t been found in the file.

Exceptions IOException

Implements IO::LCReader (p. 134).

7.156.2.9 EVENT::LCEvent ∗ SIO::SIOReader::readEvent ( int runNumber, int evtNumber, int accessMode ) throw IO::IOException, std::exception) [virtual]

Same as LCEvent∗ readEvent(int runNumber, int evtNumber allowing to set the access mode LCIO::READ_ONLY (default) or LCIO::Update.

Exceptions IOException

Implements IO::LCReader (p. 134).

7.156.2.10 LCEvent ∗ SIO::SIOReader::readNextEvent ( ) throw IO::IOException, std::exception) [virtual]

Reads the next event from the file.

Exceptions IOException

Implements IO::LCReader (p. 134).

7.156.2.11 LCEvent ∗ SIO::SIOReader::readNextEvent ( int accessMode ) throw IO::IOException, std::exception)

        [virtual]

Same as readNextRunHeader() (p. 239) but allows to set the access mode LCIO::READ_ONLY (default) or LCIO- ::Update.

Exceptions IOException

Implements IO::LCReader (p. 134).

7.156.2.12 LCRunHeader ∗ SIO::SIOReader::readNextRunHeader ( ) throw IO::IOException, std::exception)

        [virtual]

Reads the next run header from the file.

Exceptions IOException

Implements IO::LCReader (p. 134).

7.156.2.13 LCRunHeader ∗ SIO::SIOReader::readNextRunHeader ( int accessMode ) throw IO::IOException, std::exception) [virtual]

Same as readNextRunHeader() (p. 239) but allows to set the access mode LCIO::READ_ONLY (default) or LCIO- ::Update.

Exceptions IOException

Implements IO::LCReader (p. 135).

7.156.2.14 EVENT::LCRunHeader ∗ SIO::SIOReader::readRunHeader ( int runNumber ) throw IO::IOException , std::exception) [virtual]

Reads the specified runHeader from file.

Returns NULL if the specified runHeader hasn’t been found in the file.

Exceptions IOException

Implements IO::LCReader (p. 135).

7.156.2.15 EVENT::LCRunHeader ∗ SIO::SIOReader::readRunHeader ( int runNumber, int accessMode ) throw IO::IOException , std::exception) [virtual]

Same as LCEvent∗ readRunHeader(int runNumber) (p. 239) allowing to set the access mode LCIO::READ_ONLY (default) or LCIO::Update.

Exceptions IOException

Implements IO::LCReader (p. 135).

7.156.2.16 void SIO::SIOReader::readStream ( ) throw IO::IOException, std::exception) [virtual]

Reads the input stream and notifies registered listeners according to the object type found in the stream. Exceptions

IOException EndOfException

Implements IO::LCReader (p. 135).

7.156.2.17 void SIO::SIOReader::readStream ( int maxRecord ) throw IO::IOException, std::exception) [virtual]

Reads maxRecord from the input stream and notifies registered listeners according to the object type found in the stream. Throws EndOfException if less than maxRecord records are found in the stream. Exceptions

IOException EndOfException

Implements IO::LCReader (p. 135).

7.156.2.18 void SIO::SIOReader::setReadCollectionNames ( const std::vector< std::string > & colnames ) [virtual]

Limit the collection names that are going to be read to the subset given in the vector - all other collection will be ignored. This might improve the reading performance considerably in cases where only a small subset of the collections in the event is needed. Implements IO::LCReader (p. 136).

7.156.2.19 void SIO::SIOReader::skipNEvents ( int n ) throw IO::IOException, std::exception) [virtual]

Skips the next n events from the current position. In fact simply reads the next n event headers so that the next event read is the (n+1)-th event. Implements IO::LCReader (p. 136). The documentation for this class was generated from the following files:

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

• SIOReader.h

• SIOReader.cc

7.157 SIO::SIOReconstructedParticleHandler Class Reference

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of ReconstructedParticles.

#include <SIO/SIOReconstructedParticleHandler.h>

Inheritance diagram for SIO::SIOReconstructedParticleHandler:

Public Member Functions

• virtual unsigned int read (SIO_stream ∗stream, EVENT::LCObject ∗∗objP)

Reads lcio ReconstructedParticle objects from an SIO (p. 30) stream.

• virtual unsigned int write (SIO_stream ∗stream, const EVENT::LCObject ∗obj)

Writes lcio ReconstructedParticle objects to an SIO (p. 30) stream.

Additional Inherited Members

7.157.1 Detailed Description

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of ReconstructedParticles.

Author gaede

Version

Id: SIOReconstructedParticleHandler.h (p. ?? ),v 1.4 2005-04-15 08:37:43 gaede Exp

The documentation for this class was generated from the following files:

• SIOReconstructedParticleHandler.h

• SIOReconstructedParticleHandler.cc

7.158 SIO::SIORecords Class Reference

Manager class that holds instances of all known LCIO-SIO records.

#include <SIO/LCSIO.h>

242  Class Documentation

Classes

• class Unpack

Helper class that sets the unpack flag of known SIO (p. 30) records as specified by the constructor.

Public Types

• enum ID { Event =0, Header, Run, Access, Index, NumberOfRecords }

Public Member Functions

• SIO_record ∗ operator[] (size_t idx)

• void setCompress (bool flag=true)

Protected Member Functions

• void add (SIO_record ∗&rec, const char ∗name)

Protected Attributes

• SIO_record ∗ _records [NumberOfRecords]

Friends

• class LCSIO

7.158.1 Detailed Description

Manager class that holds instances of all known LCIO-SIO records. Provides Unpack (p. 311) class for setting subsets of records to be unpacked. Every new record type needs to be added here in SIORecords::ID and SIORecords::Unpack (p. 311). The documentation for this class was generated from the following files:

• LCSIO.h

• LCSIO.cc

7.159 SIO::SIORunHeaderHandler Class Reference

Handler for LCRunHeader/LCRunHeaderImpl objects.

#include <SIO/SIORunHeaderHandler.h>

Inheritance diagram for SIO::SIORunHeaderHandler:

Public Member Functions

• SIORunHeaderHandler (const std::string &name)

C’tor for writing.

• SIORunHeaderHandler (const std::string &name, IOIMPL::LCRunHeaderIOImpl ∗∗rhP)

C’tor for reading.

• virtual unsigned int xfer (SIO_stream ∗, SIO_operation, unsigned int)

• virtual unsigned int version ()

• void setRunHeader (const EVENT::LCRunHeader ∗hdr)

• void setRunHeaderPtr (IOIMPL::LCRunHeaderIOImpl ∗∗hdrP)

7.159.1 Detailed Description

Handler for LCRunHeader/LCRunHeaderImpl objects.

Author gaede

Version

Id: SIORunHeaderHandler.h (p. ?? ),v 1.8 2005-04-15 08:37:43 gaede Exp

fg 20030609 using data interface for writing

The documentation for this class was generated from the following files:

• SIORunHeaderHandler.h

• SIORunHeaderHandler.cc

7.160 SIO::SIOSimCalHitHandler Class Reference

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of SimCalorimeterHits.

#include <SIO/SIOSimCalHitHandler.h>

Inheritance diagram for SIO::SIOSimCalHitHandler:

Public Member Functions

• virtual unsigned int read (SIO_stream ∗stream, EVENT::LCObject ∗∗objP)

Reads lcio calorimeter hit objects from an SIO (p. 30) stream.

• virtual unsigned int write (SIO_stream ∗stream, const EVENT::LCObject ∗obj)

Writes lcio objects to an SIO (p. 30) stream.

244  Class Documentation

Additional Inherited Members 7.160.1 Detailed Description

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of SimCalorimeterHits.

Author gaede

Version

Id: SIOSimCalHitHandler.h (p. ?? ),v 1.6 2005-04-15 08:37:43 gaede Exp The documentation for this class was generated from the following files:

• SIOSimCalHitHandler.h

• SIOSimCalHitHandler.cc

7.161 SIO::SIOSimTrackHitHandler Class Reference

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of SimTrackerHits.

#include <SIO/SIOSimTrackHitHandler.h>

Inheritance diagram for SIO::SIOSimTrackHitHandler:

Public Member Functions

• virtual unsigned int read (SIO_stream ∗stream, EVENT::LCObject ∗∗objP)

Reads lcio tracker hit objects from an SIO (p. 30) stream.

• virtual unsigned int write (SIO_stream ∗stream, const EVENT::LCObject ∗obj)

Writes lcio objects to an SIO (p. 30) stream.

Additional Inherited Members 7.161.1 Detailed Description

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of SimTrackerHits.

Author gaede Version

Id: SIOSimTrackHitHandler.h (p. ?? ),v 1.5 2005-04-15 08:37:43 gaede Exp

The documentation for this class was generated from the following files:

• SIOSimTrackHitHandler.h

• SIOSimTrackHitHandler.cc

7.162 SIO::SIOStrVecHandler Class Reference

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of StrVecs.

#include <SIO/SIOStrVecHandler.h>

Inheritance diagram for SIO::SIOStrVecHandler:

Public Member Functions

• virtual unsigned int read (SIO_stream ∗stream, EVENT::LCObject ∗∗objP)

Reads objects from an SIO (p. 30) stream.

• virtual unsigned int write (SIO_stream ∗stream, const EVENT::LCObject ∗obj)

Writes lcio objects to an SIO (p. 30) stream.

Additional Inherited Members

7.162.1 Detailed Description

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of StrVecs.

Author vogt

Version

Id: SIOStrVecHandler.h (p. ?? ),v 1.3 2005-04-15 08:37:43 gaede Exp

The documentation for this class was generated from the following files:

• SIOStrVecHandler.h

• SIOStrVecHandler.cc

7.163 SIO::SIOTPCHitHandler Class Reference

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TPCHits.

#include <SIO/SIOTPCHitHandler.h>

Inheritance diagram for SIO::SIOTPCHitHandler:

Public Member Functions

• virtual unsigned int read (SIO_stream ∗stream, EVENT::LCObject ∗∗objP)

Reads lcio TPC hit objects from an SIO (p. 30) stream.

• virtual unsigned int write (SIO_stream ∗stream, const EVENT::LCObject ∗obj)

Writes lcio TPC hit objects to an SIO (p. 30) stream.

Additional Inherited Members

7.163.1 Detailed Description

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TPCHits.

Author gaede

Version

Id: SIOTPCHitHandler.h (p. ?? ),v 1.4 2005-04-15 08:37:43 gaede Exp

The documentation for this class was generated from the following files:

• SIOTPCHitHandler.h

• SIOTPCHitHandler.cc

7.164 SIO::SIOTrackerDataHandler Class Reference

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TrackerDatas.

#include <SIO/SIOTrackerDataHandler.h>

Inheritance diagram for SIO::SIOTrackerDataHandler:

Public Member Functions

• virtual unsigned int read (SIO_stream ∗stream, EVENT::LCObject ∗∗objP)

Reads lcio TrackerData objects from an SIO (p. 30) stream.

• virtual unsigned int write (SIO_stream ∗stream, const EVENT::LCObject ∗obj)

Writes TrackerData hit objects to an SIO (p. 30) stream.

Additional Inherited Members

7.164.1 Detailed Description

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TrackerDatas.

Author gaede

Version

Id: SIOTrackerDataHandler.h (p. ?? ),v 1.1 2005-05-31 07:43:25 gaede Exp

The documentation for this class was generated from the following files:

• SIOTrackerDataHandler.h

• SIOTrackerDataHandler.cc

7.165 SIO::SIOTrackerHitHandler Class Reference

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of SimTrackerHits.

#include <SIO/SIOTrackerHitHandler.h>

Inheritance diagram for SIO::SIOTrackerHitHandler:

Public Member Functions

• virtual unsigned int read (SIO_stream ∗stream, EVENT::LCObject ∗∗objP)

Reads lcio tracker hit objects from an SIO (p. 30) stream.

• virtual unsigned int write (SIO_stream ∗stream, const EVENT::LCObject ∗obj)

Writes lcio tracker hit objects to an SIO (p. 30) stream.

248  Class Documentation

Additional Inherited Members 7.165.1 Detailed Description

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of SimTrackerHits.

Author gaede

Version

Id: SIOTrackerHitHandler.h (p. ?? ),v 1.3 2005-04-15 08:37:43 gaede Exp The documentation for this class was generated from the following files:

• SIOTrackerHitHandler.h

• SIOTrackerHitHandler.cc

7.166 SIO::SIOTrackerHitPlaneHandler Class Reference

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TrackerHitPlane.

#include <SIO/SIOTrackerHitPlaneHandler.h>

Inheritance diagram for SIO::SIOTrackerHitPlaneHandler:

Public Member Functions

• virtual unsigned int read (SIO_stream ∗stream, EVENT::LCObject ∗∗objP)

Reads lcio object from an SIO (p. 30) stream.

• virtual unsigned int write (SIO_stream ∗stream, const EVENT::LCObject ∗obj)

Writes lcio object to an SIO (p. 30) stream.

Additional Inherited Members 7.166.1 Detailed Description

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TrackerHitPlane.

Author engels Version $Id:$

The documentation for this class was generated from the following files:

• SIOTrackerHitPlaneHandler.h

• SIOTrackerHitPlaneHandler.cc

7.167 SIO::SIOTrackerHitZCylinderHandler Class Reference

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TrackerHitZCylinder.

#include <SIO/SIOTrackerHitZCylinderHandler.h>

Inheritance diagram for SIO::SIOTrackerHitZCylinderHandler:

Public Member Functions

• virtual unsigned int read (SIO_stream ∗stream, EVENT::LCObject ∗∗objP)

Reads lcio object from an SIO (p. 30) stream.

• virtual unsigned int write (SIO_stream ∗stream, const EVENT::LCObject ∗obj)

Writes lcio object to an SIO (p. 30) stream.

Additional Inherited Members

7.167.1 Detailed Description

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TrackerHitZCylinder.

Author engels

Version $Id:$

The documentation for this class was generated from the following files:

• SIOTrackerHitZCylinderHandler.h

• SIOTrackerHitZCylinderHandler.cc

7.168 SIO::SIOTrackerPulseHandler Class Reference

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TrackerPulses.

#include <SIO/SIOTrackerPulseHandler.h>

Inheritance diagram for SIO::SIOTrackerPulseHandler:

Public Member Functions

• virtual unsigned int read (SIO_stream ∗stream, EVENT::LCObject ∗∗objP)

Reads lcio TrackerPulse objects from an SIO (p. 30) stream.

• virtual unsigned int write (SIO_stream ∗stream, const EVENT::LCObject ∗obj)

Writes TrackerPulse hit objects to an SIO (p. 30) stream.

Additional Inherited Members

7.168.1 Detailed Description

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TrackerPulses.

Author gaede

Version

Id: SIOTrackerPulseHandler.h (p. ?? ),v 1.1 2005-05-31 07:43:25 gaede Exp

The documentation for this class was generated from the following files:

• SIOTrackerPulseHandler.h

• SIOTrackerPulseHandler.cc

7.169 SIO::SIOTrackerRawDataHandler Class Reference

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TrackerRawDatas.

#include <SIO/SIOTrackerRawDataHandler.h>

Inheritance diagram for SIO::SIOTrackerRawDataHandler:

Public Member Functions

• virtual unsigned int read (SIO_stream ∗stream, EVENT::LCObject ∗∗objP)

Reads lcio TrackerRawData objects from an SIO (p. 30) stream.

• virtual unsigned int write (SIO_stream ∗stream, const EVENT::LCObject ∗obj)

Writes TrackerRawData hit objects to an SIO (p. 30) stream.

Additional Inherited Members

7.169.1 Detailed Description

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of TrackerRawDatas.

Author gaede

Version

Id: SIOTrackerRawDataHandler.h (p. ?? ),v 1.1 2005-05-31 07:43:25 gaede Exp

The documentation for this class was generated from the following files:

• SIOTrackerRawDataHandler.h

• SIOTrackerRawDataHandler.cc

7.170 SIO::SIOTrackHandler Class Reference

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of Tracks.

#include <SIO/SIOTrackHandler.h>

Inheritance diagram for SIO::SIOTrackHandler:

Public Member Functions

• virtual unsigned int read (SIO_stream ∗stream, EVENT::LCObject ∗∗objP)

Reads lcio track objects from an SIO (p. 30) stream.

• virtual unsigned int write (SIO_stream ∗stream, const EVENT::LCObject ∗obj)

Writes lcio track objects to an SIO (p. 30) stream.

Additional Inherited Members

7.170.1 Detailed Description

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of Tracks.

Author gaede

Version

Id: SIOTrackHandler.h (p. ?? ),v 1.4 2005-04-15 08:37:43 gaede Exp

The documentation for this class was generated from the following files:

• SIOTrackHandler.h

• SIOTrackHandler.cc

7.171 SIO::SIOVertexHandler Class Reference

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of Vertexes.

#include <SIO/SIOVertexHandler.h>

Inheritance diagram for SIO::SIOVertexHandler:

Public Member Functions

• virtual unsigned int init (SIO_stream ∗stream, SIO_operation op, EVENT::LCCollection ∗col, unsigned int

vers)

Initialize the handler and/or the collection.

• virtual unsigned int read (SIO_stream ∗stream, EVENT::LCObject ∗∗objP)

Reads lcio Vertex objects from an SIO (p. 30) stream.

• virtual unsigned int write (SIO_stream ∗stream, const EVENT::LCObject ∗obj)

Writes lcio Vertex objects to an SIO (p. 30) stream.

Protected Attributes

• std::map< int, std::string > imr

• std::set< std::string > _set

• EVENT::StringVec parameters

7.171.1 Detailed Description

Implementation of SIOObjectHandler (p. 229) to handle IO (p. 27) of Vertexes.

Author gaede, engels

Version

Id: SIOVertexHandler.h (p. ?? ),v 1.3 2006-12-06 19:09:18 hvogt Exp

7.171.2 Member Function Documentation

7.171.2.1 unsigned int SIO::SIOVertexHandler::init ( SIO_stream ∗ stream, SIO_operation op, EVENT::LCCollection ∗ col, unsigned int vers ) [virtual]

Initialize the handler and/or the collection. Read/write the flag and the parameters. Overwrite for classes that need specific processing. Reimplemented from SIO::SIOObjectHandler (p. 231). References EVENT::Vertex::getAlgorithmType(), EVENT::LCCollection::getElementAt(), EVENT::LCCollection- ::getFlag(), EVENT::LCCollection::getNumberOfElements(), EVENT::LCCollection::getParameters(), EVENT::LC-Parameters::getStringVals(), EVENT::LCCollection::parameters(), EVENT::LCCollection::setFlag(), and EVENT::L-CParameters::setValues(). The documentation for this class was generated from the following files:

• SIOVertexHandler.h

• SIOVertexHandler.cc

7.172 SIO::SIOWriter Class Reference

Concrete implementation of LCWriter using SIO (p. 30).

#include <SIO/SIOWriter.h>

Inheritance diagram for SIO::SIOWriter:

Public Member Functions

• SIOWriter ()

Default constructor.

• virtual ∼SIOWriter ()

Destructor.

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

• virtual void open (const std::string &filename) throw (IO::IOException, std::exception)

Opens a file for writing.

• virtual void open (const std::string &filename, int writeMode) throw (IO::IOException, std::exception)

Opens a file for writing.

• virtual void setCompressionLevel (int level)

Set the compression level - needs to be called before open() (p. 256) otherwise call will have no effect.

• virtual void writeRunHeader (const EVENT::LCRunHeader ∗hdr) throw (IO::IOException, std::exception)

Writes the given run header to file.

• virtual void writeEvent (const EVENT::LCEvent ∗evt) throw (IO::IOException, std::exception)

Writes the given event to file.

• virtual void close () throw (IO::IOException, std::exception)

Closes the output file/stream etc.

• virtual void flush () throw (IO::IOException, std::exception)

Flushes the output file/stream etc.

Protected Member Functions

• void setUpHandlers (const EVENT::LCEvent ∗evt)

Sets up the handlers for writing the current event.

• void getSIOFileName (const std::string &filename, std::string &sioFilename)

Creates a proper filename with extension ’slcio’ in sioFilename.

Protected Attributes

• SIO_stream ∗ _stream

• int _compressionLevel

• LCIORandomAccessMgr _raMgr

Friends

• class SIOReader

• class SIOUnpack

7.172.1 Detailed Description

Concrete implementation of LCWriter using SIO (p. 30).

Use LCFactory to instantiate.

See Also LCFactory

Author gaede

Version Mar 6, 2003

7.172 SIO::SIOWriter Class Reference  255

7.172.2 Member Function Documentation

7.172.2.1 void SIO::SIOWriter::close ( ) throw IO::IOException, std::exception) [virtual] Closes the output file/stream etc.

Exceptions IOException

Implements IO::LCWriter (p. 167). References SIO::LCIORandomAccessMgr::clear(), and SIO::LCIORandomAccessMgr::writeRandomAccess-Records().

7.172.2.2 void SIO::SIOWriter::flush ( ) throw IO::IOException, std::exception) [virtual]

Flushes the output file/stream etc.

Exceptions IOException

Implements IO::LCWriter (p. 167).

7.172.2.3 void SIO::SIOWriter::open ( const std::string & filename ) throw IO::IOException, std::exception) [virtual]

Opens a file for writing. If file with given name exists, an exception is thrown. Use append or new mode instead.

Exceptions IOException

Implements IO::LCWriter (p. 167).

7.172.2.4 void SIO::SIOWriter::open ( const std::string & filename, int writeMode ) throw IO::IOException, std::exception)

        [virtual]

Opens a file for writing. Possible write modes are: LCIO::WRITE_NEW (existing files are replaced) and LCIO::WRITE_APPEND.

Exceptions IOException

Implements IO::LCWriter (p. 168). References SIO::LCSIO::getValidSIOName(), and SIO::LCSIO::seekStream().

7.172.2.5 void SIO::SIOWriter::setCompressionLevel ( int level ) [virtual]

Set the compression level - needs to be called before open() (p. 256) otherwise call will have no effect. If not called the Writer will use default compression. Valid compression levels are:

• level < 0 : default compression

• level == 0 : no compression

• level > 0 : 1 (fastest) - 9 (best compression)

Experimental code - don’t use for production.

7.173 SomeClass Struct Reference  257

Parameters level

compression level

Implements IO::LCWriter (p. 168).

7.172.2.6 void SIO::SIOWriter::setUpHandlers ( const EVENT::LCEvent ∗ evt ) [protected] Sets up the handlers for writing the current event. Creates Handlers needed for writing the event on this stream. Needs to be called for every event. References EVENT::LCEvent::getCollection(), EVENT::LCEvent::getCollectionNames(), EVENT::LCCollection- ::getTypeName(), and EVENT::LCCollection::isTransient().

7.172.2.7 void SIO::SIOWriter::writeEvent ( const EVENT::LCEvent ∗ evt ) throw IO::IOException, std::exception)

        [virtual]

Writes the given event to file. Exceptions IOException Implements IO::LCWriter (p. 168).

7.172.2.8 void SIO::SIOWriter::writeRunHeader ( const EVENT::LCRunHeader ∗ hdr ) throw IO::IOException,

std::exception) [virtual]

Writes the given run header to file. Exceptions IOException Implements IO::LCWriter (p. 168). The documentation for this class was generated from the following files:

• SIOWriter.h

• SIOWriter.cc

7.173 SomeClass Struct Reference

Classes

• struct Index

Public Attributes

• double someDouble

• int someInt

• bool someBool The documentation for this struct was generated from the following file:

• test_RTextension.cc

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.174 TEST Class Reference

Public Member Functions

• TEST (const std::string &tname, std::ostream &stream=std::cout)

• void LOG (const std::string &msg)

• template<class V1 , class V2 >

void operator() (const V1 &v1, const V2 &v2, const std::string name)

• void FAILED (const std::string &msg)

The documentation for this class was generated from the following file:

• tutil.h

7.175 lcrtrel_helper::ToRelation< U > Struct Template Reference

Helper class for relations.

#include <LCRTRelations.h>

7.175.1 Detailed Description

template<class U>struct lcrtrel_helper::ToRelation< U >

Helper class for relations. The documentation for this struct was generated from the following file:

• LCRTRelations.h

7.176 EVENT::TPCHit Class Reference

Deprecated.

#include <pre-generated/EVENT/TPCHit.h>

Inheritance diagram for EVENT::TPCHit:

Public Types

• typedef TPCHit lcobject_type

Useful typedef for template programming with LCIO (p. 115).

Public Member Functions

• virtual ∼TPCHit ()

Destructor.

• virtual int getCellID () const =0

Returns the detector specific cell id.

• virtual float getTime () const =0

Returns the time of the hit.

• virtual float getCharge () const =0

Returns the integrated charge of the hit.

• virtual int getQuality () const =0

Returns a quality flag for the hit.

• virtual int getNRawDataWords () const =0

Return the number of raw data (32-bit) words stored for the hit.

• virtual int getRawDataWord (int i) const =0

Return the raw data (32-bit) word at i.

Additional Inherited Members

7.176.1 Detailed Description

Deprecated. Please use TrackerRawData (p. 294), TrackerData (p. 267) and TrackerPulse (p. 289) instead. A future version of LCIO (p. 115) will not support writing of TPCHit (p. 258) collections anymore.

Deprecated Author gaede

Version

Id: TPCHit.aid,v 1.5 2006-03-24 13:25:52 gaede Exp

7.176.2 Member Function Documentation

7.176.2.1 virtual int EVENT::TPCHit::getNRawDataWords ( ) const [pure virtual]

Return the number of raw data (32-bit) words stored for the hit. Check the flag word (bit TPCBIT_RAW) of the collection if raw data is stored at all. Implemented in IMPL::TPCHitImpl (p. 261). Referenced by UTIL::LCTOOLS::printTPCHits(), and SIO::SIOTPCHitHandler::write().

7.176.2.2 virtual int EVENT::TPCHit::getRawDataWord ( int i ) const [pure virtual]

Return the raw data (32-bit) word at i. Check the flag word (bit TPCBIT_RAW) of the collection if raw data is stored at all. Implemented in IMPL::TPCHitImpl (p. 262). Referenced by UTIL::LCTOOLS::printTPCHits(), and SIO::SIOTPCHitHandler::write(). The documentation for this class was generated from the following file:

• TPCHit.h

7.177 IMPL::TPCHitImpl Class Reference

Implementation of the real data TPCHit.

#include <IMPL/TPCHitImpl.h>

Inheritance diagram for IMPL::TPCHitImpl:

Public Member Functions

• TPCHitImpl ()

Default Constructor - initializes all data to 0’s.

• virtual ∼TPCHitImpl ()

Destructor.

• virtual int id () const

Returns an object id for internal (debugging) use in LCIO.

• virtual int getCellID () const

Returns the detector specific cell id.

• virtual float getTime () const

Returns the time of the hit.

• virtual float getCharge () const

Returns the integrated charge of the hit.

• virtual int getQuality () const

Returns a quality flag for the hit.

• virtual int getNRawDataWords () const

Return the number of raw data (32-bit) words stored for the hit.

• virtual int getRawDataWord (int i) const

Return the raw data (32-bit) word at i.

• virtual void setCellID (int cellID)

Set the cell id.

• virtual void setTime (float time)

Set the time.

• virtual void setCharge (float charge)

Set the charge.

• virtual void setQuality (int quality)

Set the cell quality word.

• virtual void setRawData (const int ∗raw, int size)

Set the raw data.

Protected Member Functions

• virtual void initRawArray (int size)

initialize the raw data array - for faster reading

Protected Attributes

• int _cellID

• float _time

• float _charge

• int _quality

• int _rawSize

• int ∗ _rawArray

Additional Inherited Members

7.177.1 Detailed Description

Implementation of the real data TPCHit.

Author gaede

Version Sep 11, 2003

7.177.2 Member Function Documentation

7.177.2.1 int IMPL::TPCHitImpl::getNRawDataWords ( ) const [virtual]

Return the number of raw data (32-bit) words stored for the hit. Check the flag word (bit TPCBIT_RAW) of the collection if raw data is stored at all. Implements EVENT::TPCHit (p. 259).

7.177.2.2 int IMPL::TPCHitImpl::getRawDataWord ( int i ) const [virtual]

Return the raw data (32-bit) word at i. Check the flag word (bit TPCBIT_RAW) of the collection if raw data is stored at all. Implements EVENT::TPCHit (p. 260). The documentation for this class was generated from the following files:

• TPCHitImpl.h

• TPCHitImpl.cc

7.178 IOIMPL::TPCHitIOImpl Class Reference

Adding stuff needed for io (friend declarations, etc.)

#include <IOIMPL/TPCHitIOImpl.h>

Inheritance diagram for IOIMPL::TPCHitIOImpl:

Friends

• class SIO::SIOTPCHitHandler

Additional Inherited Members

7.178.1 Detailed Description

Adding stuff needed for io (friend declarations, etc.)

Author gaede

Version Sep 11, 2003

The documentation for this class was generated from the following file:

• TPCHitIOImpl.h

7.179 EVENT::Track Class Reference

The LCIO (p. 115) track class.

#include <pre-generated/EVENT/Track.h>

Inheritance diagram for EVENT::Track:

Public Types

• typedef Track lcobject_type

Useful typedef for template programming with LCIO (p. 115).

Public Member Functions

• virtual ∼Track ()

Destructor.

• virtual int getType () const =0

Flagword that defines the type of track.

• virtual float getD0 () const =0

Impact paramter of the track in (r-phi).

• virtual float getPhi () const =0

Phi of the track at the reference point.

• virtual float getOmega () const =0

Omega is the signed curvature of the track in [1/mm].

• virtual float getZ0 () const =0

Impact paramter of the track in (r-z).

• virtual float getTanLambda () const =0

Lambda is the dip angle of the track in r-z at the reference point.

• virtual const FloatVec & getCovMatrix () const =0

Covariance matrix of the track parameters.

• virtual const float ∗ getReferencePoint () const =0

Reference point of the track parameters.

• virtual float getChi2 () const =0

True if the reference point is the point of closest approach.

• virtual int getNdf () const =0

Number of degrees of freedom of the track fit.

• virtual float getdEdx () const =0

dEdx of the track.

• virtual float getdEdxError () const =0

Error of dEdx.

• virtual float getRadiusOfInnermostHit () const =0

The radius of the innermost hit that has been used in the track fit.

• virtual const IntVec & getSubdetectorHitNumbers () const =0

A vector that holds the number of hits in particular subdetectors.

• virtual const TrackVec & getTracks () const =0

The tracks that have been combined to this track.

• virtual const TrackStateVec & getTrackStates () const =0

Returns track states associtated to this track.

• virtual const TrackState ∗ getClosestTrackState (float x, float y, float z) const =0

Returns track state closest to the given point.

• virtual const TrackState ∗ getTrackState (int location) const =0

Returns track state for the given location - or NULL if not found.

• virtual const TrackerHitVec & getTrackerHits () const =0

Optionaly ( check/set flag(LCIO::TRBIT_HITS)==1) return the hits that have been used to create this track.

Additional Inherited Members

7.179.1 Detailed Description

The LCIO (p. 115) track class. The helix parameterization follows that of the L3 experiment at LEP. A detailed description of the track parameters is given in the LC-Note LC-DET-2006-004 "Track Parameters in LCIO" (pdf).

Author gaede

Version

Id: Track.aid,v 1.15 2007-09-18 09:47:09 gaede Exp

7.179.2 Member Function Documentation

7.179.2.1 virtual float EVENT::Track::getChi2 ( ) const [pure virtual]

True if the reference point is the point of closest approach. Chi^{∧}2 of the track fit. Implemented in IMPL::TrackImpl (p. 301). Referenced by SIO::SIOTrackHandler::write().

7.179.2.2 virtual const TrackState∗ EVENT::Track::getClosestTrackState ( float x, float y, float z ) const [pure

        virtual]

Returns track state closest to the given point.

7.179 EVENT::Track Class Reference  265

See Also TrackState (p. 304). Implemented in IMPL::TrackImpl (p. 301). 7.179.2.3 virtual const FloatVec& EVENT::Track::getCovMatrix ( ) const [pure virtual] Covariance matrix of the track parameters. Stored as lower triangle matrix where the order of parameters is: d0, phi, omega, z0, tan(lambda). So we have cov(d0,d0), cov( phi, d0 ), cov( phi, phi), ... Information is stored in the first TrackState (p. 304) of this Track (p. 263),

See Also TrackState (p. 304). Implemented in IMPL::TrackImpl (p. 301). 7.179.2.4 virtual float EVENT::Track::getD0 ( ) const [pure virtual] Impact paramter of the track in (r-phi). Information is stored in the first TrackState (p. 304) of this Track (p. 263),

See Also TrackState (p. 304). Implemented in IMPL::TrackImpl (p. 301). 7.179.2.5 virtual float EVENT::Track::getOmega ( ) const [pure virtual] Omega is the signed curvature of the track in [1/mm]. The sign is that of the particle’s charge. Information is stored in the first TrackState (p. 304) of this Track (p. 263),

See Also TrackState (p. 304). Implemented in IMPL::TrackImpl (p. 301). 7.179.2.6 virtual float EVENT::Track::getPhi ( ) const [pure virtual] Phi of the track at the reference point.

See Also getReferencePoint (p. 266) Information is stored in the first TrackState (p. 304) of this Track (p. 263),

TrackState (p. 304). Implemented in IMPL::TrackImpl (p. 302).

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.179.2.7 virtual float EVENT::Track::getRadiusOfInnermostHit ( ) const [pure virtual]

The radius of the innermost hit that has been used in the track fit. Helps to detect V0 tracks with small impact paramters or heavy mesons. To be used as convenient information or if hits are not stored in the data set, e.g. DST or FastMC. Implemented in IMPL::TrackImpl (p. 302). Referenced by SIO::SIOTrackHandler::write().

7.179.2.8 virtual const float∗ EVENT::Track::getReferencePoint ( ) const [pure virtual]

Reference point of the track parameters. The default for the reference point is the point of closest approach. Information is stored in the first TrackState (p. 304) of this Track (p. 263),

See Also TrackState (p. 304).

Implemented in IMPL::TrackImpl (p. 302).

7.179.2.9 virtual const IntVec& EVENT::Track::getSubdetectorHitNumbers ( ) const [pure virtual]

A vector that holds the number of hits in particular subdetectors. The mapping of indices to subdetectors is implementation dependent. To be used as convenient information or if hits are not stored in the data set, e.g. DST or FastMC. Check/set collection variable TrackSubdetectorNames for decoding the indices. Implemented in IMPL::TrackImpl (p. 302). Referenced by SIO::SIOTrackHandler::write().

7.179.2.10 virtual float EVENT::Track::getTanLambda ( ) const [pure virtual]

Lambda is the dip angle of the track in r-z at the reference point.

See Also getReferencePoint (p. 266) Information is stored in the first TrackState (p. 304) of this Track (p. 263), TrackState (p. 304).

Implemented in IMPL::TrackImpl (p. 302).

7.179.2.11 virtual const TrackState∗ EVENT::Track::getTrackState ( int location ) const [pure virtual]

Returns track state for the given location - or NULL if not found.

See Also TrackState (p. 304). location can be: TrackState::AtIP, TrackState::AtFirstHit, TrackState::AtLastHit, Track-State::AtCalorimeter, TrackState::AtVertex, TrackState::AtOther

Implemented in IMPL::TrackImpl (p. 303).

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.179.2.12 virtual const TrackStateVec& EVENT::Track::getTrackStates ( ) const [pure virtual]

Returns track states associtated to this track.

See Also TrackState (p. 304).

Implemented in IMPL::TrackImpl (p. 303). Referenced by SIO::SIOTrackHandler::write().

7.179.2.13 virtual int EVENT::Track::getType ( ) const [pure virtual]

Flagword that defines the type of track. Bits 0-15 can be used to denote the subdetectors that have contributed hits used in the track fit. For the definition of the bits check/Set the collection variables TrackTypeBitNames and TrackTypeBitIndices. Bits 16-31 are used internally. Implemented in IMPL::TrackImpl (p. 303). Referenced by SIO::SIOTrackHandler::write().

7.179.2.14 virtual float EVENT::Track::getZ0 ( ) const [pure virtual]

Impact paramter of the track in (r-z). Information is stored in the first TrackState (p. 304) of this Track (p. 263),

See Also TrackState (p. 304).

Implemented in IMPL::TrackImpl (p. 303). The documentation for this class was generated from the following file:

• Track.h

7.180 EVENT::TrackerData Class Reference

TrackerData (p. 267) contains the corrected (calibrated) raw tracker data.

#include <pre-generated/EVENT/TrackerData.h>

Inheritance diagram for EVENT::TrackerData:

Public Types

• typedef TrackerData lcobject_type

Useful typedef for template programming with LCIO (p. 115).

Public Member Functions

• virtual ∼TrackerData ()

Destructor.

• virtual int getCellID0 () const =0

Returns the first detector specific (geometrical) cell id.

• virtual int getCellID1 () const =0

Returns the second detector specific (geometrical) cell id.

• virtual float getTime () const =0

Returns a time measurement associated with the adc values, e.g.

• virtual const FloatVec & getChargeValues () const =0

The corrected (calibrated) FADC spectrum.

Additional Inherited Members

7.180.1 Detailed Description

TrackerData (p. 267) contains the corrected (calibrated) raw tracker data.

See Also TrackerRawData (p. 294)

TrackerPulse (p. 289)

Author gaede

Version

Id: TrackerData.aid,v 1.2 2006-03-24 13:25:53 gaede Exp

7.180.2 Member Function Documentation

7.180.2.1 virtual int EVENT::TrackerData::getCellID1 ( ) const [pure virtual]

Returns the second detector specific (geometrical) cell id. Optional, check/set flag(LCIO::TRAWBIT_ID1)==1. Implemented in IMPL::TrackerDataImpl (p. 270). Referenced by SIO::SIOTrackerDataHandler::write().

7.180.2.2 virtual float EVENT::TrackerData::getTime ( ) const [pure virtual]

Returns a time measurement associated with the adc values, e.g.

the t0 of the spectrum for the TPC. Subdetector dependent.

Implemented in IMPL::TrackerDataImpl (p. 269).

Referenced by SIO::SIOTrackerDataHandler::write().

The documentation for this class was generated from the following file:

• TrackerData.h

7.181 IMPL::TrackerDataImpl Class Reference

Default implementation of TrackerData.

#include <IMPL/TrackerDataImpl.h>

Inheritance diagram for IMPL::TrackerDataImpl:

Public Member Functions

• TrackerDataImpl ()

Default Constructor - initializes all data to 0’s.

• virtual ∼TrackerDataImpl ()

Destructor.

• virtual int id () const

Returns an object id for internal (debugging) use in LCIO.

• virtual int getCellID0 () const

Returns the first detector specific (geometrical) cell id.

• virtual int getCellID1 () const

Returns the second detector specific (geometrical) cell id.

• virtual float getTime () const

Returns the time.

• virtual const EVENT::FloatVec & getChargeValues () const

The calibrated ADC values.

• void setCellID0 (int cellID0)

• void setCellID1 (int cellID1)

• void setTime (float time)

• void setChargeValues (const EVENT::FloatVec &charge)

Set the charge vector by copying the values.

• EVENT::FloatVec & chargeValues ()

Allows direct access to the charge vector.

Protected Attributes

• int _cellID0

• int _cellID1

• float _time

• EVENT::FloatVec _charge

Additional Inherited Members

7.181.1 Detailed Description

Default implementation of TrackerData.

Author gaede

Version

Id: TrackerDataImpl.h (p. ?? ),v 1.2 2006-08-04 16:52:46 gaede Exp

7.181.2 Member Function Documentation

7.181.2.1 FloatVec & IMPL::TrackerDataImpl::chargeValues ( )

Allows direct access to the charge vector.

Can be used e.g. to set the values: raw->chargeValues() (p. 270).clear() ; raw->chargeValues() (p. 270).resize( n ) ; raw->chargeValues() (p. 270)[i] = 42. ;

7.181.2.2 virtual int IMPL::TrackerDataImpl::getCellID1 ( ) const [inline], [virtual]

Returns the second detector specific (geometrical) cell id.

Optional, check/set flag(LCIO::TRAWBIT_ID1)==1.

Implements EVENT::TrackerData (p. 268).

The documentation for this class was generated from the following files:

• TrackerDataImpl.h

• TrackerDataImpl.cc

7.182 IOIMPL::TrackerDataIOImpl Class Reference

Adding stuff needed for io (friend declarations, etc.)

#include <IOIMPL/TrackerDataIOImpl.h>

Inheritance diagram for IOIMPL::TrackerDataIOImpl:

Friends

• class SIO::SIOTrackerDataHandler

Additional Inherited Members

7.182.1 Detailed Description

Adding stuff needed for io (friend declarations, etc.)

Author gaede

Version Sep 11, 2003

The documentation for this class was generated from the following file:

• TrackerDataIOImpl.h

7.183 EVENT::TrackerHit Class Reference

A generic tracker hit to be used by pattern recognition.

#include <pre-generated/EVENT/TrackerHit.h>

Inheritance diagram for EVENT::TrackerHit:

Public Types

• typedef TrackerHit lcobject_type

Useful typedef for template programming with LCIO (p. 115).

Public Member Functions

• virtual ∼TrackerHit ()

Destructor.

• virtual int getCellID0 () const =0

Same name as in CalorimeterHit (p. 49), even though there are no ’cells’ in this case.

• virtual int getCellID1 () const =0

Same name as in CalorimeterHit (p. 49), even though there are no ’cells’ in this case Optional, check/set flag(LCIO- ::RTHBIT_ID1)==1.

• virtual const double ∗ getPosition () const =0

The hit position in [mm].

• virtual const FloatVec & getCovMatrix () const =0

Covariance of the position (x,y,z), stored as lower triangle matrix.

• virtual float getdEdx () const =0

The dE/dx of the hit in [GeV].

• virtual float getTime () const =0

The time of the hit in [ns].

• virtual int getType () const =0

Type of raw data hit, either one of

LCIO::TPCHIT

LCIO::SIMTRACKERHIT

• virtual int getQuality () const =0

The quality bit flag of the hit.

• virtual float getEDep () const =0

The energy deposited on the hit [GeV].

• virtual float getEDepError () const =0

The error measured on EDep [GeV].

• virtual const LCObjectVec & getRawHits () const =0

The raw data hits.

Additional Inherited Members

7.183.1 Detailed Description

A generic tracker hit to be used by pattern recognition.

Author gaede

Version

Id: TrackerHit.aid,v 1.16 2010-06-17 12:28:59 engels Exp

7.183.2 Member Function Documentation

7.183.2.1 virtual const FloatVec& EVENT::TrackerHit::getCovMatrix ( ) const [pure virtual]

Covariance of the position (x,y,z), stored as lower triangle matrix. i.e. cov(x,x) , cov(y,x) , cov(y,y) , cov(z,x) , cov(z,y) , cov(z,z). Implemented in IMPL::TrackerHitPlaneImpl (p. 280), IMPL::TrackerHitZCylinderImpl (p. 286), and IMPL::- TrackerHitImpl (p. 274). Referenced by SIO::SIOTrackerHitHandler::write().

7.183.2.2 virtual float EVENT::TrackerHit::getdEdx ( ) const [pure virtual]

The dE/dx of the hit in [GeV].

Deprecated See Also getEDep() (p. 272)

Implemented in IMPL::TrackerHitPlaneImpl (p. 282), IMPL::TrackerHitZCylinderImpl (p. 287), and IMPL::- TrackerHitImpl (p. 276).

7.183.2.3 virtual const LCObjectVec& EVENT::TrackerHit::getRawHits ( ) const [pure virtual]

The raw data hits. Check getType() (p. 273) to get actual data type. Implemented in IMPL::TrackerHitPlaneImpl (p. 282), IMPL::TrackerHitZCylinderImpl (p. 288), and IMPL::- TrackerHitImpl (p. 276). Referenced by SIO::SIOTrackerHitHandler::write(), SIO::SIOTrackerHitZCylinderHandler::write(), and SIO::SIO-TrackerHitPlaneHandler::write().

7.183.2.4 virtual int EVENT::TrackerHit::getType ( ) const [pure virtual]

Type of raw data hit, either one of LCIO::TPCHIT LCIO::SIMTRACKERHIT Type of hit. Mapping of integer types to type names through collection parameters "TrackerHitTypeNames" and "TrackerHitTypeValues". Implemented in IMPL::TrackerHitPlaneImpl (p. 282), IMPL::TrackerHitZCylinderImpl (p. 288), and IMPL::- TrackerHitImpl (p. 276). Referenced by SIO::SIOTrackerHitHandler::write(), SIO::SIOTrackerHitZCylinderHandler::write(), and SIO::SIO-TrackerHitPlaneHandler::write(). The documentation for this class was generated from the following file:

• TrackerHit.h

7.184 IMPL::TrackerHitImpl Class Reference

Implementation of the generic tracker hit.

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

#include <IMPL/TrackerHitImpl.h>

Inheritance diagram for IMPL::TrackerHitImpl:

Public Member Functions

• virtual ∼TrackerHitImpl ()

Destructor.

• virtual int id () const

Returns an object id for internal (debugging) use in LCIO.

• virtual int getCellID0 () const

Same name as in CalorimeterHit, even though there are no ’cells’ in this case.

• virtual int getCellID1 () const

Same name as in CalorimeterHit, even though there are no ’cells’ in this case 0 if information is not stored - check the flag word (bit RTHBIT_ID1) of the collection.

• virtual const double ∗ getPosition () const

The hit position in [mm].

• virtual const EVENT::FloatVec & getCovMatrix () const

Covariance of the position (x,y,z)

• virtual float getdEdx () const

The dE/dx of the hit in [GeV/mm].

• virtual float getEDep () const

The deposited energy of the hit [GeV].

• virtual float getEDepError () const

The error measured on EDep [GeV].

• virtual float getTime () const

The time of the hit in [ns].

• virtual int getType () const

Type of raw data hit, either one of LCIO::TPCHIT LCIO::SIMTRACKERHIT

• virtual int getQuality () const

The quality bit flag of the hit.

• virtual const EVENT::LCObjectVec & getRawHits () const

The raw data hits.

• virtual EVENT::LCObjectVec & rawHits ()

Use to manipulate the raw hits.

• void setCellID0 (int id0)

Sets the first cell id;.

• void setCellID1 (int id1)

Sets the second cell id; Only store if the flag word (bit RTHBIT_ID1) of the collection is set.

• void setType (int type)

• void setPosition (const double pos[3])

• void setCovMatrix (const EVENT::FloatVec &cov)

• void setCovMatrix (const float cov[TRKHITNCOVMATRIX])

• void setdEdx (float dedx)

• void setEDep (float e)

• void setEDepError (float e)

• void setTime (float t)

• void setQuality (int quality)

• void setQualityBit (int bit, bool val=true)

Protected Attributes

• int _cellID0

• int _cellID1

• int _type

• double _pos [3]

• EVENT::FloatVec _cov

• float _EDep

• float _EDepError

• float _time

• int _quality

• EVENT::LCObjectVec _rawHits

Additional Inherited Members

7.184.1 Detailed Description

Implementation of the generic tracker hit.

Author gaede

Version

Id: TrackerHitImpl.h (p. ?? ),v 1.15 2010-06-17 12:29:00 engels Exp

7.184.2 Member Function Documentation

7.184.2.1 int IMPL::TrackerHitImpl::getCellID1 ( ) const [virtual]

Same name as in CalorimeterHit, even though there are no ’cells’ in this case 0 if information is not stored - check the flag word (bit RTHBIT_ID1) of the collection. Default is to store only cellid0. Implements EVENT::TrackerHit (p. 272).

7.184.2.2 float IMPL::TrackerHitImpl::getdEdx ( ) const [virtual]

The dE/dx of the hit in [GeV/mm].

Deprecated See Also getEDep() (p. 274)

Implements EVENT::TrackerHit (p. 273). References UTIL::LCWarning::printWarning().

7.184.2.3 const EVENT::LCObjectVec & IMPL::TrackerHitImpl::getRawHits ( ) const [virtual]

The raw data hits. Check getType() (p. 276) to get actual data type. Implements EVENT::TrackerHit (p. 273).

7.184.2.4 float IMPL::TrackerHitImpl::getTime ( ) const [virtual]

The time of the hit in [ns]. Is this needed ? Implements EVENT::TrackerHit (p. 272).

7.184.2.5 int IMPL::TrackerHitImpl::getType ( ) const [virtual]

Type of raw data hit, either one of LCIO::TPCHIT LCIO::SIMTRACKERHIT Type of hit. Mapping of integer types to type names through collection parameters "TrackerHitTypeNames" and "TrackerHitTypeValues". Implements EVENT::TrackerHit (p. 273).

7.184.2.6 void IMPL::TrackerHitImpl::setCellID1 ( int id1 )

Sets the second cell id; Only store if the flag word (bit RTHBIT_ID1) of the collection is set. Default is to store only cellid0. The documentation for this class was generated from the following files:

• TrackerHitImpl.h

• TrackerHitImpl.cc

7.185 IOIMPL::TrackerHitIOImpl Class Reference

Adding stuff needed for io (friend declarations, etc.)

#include <IOIMPL/TrackerHitIOImpl.h>

Inheritance diagram for IOIMPL::TrackerHitIOImpl:

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Friends

• class SIO::SIOTrackerHitHandler

Additional Inherited Members

7.185.1 Detailed Description

Adding stuff needed for io (friend declarations, etc.)

Author gaede

Version Mar 12, 2003

The documentation for this class was generated from the following file:

• TrackerHitIOImpl.h

7.186 EVENT::TrackerHitPlane Class Reference

A tracker hit on a planar surface - the hit is stored as: x,y,z, u(theta, phi), v(theta,phi), du, dv where (x,y,z) is the position in global cartesian coordinates, u,v are the the meassurement directions (unit vectors, spanning the plane) and du,dv are the measurement errors along these directions.

#include <pre-generated/EVENT/TrackerHitPlane.h>

Inheritance diagram for EVENT::TrackerHitPlane:

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Public Types

• typedef TrackerHitPlane lcobject_type

Useful typedef for template programming with LCIO (p. 115).

Public Member Functions

• virtual ∼TrackerHitPlane ()

Destructor.

• virtual int getCellID0 () const =0

Same name as in CalorimeterHit (p. 49), even though there are no ’cells’ in this case.

• virtual int getCellID1 () const =0

Same name as in CalorimeterHit (p. 49), even though there are no ’cells’ in this case Optional, check/set flag(LCIO- ::RTHPBIT_ID1)==1.

• virtual const float ∗ getU () const =0

Direction of first measurement - given as (theta, phi).

• virtual const float ∗ getV () const =0

Direction of second measurement - given as (theta, phi).

• virtual float getdU () const =0

Measurement error along u.

• virtual float getdV () const =0

Measurement error along v.

Additional Inherited Members

7.186.1 Detailed Description

A tracker hit on a planar surface - the hit is stored as: x,y,z, u(theta, phi), v(theta,phi), du, dv where (x,y,z) is the position in global cartesian coordinates, u,v are the the meassurement directions (unit vectors, spanning the plane) and du,dv are the measurement errors along these directions. The order of u and v should follow the global cartesian coordinates x, y and z, e.g. in case of a plane parallel to the z-axis, u lies in the x-y plane and v is along z.

Author gaede

Version $Id:$

7.186.2 Member Function Documentation

7.186.2.1 virtual const float∗ EVENT::TrackerHitPlane::getU ( ) const [pure virtual]

Direction of first measurement - given as (theta, phi). Defines spanning vector of measurement plane. Implemented in IMPL::TrackerHitPlaneImpl (p. 282). Referenced by SIO::SIOTrackerHitPlaneHandler::write().

7.186.2.2 virtual const float∗ EVENT::TrackerHitPlane::getV ( ) const [pure virtual]

Direction of second measurement - given as (theta, phi). Defines spanning vector of measurement plane. Implemented in IMPL::TrackerHitPlaneImpl (p. 282). Referenced by SIO::SIOTrackerHitPlaneHandler::write(). The documentation for this class was generated from the following file:

• TrackerHitPlane.h

7.187 IMPL::TrackerHitPlaneImpl Class Reference

Implementation of the planar tracker hit.

#include <IMPL/TrackerHitPlaneImpl.h>

Inheritance diagram for IMPL::TrackerHitPlaneImpl:

Public Member Functions

• virtual ∼TrackerHitPlaneImpl ()

Destructor.

• virtual int id () const

Returns an object id for internal (debugging) use in LCIO.

• virtual int getCellID0 () const

Same name as in CalorimeterHit, even though there are no ’cells’ in this case.

• virtual int getCellID1 () const

Same name as in CalorimeterHit, even though there are no ’cells’ in this case 0 if information is not stored - check the flag word (bit RTHPBIT_ID1) of the collection.

• virtual const double ∗ getPosition () const

The hit position in [mm].

• virtual const float ∗ getU () const

Direction of first measurement - given as (theta, phi).

• virtual const float ∗ getV () const

Direction of second measurement - given as (theta, phi).

• virtual float getdU () const

Error along u.

• virtual float getdV () const

Error along v.

• virtual const EVENT::FloatVec & getCovMatrix () const

Covariance of the position (x,y,z)

• virtual float getdEdx () const

The dE/dx of the hit in [GeV/mm].

• virtual float getEDep () const

The deposited energy of the hit [GeV].

• virtual float getEDepError () const

The error measured on EDep [GeV].

• virtual float getTime () const

The time of the hit in [ns].

• virtual int getType () const

Type of hit.

• virtual int getQuality () const

The quality bit flag of the hit.

• virtual const EVENT::LCObjectVec & getRawHits () const

The raw data hits.

• virtual EVENT::LCObjectVec & rawHits ()

Use to manipulate the raw hits.

• void setCellID0 (int id0)

Sets the first cell id;.

• void setCellID1 (int id1)

Sets the second cell id; Only store if the flag word (bit RTHPBIT_ID1) of the collection is set.

• void setType (int type)

• void setPosition (const double pos[3])

• void setU (const float u[2])

• void setU (float theta, float phi)

• void setV (const float v[2])

• void setV (float theta, float phi)

• void setdU (float du)

• void setdV (float dv)

• void setEDep (float e)

• void setEDepError (float e)

• void setTime (float t)

• void setQuality (int quality)

• void setQualityBit (int bit, bool val=true)

Protected Attributes

• int _cellID0

• int _cellID1

• int _type

• double _pos [3]

• float _u [2]

• float _v [2]

• float _du

• float _dv

• float _EDep

• float _EDepError

• float _time

• int _quality

• EVENT::FloatVec _cov

• EVENT::LCObjectVec _rawHits

Additional Inherited Members

7.187.1 Detailed Description

Implementation of the planar tracker hit.

Author gaede

Version

Id:

7.187.2 Member Function Documentation

7.187.2.1 virtual int IMPL::TrackerHitPlaneImpl::getCellID1 ( ) const [inline], [virtual]

Same name as in CalorimeterHit, even though there are no ’cells’ in this case 0 if information is not stored - check the flag word (bit RTHPBIT_ID1) of the collection. Default is to store only cellid0. Implements EVENT::TrackerHitPlane (p. 278).

7.187.2.2 float IMPL::TrackerHitPlaneImpl::getdEdx ( ) const [virtual]

The dE/dx of the hit in [GeV/mm].

Deprecated See Also getEDep() (p. 280)

Implements EVENT::TrackerHit (p. 273). References UTIL::LCWarning::printWarning().

7.187.2.3 virtual const EVENT::LCObjectVec& IMPL::TrackerHitPlaneImpl::getRawHits ( ) const [inline],

        [virtual]

The raw data hits. Check getType() (p. 282) to get actual data type. Implements EVENT::TrackerHit (p. 273).

7.187.2.4 virtual float IMPL::TrackerHitPlaneImpl::getTime ( ) const [inline], [virtual]

The time of the hit in [ns]. Is this needed ? Implements EVENT::TrackerHit (p. 272).

7.187.2.5 virtual int IMPL::TrackerHitPlaneImpl::getType ( ) const [inline], [virtual]

Type of hit. Mapping of integer types to type names through collection parameters "TrackerHitTypeNames" and "TrackerHitType-Values". Implements EVENT::TrackerHit (p. 273).

7.187.2.6 virtual const float∗ IMPL::TrackerHitPlaneImpl::getU ( ) const [inline], [virtual]

Direction of first measurement - given as (theta, phi). Defines spanning vector of measurement plane. Implements EVENT::TrackerHitPlane (p. 279).

7.187.2.7 virtual const float∗ IMPL::TrackerHitPlaneImpl::getV ( ) const [inline], [virtual]

Direction of second measurement - given as (theta, phi). Defines spanning vector of measurement plane. Implements EVENT::TrackerHitPlane (p. 279).

7.187.2.8 void IMPL::TrackerHitPlaneImpl::setCellID1 ( int id1 )

Sets the second cell id; Only store if the flag word (bit RTHPBIT_ID1) of the collection is set. Default is to store only cellid0. The documentation for this class was generated from the following files:

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

• TrackerHitPlaneImpl.h

• TrackerHitPlaneImpl.cc

7.188 IOIMPL::TrackerHitPlaneIOImpl Class Reference

Adding stuff needed for io (friend declarations, etc.)

#include <IOIMPL/TrackerHitPlaneIOImpl.h>

Inheritance diagram for IOIMPL::TrackerHitPlaneIOImpl:

Friends

• class SIO::SIOTrackerHitPlaneHandler

Additional Inherited Members

7.188.1 Detailed Description

Adding stuff needed for io (friend declarations, etc.)

Author engels

Version 06-2011

The documentation for this class was generated from the following file:

• TrackerHitPlaneIOImpl.h

7.189 EVENT::TrackerHitZCylinder Class Reference

A tracker hit on a cylindrical surface that runs parallel to the z-axis - the hit is stored as: x,y,z, center(x,y), dRPhi, dz

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

where (x,y,z) is the position in global cartesian coordinates, center(x,y) is the center of the cylinder and Rphi and z are the measurement errors.

#include <pre-generated/EVENT/TrackerHitZCylinder.h>

Inheritance diagram for EVENT::TrackerHitZCylinder:

Public Types

• typedef TrackerHitZCylinder lcobject_type

Useful typedef for template programming with LCIO (p. 115).

Public Member Functions

• virtual ∼TrackerHitZCylinder ()

Destructor.

• virtual int getCellID0 () const =0

Same name as in CalorimeterHit (p. 49), even though there are no ’cells’ in this case.

• virtual int getCellID1 () const =0

Same name as in CalorimeterHit (p. 49), even though there are no ’cells’ in this case Optional, check/set flag(LCIO- ::RTHZBIT_ID1)==1.

• virtual const float ∗ getCenter () const =0

Radius of cylinder.

• virtual float getdRPhi () const =0

Measurement error along RPhi.

• virtual float getdZ () const =0

Measurement error along z.

Additional Inherited Members

7.189.1 Detailed Description

A tracker hit on a cylindrical surface that runs parallel to the z-axis - the hit is stored as: x,y,z, center(x,y), dRPhi, dz where (x,y,z) is the position in global cartesian coordinates, center(x,y) is the center of the cylinder and Rphi and z are the measurement errors. Note the radius of the cylinder is given by R=sqrt( dx^{∧}2c+ dy^{∧}2 ) and dx=x-x_c, dy=y-y_c.

Author gaede

Version $Id:$

7.189.2 Member Function Documentation

7.189.2.1 virtual const float∗ EVENT::TrackerHitZCylinder::getCenter ( ) const [pure virtual]

Radius of cylinder. Center of cylinder in RPhi-plane - given as (x,y). Implemented in IMPL::TrackerHitZCylinderImpl (p. 287). Referenced by SIO::SIOTrackerHitZCylinderHandler::write(). The documentation for this class was generated from the following file:

• TrackerHitZCylinder.h

7.190 IMPL::TrackerHitZCylinderImpl Class Reference

Implementation of the tracker hit on a cylindrical surface parallel to z.

#include <IMPL/TrackerHitZCylinderImpl.h>

Inheritance diagram for IMPL::TrackerHitZCylinderImpl:

Public Member Functions

• virtual ∼TrackerHitZCylinderImpl ()

Destructor.

• virtual int id () const

Returns an object id for internal (debugging) use in LCIO.

• virtual int getCellID0 () const

Same name as in CalorimeterHit, even though there are no ’cells’ in this case.

• virtual int getCellID1 () const

Same name as in CalorimeterHit, even though there are no ’cells’ in this case 0 if information is not stored - check the flag word (bit RTHZBIT_ID1) of the collection.

• virtual const double ∗ getPosition () const

The hit position in [mm].

• virtual const float ∗ getCenter () const

Radius of cylinder.

• virtual float getdRPhi () const

Measurement error along RPhi.

• virtual float getdZ () const

Measurement error along z.

• virtual const EVENT::FloatVec & getCovMatrix () const

Covariance of the position (x,y,z)

• virtual float getdEdx () const

The dE/dx of the hit in [GeV/mm].

• virtual float getEDep () const

The deposited energy of the hit [GeV].

• virtual float getEDepError () const

The error measured on EDep [GeV].

• virtual float getTime () const

The time of the hit in [ns].

• virtual int getType () const

Type of hit.

• virtual int getQuality () const

The quality bit flag of the hit.

• virtual const EVENT::LCObjectVec & getRawHits () const

The raw data hits.

• virtual EVENT::LCObjectVec & rawHits ()

Use to manipulate the raw hits.

• void setCellID0 (int id0)

Sets the first cell id;.

• void setCellID1 (int id1)

Sets the second cell id; Only store if the flag word (bit RTHZBIT_ID1) of the collection is set.

• void setType (int type)

• void setPosition (const double pos[3])

• void setCenter (const float c[2])

• void setCenter (float xc, float yc)

• void setdRPhi (float drphi)

• void setdZ (float dz)

• void setEDep (float e)

• void setEDepError (float e)

• void setTime (float t)

• void setQuality (int quality)

• void setQualityBit (int bit, bool val=true)

Protected Attributes

• int _cellID0

• int _cellID1

• int _type

• double _pos [3]

• float _center [2]

• float _drphi

• float _dz

• float _EDep

7.190 IMPL::TrackerHitZCylinderImpl Class Reference  287

• float _EDepError

• float _time

• int _quality

• EVENT::FloatVec _cov

• EVENT::LCObjectVec _rawHits

Additional Inherited Members 7.190.1 Detailed Description

Implementation of the tracker hit on a cylindrical surface parallel to z.

Author gaede

Version

Id:

7.190.2 Member Function Documentation

7.190.2.1 virtual int IMPL::TrackerHitZCylinderImpl::getCellID1 ( ) const [inline], [virtual] Same name as in CalorimeterHit, even though there are no ’cells’ in this case 0 if information is not stored - check the flag word (bit RTHZBIT_ID1) of the collection. Default is to store only cellid0. Implements EVENT::TrackerHitZCylinder (p. 284).

7.190.2.2 virtual const float∗ IMPL::TrackerHitZCylinderImpl::getCenter ( ) const [inline], [virtual] Radius of cylinder. Center of cylinder in RPhi-plane - given as (x,y). Implements EVENT::TrackerHitZCylinder (p. 285).

7.190.2.3 float IMPL::TrackerHitZCylinderImpl::getdEdx ( ) const [virtual] The dE/dx of the hit in [GeV/mm].

Deprecated See Also getEDep() (p. 286) Implements EVENT::TrackerHit (p. 273). References UTIL::LCWarning::printWarning().

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.190.2.4 virtual const EVENT::LCObjectVec& IMPL::TrackerHitZCylinderImpl::getRawHits ( ) const [inline],

        [virtual]

The raw data hits. Check getType() (p. 288) to get actual data type. Implements EVENT::TrackerHit (p. 273).

7.190.2.5 virtual float IMPL::TrackerHitZCylinderImpl::getTime ( ) const [inline], [virtual]

The time of the hit in [ns]. Is this needed ? Implements EVENT::TrackerHit (p. 272).

7.190.2.6 virtual int IMPL::TrackerHitZCylinderImpl::getType ( ) const [inline], [virtual]

Type of hit. Mapping of integer types to type names through collection parameters "TrackerHitTypeNames" and "TrackerHitType-Values". Implements EVENT::TrackerHit (p. 273).

7.190.2.7 void IMPL::TrackerHitZCylinderImpl::setCellID1 ( int id1 )

Sets the second cell id; Only store if the flag word (bit RTHZBIT_ID1) of the collection is set. Default is to store only cellid0. The documentation for this class was generated from the following files:

• TrackerHitZCylinderImpl.h

• TrackerHitZCylinderImpl.cc

7.191 IOIMPL::TrackerHitZCylinderIOImpl Class Reference

Adding stuff needed for io (friend declarations, etc.)

#include <IOIMPL/TrackerHitZCylinderIOImpl.h>

Inheritance diagram for IOIMPL::TrackerHitZCylinderIOImpl:

Friends

• class SIO::SIOTrackerHitZCylinderHandler

Additional Inherited Members 7.191.1 Detailed Description

Adding stuff needed for io (friend declarations, etc.)

Author engels

Version 2011-06 The documentation for this class was generated from the following file:

• TrackerHitZCylinderIOImpl.h

7.192 EVENT::TrackerPulse Class Reference

Tracker pulses as computed from TrackerData (p. 267) objects or as directly measured by a specific subdetector.

#include <pre-generated/EVENT/TrackerPulse.h>

Inheritance diagram for EVENT::TrackerPulse:

Public Types

• typedef TrackerPulse lcobject_type

Useful typedef for template programming with LCIO (p. 115).

Public Member Functions

• virtual ∼TrackerPulse ()

Destructor.

• virtual int getCellID0 () const =0

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Returns the first detector specific (geometrical) cell id.

• virtual int getCellID1 () const =0

Returns the second detector specific (geometrical) cell id.

• virtual float getTime () const =0

The time of the pulse - arbitrary units.

• virtual float getCharge () const =0

The integrated charge of the pulse - arbitrary units.

• virtual const FloatVec & getCovMatrix () const =0

Covariance matrix of the charge (c) and time (t) measurements.

• virtual int getQuality () const =0

The quality bit flag of the pulse - check/set collection parameters TrackerPulseQualityNames and TrackerPulseQuality-NamesValues.

• virtual TrackerData ∗ getTrackerData () const =0

Optionally the TrackerData (p. 267) that has been used to create the pulse can be stored with the pulse - NULL if none.

Additional Inherited Members

7.192.1 Detailed Description

Tracker pulses as computed from TrackerData (p. 267) objects or as directly measured by a specific subdetector. Typically TrackerHIts are created from TrackerPulses by some sort of clustering.

See Also TrackerRawData (p. 294)

TrackerData (p. 267)

Author gaede

Version

Id: TrackerPulse.aid,v 1.5 2010-11-11 09:35:27 engels Exp

7.192.2 Member Function Documentation

7.192.2.1 virtual int EVENT::TrackerPulse::getCellID1 ( ) const [pure virtual]

Returns the second detector specific (geometrical) cell id. Optional, check/set flag(LCIO::TRAWBIT_ID1)==1. Implemented in IMPL::TrackerPulseImpl (p. 292). Referenced by SIO::SIOTrackerPulseHandler::write().

7.192.2.2 virtual const FloatVec& EVENT::TrackerPulse::getCovMatrix ( ) const [pure virtual]

Covariance matrix of the charge (c) and time (t) measurements. Stored as lower triangle matrix, i.e. cov(c,c) , cov(t,c) , cov(t,t). Optional, check/set flag(LCIO::TRAWBIT_CM)==1. Implemented in IMPL::TrackerPulseImpl (p. 293). Referenced by SIO::SIOTrackerPulseHandler::write().

7.192.2.3 virtual TrackerData∗ EVENT::TrackerPulse::getTrackerData ( ) const [pure virtual]

Optionally the TrackerData (p. 267) that has been used to create the pulse can be stored with the pulse - NULL if none. Check the quality bits for reason why the spectrum has been stored for the pulse. Implemented in IMPL::TrackerPulseImpl (p. 293). Referenced by SIO::SIOTrackerPulseHandler::write(). The documentation for this class was generated from the following file:

• TrackerPulse.h

7.193 IMPL::TrackerPulseImpl Class Reference

Default implementation of TrackerPulse.

#include <IMPL/TrackerPulseImpl.h>

Inheritance diagram for IMPL::TrackerPulseImpl:

Public Member Functions

• TrackerPulseImpl ()

Default Constructor - initializes all data to 0’s.

• virtual ∼TrackerPulseImpl ()

Destructor.

• virtual int id () const

Returns an object id for internal (debugging) use in LCIO.

• virtual int getCellID0 () const

Returns the first detector specific (geometrical) cell id.

• virtual int getCellID1 () const

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Returns the second detector specific (geometrical) cell id.

• virtual float getTime () const

The time of the pulse.

• virtual float getCharge () const

The integrated charge of the pulse // FIXME: unit ?.

• virtual const EVENT::FloatVec & getCovMatrix () const

Covariance matrix of the charge (c) and time (t) measurements.

• virtual int getQuality () const

The quality bit flag of the pulse - use the defined constants for referring to the bits.

• virtual EVENT::TrackerData ∗ getTrackerData () const

Optionally the TrackerData that has been uesed to create the pulse can be stored with the pulse - NULL if none.

• void setCellID0 (int cellID0)

• void setCellID1 (int cellID1)

• void setTime (float time)

• void setCharge (float charge)

• void setCovMatrix (const float ∗cov)

• void setCovMatrix (const EVENT::FloatVec &)

• void setQuality (int quality)

• void setQualityBit (int bit, bool val=true)

• void setTrackerData (EVENT::TrackerData ∗corrData)

Protected Attributes

• int _cellID0

• int _cellID1

• float _time

• float _charge

• int _quality

• EVENT::FloatVec _cov

• EVENT::TrackerData ∗ _corrData

Additional Inherited Members

7.193.1 Detailed Description

Default implementation of TrackerPulse.

Author gaede

Version

Id: TrackerPulseImpl.h (p. ?? ),v 1.4 2010-05-12 14:50:00 engels Exp

7.193.2 Member Function Documentation

7.193.2.1 virtual int IMPL::TrackerPulseImpl::getCellID1 ( ) const [inline], [virtual]

Returns the second detector specific (geometrical) cell id.

Optional, check/set flag(LCIO::TRAWBIT_ID1)==1.

Implements EVENT::TrackerPulse (p. 290).

7.193.2.2 virtual const EVENT::FloatVec& IMPL::TrackerPulseImpl::getCovMatrix ( ) const [inline], [virtual]

Covariance matrix of the charge (c) and time (t) measurements. Stored as lower triangle matrix, i.e. cov(c,c) , cov(t,c) , cov(t,t). Optional, check/set flag(LCIO::TRAWBIT_CM)==1. Implements EVENT::TrackerPulse (p. 291).

7.193.2.3 virtual EVENT::TrackerData∗ IMPL::TrackerPulseImpl::getTrackerData ( ) const [inline], [virtual]

Optionally the TrackerData that has been uesed to create the pulse can be stored with the pulse - NULL if none. Check the quality bits for reason why the spectrum has been stored for the pulse. Implements EVENT::TrackerPulse (p. 291). The documentation for this class was generated from the following files:

• TrackerPulseImpl.h

• TrackerPulseImpl.cc

7.194 IOIMPL::TrackerPulseIOImpl Class Reference

Adding stuff needed for io (friend declarations, etc.)

#include <IOIMPL/TrackerPulseIOImpl.h>

Inheritance diagram for IOIMPL::TrackerPulseIOImpl:

Friends

• class SIO::SIOTrackerPulseHandler

Additional Inherited Members

7.194.1 Detailed Description

Adding stuff needed for io (friend declarations, etc.)

Author gaede Version Sep 11, 2003

The documentation for this class was generated from the following file:

• TrackerPulseIOImpl.h

7.195 EVENT::TrackerRawData Class Reference

Generic class for raw tracker data.

#include <pre-generated/EVENT/TrackerRawData.h>

Inheritance diagram for EVENT::TrackerRawData:

Public Types

• typedef TrackerRawData lcobject_type

Useful typedef for template programming with LCIO (p. 115).

Public Member Functions

• virtual ∼TrackerRawData ()

Destructor.

• virtual int getCellID0 () const =0

Returns the first detector specific (geometrical) cell id.

• virtual int getCellID1 () const =0

Returns the second detector specific (geometrical) cell id.

• virtual int getTime () const =0

Returns a time measurement associated with the adc values, e.g.

• virtual const ShortVec & getADCValues () const =0

The actual FADC spectrum.

Additional Inherited Members

7.195.1 Detailed Description

Generic class for raw tracker data. It can be used to store the full FADC spectrum as it comes out of the tracker DAQ or just one ore more single FADC readout values.

See Also TrackerData (p. 267)

TrackerPulse (p. 289)

Author gaede

Version

Id: TrackerRawData.aid,v 1.2 2006-03-24 13:25:53 gaede Exp

7.195.2 Member Function Documentation

7.195.2.1 virtual int EVENT::TrackerRawData::getCellID1 ( ) const [pure virtual]

Returns the second detector specific (geometrical) cell id. Optional, check/set flag(LCIO::TRAWBIT_ID1)==1. Implemented in IMPL::TrackerRawDataImpl (p. 297). Referenced by SIO::SIOTrackerRawDataHandler::write().

7.195.2.2 virtual int EVENT::TrackerRawData::getTime ( ) const [pure virtual]

Returns a time measurement associated with the adc values, e.g. the t0 of the spectrum for the TPC. Subdetector dependent. Implemented in IMPL::TrackerRawDataImpl (p. 296). Referenced by SIO::SIOTrackerRawDataHandler::write(). The documentation for this class was generated from the following file:

• TrackerRawData.h

7.196 IMPL::TrackerRawDataImpl Class Reference

Default implementation of TrackerRawData.

#include <IMPL/TrackerRawDataImpl.h>

Inheritance diagram for IMPL::TrackerRawDataImpl:

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Public Member Functions

• TrackerRawDataImpl ()

Default Constructor - initializes all data to 0’s.

• virtual ∼TrackerRawDataImpl ()

Destructor.

• virtual int id () const

Returns an object id for internal (debugging) use in LCIO.

• virtual int getCellID0 () const

Returns the first detector specific (geometrical) cell id.

• virtual int getCellID1 () const

Returns the second detector specific (geometrical) cell id.

• virtual int getTime () const

Returns the time.

• virtual const EVENT::ShortVec & getADCValues () const

The measured ADC values.

• void setCellID0 (int cellID0)

• void setCellID1 (int cellID1)

• void setTime (int time)

• void setADCValues (const EVENT::ShortVec &adc)

Set the ADC vector by copying the values.

• EVENT::ShortVec & adcValues ()

Allows direct access to the adc vector.

Protected Attributes

• int _cellID0

• int _cellID1

• int _channelID

• int _time

• EVENT::ShortVec _adc

Additional Inherited Members

7.196.1 Detailed Description

Default implementation of TrackerRawData.

Author gaede

Version

Id: TrackerRawDataImpl.h (p. ?? ),v 1.2 2006-08-04 16:52:47 gaede Exp

7.196.2 Member Function Documentation

7.196.2.1 ShortVec & IMPL::TrackerRawDataImpl::adcValues ( )

Allows direct access to the adc vector. Can be used e.g. to set the values: raw->adcValues() (p. 297).clear() ; raw->adcValues() (p. 297).resize( n ) ; raw->adcValues() (p. 297)[i] = 42 ;

7.196.2.2 virtual int IMPL::TrackerRawDataImpl::getCellID1 ( ) const [inline], [virtual]

Returns the second detector specific (geometrical) cell id. Optional, check/set flag(LCIO::TRAWBIT_ID1)==1. Implements EVENT::TrackerRawData (p. 295). The documentation for this class was generated from the following files:

• TrackerRawDataImpl.h

• TrackerRawDataImpl.cc

7.197 IOIMPL::TrackerRawDataIOImpl Class Reference

Adding stuff needed for io (friend declarations, etc.)

#include <IOIMPL/TrackerRawDataIOImpl.h>

Inheritance diagram for IOIMPL::TrackerRawDataIOImpl:

Friends

• class SIO::SIOTrackerRawDataHandler

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Additional Inherited Members

7.197.1 Detailed Description

Adding stuff needed for io (friend declarations, etc.)

Author gaede

Version Sep 11, 2003

The documentation for this class was generated from the following file:

• TrackerRawDataIOImpl.h

7.198 IMPL::TrackImpl Class Reference

Implementation of the LCIO track class.

#include <IMPL/TrackImpl.h>

Inheritance diagram for IMPL::TrackImpl:

Public Member Functions

• TrackImpl ()

Default constructor, initializes values to 0.

• TrackImpl (const TrackImpl &)

Copy constructor - creates shallow copy, i.e.

• const TrackImpl & operator= (const TrackImpl &o)

• virtual ∼TrackImpl ()

Destructor.

• virtual int id () const

Returns an object id for internal (debugging) use in LCIO.

• virtual int getType () const

Type of track, i.e.

• virtual float getD0 () const

Returns true if the corresponding bit in the type word is set.

• virtual float getPhi () const

Phi of the track at reference point.

• virtual float getOmega () const

Omega is the signed curvature of the track in [1/mm].

• virtual float getZ0 () const

Impact paramter of the track in (r-z).

• virtual float getTanLambda () const

Lambda is the dip angle of the track in r-z at the reference point.

• virtual const EVENT::FloatVec & getCovMatrix () const

Covariance matrix of the track parameters.

• virtual const float ∗ getReferencePoint () const

Reference point of the track parameters.

• virtual float getChi2 () const

True if the reference point is the point of closest approach.

• virtual int getNdf () const

Number of degrees of freedom of the track fit.

• virtual float getdEdx () const

dEdx of the track.

• virtual float getdEdxError () const

Error of dEdx.

• virtual float getRadiusOfInnermostHit () const

The radius of the innermost hit that has been used in the track fit.

• virtual const EVENT::IntVec & getSubdetectorHitNumbers () const

A vector that holds the number of hits in particular subdetectors.

• virtual const EVENT::TrackVec & getTracks () const

The tracks (as Track objects) that have been combined to this track.

• virtual const

EVENT::TrackStateVec & getTrackStates () const

Returns track states associtated with this track.

• virtual const EVENT::TrackState ∗ getClosestTrackState (float x, float y, float z) const

Returns track state closest to the given point.

• virtual const EVENT::TrackState ∗ getTrackState (int location) const

Returns track state for the given location - or NULL if not found.

• virtual const

EVENT::TrackerHitVec & getTrackerHits () const

Optionaly ( check/set flag(LCIO::TRBIT_HITS)==1) return the hits that have been used to create this track.

• virtual void setTypeBit (int index, bool val=true)

• virtual void setD0 (float d0)

• virtual void setPhi (float phi)

• virtual void setOmega (float omega)

• virtual void setZ0 (float z0)

• virtual void setTanLambda (float tanLambda)

• virtual void setCovMatrix (const float ∗cov)

• virtual void setCovMatrix (const EVENT::FloatVec &cov)

• virtual void setReferencePoint (const float ∗rPnt)

• virtual void setChi2 (float chi2)

• virtual void setNdf (int ndf)

• virtual void setdEdx (float dEdx)

• virtual void setdEdxError (float dEdxError)

• virtual void addTrack (EVENT::Track ∗trk)

• virtual void addTrackState (EVENT::TrackState ∗trkstate)

• virtual void addHit (EVENT::TrackerHit ∗hit)

• virtual EVENT::TrackStateVec & trackStates ()

• virtual void setRadiusOfInnermostHit (float r)

• virtual EVENT::IntVec & subdetectorHitNumbers ()

Allows modification of the subdetectorHitNumbers, e.g.

Protected Member Functions

• virtual void setType (int type)

Protected Attributes

• std::bitset< 32 > _type

• float _chi2

• int _ndf

• float _dEdx

• float _dEdxError

• float _radiusOfInnermostHit

• EVENT::IntVec _subdetectorHitNumbers

• EVENT::TrackVec _tracks

• EVENT::TrackerHitVec _hits

• EVENT::TrackStateVec _trackStates

Additional Inherited Members

7.198.1 Detailed Description

Implementation of the LCIO track class.

The helix parameterization follows that of the L3 experiment at LEP. A detailed description of the track parameters is given in the LC-Note LC-DET-2006-004 "Track Parameters in LCIO" (pdf).

See Also Track

Author gaede

Version

Id: TrackImpl.h (p. ?? ),v 1.15 2007-09-18 09:47:10 gaede Exp

7.198.2 Constructor & Destructor Documentation

7.198.2.1 IMPL::TrackImpl::TrackImpl ( const TrackImpl & o )

Copy constructor - creates shallow copy, i.e.

all data members are copied but pointers to other LCObjects i.e. TrackerHits and Tracks are preserved.

7.198 IMPL::TrackImpl Class Reference  301

7.198.3 Member Function Documentation

7.198.3.1 float IMPL::TrackImpl::getChi2 ( ) const [virtual] True if the reference point is the point of closest approach. Chi^{∧}2 of the track fit. Implements EVENT::Track (p. 264).

7.198.3.2 const TrackState ∗ IMPL::TrackImpl::getClosestTrackState ( float x, float y, float z ) const [virtual] Returns track state closest to the given point.

See Also TrackState. Implements EVENT::Track (p. 264).

7.198.3.3 const FloatVec & IMPL::TrackImpl::getCovMatrix ( ) const [virtual] Covariance matrix of the track parameters. Stored as lower triangle matrix where the order of parameters is: d0, phi, omega, z0, tan(lambda). So we have cov(d0,d0), cov( phi, d0 ), cov( phi, phi), ... Information is stored in the first TrackState of this Track,

See Also TrackState. Implements EVENT::Track (p. 265). References IMPL::TrackStateImpl::getCovMatrix().

7.198.3.4 float IMPL::TrackImpl::getD0 ( ) const [virtual] Returns true if the corresponding bit in the type word is set. Impact paramter of the track in (r-phi). Information is stored in the first TrackState of this Track,

See Also TrackState. Implements EVENT::Track (p. 265).

7.198.3.5 float IMPL::TrackImpl::getOmega ( ) const [virtual] Omega is the signed curvature of the track in [1/mm]. The sign is that of the particle’s charge. Information is stored in the first TrackState of this Track,

See Also TrackState. Implements EVENT::Track (p. 265).

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.198.3.6 float IMPL::TrackImpl::getPhi ( ) const [virtual]

Phi of the track at reference point. Information is stored in the first TrackState of this Track,

See Also TrackState.

Implements EVENT::Track (p. 265).

7.198.3.7 float IMPL::TrackImpl::getRadiusOfInnermostHit ( ) const [virtual]

The radius of the innermost hit that has been used in the track fit. Helps to detect V0 tracks with small impact paramters or haevy mesons. To be used as convenient information or if hits are not stored in the data set, e.g. DST or FastMC. Implements EVENT::Track (p. 266).

7.198.3.8 const float ∗ IMPL::TrackImpl::getReferencePoint ( ) const [virtual]

Reference point of the track parameters. The default for the reference point is the point of closest approach. Information is stored in the first TrackState of this Track,

See Also TrackState.

Implements EVENT::Track (p. 266). References IMPL::TrackStateImpl::getReferencePoint().

7.198.3.9 const IntVec & IMPL::TrackImpl::getSubdetectorHitNumbers ( ) const [virtual]

A vector that holds the number of hits in particular subdetectors. The mapping of indices to subdetectors is implementation dependent. To be used as convenient information or if hits are not stored in the data set, e.g. DST or FastMC. TODO: Provide way to store mapping in event/run header. Implements EVENT::Track (p. 266).

7.198.3.10 float IMPL::TrackImpl::getTanLambda ( ) const [virtual]

Lambda is the dip angle of the track in r-z at the reference point. Information is stored in the first TrackState of this Track,

See Also TrackState.

Implements EVENT::Track (p. 266).

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.198.3.11 const TrackState ∗ IMPL::TrackImpl::getTrackState ( int location ) const [virtual]

Returns track state for the given location - or NULL if not found.

See Also TrackState. location can be set to: AtIP, AtFirstHit, AtLastHit, AtCalorimeter, AtVertex, AtOther

Implements EVENT::Track (p. 266).

7.198.3.12 const TrackStateVec & IMPL::TrackImpl::getTrackStates ( ) const [virtual]

Returns track states associtated with this track.

See Also TrackState.

Implements EVENT::Track (p. 267).

7.198.3.13 int IMPL::TrackImpl::getType ( ) const [virtual]

Type of track, i.e. the subdetector(s) used to derive the track paramters, e.g. "TPC", "COMB", etc. In order to save disc space working groups should try and define three to four letter acronyms for their tracking detectors.Flagword that defines the type of track. Bits 0-15 can be used to denote the subdetectors that have contributed hits used in the track fit. The definition of the hits has to be done elsewhere, e.g. in the run header. Before LCIO 2.0 bit 31 was used to encode isReferencePointPCA (now deprecated). Implements EVENT::Track (p. 267).

7.198.3.14 float IMPL::TrackImpl::getZ0 ( ) const [virtual]

Impact paramter of the track in (r-z). Information is stored in the first TrackState of this Track,

See Also TrackState.

Implements EVENT::Track (p. 267).

7.198.3.15 IntVec & IMPL::TrackImpl::subdetectorHitNumbers ( ) [virtual]

Allows modification of the subdetectorHitNumbers, e.g. track->subdetectorHitNumbers() (p. 303).resize(5) ; track->subdetectorHitNumbers() (p. 303)[4] = 42 ; Referenced by SIO::SIOTrackHandler::read(). The documentation for this class was generated from the following files:

• TrackImpl.h

• TrackImpl.cc

7.199 IOIMPL::TrackIOImpl Class Reference

Adding stuff needed for io (friend declarations, etc.)

#include <IOIMPL/TrackIOImpl.h>

Inheritance diagram for IOIMPL::TrackIOImpl:

Friends

• class SIO::SIOTrackHandler

Additional Inherited Members

7.199.1 Detailed Description

Adding stuff needed for io (friend declarations, etc.)

Author gaede

Version Mar 15, 2004

The documentation for this class was generated from the following file:

• TrackIOImpl.h

7.200 EVENT::TrackState Class Reference

The LCIO (p. 115) TrackState (p. 304) class.

#include <pre-generated/EVENT/TrackState.h>

Inheritance diagram for EVENT::TrackState:

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Public Types

• typedef TrackState lcobject_type

Useful typedef for template programming with LCIO (p. 115).

Public Member Functions

• virtual ∼TrackState ()

Destructor.

• virtual int getLocation () const =0

The location of the track state.

• virtual float getD0 () const =0

Impact paramter of the track in (r-phi).

• virtual float getPhi () const =0

Phi of the track at the reference point.

• virtual float getOmega () const =0

Omega is the signed curvature of the track in [1/mm].

• virtual float getZ0 () const =0

Impact paramter of the track in (r-z).

• virtual float getTanLambda () const =0

Lambda is the dip angle of the track in r-z at the reference point.

• virtual const FloatVec & getCovMatrix () const =0

Covariance matrix of the track parameters.

• virtual const float ∗ getReferencePoint () const =0

Reference point of the track parameters, e.g.

Static Public Attributes

• static const int AtOther = 0

• static const int AtIP = 1

• static const int AtFirstHit = 2

• static const int AtLastHit = 3

• static const int AtCalorimeter = 4

• static const int AtVertex = 5

• static const int LastLocation = AtVertex

Additional Inherited Members

7.200.1 Detailed Description

The LCIO (p. 115) TrackState (p. 304) class.

Author gaede, engels

Version $Id:$

7.200.2 Member Function Documentation

7.200.2.1 virtual const FloatVec& EVENT::TrackState::getCovMatrix ( ) const [pure virtual]

Covariance matrix of the track parameters. Stored as lower triangle matrix where the order of parameters is: d0, phi, omega, z0, tan(lambda). So we have cov(d0,d0), cov( phi, d0 ), cov( phi, phi), ... Implemented in IMPL::TrackStateImpl (p. 309).

7.200.2.2 virtual int EVENT::TrackState::getLocation ( ) const [pure virtual]

The location of the track state. Location can be set to: TrackState::AtIP, TrackState::AtFirstHit, TrackState::AtLastHit, TrackState::AtCalorimeter, TrackState::AtVertex, TrackState::AtOther Implemented in IMPL::TrackStateImpl (p. 309). Referenced by IMPL::TrackStateImpl::TrackStateImpl().

7.200.2.3 virtual float EVENT::TrackState::getOmega ( ) const [pure virtual]

Omega is the signed curvature of the track in [1/mm]. The sign is that of the particle’s charge. Implemented in IMPL::TrackStateImpl (p. 309).

7.200.2.4 virtual float EVENT::TrackState::getPhi ( ) const [pure virtual]

Phi of the track at the reference point.

See Also getReferencePoint (p. 306)

Implemented in IMPL::TrackStateImpl (p. 308).

7.200.2.5 virtual const float∗ EVENT::TrackState::getReferencePoint ( ) const [pure virtual]

Reference point of the track parameters, e.g. the origin at the IP, or the position of the first/last hits or the entry point into the calorimeter.

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Implemented in IMPL::TrackStateImpl (p. 309).

Referenced by IMPL::TrackStateImpl::TrackStateImpl().

7.200.2.6 virtual float EVENT::TrackState::getTanLambda ( ) const [pure virtual]

Lambda is the dip angle of the track in r-z at the reference point.

See Also getReferencePoint (p. 306)

Implemented in IMPL::TrackStateImpl (p. 308).

The documentation for this class was generated from the following files:

• TrackState.h

• TrackStateImpl.cc

7.201 IMPL::TrackStateImpl Class Reference

Implementation of the LCIO TrackState class.

#include <IMPL/TrackStateImpl.h>

Inheritance diagram for IMPL::TrackStateImpl:

Public Member Functions

• TrackStateImpl ()

Default constructor, initializes values to 0.

• TrackStateImpl (int location, float d0, float phi, float omega, float z0, float tanLambda, const float ∗covMatrix, const float ∗reference)

• TrackStateImpl (int location, float d0, float phi, float omega, float z0, float tanLambda, const EVENT::Float-Vec &covMatrix, const float ∗reference)

• TrackStateImpl (const EVENT::TrackState &p)

Copy constructor which takes as an argument an EVENT::TrackState (p. 304) reference.

• virtual ∼TrackStateImpl ()

Destructor.

• virtual int id () const

Returns an object id for internal (debugging) use in LCIO.

• virtual int getLocation () const

The location of the track state.

• virtual float getD0 () const

Impact paramter of the track in (r-phi).

• virtual float getPhi () const

Phi of the track at reference point.

• virtual float getOmega () const

Omega is the signed curvature of the track in [1/mm].

• virtual float getZ0 () const

Impact paramter of the track in (r-z).

• virtual float getTanLambda () const

Lambda is the dip angle of the track in r-z at the reference point.

• virtual const EVENT::FloatVec & getCovMatrix () const

Covariance matrix of the track parameters.

• virtual const float ∗ getReferencePoint () const

Reference point of the track parameters, e.g.

• virtual void setLocation (int location)

• virtual void setD0 (float d0)

• virtual void setPhi (float phi)

• virtual void setOmega (float omega)

• virtual void setZ0 (float z0)

• virtual void setTanLambda (float tanLambda)

• virtual void setCovMatrix (const float ∗cov)

• virtual void setCovMatrix (const EVENT::FloatVec &cov)

• virtual void setReferencePoint (const float ∗rPnt)

Protected Attributes

• int _location

• float _d0

• float _phi

• float _omega

• float _z0

• float _tanLambda

• EVENT::FloatVec _covMatrix

• float _reference [TRKSTATENREFSIZE]

Additional Inherited Members

7.201.1 Detailed Description

Implementation of the LCIO TrackState class.

See Also TrackState

Author gaede, engels

Version $Id:$

7.201.2 Member Function Documentation

7.201.2.1 const FloatVec & IMPL::TrackStateImpl::getCovMatrix ( ) const [virtual]

Covariance matrix of the track parameters. Stored as lower triangle matrix where the order of parameters is: d0, phi, omega, z0, tan(lambda). So we have cov(d0,d0), cov( phi, d0 ), cov( phi, phi), ... Implements EVENT::TrackState (p. 306). Referenced by IMPL::TrackImpl::getCovMatrix().

7.201.2.2 int IMPL::TrackStateImpl::getLocation ( ) const [virtual]

The location of the track state. Location can be set to: AtIP, AtFirstHit, AtLastHit, AtCalorimeter, AtVertex, AtOther Implements EVENT::TrackState (p. 306).

7.201.2.3 float IMPL::TrackStateImpl::getOmega ( ) const [virtual]

Omega is the signed curvature of the track in [1/mm]. The sign is that of the particle’s charge. Implements EVENT::TrackState (p. 306).

7.201.2.4 const float ∗ IMPL::TrackStateImpl::getReferencePoint ( ) const [virtual]

Reference point of the track parameters, e.g. the origin at the IP, or the position of the first/last hits or the entry point into the calorimeter. Implements EVENT::TrackState (p. 306). Referenced by IMPL::TrackImpl::getReferencePoint(). The documentation for this class was generated from the following files:

• TrackStateImpl.h

• TrackStateImpl.cc

7.202 IOIMPL::TrackStateIOImpl Class Reference

Adding stuff needed for io (friend declarations, etc.)

#include <IOIMPL/TrackStateIOImpl.h>

Inheritance diagram for IOIMPL::TrackStateIOImpl:

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Friends

• class SIO::SIOTrackHandler

Additional Inherited Members 7.202.1 Detailed Description

Adding stuff needed for io (friend declarations, etc.)

Author gaede, engels

Version Mar 15, 2004 The documentation for this class was generated from the following file:

• TrackStateIOImpl.h

7.203 TrkCluLink Struct Reference

Inheritance diagram for TrkCluLink:

The documentation for this struct was generated from the following file:

• lcrtrelation.cc

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.204 UTIL::UnknownAlgorithm Class Reference

Exception for unknown algorithms.

#include <UTIL/PIDHandler.h>

Inheritance diagram for UTIL::UnknownAlgorithm:

Public Member Functions

• UnknownAlgorithm (std::string text)

Additional Inherited Members

7.204.1 Detailed Description

Exception for unknown algorithms.

The documentation for this class was generated from the following file:

• PIDHandler.h

7.205 SIO::SIORecords::Unpack Class Reference

Helper class that sets the unpack flag of known SIO (p. 30) records as specified by the constructor.

#include <SIO/LCSIO.h>

Public Member Functions

• Unpack (unsigned recordFlag)

Static Public Attributes

• static const unsigned Event = 0x0001 << SIORecords::Event

• static const unsigned Header = 0x0001 << SIORecords::Header

• static const unsigned Run = 0x0001 << SIORecords::Run

• static const unsigned Access = 0x0001 << SIORecords::Access

• static const unsigned Index = 0x0001 << SIORecords::Index

• static const unsigned All = 0xFFFFFFFF

Protected Attributes

• bool _flags [NumberOfRecords]

7.205.1 Detailed Description

Helper class that sets the unpack flag of known SIO (p. 30) records as specified by the constructor. After an object of this class goes out of scope the original setting is restored. This is needed to have multiple read streams (event overlay). The documentation for this class was generated from the following files:

• LCSIO.h

• LCSIO.cc

7.206 UserClass Struct Reference

Public Attributes

• int someInt

• float aFloat

The documentation for this struct was generated from the following file:

• lcrtrelation.cc

7.207 UserStruct Struct Reference

Inheritance diagram for UserStruct:

The documentation for this struct was generated from the following file:

• test_RTextension.cc

7.208 EVENT::Vertex Class Reference

The Vertex (p. 312) class for LCIO (p. 115).

#include <pre-generated/EVENT/Vertex.h>

Inheritance diagram for EVENT::Vertex:

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

7.208 EVENT::Vertex Class Reference  313

Public Types

• typedef Vertex lcobject_type

Useful typedef for template programming with LCIO (p. 115).

Public Member Functions

• virtual ∼Vertex ()

Destructor.

• virtual bool isPrimary () const =0

Checks if the Vertex (p. 312) is the primary vertex of the event.

• virtual const std::string & getAlgorithmType () const =0

Type code for the algorithm that has been used to create the vertex - check/set the collection parameters Algorithm-Name and AlgorithmType.

• virtual float getChi2 () const =0

Chi squared of the vertex fit.

• virtual float getProbability () const =0

Probability of the vertex fit.

• virtual const float ∗ getPosition () const =0

Position of the vertex.

• virtual const FloatVec & getCovMatrix () const =0

Covariance matrix of the position (stored as lower triangle matrix, i.e.

• virtual const FloatVec & getParameters () const =0

Additional parameters related to this vertex - check/set the collection parameter "VertexParameterNames" for the parameters’ meaning.

• virtual ReconstructedParticle ∗ getAssociatedParticle () const =0

Returns Reconstructed Particle associated to the Vertex (p. 312).

Additional Inherited Members

7.208.1 Detailed Description

The Vertex (p. 312) class for LCIO (p. 115). UNDER DEVELOPMENT!

see: here

UNDER DEVELOPMENT!

Author gaede, engels

Version

Id: Vertex.aid,v 1.3 2006-09-21 06:10:36 gaede Exp

7.208.2 Member Function Documentation

7.208.2.1 virtual const FloatVec& EVENT::Vertex::getCovMatrix ( ) const [pure virtual]

Covariance matrix of the position (stored as lower triangle matrix, i.e. cov(xx),cov(y,x),cov(y,y) ). Implemented in IMPL::VertexImpl (p. 316). Referenced by SIO::SIOVertexHandler::write().

7.208.2.2 virtual bool EVENT::Vertex::isPrimary ( ) const [pure virtual]

Checks if the Vertex (p. 312) is the primary vertex of the event. Only one primary vertex per event is allowed Implemented in IMPL::VertexImpl (p. 316). Referenced by SIO::SIOVertexHandler::write(). The documentation for this class was generated from the following file:

• Vertex.h

7.209 IMPL::VertexImpl Class Reference

Implementation of Vertex Class.

#include <IMPL/VertexImpl.h>

Inheritance diagram for IMPL::VertexImpl:

Public Member Functions

• VertexImpl ()

Default constructor, initializes values to 0.

• virtual int id () const

Returns an object id for internal (debugging) use in LCIO.

• virtual bool isPrimary () const

Checks if the Vertex is the primary vertex of the event.

• virtual const std::string & getAlgorithmType () const

Type code for the algorithm that has been used to create the vertex - check/set the collection parameters Algorithm-Name and AlgorithmType.

• virtual float getChi2 () const

Chi squared of the vertex fit.

• virtual float getProbability () const

Probability of the vertex fit.

• virtual const float ∗ getPosition () const

Position of the vertex.

• virtual const EVENT::FloatVec & getCovMatrix () const

Covariance matrix of the position (stored as lower triangle matrix, i.e.

• virtual const EVENT::FloatVec & getParameters () const

Additional parameters related to this vertex - check/set the collection parameter "VertexParameterNames" for the parameters’ meaning.

• virtual

EVENT::ReconstructedParticle ∗ getAssociatedParticle () const

Returns Reconstructed Particle associated to the Vertex.

• void setPrimary (bool primary)

• void setAlgorithmType (std::string type)

• void setChi2 (float chi2)

• void setProbability (float probability)

• void setPosition (const float vpos[3])

• void setPosition (float px, float py, float pz)

• void setCovMatrix (const float ∗cov)

• void setCovMatrix (const EVENT::FloatVec &)

• void setAssociatedParticle (EVENT::ReconstructedParticle ∗aP)

• void addParameter (float p)

Protected Attributes

• int _primary

• std::string _type

• float _chi2

• float _probability

• float _vpos [3]

• EVENT::FloatVec _cov

• EVENT::FloatVec _par

• EVENT::ReconstructedParticle ∗ _aParticle

Additional Inherited Members

7.209.1 Detailed Description

Implementation of Vertex Class.

See Also Vertex

Author gaede, engels

Version Aug 09, 2006

7.209.2 Member Function Documentation

7.209.2.1 const EVENT::FloatVec & IMPL::VertexImpl::getCovMatrix ( ) const [virtual]

Covariance matrix of the position (stored as lower triangle matrix, i.e. cov(xx),cov(y,x),cov(y,y) ). Implements EVENT::Vertex (p. 314).

7.209.2.2 bool IMPL::VertexImpl::isPrimary ( ) const [virtual]

Checks if the Vertex is the primary vertex of the event. Only one primary vertex per event is allowed Implements EVENT::Vertex (p. 314). The documentation for this class was generated from the following files:

• VertexImpl.h

• VertexImpl.cc

7.210 IOIMPL::VertexIOImpl Class Reference

Adding stuff needed for io (friend declarations, etc.)

#include <IOIMPL/VertexIOImpl.h>

Inheritance diagram for IOIMPL::VertexIOImpl:

Generated on Fri Dec 2 2016 11:44:41 for LCIO by Doxygen

Friends

• class SIO::SIOVertexHandler

Additional Inherited Members 7.210.1 Detailed Description

Adding stuff needed for io (friend declarations, etc.)

Author gaede, engels

Version Aug 15, 2006 The documentation for this class was generated from the following file:

• VertexIOImpl.h